A Bibliography of Publications in International Journal of Computers and Applications

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

21 February 2025 Version 2.01

Title word cross-reference

(1, N) [RS16]. (K, N) [Bai10]. 1 [MM00, NPY07]. 2 [JJY^{+}10, MM00, MMM13, WLZZ12, ZNLA08]. 2 × 2 [SYJ11]. 3 [HC11, Man15, MKMA07, Wan06, WLZZ12, Wer09, Yut18, ZMY^{+}00]. 5 [BKK17]. 8 [APSK14]. _{2} [EMN17]. _{3} [SYJ11]. GF(2^{m}) [LC07]. K [AWA20, JWC^{+}09, XZ18, AAH25, JN17]. l [VLK19]. N [Kat07]. π [IIDI01]. ^{3} \sqrt{2} [LHJ07].

-arithmetic [RS16]. -Bit [APSK14]. -Calculus [IIDI01]. -D [MM00, MMM13, NPY07, WLZZ12, Yut18, ZNLA08, ZMY^{+}00]. -Dimensional [Kat07]. -DOF [BKK17]. -means [JN17, AWA20, XZ18]. -nearest [AAH25]. -Nearest-Neighbour [JWC^{+}09]. -Subdivision [LHJ07].

1.1 [KJ03]. 19 [Pat20, Pat22a]. 1C

[PKM18].

2 [CJNY^{+}07, GG05]. 2000 [Ano00a]. 2D [GBBZ19, SAM23].

3D [AA18, Bay18, GBBZ19, Rui19].

4 [CJNY^{+}07]. 4-bit [DG18].

5G [ARAAMD24].

6-generator [SSMM16]. 6TiSCH [PC22].

Abjad [HEZ05]. abnormal [Ano21c, FCW^{+}20]. Abstract [SHP25]. accelerate [Man15]. Accelerating [EHA16]. accelerator [SHDY10]. accelerometer [WT15]. Access [Amm07, GM01, Pat20, RDEDAE09, SGD07, YCYL04, ZYS13, ZB08, ASM21, BYBM21, DU19, DM20, HDS17, HZ19, KPMR20, PCK21, SPKS24]. Accuracy [JWC^{+}09, TD23]. Accurate [DS10, HZ19]. achieve [DD16, MM24]. achievement [Tia18]. ACO [SSMM16, SK17, VS24]. Acoustic [KO00]. acquaintance [JMMS18]. acquisition [WDW17, ZP19]. Across [GYQC14]. action [Yao17]. Active [Lee07, ZKS05, FD23]. Activities [CLLC10]. Activity [FR00, KLRH13, Wan11, AUHH^{+}18, DZ23, SLSS24, SMAM25]. Ad [AJS08, CL11b, GT10, HL10, JR12, LH12, NS10b, SN12, WTLW07, BYBM21, Ogu18, VR19, WZ05]. ad-hoc [BYBM21, Ogu18]. Adams [FO09]. Adaptation [Alm11, Din11, LM02, RABB08, ZJH08, MSH21, SN23]. adaptation-based [MSH21]. Adapter [JYHS01]. Adapting [TG24]. Adaptive [ASM21, AEA19, CS22, DP17, HS24, HP04, HJR08, IAD10, JS19, Jin10, KSB19, LHK03, LS00, MK11b, PC13, PCS08, SAD09, ZXF07, ZKNL10, Bai21, BYBM21, Day17, DRF18, HPC18, JXG25, KC20, KKKV19, KR19b, Man15, NM16, SS20a, Yut18]. Adaptively [LHJ07]. addictions [ABK24]. Address [DLL05]. Addressing [BB23a]. adenocarcinoma [MSG18, PBS19]. Adhoc [KM09a]. Adjustable [LZXB24]. adjusted [CLTY17]. Admissible [WF06]. Admission [ACA07]. adopting [SJZ^{+}18]. Ads [dMRF^{+}14]. Advanced [BK24, SK18, KKS18, YWH19]. advancements [Pat18]. adversarial [AASL23, SS21]. AES [BM18]. affective [DZ18]. Affinity [WZTH18, AJP18, WWHW18]. Affordable [MBS01]. Against [CEJ12, DVT20, VLK19]. age [DK23, YB18]. age-related [DK23]. Agent [AEDA03, HAA05, IIDI01, Kim01, LL03, MS03, WLC05]. Agent-Based [MS03, WLC05]. Agents [AD00, LL03, PS12, SV03, VNR19]. Agentservice [VGPB09]. Aggregate [PBA10]. Aggregation [Dua07, CDR24, Dua23, Ogu19, VV19]. aggression [CRAV24]. Agile [AAB19]. agility [KKKV19]. Agreement [Pop05, CKC15]. agricultural [Ano21d, ZSWF20]. AI [AFLdS25, KK24, Sen24]. AI-driven [Sen24]. aid [KO00]. aided [ARAAMD24, SR20]. Aids [BSB10, BH21]. AIIMLSDSPS [NN17]. aircraft [BTZA18]. alarming [RWS17]. alert [Ali24]. Algebra [Kar12, PG04]. Algebraic [EA05]. Algorithm [AKAJ23, ACA07, AAAZ00, ABR10, AJS08, ASA05, AEDS07, BY03, CT13, CX03, CC06, CL07, CGL12, Dar09, DNP19, DDL08, GYQC14, JR12, JS08, JH00, Kat07, KB09, LC07, LS10, LV06, Loo04, Nom12, PTLH08, PKW^{+}11, PR19, PYMD09, PS06, SVM^{+}00, SSP07, SAAAOH10, SWL07, SS12, TBS00, VR08, WLL10, YLY07, YYD10, ZMY^{+}00, AAH25, AKAJ24, AM16, AASA22, Ano21c, Ano21i, Ano22b, BFB^{+}20, Bay18, BBL23, BP17, CXJ24, CGI18, Dai19, DS19a, DWT23, DRF18, DVT20, EK23, Fan18, FMP^{+}19, FCW^{+}20, Fu25, GBBZ19, GGB23, GAS23, HLW^{+}18, HZW^{+}19, HNB24, Hua18b, IAT20, JNS18, JS16, JN17, KS17, LL18, Li20a, Li20b, LC20, Mal24a, MGT19, MR19b, MMS20, NN17, NN21, NX18, NA19, NZT^{+}20, PR17, PV18, PW19, RR17, RDYGE18, SMSI23, SS18, SG16, SPC19]. algorithm [TYBwY17, TLC18, TW19, TA23, WZTH18, WWHW18, Wan19, fWfH19, Wu18, WG19, XZ18, Yan19, Yin18b, YxLmLx17, Yu19, Zha17b, ZP19, Zhe19, ZLHW18, ZZ20]. algorithm-based [AASA22]. Algorithms [AZP03, CKC09, CHL01, CGH11, DJQS04, HH07, JJY^{+}10, LE01, MM00, SMF12, WCC06, ABK24, BKXK19, Li19, Man15, RN17, RH23, SS20c, SV23]. alignment [MSH21, PW19, ZX24]. Allan [LLH08]. Allocation [CL11a, CL11b, Dua07, FA06, LV06, MKMA07, SD03, SWL07, VTS06, AKB^{+}19, AZK23, JMMS18, KBL21, RVS24]. Allocations [ANB09]. alone [SD17]. Alternative [KMR13]. Alzheimer [HHU24]. Ambient [JPJ^{+}19]. AMBTC [DS11]. AMPL [GB20]. AMVRGO [CS22]. Analyses [INyN11]. Analysis [AAY02, BW07, BBAR08, BD11, CS03, FJK01, GS13, GS11, HL10, JV12, Jij15, Kur07, LB04b, Loo04, Mag05, MBS01, MK11a, MMHJ10, SML^{+}13, SEB06, SBS12, SWLH08, Tia18, VHL11, Vee05, WMB02, Xia18, Yan17, AR19, BK24, BS19, BS25, BR17, BP17, BM17, CDE20, CSFS17, CU19, CPKT18, DR15, DS19b, DG19, Don18, DWT23, GR19, GGMM23, HDS17, Hou18, KKXB19, KD19, Li17, Li20a, Lin17, LHZ18, MS19, MPS^{+}23, MMH18, NX18, OMA24, PKD18, PN20, RMG19, RK15, Ren18, SD17, SAM23, SS20b, VP17, VP19a, Wan19, YMRLJ17, Yin18a, YYM17, Zha17a]. analytical [AEEO23, GMD16]. Analytics [RWL19, GT19]. analytics-based [GT19]. Analyzing [KK18, ACM20]. anchor [WGE17]. AND-XOR [PC11]. Android [Ano22b, Li20b]. ANFIS [CBXh18]. angle [QD18]. ANN [ABEAA10]. Annealing [AAA04, YYD10]. anomalies [DND25]. anomaly [BJBA24, IQ22, MMS18, Men18, MSYG18, TAUK25, YWDC24]. Anonymity [NN17]. Anonymity-based [NN17]. anonymous [KKA20]. Anser [ZXF07]. answering [BJ21, SHI24, SJTl25]. ant [MR23]. antenna [Gho17]. Anti [SS21]. Anti-malware [SS21]. API [SS20c]. Applicable [KM06]. Application [BL01, CXJ24, DG09, Fu25, JS08, JSZ09, KB10, MMH01, OE12, RCJJ06, Sam10, SIT11, Yao17, YWH19, ZFV04, ZXT^{+}06, ZMY^{+}00, ZLC06, Bai21, BR17, CT15, DG19, Guo19, KB22, MSG18, RR17, WLHZ18]. Applications [AP01, BA03, FR00, HP04, LH13, Lic06, LE01, MK11a, OA01, SMF12, SR06, Shi12a, SS19b, YP05, Ali24, Dar20, DS19b, FM18, JCYZ18, KSD23, KK18, KKR25, KK24, MT24, MdCdAdS24, MV18, RPZW17, SSS21a, SA24, SMAM25, SK18, ST16, WWHW18]. applied [AUHH^{+}18]. Applying [LM02, CDE20]. APPR [PYMD09]. Approach [Abd11, ACA07, AMJ12, Che04, Chi08, FWQ00, FRP07, FMI11, GA11, KS14, KNS07, KLRH13, KM09a, Lap05, LWH12, MDY14, MB10, Nya10, PMR^{+}03, PT14, PKDP12, Rus06, SVM^{+}00, SM10, SOC07, Sha11, SHS^{+}02, UU03, YP05, APS^{+}21, AYI23, AKB^{+}19, Bad19, BJBA24, BA16, BS19, BA23, BS17, BTZA18, DR15, DSBR19, DS19a, DZ23, ECBR20, GKMR19, GS19, GS22, GMD16, GSP19, HDS17, JS19, JMMS18, JSYT21, KBD21, KC19, KR18, KR19b, Mal24a, MB19, MMS20, NB24, NZT^{+}20, PS19, PR18, PK20, RMF19, SK17, SSS20, SK24, TS15, WDW17, WGE17, ZEZ17, ZL18]. Approaches [GG04, KM09c, EhMZ24, KKXB19, MV18, MS16, MMH19, YNA23]. Approximation [YWZJ06]. apps [Dud19]. APSSNMP [WBV04]. Aquila [INR24]. Arabic [AAM^{+}18, AAD06, ECBR20, HEZ05, KAKF11, KB09, SS05]. ARCH [OW11]. archetype [NM16]. Architectural [KHVW04, Kur07, SAN24]. Architecture [AA12, AP01, HY00, JSHB10, KB09, LS03, MMM13, MK11a, MR12, MK11b, SAD09, WPJZ08, WLC05, Wu08, ZNLA08, ASM21, Aa20, BTAB21, DE17, Dar20, HT19, PR17, PV18, SA24, ST16, XM23]. architecture-based [PV18]. Architecture-Independent [LS03]. Architectures [CB05, EGL05, MM00, MBS01, MB10, YW08]. Area [Bha00, BMM^{+}08, MMHJ10, MSK24, ND18, PN20, RDYGE18]. areas [PKM18]. Arithmetic [APSK14, DD16, RS16]. Arnold [AKAJ23]. Array [CBXh18, ITK08, JH00, RDB07]. arrival [NN21]. Art [BSB10, SMAM25, Yao17]. Artefact [FO09]. arthritis [PV18]. Article [CZC20, FCW^{+}20, Gen19, HZ20, Li20b, LC20, WZW20, YZ19a, ZSWF20, ZY19a, ZY19b, ZB20, Ano19, Ano21j, Ano21g, Ano21a, Ano21c, Ano21i, Ano21b, Ano21d, Ano21h, Ano21f, Ano21e]. Artificial [HSR20, WG19, ZXF07, ARAAMD24, AA24, Ano21j, Ano21a, Gen19, JN17, PK20, ZY19b]. artificial-noise-aided [ARAAMD24]. ASCII [EHA16]. aspect [GB20]. Asphyxia [SML^{+}13]. Assessing [ASGW20, Dud19, BRR17]. Assessment [LSMI11, LM02, PMR^{+}03, SS20c, HRMA17, KB22]. Assignment [CHC01, KNS07, SS10, TZ07, Di18, Mal24b]. assist [TS19]. assisted [OHZ^{+}23]. association [BKXK19, BD24, Hua18a, RN17]. Asymmetric [SYJ11]. Asynchronous [AAA02, Juh06, MYA16, PS16]. AT-densenet [SR23a]. Atmospheric [BHB09]. Atoms [UU03]. Attachable [Shi13]. attack [DVT20, KSS19, MB19, NCGA24, YxLmLx17, Zhe19]. Attacks [CEJ12, NK24, BMSP24, DT20, HRMA17, SS21, VLK19]. Attendance [TS11]. attention [HFG22, MBP23, NMK23, SAM23, TG24, XM23, YZ21, ZDG^{+}23]. attitude [RS17]. attribute [CG23, DM20]. attribute-based [DM20]. attribute-level [CG23]. Attributes [MS03, MS16]. Attribution [SRCXLC10]. auction [WQX^{+}15]. Audio [AS09, SS19a]. auditability [SS15]. augmentation [IB23, SD23]. Augmentative [BB10]. Augmented [dMRF^{+}14]. augmenting [GMD^{+}24]. Australia [RWS17]. Authenticated [Tan11, JCYZ18, RRA^{+}23]. Authenticating [Chi08]. Authentication [LH12, PBS22, WL07, DG19, GMD^{+}24, GAS23, JCI18, JXLZ15, JC19, KC19, KPMR20, KKA20, LTY^{+}24, MKA^{+}18, NNDP22, Ogu18, SK18, SCM18, TK24]. author [LLL^{+}15]. Authoring [DAC03]. Auto [LH13, LBYB17, SS20a, IQ22, SR23b]. Auto-adaptive [SS20a]. Auto-Design [LH13]. auto-encoders [IQ22]. Auto-establishing [LBYB17]. auto-solving [LBYB17]. autocorrelation [BFB^{+}20]. autoencoder [DND25, TN24]. autoencoder-based [TN24]. automata [GMD16, HC17]. Automated [CH10, CS09, KKSR22, PKM18, Shi12a, SAAAOH10, AFLdS25, SC25]. Automatic [KMR13, Shi12b, AAA17, BHZGG22, HKE20, KB22, NB24, PGK19, RPAU17, ZJG17]. Automatically [DGR09]. Automating [CPKT18]. Automation [GSV08, JSZ09, Yu19]. Autonomous [LGD^{+}08a, MMCM06, Ogu18, PC22, RS17]. Available [HM07]. AVC [YMTbB15]. Aware [AJS08, CL11a, KHVW04, MK11b, NS10b, PC11, SS04, WTLW07, AASA22, Dar20, GBBZ19, HKS19, KV17, KK23b, MSK24, NN17, PC22, RS19, RK15]. Awareness [SG05, JPJ^{+}19, MJ19a].

B [GG10]. B2B [FH05]. Backbone [SN12]. Background [GGMM23, HY18, Lou19, SLSS24]. backgrounds [EMN17]. Backoff [RDEDAE09]. Backpropagation [OS08, Sar13, EHA16]. backs [OB15]. Backup [HCSC06]. Balanced [JJY^{+}10, TN04]. Balancing [BW07, GT10, Vee05, GP23, HKS19, KS17, SP21]. balking [AJJ18]. ball [SSS21b]. Band [Wan11, Gho17, SCM18]. Band-Splitting [Wan11]. Bandwidth [ANB09, Ani14, CLBP09, HJR08, SSP07, SLH08, TT08]. Base [GM01, AMM^{+}17, CSFS17, VR19]. Based [Abd11, AKAJ23, AA12, AH02, AFbMSA10, ABR10, AJAk00, AR05, AAED06, AEDS07, AP01, AB09, APSK14, AK06, BL01, BW07, Beg06, BSB10, CT13, CHC01, CCZC13, CGY^{+}13, CX03, CL11b, CBL06, CLBP09, DK05, Dar09, DP03, EHA14, FL04, FRP07, FH05, GT10, GHRK11, HL11, HCH12, JKRS02, JS08, JYYJ08, Kar03, KB09, KZ07, LL14, Lap05, LC07, LLC11, LSMI11, LXH07, LXH08, LWH12, LLRW07, LSC03, LQLX11, LXM^{+}12, LM02, LD13, MS03, MD14, MKH02, NS10a, NC10, OSKC02, PC13, PF10, PTLH08, PS06, RHR06, RGNMPM12, RDEDAE09, SLZZ10, SL03, SN12, SSP07, SM10, Sha11, SX12, Shi12b, SHS^{+}02, SN11, SWL07, SRJS08, SS12, TAD03, TS11, UU03, VT01, VG14, VTS06, Wan06, WVK07, WPJZ08, WZW20, Wer09, WLC05, WLC10, XLQ09, XGC12, XCCL13, YLY07]. Based [YHL04, YHK^{+}10b, YTW^{+}06, YYD10, ZC07, AL19, ASGW20, AESBE17, ASM18, AA18, AGM23, AASA22, ADM21, AS24a, AKB^{+}19, ABB19, AB24b, Ano21g, Ano21i, Ano21b, Ano21d, Ano21f, Ano21e, AMM^{+}17, AEA19, BKXK19, Bad19, Bai21, Bay18, BR17, BA23, BBTH18, BB23b, CKC15, CG23, CM20, CLTY17, Che20, CXJ24, CBXh18, CDR24, CS22, CRAV24, CDM^{+}21, DY18, Dai19, DD16, DMB23, DNP24, DSBR19, DSP^{+}18, DND25, DP17, DS19a, Di18, DZ18, DZ23, Don18, DWT23, Dua23, DM20, DRF18, EMM24, EBA19, Fan18, FR18, FSL17, GMS19, GMS18, GT19, GR19, GMD16, GP17, GSP19, HKS19, HLW^{+}18, HZW^{+}19, HFG22, HNB24, HNH19, HZ20, HMP20, HY18, HA20, HCC16, HC17, Hu17, Hua18a, Hua18b, IAT20, IG23, INR24, IQ22, IB23, JNS18, JPJ^{+}19, JXG25, JSM19, JSYT21, JC19, Kao17, KSD23]. based [KSB19, KC20, KKR25, KB22, KJ22, KV17, KKM17, KR19a, KR19b, KK23b, KKA20, Li17, LZL^{+}17, LWG17, LL18, LC20, LZXB24, LGD^{+}08a, LCZA17, LZ18, LTY^{+}24, LCZD02, Lou19, MSH21, MBP23, MR19a, MPS^{+}19, MSG18, Mal24b, MP18, MB19, MSD24, MMH18, MMH19, MS17, MMS20, NN17, NN21, NMK23, ND18, NX18, NNDP22, NZT^{+}20, NBS21, PBS19, PV18, PSP19, Pat20, Pei18, PCK21, PR19, PBS22, PNAB18, QDQ24, QD18, RPAU17, RRA^{+}23, RN17, RR17, RHJ24, RJS18, RDYGE18, Rui19, SMSI23, SA24, SR20, SS20a, SMAM25, SKS19, SS18, Sha20, SJTl25, SHLA17, SS20b, SPC19, ST16, SN20, SD23, SGZL20, SPKS24, TTW15, TYBwY17, TLC18, TKY23, TN24, TW19, Tia18, TA23, VP19b, VS24, VV19, WC24, WZTH18, WLHZ18, WWHW18, Wan19, WYZ^{+}19, fWfH19, WG17, Wu18, Xia18]. based [XZ18, XZCL17, XM23, YMRLJ17, YLZ^{+}15, Yan19, Ye17, Yin18a, Yin18b, YxLmLx17, YYM17, YZ19a, YWDC24, YWH19, Yut18, ZEZ17, Zha17a, ZSWF20, ZDG^{+}23, ZX24, ZWC24, ZLHW18, Zhu17, ZDW^{+}21, ZB20]. Based-NoC [JS08]. Basic [BK01, KO00, GMS18]. Basis [KJ03, LE02]. BAT [MGT19]. Batch [HB01, SNS18, MM24]. Bayes [AEDS07, JWC^{+}09, JCW10]. Bayesian [BBL23, DRF18, KLRH13, LLR05, SG16, Yan18, Yan19]. BDD [ZWC24]. BDD-based [ZWC24]. BDT [Li20a]. BDT-SVM [Li20a]. Bed [EKAO08]. bee [JN17, PK20]. behavior [AMM^{+}17, BO18, Dua23, LZL^{+}17, MS19]. Behaviour [Sat08]. being [Pat23]. better [APZ17, DD16]. Between [MU09, JPJ^{+}19]. Beyond [BB23b]. Bézier [MU09]. Bi [GK23]. Bi-GRU [GK23]. bibliographic [yZ19b]. Bidirectional [RCJJ06]. big [Ano21c, CGC17, Che20, DY18, FCW^{+}20, HZ19, Li17, LZL^{+}17, LGL20, MR19a, SK17, Wan19, Zha17b, Zhe19]. Biharmonic [MU09]. Bilateral [LXM^{+}12]. Bilinear [LXH08]. Billing [AA06]. Binary [MD14, HNH19, IAT20, IG23, TN24]. Binding [BO12]. Biodiesel [KB10]. biological [Sop23]. biological-like [Sop23]. biomarker [PBS19]. biomarkers [MSG18]. biomedical [GK23, HNH19]. Biometric [DS10, PBS22]. Biometrics [Nom11, TS11, HA20, JCI18, SCM18]. Biometrics-Based [TS11]. biomodel [AA18]. Bionic [GS19]. BIST [HH07]. Bit [ABR10, APSK14, DS11, EQSMB07, LC07, DG18, EHA16]. Bit-Counting [EQSMB07]. Bit-Level [ABR10]. bit-mapped [EHA16]. Bit-Parallel [LC07]. Bit-Rate [DS11]. Bitrate [CCLC10]. bitwise [LWG17]. Bivariate [AC06]. black [KK23a]. Blended [KH19]. blends [KB10]. Blind [CLX10, CX03, FL04, GMD^{+}24, DNP24, PR18]. Block [IG23, LLRW07, PCK21, XXR09, BK23, MJ19b]. Block-Based [LLRW07]. Blockchain [OHZ^{+}23, AJK23, CDR24, CDM^{+}21, HL24, KSD23, MMPM21, NNDP22, SA24]. Blockchain-assisted [OHZ^{+}23]. blockchain-based [CDR24, CDM^{+}21, NNDP22]. blockchain-enhanced [HL24]. blockchain-layered [SA24]. blocking [Pat20]. Blocks [CTW12]. blog [DZ18, Yin18a]. Bloom [MK11a]. Board [Ano17b, LLC11]. BOCA [NNDP22]. body [CZZ18, MSK24, ND18]. Boolean [DG18]. Booth [LC07]. bots [VLK19]. bound [ASM21]. boundaries [SN23]. Bounded [AAAZ00]. box [KK23a]. boxes [DG18]. BPO [JN17]. Bragg [AFbMSA10]. brain [HHU24, Kao17, QDQ24]. brainstorm [AD18]. Breadth [OL00]. Breadth-First [OL00]. breast [MBP23, SJ17]. Bregman [CCZC13]. Broadcast [Bha00, CEJ12, XYSX10]. Browser [SGD07]. BSS [WLL10]. Bucket [SR06]. Buffer [GH06, Tsu07, CSB19]. Buffered [BLDD06, Dua07]. Buffers [AKK10]. bug [HE20]. Building [AEDA03, CW03, TOG^{+}05, ENER23]. Buildings [LM02, GT19]. Bus [LGD^{+}08a, OB15]. Butterworth [GS19].

C [EMN17, KMR13, KC10]. C-Means [KMR13]. Cable [FJZ06]. Cache [AL06, ASA04, DLDB04, JH00, CSB19, OB15]. Caches [LHK03]. Caching [SSP07]. CAD [LBYB17]. Calculation [LLH08]. Calculus [IIDI01]. Call [Ano00a, Ano00b, Ano02a, YM01, YYM17]. calls [SS20c]. calories [LGSS18]. Camera [KMGS11, MC07, WLC10]. can [ABK24, GHRK11, LGD^{+}08a]. CAN-Based [GHRK11]. CAN-Bus [LGD^{+}08a]. cancellation [RVS24]. cancer [GGK^{+}24, MBP23, SR19, SJ17, WYZ^{+}19]. Capability [DTDE06, HYT^{+}05]. capacitor [SD17]. Capacity [AB09, HKS19, KNS07, ZYM03, CPJ17, FR18, MFK^{+}15]. CAPTCHA [VLK19]. caption [ZFH24]. capture [Ye17]. Capturing [MBS^{+}24]. Carbon [OIB17]. Cards [dSTE05]. Carlo [TLC18]. Case [ASA04, DGR09, LH13, Loo04, LM02, RW07, EHA16, MT24, MP17, RMG19, SJZ^{+}18]. Case-Based [LM02]. Cases [LH00]. categorization [CM20, KKSR22]. Cell [AAAZ00, BLDD06, SR19]. Cellular [CGC01, SBS12, SD03, GMD16, HC17]. cement [GKMR19]. Cemo [ZLF19]. center [CT15, HNH19]. center-surround [CT15]. center-symmetric [HNH19]. centers [BZT23, DM20]. Centralized [LV06]. centric [Dar20, OHZ^{+}23]. CEP [LGSS18]. Certain [AS19, Pat23]. Certificate [LXH08]. Certificate-Based [LXH08]. certificateless [JXLZ15]. certificates [NNDP22]. Certified [LXH07, XGC12]. CF [CS22]. CF-AMVRGO [CS22]. CGSA [SR23b]. Chain [YLY07, BK23, KSD23, MMPM21, PCK21, RO24]. chain-based [PCK21]. Chains [FL04]. challenges [EhMZ24, GR19, YNA23, udQ22]. change [DR15, DS19b, GS22, PK19, YYM17]. changeability [ASGW20]. Changing [IIDI01]. Channel [CHC01, GT09, HHGG11, KJ22, SBS12, SD03, SRJS08, LSDJ23]. Channels [AMJ12, Bab11, SdSNL06]. Chaos [AKAJ23, Hou18, KR19a, AKAJ24, DNP24]. Chaos-based [KR19a]. Chaotic [DG19, IAD10, AS24b, CKC15, FMP^{+}19, GAS23]. character [EHA16]. Characteristics [MS03, NvV10, SR20]. Characterization [RS17]. Characterizing [LLH08]. Characters [AAD06, DJQS04, FR18]. Chat [OE12]. Chebyshev [DG19]. Checker [LYR09]. Checking [INyN11]. Checkpointing [AMJ12]. Chief [PF18, Pat24]. Chikungunya [KC20]. child [AGM23]. children [BBL23]. China [WXYN13]. Chinese [LLL^{+}15]. Chip [CJNY^{+}07, LHK03, GBBZ19]. choices [NCGA24]. Chopper [EHA14]. Chromosome [PTLH08]. chronic [JS19, SV23]. chronological [ASM18]. chronological-whale [ASM18]. cipher [MJ19b]. Circuit [KS14, MMCM06, SS24]. Circular [ASM18, Tsu07, VR08]. cities [SK18]. citizens [CDM^{+}21]. City [ZW08, Ano21f, HZ20, ZTSL08]. Class [LC07, LQLX11, MS01, RW07, ZB08]. Classes [DTDE06, Shi13]. Classical [SBS12]. Classification [ANB09, GMS19, JCW10, KAKF11, KM09c, ZXT^{+}06, AS20, BFB^{+}20, BJ17, CLTY17, CMG20, DMB23, GH20, GR19, GK23, Hou18, Hua18a, JS19, JGM25, Kao17, KJ22, KR18, KJ19, Li20a, MPS^{+}23, Mal24a, NM16, PK20, RPS18, SR20, SP18, SHLA17, SGZL20, SCS19, SKN18, SK24, VS24, Yin18a]. classification-based [DMB23]. classification-oriented [CMG20]. classifications [ZA18]. Classified [LS00, Zha17b, dMRF^{+}14]. Classifier [AEDS07, Sam10, AR19, BJ17, CU19, GMS19, KH19, MBN^{+}24, NCGA24, SP18, SKS19, Yan19]. Classifiers [JWC^{+}09]. classify [JSYT21]. Classifying [SJTl25]. Client [TZ07, ZC07]. clinical [FSNK22, MZZC18]. Clock [KS14, LLH08].

Clock-Gating [KS14]. clones [BS20]. CLOS [Dai19]. Cloth [ZFV04]. Cloud [KRMS19, PSP19, PBS22, RWL19, RSS24, ASM18, AKB^{+}19, AA24, Ano21f, Ano21e, AEA19, BM18, BSG19, BTAB21, BTZA18, CG23, CZL^{+}17, DE17, DM20, GP23, GGB23, GAS23, HKS19, HZ20, HFK19, HA20, KKKV19, KS17, KKA20, MPS^{+}19, Mal24b, NY17, RG23, RMF19, SZT18, SS20a, SS15, Sha20, Yan17, Zha17b, Zha17a, Zhu17, ZB20]. Cloud-based [PSP19, MPS^{+}19]. cloud-medical [KKA20]. cloudlet [RS19]. Cloudsim [BZT23]. Clues [RCNM03]. Cluster [Ano02a, BW07, GG05, HWL07, Juh06, KYL11, MPDK13, NS10a, PR19, SAAAOH10, VTS06, ZFV04, AWA20, BR17, JC19, VV19]. Cluster-Based [NS10a, VTS06, JC19]. Cluster/Grid [Ano02a]. Clustered [YWZJ06]. Clustering [JR12, KYL11, KM09a, KMR13, LQLX11, LHZ18, MR19a, SAD09, YLY07, YA06, AWA20, AM16, Ano21c, Bad19, CM20, CLTY17, DSBR19, FCW^{+}20, INR24, IB23, JNS18, JYN19, MMS20, NB24, NA19, RN17, RH23, RHJ24, WZTH18, WLHZ18, XZ18, Yan19, ZLHW18]. clustering-based [INR24]. Clusters [HR04, Vee05]. cluttered [EMN17, SLSS24]. CMA [DSBR19]. CMA-ES [DSBR19]. CNN [AS24a, GK23, HS24, KJ22, MPS^{+}23, SAM23]. CNN-GRU [AS24a]. Co [AJAk00, Ano00b, HNH19]. co-occurrence [HNH19]. Co-Operative [Ano00b]. Co-Scheduling [AJAk00]. coalition [PNAB18]. cocaine [BB23b]. Code [AFbMSA10, PS12, SS10, YCYL04, YWZJ06, AS24a, GMD^{+}24, Kar19, SHP25, ZJG17]. Codebook [CKC09, CM20]. Codecs [LLRW07]. Coded [VR08, KSB19]. codes [MFK^{+}15, RMG19]. Coding [Kao10, LS00, DD16, SN20]. Coefficients [BDB14]. Cognitive [KMCV09, ARAAMD24, Ano21j, DU19, EhMZ24, Gen19]. Coherence [AL06, ASA04, DLDB04, GCMS20]. Coherence-Replacement [AL06]. cold [BB23a]. Collaboration [DHSR05, VCP00, VNBA03, YWZ^{+}06]. Collaborative [CW03, CS22, FCW^{+}20, KK13, KZ07, KL03, AJP18, Ano21c, CXJ24, DZ18, Li19, QD18]. Collaborative-Learning [OSKC02]. collection [AWO^{+}18, RK24]. colony [JN17, MR23, PK20]. color [CPJ17, CJG18, HS24, IAT20, IG23, PBM16, ZEZ17]. Colour [KC10, Sam10, HMP20]. Combination [NC10, BM17]. combinations [JSM19]. Combinatorial [WCC06, Bar18]. Combined [FWQ00, DS19a, KJ19, ZLF19]. Combining [AM14, FMP^{+}19, GSJ24, PS19]. Command [MMH01, Shi12b, RS17]. Comments [Tan11]. Commerce [AP04b, AP01, CLX10, CB05, FH05, CXJ24, Li19, TW19, ZLF19]. Commercial [Lap05]. Commissioning [GSV08]. Communication [Beg06, BSB10, HAA05, JYHS01, KYL11, NPY07, PLK^{+}04, RCJJ06, SdSNL06, TAD03, VCP00, CZL^{+}17, JJS18, JC19]. Communications [CHC01, MJ19b, Pat22a]. communities [BA16]. Community [Pei18, SHI24]. companies [Pat18]. Comparative [GT09, JV12, MU09, PBA10, GH20, MPS^{+}23]. Comparing [SEH05]. Comparison [AAA02, BKK17, GTY14, Oh10, GBBZ19, MS16, QDQ24, SG16, TYBwY17]. competence [LGL20]. competency [BHZGG22]. Compiler [TOG^{+}05]. Compiling [DDL08]. Complex [Kur07, Men18, BA16, CSFS17, Lou19, OOG^{+}18, Pei18, ZDG^{+}23]. Complexity [LC07, YF18]. Comply [ÖZ11]. Component [AA12, LLC11, PW02, SML^{+}13, SIT11, ASGW20, DS19b, GGMM23]. Component-Based [AA12, LLC11, ASGW20]. Components [Har03, SV03, SPW02, BBTH18]. Composed [LL03]. Composite [GP11]. Composition [KM09c, PW02, SPW02, BBTH18, Sha20]. COMPOW [HL11]. COMPOW-Based [HL11]. comprehensive [Mal24a, MG19, OMA24, SC19]. Compressed [CHL01, FJK01, LHK03, Oh10, CJG18]. Compressing [GA11]. Compression [ABR10, CKC09, CC06, DS11, LE01, LE02, LD13, MAL10, PC13, PS06, FSL17, KSB19, KR19b, MGT19, NM16, SSP19]. Compressive [CCZC13]. Compressor [GA11, TN24]. Computation [BA03, FJZ06, Jin10, YYZ08, DM20, HFK19, KH19, MMPM21, SC19]. Computational [Ano00b, JJS18, SNS18]. Computationally [AK06]. Computations [WR00]. Computer [AO09, EHA14, Hag03, LSMI11, MLT09, SR20, Shi12b, VNBA03, Ano21d, Ano21h, CZC20, CPKT18, Kao17, RJS18, Yan17, Yao17, Zha17a, ZSWF20]. Computer-aided [SR20]. Computer-Based [LSMI11, Ano21d, ZSWF20]. Computer-Mediated [Hag03, VNBA03]. Computerized [AEDA03, SJ17]. Computers [KYPS06]. Computing [Ano02a, AMJ12, AAA^{+}11, CB05, Dar09, HP04, KBL21, Lam13, LB04a, SS19b, TN04, VTS06, XY02, YW08, ZKNL10, ASM21, AA24, Ano21f, Ano21e, AZK23, BP17, CZL^{+}17, CSB19, HKS19, HZ20, HWL^{+}18, KKK22, KRMS19, LJL^{+}19, MV18, NN21, NY17, Oka19, RSS24, RMF19, SZT18, SKS19, Sha20, Yan17, YWH19, Zha17b, Zha17a, Zhu17, ZB20]. computing-based [Ano21e, SKS19, ZB20]. concatenated [JGM25]. Concealment [SMF12]. concept [AB24a, IB23, PRR20]. concept-drift [PRR20]. Conceptual [SCCB10]. concerned [Pat18]. concerns [RWS17]. Concrete [GA08]. Concurrency [XYSX10]. Concurrent [GT10, Haq06, HH07, WHBS01, Ano21f, HZ20]. Conditional [MR19a, Ogu18]. cone [TW19]. confidentiality [BM18]. Configurable [SKP09]. Configuration [MRU04, LJL^{+}19]. confirmation [ZJG17]. conflict [GGB23, KKS18]. conflicts [ZWC24]. congestion [Dar20, GBBZ19]. congestion-aware [Dar20, GBBZ19]. Conjugate [EA05]. Conjugate-Gradient [EA05]. conjunction [MBP23]. Connected [HY00, MPDK13, Rah19]. Connectedness [SWLH08]. connection [KJ22]. Connections [SVM^{+}00, TT08]. Connectivity [SLZZ10, BK24, SS20d]. Connectivity-Based [SLZZ10]. conserved [PW19]. Considerations [MMCM06, BS24]. considering [PKD18]. constant [OHZ^{+}23]. constant-size [OHZ^{+}23]. Constrained [KM09b]. Constraint [LS10, CJW^{+}23, SSMM16]. Constraints [Sat08]. Constraints1 [ZKNL10]. Constructing [AAA17, ZYM03]. Construction [Sat08, HZ19, Li17]. Constructivist [Kom03]. constructor [SS24]. consuming [DY18]. consumption [Ano22a, JNS18, ZY19a]. contacts [BO18]. Content [JKRS02, MS03, Rad04, SHS^{+}02, TVV^{+}02, Wan06, AEEO23, VP19b]. Content-Based [JKRS02, SHS^{+}02, Wan06, VP19b]. contention [SN23]. Contents [Fur07]. Context [CL11a, MK11b, WR00, ZWC24]. Context-Aware [CL11a, MK11b]. Context-Driven [WR00]. contexts [Sen24]. contextual [DR15, JLA20, ZWC24]. Continuous [AEDS07, LCHR06, VV19]. contour [DRF18]. contracts [SMT^{+}23]. Control [AAA02, ACA07, CL10, Din11, KZ07, LV06, LCZD02, SGD07, XYSX10, ZB08, ASM21, BYBM21, BKK17, CPKT18, DM20, HDS17, KKS18, MPĆ15, PCK21, RS17]. Controlled [CCLC10, LB04a, RDEDAE09, ASGW20].

Controller [ANB09, EHA14, PS16]. controllers [BKK17]. Controlling [MLT09]. convergence [EHA16, Man15]. Convergency [YTW^{+}06]. Converging [DNP19]. conversational [MBS^{+}24]. Conversion [ZLC06, EHA16, SD17, SMT19]. converting [MYA16]. ConvNet [XM23]. Convolution [ZNLA08, MBP23, YZ21]. Convolutional [AJK23, SN11, AB24b, ISM25, JSYT21, LSDJ23, SCS19, SK24, WYZ^{+}19]. Cooperating [LL03]. Cooperation [INyN11]. Cooperative [DAC03, IM19, AASA22, Yu19]. Coordinated [AMJ12, BBAR08]. Coordinating [CW03]. Coordination [XY02]. Copy [SSS21b]. copyright [DNP24]. CORBA [OA01]. Core [LL14, HS24]. Corpus [KAKF11]. Corrected [CT15]. Correctness [CS03, SP14]. Correlation [LL11a, LD13, PC13, Ano21c, CGI18, FCW^{+}20, LSDJ23, MSH21, NX18, RG23]. Corrigendum [Ano17a]. COSMIC [DG09]. COSMIC-FFP [DG09]. Cost [INyN11, KLB06, LM02, YYZ08, NCGA24, Rah19]. cost-effective [Rah19]. Cost-Modelling [KLB06]. counterfeit [DR15]. Counting [EQSMB07, BFB^{+}20]. countries [Pat23]. Coupled [LXM^{+}12]. Coupling [AM14]. Course [MS03, RHR06]. Courseware [DAC03]. Courseware-Authoring [DAC03]. cover [DS19b, GMS19, GR19]. coverage [SC19, SS20d]. covert [MFK^{+}15]. COVID [Pat20, Pat22a]. COVID-19 [Pat20, Pat22a]. Covid’19 [JSYT21]. CPABE [CG23]. CPDALB [HKS19]. CPU [SEB06]. crawling [DS19a]. create [Al 18, SK24]. creating [RWS17]. Creation [PBS22, PGK19]. credit [DM20]. criminal [RMG19]. crisisMMD [JGM25]. Criteria [Lap05, FSL17, SMT19]. Critical [SD17, ND18, SAN24, ST16]. cropping [RPAU17]. Cross [WLL10, LB04b].

Cross-Tree [LB04b]. Crossbar [Dua07]. Crowd [DM20]. crowdsourcing [Pat20]. cruise [HC17]. Cry [SML^{+}13]. Cryptanalysis [KPMR20, YHK^{+}10a, DG18]. crypto [DG18]. cryptocurrency [Pat23]. cryptographic [WC24]. Cryptography [WWL^{+}14, AEH17, IAT20, PC19, RRA^{+}23]. Cryptosystem [SN11, HA19]. CS [Kur07]. CSP [AKB^{+}19]. CT [BRB22, WMB02, WLZZ12]. Cube [ASA05, HY00]. Cube-Connected [HY00]. Cubic [ITK08]. Cuckoo [DNP19]. Cumulative [MS01]. CUR [PS19]. Current [DP03, LL07]. Cursive [SS05]. curve [RRA^{+}23]. curvelet [SHLA17]. Curves [Pop05]. custom [TG24]. Customer [KZ07, Lin11, ZLF19]. customized [FSNK22]. Cut [HR04, RRRR17]. Cutting [CGH11]. CVE [Hsi10]. CVE-ES [Hsi10]. cyber [Mal24a, Oka19]. cyber-physical [Oka19]. Cyberbullying [AEEO23]. Cybersecurity [Pat22b]. Cyberspace [Pat22b]. cycle [Xia18]. Cycles [HY00]. Cycling [Smi03, DNP24].

D [JJY^{+}10, MM00, MMM13, Man15, MKMA07, NPY07, Wan06, WLZZ12, Wer09, Yut18, ZNLA08, ZMY^{+}00]. D2D [VP17]. DAG [ZDW^{+}21]. DAG-SVM [ZDW^{+}21]. DAM [GSP19]. DAM-FL [GSP19]. Damaged [DJQS04]. Data [AAA02, AAY02, AWO^{+}18, CGC01, CHL01, CCLC10, CGL12, GHRK11, KC10, LCHR06, LLRW07, LB04b, LE02, MDY14, PT14, Sam10, SCCB10, SP14, Tia18, VHL11, YYM17, YTW^{+}06, ZWLH14, dMRF^{+}14, AM16, Ano21c, Ano21f, Ano21e, AS24b, BMS24, BD24, BK23, BS24, BZT23, CGC17, Che20, CPJ17, DY18, Dai19, DM20, FCW^{+}20, GH20, GGMM23, He18, HZ19, HZ20, HA20, Hou18, HPC18, Hua18a, KS21, KR18, KSS19, KJ19, KPMR20, Li17, LZL^{+}17, LGL20, LHZ18, LZ18, MYA16, MR19a, Mal24a, MP18, Men18, MBS^{+}24, NZT^{+}20, Ogu19, Pei18, PRR20, RMG19, RN17, RG23, RK24, RWS17, SK17, SP18, SS15, SD23, TAUK25, VR19, VNR19, VV19, WZTH18, Wan19, fWfH19, WG19, XF18, XZCL17, YNA23, Ye17, Yin18a, Yu19, YWDC24, ZLF19, Zha17b, ZP19, Zhe19, Zhu17, ZB20, AEH17]. Data-Dependency [KC10]. Data-Dependent [Sam10]. data-intensive [LHZ18]. Data-Link [AAA02]. data-set [WZTH18]. Database [Bha00, Lee01, Loo04, PG04, Rad04, MR23]. Datahiding [IEZA10]. dataset [TG24]. datasets [Bad19, BS24, EhMZ24, KBD21, SK21]. day [KH19]. DBSCAN [RH23]. DCT [BDB14, CJG18, LSC03]. DCT-Based [LSC03]. DDC [YPY20]. DDoS [BMSP24]. DDPT [MP18]. de-constructor [SS24]. DEA [Tia18]. deadline [HKS19]. Deadlock [Haq06]. deblocking [KSB19]. decentralized [MMPM21]. Decision [AAD06, SSG06, TCC^{+}14, Dar20, DMB23, FSNK22, He18, QDQ24, SP18]. decision-based [QDQ24]. decision-driven [Dar20]. Decision-Making [TCC^{+}14]. Decoder [AFbMSA10, MR12]. Decommissioning [LCZD02]. Decomposition [EGL05, KNS07, KC10, PS06, ADM21, CJW^{+}23, GMD^{+}24, GGMM23, HNB24, PS19]. Decompression [LHK03]. Decontamination [LCZD02]. Deductive [OL00, OL04]. Deep [AKK10, HE20, AYI23, AB24b, Ano21b, DZ23, EhMZ24, ECBR20, FSNK22, FD23, GGK^{+}24, GSJ24, HT19, HHU24, JGM25, LSDJ23, MBN^{+}24, NMK23, PN20, RPAU17, SN23, SV23, SGMT22, SCS19, TAUK25, TG24, WZW20, udQ22]. Defect [GS11, FD23, VJTG24]. Defense [DVT20]. Defined [FA06, SP21, TKY23, VP17]. Definition [BA03, ÖZ11]. deformable [CZZ18, ZDG^{+}23]. deformation [AA18]. degeneration [DK23]. dehazing [TTW15]. Delay [BW07, MS01, SN12, SLH08, VNR19, MSK24]. delay-aware [MSK24]. delayed [Hao17]. delivery [KS21]. demographics [BB23b]. Demystifying [KK23a]. Deniable [JCYZ18, JXLZ15]. Denoising [Abd11, SYZS10, SS12, AB24b]. dense [AUHH^{+}18]. densenet [SR23a]. Density [SWL07, Bad19, JNS18, WLHZ18, ZLHW18]. density-based [JNS18, WLHZ18]. Dental [Nom12, BM17]. Dependable [Alm11]. Dependence [AAY02, Man17]. Dependencies [ADM21]. Dependencies-based [ADM21]. Dependency [KC10]. Dependent [BCP09, LCHR06, Sam10, Sha11, ZC07, He18, NN17]. Deployment [SSM12]. Depth [AAA^{+}11, CGI18, DWT23, TW19]. depths [BS24]. dereferences [Don18]. derivative [Rah19]. Dermatology [JS14]. Description [AAD06, Fur07]. Descriptions [CS03, Kur07]. descriptor [HNH19, VP19a]. Descriptors [SLZZ10, KB16, SS19a]. Design [AEHSES08, ACGP10, ANB09, AP01, APSK14, BB10, BSB10, DLL05, DDL08, FWQ00, GA08, Gho17, GHRK11, Hao17, Har03, Kur07, LH13, LYY^{+}08, PR08, PLK^{+}04, PBS22, RDB07, RS08, SEH05, SYJ11, VP19a, Vee05, WHBS01, ZKNL10, ASGW20, AJP18, Ano22b, DNP19, GMD16, Li20b, MR19b, udQ22]. Designated [CLX10]. Designated-Verifier [CLX10]. Designing [GP11, RKS^{+}25]. despite [Pat23]. detect [GS22, OMA24]. Detecting [BA16, CTW12, Don18, MC07, NK24, LTS25, ZWC24]. Detection [AFbMSA10, ABMKO14, BMSP24, CH10, DS10, DT20, Haq06, Lou19, SL03, TS08, WLC10, AB24a, AYI23, AJK23, AEEO23, BK24, BMS24, BJBA24, BB18, BBL23, CDE20, CGC17, CRAV24, DMB23, DR15, DK23, DS19b, DVT20, EMM24, EMN17, GMS18, GGK^{+}24, HT19, HY18, IQ22, Kar19, KKS18, MBN^{+}24, MSYG18, NCGA24, PR18, PRR20, PN24, RPAU17, RSS24, RPS18, RvW20, RDYGE18, SJ17, SS20c, SK21, SSS21b, SGMT22, SD23, SLQR20, SR23b, TAUK25, TLC18, WT15, YxLmLx17, YWDC24, ZDW^{+}21]. Detector [Wan11, ZZ20]. Detectors [KB16]. determinant [HFK19]. Determination [CLBP09, YA06, RS17]. Deterministic [OA01, SSO05]. detrimental [Pat18]. developer [AAB19]. Developing [LLC11, TS19, YWZ^{+}06]. Development [CS09, DG09, GSV08, Hag03, LZ07, MJ19a, MRU04, MMHJ10, SV03, VGPB09, ZSWF20, Ano21d, Bar18]. Deviation [LLH08]. Device [CL11a]. Devices [AC06, MMCM06, Ano22b, BSG19, Li20b, RWS17, SCM18]. DFMCloudsim [BZT23]. diabetic [MBN^{+}24]. Diagnosing [KM09a]. Diagnosis [GG10, Shi12b]. Diagrams [XXR09, WDW17]. Diesel [KB10]. difference [BRB22, RDYGE18, ZZ20]. Different [DA16, GG04, GT09, NK24, PGK19]. Differential [EJ05, JV12, JS16]. Differentially [SSS21a]. Differentiation [BJ07, SLH08]. differentiator [MR19b]. diffusibility [AS20]. Diffusion [BHB09]. Digit [NC10]. Digital [Jin10, Pat22a, RDB07, Ano21g, DNP24, HL24, Pat23, Ren18, SJ17, Yao17, YZ19a, yZ19b]. dilemma [Pat20]. Dimensional [Kat07, SYZS10, ZWLH14, GMD16, KD19, VV19]. Dimensionality [WC24]. Dimensioning [SR06]. directed [KLT15, RR17]. Disabilities [AR05]. Disabled [YCYL04]. Disaster [ZWLH14]. disc [BO18]. disc-pads [BO18]. discharge [RRRR17]. Discontinuities [SYZS10]. Discovery [Amm07, KMCV09, MR23, JYN19, NX18, Pei18, YxLmLx17]. Discrete [HWS09, MM00, BA16]. Discrete-Sized [HWS09]. discriminate [SKS19]. Discrimination [AS09, AAA17, NCGA24, PKM18]. discriminative [LLL^{+}15]. discussion [NY17]. Disease [GG10, RN17, HHU24, JS19, KC20, PV18, SV23, TS19, ZDG^{+}23]. diseases [GCMS20, NMK23]. Disk [JH00, YLZ^{+}15]. disorder [BB23b]. dispatch [SSMM16]. Display [Shi13]. Dispo [MGK^{+}00]. Disputes [MLT09]. Distance [Kar03, OSKC02, VP19b]. distance-based [VP19b]. Distance-Learning [Kar03]. distillation [ISM25]. distorted [KB16]. Distortion [CTW12]. Distributed [AESBE17, AH02, AA06, AD00, AMJ12, BBAR08, Bha00, BMM^{+}08, CHY00, DB04, DLL05, FA06, HYT^{+}05, INyN11, KB09, KL03, LVK03, LV06, MGK^{+}00, OA01, PKDP12, Rad04, SP14, TZ07, TN04, VGPB09, VTS06, AWO^{+}18, BZT23, DVT20, NN21, XF18]. Distributed/Parallel [AH02]. Distribution [GHRK11, GSV08, MB09]. Distributions [LGD08b, Wer09]. disturbances [SHLA17]. diversity [Bay18, GSP19]. diversity-based [GSP19]. Divide [Pat22a]. Divisible [FL04, Vee05]. division [Dua23]. DMCMPRA [LS10]. DNA [AEH17, Dar09, GA11, NK24]. DNN [JSYT21]. DNS [BW07]. DNS-Based [BW07]. do [Pat23]. Docker [LCZA17]. Doctor [OE12]. Document [INR24, IB23, MYA16]. document-elements [IB23]. documents [KB23]. Does [DP03]. DOF [BKK17]. Domain [AKAJ23, BS19, JSHB10, PS06, TT08, AKAJ24, AAA17, GSJ24, MSH21, MP18, MKA^{+}18]. Domain-specific [BS19, AAA17]. domains [LTS25]. Domination [MSD24]. Domino [KS14]. DORB [OA01]. Dot [GMD16]. Double [JR12, BFB^{+}20]. Downloading [Ani14]. Downstream [AAA^{+}11]. DPOS [BK23]. DPOS-hyper [BK23]. dragonfly [IAT20]. Drbd [XXR09]. drift [AB24a, PRR20]. drive [PS16]. Driven [Che07, DSV12, Kao10, MMH01, SH10, WR00, Ano21j, Ano21a, BS19, Dar20, EMM24, Gen19, LL18, Sen24, ZY19b]. driver [NPN^{+}23]. Drones [PCK21]. dry [BO18]. DSF [PCD25]. DSF-GAM [PCD25]. DSP [Ano21g, YZ19a]. DT [SZT18]. DT-MG [SZT18]. DTW [KB09]. Dual [CT13, HPC18, TYBwY17, Xia18]. dual-cycle [Xia18]. Dual-Exposure [CT13]. dual-frequency [TYBwY17]. ductal [PBS19]. due [PN20]. During [BO12, Pat20]. DWT [BK23, LD13]. Dyad [CLLC10]. Dynamic [AEHSES08, CL07, GP11, HYT^{+}05, JV12, JOYB13, KS17, Lee01, LS10, LV06, Oka19, PLK^{+}04, PK07, PSP19, SAD09, SNS18, SD03, SN23, SPW02, Wan11, XXR09, ZFV04, ZB08, AEA19, CT15, DU19, HKS19, HZW^{+}19, OOG^{+}18, SS20a, SS20c, SPKS24, XZ18, Yan18]. Dynamically [IIDI01]. dynamics [PK19].

E-Commerce [AP04b, AP01, CLX10, CB05, FH05, TW19, CXJ24, ZLF19]. E-governance [SK18]. E-Learning [KL03, MS03, RCNM03]. e-mail [JCYZ18]. EAAF [KKA20]. Each [Sha11]. Early [TCC^{+}14, JXG25]. ECC [BK23, KKA20]. ECC-based [KKA20]. ecological [Lin17]. economic [SSMM16]. Edge [AAAZ00, CH10, KBL21, RABB08, SL03, ZJZ15, DR15, KLT15, LJL^{+}19, Oka19, SSP19]. edge-directed [KLT15]. Edge-preserving [ZJZ15]. Edge-Width [AAAZ00]. edges [HNH19]. Editor [Pat24]. Editor-in-chief [Pat24]. Editorial [Ano17b, RWL19]. Editors [Ano05, PF18]. Editors-in-Chief [PF18]. Education [BK01, DP03, HWH03, JSZ09, PMR^{+}03, SHS^{+}02, UU03, Al 18]. Educational [Har03, NNDP22, Tia18]. EEG [BK24, BK24]. Effect [AO09, NvV10, TAD03, AS19, OIB17, PKD18]. Effective [PF10, PN20, TS11, HNH19, Rah19, RKS^{+}25]. effectively [KS17]. Effectiveness [NvV10, FD23]. Effects [MAL10, ABK24, KSB19]. efficacious [AS24b]. Efficiency [EQSMB07, GH06, HB01, JN17, LB04a, NM16, KK18, MM24, SPKS24]. Efficient [Ani14, AK06, AMJ12, Chi08, DS10, FA06, IAD10, JH00, JSM19, LXH08, MMM13, MR10, NY17, OS08, PT14, RS08, SAD09, SVM^{+}00, SHDY10, SS12, Wan06, WCC06, YP05, AL19, BSG19, BTZA18, CG23, DZ23, EBA19, GKMR19, HA19, IM19, IG23, KPSK23, KC19, KSD23, KR18, MMPM21, MT24, MG19, MBN^{+}24, MV18, MR19b, MSK24, ND18, PNAB18, RVS24, RHJ24, SS18, SS20d]. Effort [AB09, DG09, GS11, LYY04, Bar18, RKS^{+}25, ST16]. Egophobic [AZP03]. EHR [BM18]. Eigenproblems [GG04]. Elastic [SS24]. elderly [ZDW^{+}21]. Electric [ABEAA10, RRRR17, Zhu17]. Electricity [DK05]. electrocardiograms [MR19b]. electroencephalographic [Kao17]. Electron [GMD16]. electronic [AESBE17, HWL^{+}18, JCI18, Li19, OHZ^{+}23, TS19, WQX^{+}15, YWH19]. electronics [OIB17]. element [Yin18a]. elements [IB23, Lin17, Lou19]. Elevator [HLW04]. Eleview [HLW04]. ElGamal [HA19]. elimination [ZJG17]. Elliptic [Pop05, RRA^{+}23]. email [DA16, DMB23, FR18]. email-based [FR18]. Embedded [JR12, PR08, OOG^{+}18]. Embedding [LSC03, NK24, CPJ17, PSP19, PN24, Wu18, Yin18b]. embeddings [JGM25, SAM23]. emergency [Ali24, FT20, MJ19a, ND18]. emerging [NPN^{+}23, YNA23]. emotion [BK24, ECBR20, GT19]. Emotions [AAM^{+}18, MBS^{+}24]. Empirical [AM14, GS11, JV12, RKS^{+}25, BS25, KKXB19, KKS18, SPC19, SKN18, VJTG24]. Empowering [CDM^{+}21]. Enabled [RS04, YHL04, MKN^{+}18, SP18]. Encoded [SMF12]. Encoder [AFbMSA10, SR23b]. Encoders [CJNY^{+}07, IQ22]. encoding [KPSK23]. Encryption [AKAJ23, AEH17, ABB19, IAD10, LLRW07, Tan11, AKAJ24, AS24b, EK23, FMP^{+}19, HNB24, HA20, JCYZ18, MB19, RRA^{+}23]. Encryption-based [ABB19]. end [HFG22]. end-to-end [HFG22]. endless [GMS18]. Energy [AYV10, AA06, EBA19, KYL11, Kar12, KV17, LB04a, RVS24, SAD09, SS20d, AASA22, Ano22a, GMD16, Gho17, IM19, JNS18, MG19, MV18, MSK24, ND18, PR18, PNAB18, RK15, RHJ24, SPKS24, ZY19a]. energy-aware [AASA22]. Energy-Efficient [SAD09, EBA19, RVS24, SS20d, MG19, MV18, ND18, PNAB18]. Engine [Abd11, DB04]. Engineered [BD11]. Engineering [ACGP10, BL01, Har03, Kar03, DF09, TVV^{+}02, YWZ^{+}06, SHI24, Sop23]. engines [BTZA18, SS21]. English [OE12, WQX^{+}15]. enhance [SS24]. Enhanced [AM16, AASA22, MS17, OB15, PK20, SCS19, BD24, BS24, CG23, FSNK22, FMP^{+}19, GMD^{+}24, HL24, IAT20, RH23, SAM23, SHP25]. Enhancement [CKC15, CT13, Pon05, RDEDAE09, BK24, GS19, Guo19, GAS23, PK20, SKS19, WL18]. Enhancement-Based [RDEDAE09]. Enhancing [BMS24, DNP24, ENER23, HAA05, Lin11, Mal24a, Sen24, VLK19, EhMZ24, KC19, VV19, XZCL17]. Enlargement [Wu10]. Ensemble [HHU24, PBS19, PRR20, AAH25, BB23b, CU19, CRAV24, KJ19, SP18, SV23]. ensemble-based [BB23b]. ensembled [FSNK22]. ensure [SS15]. entailment [PN22]. entangled [KK18]. Entropy [MSG18, WLL10, Wan11, He18]. Entropy-based [MSG18]. Envelopment [Tia18]. Environment [BCP09, BK01, KL03, MRU04, OL04, PBS22, SKP09, VBM02, ZTSL08, Al 18, AWO^{+}18, AEA19, GGB23, GAS23, HKS19, KRMS19, KS17, LJL^{+}19, MPS^{+}19, MM24, NY17, RS19, SKM20, SLSS24, SS20a, TK24, XF18, ZDW^{+}21]. environment* [BM18]. environmental [RH23]. Environments [CL10, HP04, Juh06, RCNM03, Sha11, TZ07, XY02, XYSX10, MPĆ15, Sop23]. epidemic [TS19]. Equal [CGH11]. Equation [SHDY10, He18]. Equations [EA05]. equivalence [BR17]. era [Pat20, Xia18]. Ergodic [IAD10]. Erratum [Ano02b]. Error [SMF12]. Errors [GS13, NK24]. Essential [SJZ^{+}18]. establishing [LBYB17]. Establishment [MC07]. Estate [dMRF^{+}14]. Estimate [LGD08b, Wan11]. Estimating [AB09, DG09, KB10, LWH12, ZXF07]. Estimation [CDB13, FMI11, GG05, KHVW04, Li17, Wan11, Bar18, FSL17, LGSS18, RKS^{+}25, ST16, Wan19]. Estimator [WLL10]. Ethernet [CBL06, Vee05]. Ethernet-Based [CBL06]. Euclidean [LL11a]. Euclidean-Overlap [LL11a]. Evacuation [HC17]. Evaluating [NCGA24, dMRF^{+}14]. Evaluation [AM14, Hag03, OL00, SSP07, SCCB10, VNBA03, dMM03, AJP18, KKS18, MJ19a, PGK19, QD18, SAN24, SPC19, SKN18, Yan19]. Event [FRP07, LL18, WT15]. Everything [SS19b]. evidence [JSM19]. evidential [KB22]. Evidentiary [HRMA17]. Evolution [JV12, JS16, YxLmLx17]. Evolutionary [AS09, PTLH08, SOC07, XM23, BKXK19, BA16, DVT20]. examine [ABK24]. Exchange [BMM^{+}08, DK05, XCCL13, Dai19, MJ21, SMT^{+}23]. Execution [AB09, HB09, KM06, MGK^{+}00, SSO05]. Expandable [AEDA03]. Expansion [Hsi10, INR24]. Experience [EKAO08, GSV08, KKKV19]. Experiential [Kom03]. experiment [ASGW20]. Experimental [LGD^{+}08a, MS16]. Expert [Hsi10, MP18]. explainability [KK23a]. explainable [KK24, MBN^{+}24]. explicit [DRF18]. Exploitation [AEDA03]. Exploiting [CTLT04, LSDJ23, MMH01, SGMT22, DT20]. Exploration [DDL08, KK13]. Exploring [MKA^{+}18, PN22]. Exposure [CT13]. Expression [SRCXLC10, YTW^{+}06, AI18, GP17]. expressions [PBS19]. extended [JLA20]. Extending [SAAAOH10]. Extensible [DHSR05, Fur07]. Extension [SWLH08, BZT23]. Extensions [DHSR05]. External [CB05]. Extracting [CGI18, SC25]. Extraction [LYR09, AAM^{+}18, AWA20, CMG20, ENER23, HKE20, Kao17, LL18, MS16, PBM16, PKM18, RPS18]. Extractive [MLM^{+}24]. Extractors [VR08]. Extremal [MU09]. eye [GCMS20].

Fabric [MMHJ10, BK23]. Face

[KMGS11, MD14, Sar13, EMN17, HNH19, HCC16, TG24]. faces [DR15]. facial [GT19, GP17, HCC16, MS16]. factoid [BJ21]. Factor [OS08, SK18]. factors [LGL20, SJZ^{+}18]. Fading [Bab11, GT09, Zhe19]. FAHP [Mal24b]. failure [SHE^{+}23, SS20b, YLZ^{+}15]. failures [TKY23, TK24]. fair [Dai19]. Fairness [GT10, NvV10, ÖZ11, RDEDAE09, MM24]. Fall [WT15, ZDW^{+}21]. False [ZJG17]. Farming [Hsi10]. Fast [AC06, CC06, DNP19, DSV12, DS11, EK23, EKAO08, FR00, KGEL06, ZP19, CGI18, FM18, LL18, RPAU17, Xia18, NPY07]. fastDTW [LTY^{+}24]. Fault [AH02, ASA05, CBL06, HY00, JJY^{+}10, KRMS19, KM09b, LL07, WL07, ZJR08, GSP19, RR17, RMF19, ZJG17]. Fault-Tolerant [AH02, ASA05, CBL06, WL07]. Favourable [ÖZ11]. FAX [CHY00, JYHS01]. FAX-to-FAX [JYHS01]. Faxportal [CHY00]. FBGs [AFbMSA10]. FC [AZK23]. FCM [CBXh18]. feasible [OHZ^{+}23]. Feature [KSS19, LGD08b, SHI24, SMSI23, VR08, Zhe19, AASL23, AR19, CMG20, DWT23, Dua23, FSNK22, Hua18a, JS19, JXG25, Kao17, LZXB24, LLL^{+}15, MS16, NBS21, RPS18, Ren18, RDYGE18, Rui19, SHE^{+}23, SV23, TD23, VS24]. features [EMN17, HCC16, KD19, SS19a, SR19, SKN18, ZDW^{+}21]. Federated [CDR24, Wu08]. feedback [INR24]. FFP [DG09]. Fibre [AFbMSA10]. Field [Nom11, OIB17]. Figure [AB24b]. Figure-ground [AB24b]. File [Tsu07, ZJH08, LZ18, SS18, SS20b, XF18]. file-based [SS18]. filling [Yu19]. Filter [MK11a, RS08, RABB08, FM18, GS19, RPAU17, TLC18, ZJZ15]. Filtering [CS22, Kim01, Lee07, LXM^{+}12, SS12, Ano21g, CXJ24, DA16, DZ18, Li19, QDQ24, QD18, YZ19a]. filters [DNP19]. Financial [KLB06, KLRH13, Ano21h, CZC20, Ye17]. Finding [SVM^{+}00, ACM20, WGE17]. Fine [AYV10, FO09, Kao10, SGD07, CRAV24, MPS^{+}23, MBS^{+}24, MLM^{+}24, SCOD17]. Fine-Grained [FO09, SGD07, SCOD17]. fine-tuned [CRAV24, MPS^{+}23, MLM^{+}24]. fingerprint [JCI18]. Finite [AJJ18, LS00]. Finite-State [LS00]. FIR [DNP19]. Fireworks [HZW^{+}19]. First [MS17, OL00, DSD^{+}13]. Fix [GS11]. Fix-Effort [GS11]. Fixed [Lic06]. FL [GSP19]. fleet [BTZA18]. Flexible [AEDA03, TOG^{+}05, BYBM21]. FLFNN [Ye17]. flooding [Ali24, KSS19]. floor [MPS^{+}19]. Flow [GT10, HC05, LGD08b, WF06, Ano21c, Di18, FCW^{+}20, TKY23]. Flow-Shop [WF06]. flower [AM16]. Flows [Beg06, GT10]. Flows-Based [Beg06]. Fly [PR19, HLW^{+}18, ZLHW18]. focused [DS19a]. fog [ASM21, AZK23, KKK22, OOG^{+}18]. folding [AD18, PR17]. food [LGSS18]. foot [AA18]. Force [Nom11]. forecast [SG16]. Forecasting [OW11, BN24, CZZ^{+}23, Hao17, SS20a, TS19]. forest [AR19, LTY^{+}24]. forest-based [LTY^{+}24]. Foreword [PF18, Pat24]. forgery [SSS21b]. Formal [BL01, BA03, CS03, Kur07, LYY04, HDS17].

Formalism [HB09]. Formation [JJY^{+}10]. Formulation [SSP07, DRF18]. Forward [HYLS12, SS10, YHK^{+}10a, YHK^{+}10b, Yut18]. Forwarding [AAED06]. four [LWG17, TK24]. four-valued [LWG17]. Fourier [NPY07]. fox [Mal24a]. FPGA [LH13, SN11, SHDY10]. FPGA-Based [SN11]. fractal [MGT19, NM16]. fractional [Guo19, MR19b]. Fragmentation [BHB09]. fragments [BZT23]. Frame [BJ07, PR18, RDYGE18, TLC18]. frameless [Day17]. Framework [DSD^{+}13, DLL05, LLC11, LLR05, VGPB09, ZWLH14, BS19, Bar18, BSG19, BB23b, CG23, DSP^{+}18, HFG22, KSD23, KM09c, KJ19, KKA20, MYA16, MS19, MSG18, MT24, MMH18, MMH19, OHZ^{+}23, PRR20, RPAU17, SK17, Sop23, SS20d, SCOD17, SGZL20, ZLF19, ZDW^{+}21]. fraud [BMS24]. Free [MDY14, LTS25, LTS25]. Free-Phish [LTS25]. freelancers [Pat23]. frequency [TYBwY17]. frequent [MR23]. frog [MR19b]. frontend [KKR25]. Fruit [PR19, HLW^{+}18, ZLHW18]. FSBMA [YMTbB15]. Function [AC06, SSM12, FMP^{+}19, MR19a, YYM17, ZA18]. Functional [BO12, FWQ00, KK13, MB10, SIT11, BK24, SAN24]. functionality [Jij15]. Functions [KLB06, RDB07, WF06, YWZJ06, DG18, Man17]. funneling [AS19]. fused [GGK^{+}24]. Fusion [GK23, RDYGE18, SX12, BS17, Pei18, fWfH19]. FusionEXNet [GGK^{+}24]. future [HL24, OIB17, PJ19, SAN24]. Fuzzy [Abd11, ANB09, CGC01, Che04, Din11, HC05, KMR13, PG04, RDEDAE09, SSM12, VG14, AESBE17, AKB^{+}19, AEA19, BR17, BKK17, BS17, BM17, DSBR19, KC20, MSG18, NB24, Ren18, RMF19, SKM20, YMRLJ17].

G [HC11]. GA [BKK17]. Gabor [VP19a]. GAM [PCD25]. Game [CLLC10, JKRS02, PNAB18, SZT18].

Games [FRP07, LLC11, AAB19, KK18]. GAN [BK24]. gastric [WYZ^{+}19]. Gating [KS14]. Gauss [TLC18]. Gaussian [BHB09, CTW12, OW11, RPAU17]. Gender [Sha11]. Gender-Dependent [Sha11]. Gene [SRCXLC10, YTW^{+}06, MSG18, PBS19]. Generalized [WLL10, ZA18]. Generate [DJQS04]. Generated [DGR09, AMM^{+}17]. Generating [APZ17, LH00]. Generation [SIT11, TVV^{+}02, WXYN13, AA18, AASL23, BHZGG22, CM20, CZL^{+}17, LWG17, PR17, PN22]. Generative [AASL23]. Generator [BB10, AS24a, SSMM16]. Generic [LLR05, MMH18]. Genetic [HH07, JS08, NN21, RABB08, SVM^{+}00, TBS00, VR08, YTW^{+}06, ZMY^{+}00, FMP^{+}19, MMS20, RR17, RH23, SMSI23, SPC19, TA23]. geographic [HZ19, Yut18]. Geographical [AP04a]. geographically [Ali24]. Geometric [HCH12, SYZS10, Wer09]. geometry [LBYB17]. Geospatial [YWZ^{+}06]. gesture [LTY^{+}24]. Getresponse [WBV04]. GisTool [AP04a]. Glass [NK24]. Global [GHRK11, AL19, JS16, MJ19b, Wu18]. globally [KB16]. glyph [BB18]. Gnutella [SWLH08]. goals [ZWC24]. golden [SSS21b]. good [WGE17]. Google [DB04]. Goql [KPP02]. governance [Pat20, SK18]. GPGPU [HMP20]. GPGPU-based [HMP20]. GPS [AGM23, AMM^{+}17]. Gradient [CS22, EA05, PKW^{+}11, SSP19]. gradients [VP19b]. Grain [HWL07]. Grained [FO09, SGD07, SCOD17]. Granularity [CL10, Kao10]. Graph [EMM24, SSP07, WPJZ08, Bai21, Fu25, ISM25, MSD24, udQ22]. Graph-Based [SSP07, EMM24]. Graphcut [Nom12]. Graphical [KPP02]. graphics [Bay18]. graphs [APS^{+}21, RS16]. Gratings [AFbMSA10]. grayscale [HMP20]. Greedy [SJMO02, SWL07, KBD21]. GreedyBigVis [KBD21]. green [DE17, MV18]. Grid [Ano02a, BH21, CL07, CL10, CL11a, CL11b, CGL12, KB09, PLK^{+}04, YYZ08, HS24, JPJ^{+}19, KR19a, MJ21, SNS18, TA23, YWH19]. Grids [HR04, LYY04]. ground [AB24b]. Group [AK06, CGY^{+}13, HM07, AJP18, XM23]. grouping [Yan19]. Groupware [FR00, Lic06]. growing [ZEZ17]. GRU [AS24a, GK23]. GRVSS [HMP20]. GSM [Gho17]. GTW [BK24]. GTW-EEG-GAN [BK24]. guarantee [ZL18]. Guaranteed [CL07, TT08]. Guarantees [SSO05]. Guard [SD03]. GUI [MMH01]. Guidance [ZW08]. Guidance1 [JSHB10]. Guided [DB04]. Gupta [HC05]. gyroscopic [WT15].

H.264 [YMTbB15]. H.264/AVC

[YMTbB15]. Haar [GP17]. Hadamard [DNP24]. Hadoop [SS20b, Zhu17]. Haematoma [ABMKO14]. Hamming [GMD^{+}24]. Hand [AAD06]. Hand-Printed [AAD06]. handling [MR19b]. Handoff [HC11, DU19]. Handover [EKAO08, VP17]. Handwriting [HEZ05]. Handwritten [NC10]. Hankel [CJW^{+}23]. Hard [AYV10, SSO05]. Hardware [AJAk00, DLDB04, DDL08, RS08, YMTbB15]. Hartley [AKAJ24, MM00, MMM13]. Hartley-Like [MMM13]. Hash [FL04, INR24]. hash-Q [INR24]. Hashing [MK11a, RGNMPM12]. Hawwaz [HEZ05]. HAZOP [MWC^{+}17]. HCDC [ABR10]. HCR [HHGG11]. Head [KYL11, PR19]. heading [FSL17]. headlines [ECBR20]. health [AESBE17, OHZ^{+}23, TS19]. healthcare [BMS24, JCI18, SSS21a, SA24]. Heavy [LGD08b]. Heavy-Tailed [LGD08b]. Hepatitis [GG10]. Hermite [ZA18]. heroin [BB23b]. Heterogeneous [HP04, Lam13, LL11a, SRJS08, KB23, KSD23]. Heuristic [AAA04, SN20, Vee05, WF06, AASA22, MMS20, SPKS24]. Heuristic-based [SN20, SPKS24]. HEVC [TCC^{+}14]. HFDI [BP17]. Hidden [dMM03, Hua18b, RST23].

Hide [CC06]. Hiding [AK06, CCLC10, CPJ17, DD16, HPC18]. Hierarchical [EKAO08, KM09a, MR10, SSP07, HZ19, PNAB18, SK17, WWHW18]. Hierarchies [ZB08]. Hierarchy [Kur07, Lee01, PTLH08, ZC07]. High [BK12, BO12, MR12, PKW^{+}11, Rad04, SEB06, TYBwY17, Tsu07, YW08, ZNLA08, ZYM03, CT15, CSB19, FR18, HWL^{+}18, QDQ24, RRRR17, VV19]. High-Capacity [ZYM03]. high-dimensional [VV19]. High-Level [BO12]. High-Performance [Rad04, YW08, CSB19, HWL^{+}18]. High-Speed [SEB06, ZYM03, RRRR17]. higher [Al 18]. Highly [HM07, MPDK13, CJG18]. histogram [VP19b]. HIV [BH21]. HIV/AIDS [BH21]. HMGA [SK17]. HMGA-ACO [SK17]. HMM [XZCL17]. HMM-based [XZCL17]. HMMN [Rah19]. Hoc [AJS08, CL11b, GT10, HL10, LH12, NS10b, BYBM21, Ogu18, VR19, WZ05, JR12, SN12, WTLW07]. holoentropy [SP18]. holoentropy-enabled [SP18]. home [ZP19]. Hop [Amm07, HL11, TBS00, SN20]. Hopfield [Sam10]. Horn [CDB13]. hospital [SJZ^{+}18]. hosted [LTS25]. Hosting [CB05, LTS25]. hotspots [DZ18]. HSS [RRRR17]. HTTP [KKKV19, SRJS08]. Human [KK13, MLT09, MD14, Nom11, Nom12, AA18, AUHH^{+}18, CDE20, DZ23, MSG18, SMAM25]. Hybrid [ACA07, AAA04, HHGG11, JWC^{+}09, Kim01, KK23b, LL03, PYMD09, SVM^{+}00, SSMM16, SHS^{+}02, AKAJ24, AH18, AS24a, AD18, BA16, BBTH18, BS24, CM20, CXJ24, GMS19, GS19, GS22, GP23, GAS23, HA20, HCC16, IM19, JS19, JN17, JSM19, KR19b, MSG18, MR19b, MMH19, PS19, PW19, PC19, RPS18, SMSI23, SS20d, SPKS24]. hybridized [DSP^{+}18]. Hyfi [LS03]. hyper [BK23]. Hypermap [Nya10]. Hypermedia [Hag03]. hyperparameter [BS25, BH21]. Hypervideo [CCG01].

I/O [Tsu07, VT01]. ICA [CX03]. Icon [AO09]. ICT [Pat20, Pat22a]. IDA [Al 18]. IDE [BS20]. Identification [Nom11, Nom12, Sar13, Sha11, BD24, HZ19, KB22, PBS19, SS19a, VV19]. Identifier [SAAAOH10]. identifiers [KB23]. identify [MSG18]. Identifying [NBS21, NMK23, PW19]. Identity [CGY^{+}13, XLQ09, YHK^{+}10b]. Identity-Based [CGY^{+}13, YHK^{+}10b]. IEEE [BJ07]. IFFOA [PR19]. III [PKM18, PK19]. illegal [Pat23]. illumination [KB16, SLSS24, VP19a]. ILP [IB23]. Image [Abd11, AKAJ23, AP04a, Bai10, BDB14, BJ17, CJW^{+}23, CKC09, CT13, CC06, DS11, HCH12, IAD10, LVK03, LS00, LSC03, LXM^{+}12, MAL10, Oh10, PS06, RGNMPM12, SL03, SX12, SHS^{+}02, SS12, WLZZ12, Wu10, XZ18, AL19, AKAJ24, AB24b, Ano21g, Ano21b, Ano22b, AS24b, BS17, BRB22, CPJ17, CJG18, DD16, DK23, EHA16, EK23, FMP^{+}19, GR19, GMD^{+}24, GSJ24, Guo19, HFG22, HNB24, HPC18, IG23, KB22, KJ22, KR19b, Li20b, MBP23, MPS^{+}23, MB19, MGT19, MS17, NM16, NB24, PGK19, PN24, RPAU17, RRA^{+}23, Ren18, SKS19, SSP19, TTW15, TN24, VP19a, VP19b, WZW20, YZ21, Yao17, YZ19a, ZEZ17, ZFH24]. image-based [GR19]. imagery [CDE20, PKM18, PK19]. Images [LSMI11, LWH12, Sam10, WMB02, ZXT^{+}06, ABB19, BM17, CT15, CGI18, DD16, DR15, DSBR19, ENER23, GMS19, GCMS20, HNH19, HMP20, KB16, KSB19, PSP19, PK20, QDQ24, RDYGE18, SSS21b, SKN18]. imaging [KK24]. imbalanced [BS24, KJ19]. immune [Man15]. impact [LGL20]. imperceptible [IG23]. imperfect [RVS24]. Implementation [AP01, APSK14, BMM^{+}08, CDB13, DB04, GHRK11, PLK^{+}04, PS06, RS08, TS11, WCC06, ZNLA08, AJP18, HFK19, LHZ18, YMTbB15, udQ22]. Implementations [NPY07]. Implementing [WR00, SK24]. implication [PK19]. implications [KK24]. Improve [CTLT04, LLB17, MGT19]. Improved [AAAZ00, CKC09, CPJ17, JJY^{+}10, KLT15, Kao17, LWH12, YLY07, ZLC06, BFB^{+}20, GSJ24, GGB23, HLW^{+}18, PS16, RG23, SGZL20, TD23, WLHZ18, Ye17, PR19]. Improvement [Din11, LXZ08, RGNMPM12, KKR25, KPMR20]. Improving [AA12, CZL^{+}17, CSB19, Day17, HL10, HA20, JSHB10, JCW10, SHP25, AA24, BYBM21]. Impulse [RABB08, ZZ20]. IMSA [Ano00a]. incentives [CDR24]. incidents [Mal24a]. Inconsistent [GTY14]. Incorporating [SOC07]. incorporation [JYN19]. increase [RMF19]. increased [MFK^{+}15]. Increasing [AEH17, EQSMB07]. Incremental [MDY14, WG17, Bai21]. Independent [LS03, AAB19, SSP19]. Indexing [Rad04, BRB22]. Indicator [PF10]. individual [LZL^{+}17, LLL^{+}15]. individual-author [LLL^{+}15]. indoor [MPS^{+}19, SLSS24]. induced [LSDJ23]. industrial [HL24, KPMR20, MMS18]. industries [JN17]. inertial [DZ23, SMAM25]. Infant [SML^{+}13]. inference [KC20, PCD25]. Influence [MS01]. influenza [TS19]. Information [Bai10, CX03, CC06, DK05, JSHB10, Lap05, Lee07, MAL10, PF10, SHS^{+}02, Smi03, SAS07, UU03, VBM02, AL19, Ano21d, AMM^{+}17, CGI18, Dar20, DS19a, FSL17, He18, HZ19, HWL^{+}18, JLA20, Kar19, KR18, KKM17, KR19a, Li19, MZZC18, MP18, Pat20, Pat22a, SJZ^{+}18, VV19, yZ19b, ZSWF20]. information-centric [Dar20]. Information-Theoretic [CX03]. Informative [JGM25]. infrared [RDYGE18]. Infrastructures [MR10, dSTE05, MMS18]. Inheritance [Lee01]. Initial [SJMO02]. Initialization [Sam10, GGMM23]. inland [CSFS17]. innovations [Pat18]. Inpainting [DSV12, HCH12, CJW^{+}23, YZ21]. Input [BSB10, BLDD06, SRJS08, LWG17]. Input-Buffered [BLDD06]. insights [DND25, MLM^{+}24]. inspired [FT20, MGT19]. Instantiation [DB20]. Instruct [CTLT04]. Instruction [YWZJ06]. Instruction-Level [YWZJ06]. Instructional [CLLC10]. Insulators [BK12]. insurance [CMG20]. integral [DRF18]. Integrate [DSD^{+}13, MMH01]. Integrated [HC11, MLT09, SSM12, ZWLH14, BK24, CMG20, NMK23]. Integrating [LH12, SHS^{+}02, AEH17, Men18, RH23]. Integration [CCG01, CGL12, yZ19b]. integrity [BMS24, KSD23, KR19a, SS15]. Intel [SKP09]. Intelligence [Ano00b, PF10, AA24, Ano21j, Ano21a, DU19, Gen19, HSR20, JJS18, MSYG18, SHE^{+}23, WG19, ZY19b]. intelligence-driven [Ano21j, Gen19]. Intelligent [AEDA03, Ano22a, AMM^{+}17, DU19, GG10, HLW04, JS14, LL03, MK11b, SV03, ZY19a, APS^{+}21, Ano21d, BMSP24, BP17, Che20, DVT20, Fu25, Hu17, MPS^{+}19, PN22, PR18, WYZ^{+}19, Yan17, YWH19, Zha17a, ZSWF20]. intensive [LHZ18]. Inter [VNBA03, NN17]. Inter-Rater [VNBA03]. interactions [ABK24]. Interactive [MPĆ15, Smi03, RR17]. Interconnection [AKK10, MKMA07]. interest [DY18, KSS19, KR19b]. interest-based [KR19b]. interesting [MR23]. Interface [Rus06, AS24a, BO18]. interfaces [Kao17]. Interference [WTLW07, RVS24, Zhe19]. Interference-Aware [WTLW07]. Interferometric [SYJ11]. Interleaving [JS16]. internal [HRMA17]. Internet [YZ19a, ASM21, AZK23, BJBA24, BRR17, CMG20, HSR20, Tia18, Yin18b, Amm07, Ano21g, BK01, CHY00, DY18, Dar20, DRF18, FT20, FJK01, HWH03, HY18, HL24, JYHS01, KC19, Lee07, LL18, Mag05, Ogu19, PCK21, PR19, RS08, Sen24, SS19b, TVV^{+}02, Xia18, YCYL04, Ye17]. Internet-of-Things [Dar20, Ogu19]. interoperability [AESBE17]. Interpolation [MM00, Wu10, ASM18]. interpretable [GGK^{+}24]. Interprocessor [NPY07]. Interstate [MLT09]. Interval [RDEDAE09]. intra [KH19, NN17]. intra-day [KH19]. Introducing [MGT19]. Intrusion [BJBA24, TS08, AYI23, AJK23, SK21, YxLmLx17]. invariant [VP19a]. inventive [HWL^{+}18]. inverted [DD16]. investigation [AFLdS25, KKKV19, RWS17]. investigations [AS19, RMG19]. IoENACS [SS19b]. iOS [SK24]. Iostreams [MMH01]. IOT [Li20a, TLC18, AGM23, AASA22, Ano21e, AZK23, MV18, MJ18, MKN^{+}18, NX18, NCGA24, PV18, QD18, RWL19, RH23, RHJ24, RJS18, TK24, ZB20]. IoT-based [AGM23]. IoT-enabled [MKN^{+}18]. IOTs [Al 18]. IoV [FT20]. IP [AAED06, CBL06, KGGJ01, MWC^{+}17, SR06, Sha10, WLC05]. IPv6 [CLBP09, EKAO08]. IPv6-Based [CLBP09]. IPVPNs [HJR08]. Iris [DS10, Nom11, KB22]. irradiance [PS16]. IRS [PKM18]. IRS-1C [PKM18]. ISMAC [RWL19]. isolation [LTY^{+}24]. ISSE [HKE20]. Issue [Ano00b, Ano02a, DF09, YW08]. Issues [BSB10, RCNM03, RPZW17]. item [MR23]. Iteration [CCZC13]. Iterative [LZW07, YYD10, HKE20, Wu18]. Itinerary [PS12]. Ixp2400 [SKP09].

Java [SRJS08, TOG^{+}05, VT01, WHBS01, YHL04]. Java-Based [SRJS08, VT01, YHL04]. Jaya [KR19b]. jitter [MMS18]. Job [MKMA07, SS20b]. Joint [TS15, CDE20, SHE^{+}23]. JPEG [KSB19, MAL10, Oh10]. JPEG2000 [Oh10, CCLC10]. Jump [AAA^{+}11]. Jumps [OW11]. just [VJTG24, ZZ20]. just-in-time-defect [VJTG24].

K2 [BBL23, SG16]. K2PC [BBL23].

KASUMI [MJ19b]. kernel

[PS19, XM23, ZA18]. Kernelized [MR19a]. Kernels [ZNLA08]. Key [Ano21e, AK06, Fan18, HZ20, HM07, MB09, MR10, Pop05, XCCL13, ZB20, ZB08, dSTE05, Ano21f, CKC15, JC19, KC19, TD23, ZLF19]. Key-Hiding [AK06]. Keyboards [GS13]. Keying [AK06]. Keys [LXH07, XGC12, WC24]. Keyword [PK07, XF18, RN17]. KF [BS18]. KF-PSO [BS18]. kidney [SV23]. kinetics [AD18]. Knapsack [HWS09, AH18, IG23]. knee [CDE20]. KNN [AAA17, BJ17]. Knowledge [CW03, ISM25, Lin11, SOC07, APS^{+}21, Bai21, HLW^{+}18, MJ19a, QD18, SJTl25, Yin18a, ZLHW18]. knowledge-based [SJTl25].

L2 [BK23, BKK17]. L2-DWT [BK23]. LAAPS [SS18]. label [VS24]. Laboratory [HWH03, Kar03]. Lack [DP03]. Lagrange [Wu10]. Land [DS19b, SSG06, GMS19, GR19, KR18, PKM18, PK19]. Landslide [ZWLH14]. Lane [WLC10]. Language [Fur07, HWL07, KPP02, KM06, OE12, PG04, Pon04, TN04, CRAV24, KB23]. Languages [SEH05]. Lans [MKH02, TS08, BJ07]. Large [BBAR08, KM09a, MMHJ10, ZFV04, Bad19, He18, HFK19, HM07, KBD21, Men18, WZTH18, XF18, XM23, Yin18a]. large-scale [HFK19, WZTH18]. Latency [AJAk00, HB01, SJMO02, BB18, MMS18]. lattice [WC24]. lattice-based [WC24]. Layer [Bab11, RCJJ06, ARAAMD24, NN17]. Layered [PCS08, WWL^{+}14, MFK^{+}15, SA24]. layout [HC17]. LDPC [PN24]. LDVP [YWDC24].

LDVP-based [YWDC24]. leaders [OMA24]. leaf [PS19]. leaping [MR19b]. Learner [MS03]. Learning [AEDS07, AAD06, CLLC10, Hag03, Kar03, Kom03, KL03, LL11b, MS03, OSKC02, RCNM03, SV03, UU03, VNBA03, dMM03, ABK24, AB24a, AYI23, Ano21b, Bai21, BMS24, BH21, BS24, BBL23, CU19, CDR24, CMG20, DP17, DS19a, DZ23, EhMZ24, ENER23, EMM24, ECBR20, FD23, GGK^{+}24, GSJ24, GP17, HT19, HHU24, IB23, JGM25, JSYT21, KKXB19, KK23a, KKM17, KH19, MP18, MMH19, NMK23, PN20, RST23, SHI24, SMSI23, SS20a, SR19, SPC19, SSS20, SS20c, SK21, SV23, SGMT22, SLQR20, SHP25, SK24, TAUK25, TW19, VJTG24, WZW20, ZX24, ZLHW18]. learning-based [DZ23, GP17, NMK23, SS20a, SPC19, ZLHW18]. learning-to-rank [SHI24]. ledger [BK23]. left [Jij15]. legs [LBYB17]. Leisure [Hsi10]. lending [SMSI23]. Length [LGD08b]. lesion [SK24]. Leukocyte [SR20]. Level [ABR10, BO12, PC11, SCCB10, YWZJ06, AL19, AYI23, CG23, ECBR20, GP17, Kar19, KH19, Xia18]. levels [PN20]. leveraging [SV23, TG24]. lexicon [AAA17, MMH18, MMH19]. lexicon-based [MMH18]. Library [LYR09, PLK^{+}04, yZ19b]. Life [VG14]. lifetime [YPY20]. lightweight [LTY^{+}24, Ogu18, Ogu19, ZFH24]. Like [MMM13, Sop23]. likelihood [BB23b]. Limited [JOYB13]. Limiting [OW11]. LiNbO [SYJ11]. Linear [ABT09, OW11, SHDY10, WXYN13, LBYB17, SKS19, Wu18, Yin18b]. Lines [ABEAA10, GB20]. Link [AAA02, SS10, WVK07, MSK24, TKY23]. Links [SSP07]. lion [AKAJ24, KR19b]. LISS [PKM18, PK19]. list [CDM^{+}21]. literature [AFLdS25, BB23a, BS20, GR19, MdCdAdS24, RvW20, RO24]. live [AA24]. Livelock [SEB06]. Liver [KMR13]. Lmcgrid [YYZ08]. Load [BW07, BBAR08, SP21, Vee05, VTS06, GP23, HKS19, KS17, SSMM16, TA23]. Load-Balancing [Vee05, GP23]. Loads [Vee05]. Local [Bha00, MBS01, MD14, Wu18, BRB22, CZZ18, HFG22, HNH19, MKN^{+}18, PR18, ZX24]. Local-global [Wu18]. Localization [KO00, Sar13, GSP19, TYBwY17, ZX24]. Location [DY18, EJ05, LCHR06, YA06, AMM^{+}17, PCD25, Yut18]. Location-Clustering [YA06]. Location-Dependent [LCHR06]. log [HFK19, SS20b]. log-determinant [HFK19]. Logarithmic [BA07]. Logic [ANB09, APSK14, PKDP12, RDB07, BKK17, LWG17, RMF19]. Logistics [Hu17]. loneliness [YB18]. Lookup [AAED06]. loop [GMS18]. Loss [SLH08, SMT19, ZLF19]. Lossless [LE01, LE02, HNB24, SSP19]. Low [Bai10, CJW^{+}23, DS11, KS14, KYL11, LC07, YYZ08, Kar19, MSH21, PK19]. Low-Complexity [LC07]. Low-Energy [KYL11]. low-level [Kar19]. Low-rank [CJW^{+}23, MSH21]. low-resolution [PK19]. loyalty [SMT^{+}23]. LPM [LD13]. LRU [JH00]. LSB [DD16]. LSK [PC13]. LST [Lin17]. LSTM [SR23b]. LTE [JC19]. lung [MSG18, SR19].

MAC [HHGG11, VG14]. Machine

[AB24a, Che07, FMI11, SML^{+}13, SHE^{+}23, ABK24, AA24, BH21, CU19, CMG20, EMM24, GP17, JC19, KKXB19, KK23a, KH19, LL18, MMH19, SMSI23, SR19, SSS20, SS20c, SK21, SGMT22, SLQR20, SK24, TAUK25, VJTG24, WYZ^{+}19, ZA18]. machine-type [JC19]. Machinery [GA08]. Machines [HC05, LZW07, ZXT^{+}06]. machining [RRRR17]. Machzehnder [SYJ11]. macular [DK23]. Made [NK24]. Magnetic [PKW^{+}11, ZXT^{+}06]. mail [JCYZ18]. Maintainability [KKR25]. Maintenance [CHL01, SH10]. Majority [FSNK22]. makespan [KS17]. Making [TCC^{+}14]. malarial [SCS19]. Malicious [SdSNL06, MWC^{+}17]. Malignancy [LSMI11]. Malware [MPS^{+}23, SS21, SS20c]. mammograms [SJ17]. Man [TS08]. Man-in-the-Middle [TS08]. Management [CO00, FO09, HJR08, HB01, HM07, LYY04, LL07, Lin11, MRU04, PF10, RCNM03, SSG06, YYZ08, ZB08, Ano21h, CZC20, JC19, KSD23, MMPM21, NY17, SBA^{+}22, SCOD17]. Managementin [FO09]. Managing [GM01, BS20]. Manet [HL11, NS10a]. Manhattan [BY03, VP19b]. Manifold [ZX24]. manipulators [LBYB17]. Manual [AB09]. Manufacturing [DJQS04, GA08]. Many [LL14, HS24, SZT18]. Many-Core [LL14, HS24]. many-to-one [SZT18]. Map [Bad19, Din11, SAAAOH10, AUHH^{+}18, DG19, EMN17, JPJ^{+}19, KK23b, RPAU17]. map-based [JPJ^{+}19]. Map-Scan-Reduce [Bad19]. Map-VFS [Din11]. mapped [EHA16]. Mapping [JS08, KK13, MS17, KKM17, RvW20]. Mappings [Rus06]. MapReduce [LHZ18, MR19a, MDY14]. maps [CKC15]. Market [CL11b, DK05, KH19]. Market-Based [CL11b]. Markov [Hua18b, KV17, RST23, VG14, YLY07, dMM03]. Markovian [Che04]. martial [Yao17]. masked [TG24]. massive [Ano21f, HZ20]. Massively [EGL05, FRP07, OL04]. Match [LS00]. Matching [MDY14, MS03, Wer09, Bay18, BA23, CGI18, Di18, GP17, LZ18, Ren18, SZT18]. Materials [SV03, IB23]. mathematical [KR19b]. Matrices [CGC01, GTY14]. Matrix [IAD10, CPJ17, CJG18, HNB24, PS19]. Matters [BA07]. Max [ÖZ11, AH18]. Max-Min [ÖZ11, AH18]. Maximization [VG14]. Maximum [MB09]. MaxSAT [KPSK23]. MCDL [Fur07]. MCFS [SGZL20]. MCMC [SG16]. MDL [FSL17]. meaningful [SC25]. Means [AWA20, KMR13, JN17, XZ18]. measure [BR17]. measurement [MWC^{+}17, SMAM25, ZJZ15]. Measurements [EJ05, MKH02, SR06]. measures [AB24a, VV19, YB18]. Measuring [FM06]. MEC [RVS24]. mechanical [BO18]. Mechanism [AH02, ZYS13, AEA19, Bay18, DM20, HFG22, HL24, MBP23, MKN^{+}18, NMK23, VP17, YZ21, ZDG^{+}23]. Media [GP11, WLZZ12, KKKV19, Xia18]. median [QDQ24]. Mediated [Hag03, VNBA03, YHK^{+}10b]. Medical [BS17, Oh10, PF10, SHS^{+}02, AB24b, BK23, CMG20, KK24, KKA20, MLM^{+}24, NZT^{+}20, PN22, PSP19, SSP19, SSS21b, VP19b]. medium [BYBM21]. meets [JLA20]. Mellin [VR08]. membership [Man17]. Memory [Che07, CJNY^{+}07, MBS01, HLW^{+}18]. Merge [MR10]. Merging [BB08, NB24, SGZL20]. Mesh [JS08, JJY^{+}10, KNS07, MKMA07, ZC07, AA18, Rah19, SPKS24]. Meshes [LHJ07]. Message [Juh06, FT20]. Messages [NK24, DT20]. Messaging [YHL04]. Meta [DB04]. Metadata [GM01]. Metaheuristic [MMS18, KK23b]. metallurgical [Lou19]. Method [CDB13, CTW12, CLBP09, DS10, DG09, HC11, JOYB13, KD19, LD13, MPDK13, MR10, NK24, WXYN13, AL19, ADM21, Ano21g, Ano21a, CM20, Di18, FSL17, GMS18, HY18, Hu17, Hua18a, KJ19, LHZ18, LZ18, Pei18, Ren18, SC25, WYZ^{+}19, fWfH19, XF18, Yan17, YZ19a, YWDC24, Zha17a, ZY19b, ZZ20]. methodologies [AAB19]. Methodology [ZLC06]. Methods [BL01, DK05, Har03, Kao10, LHK03, PYMD09, IB23, Kao17, KB22, KK23a, PGK19, SR19, TD23, ZDG^{+}23]. Metric [AM14, LL11a, YF18]. Metrics [Cur02, GS11, NvV10, SS04]. MG [SZT18]. Micro [MB10, WLZZ12, DZ18, KKR25, Yin18a]. Micro-Architectures [MB10]. micro-blog [DZ18, Yin18a]. Micro-CT [WLZZ12]. Micronetwork [WPJZ08]. Microprocessor [KC10]. microservices [ADM21, KKR25, SS24]. Microsoft [BMM^{+}08]. Middle [TS08]. Middleware [AA12]. midimew [Rah19]. midimew-connected [Rah19]. Migration [Sat08, AA24, AZK23, BZT23]. Min [ÖZ11, AH18]. Minimization [AYV10]. Minimize [LHK03, KS17]. Minimum [SN12]. Mining [YTW^{+}06, APS^{+}21, Ano21e, BKXK19, BD24, HE20, Hou18, Mal24a, MKN^{+}18, MMH19, NZT^{+}20, OMA24, RN17, SC25, TAUK25, Tia18, VV19, Zhu17, ZB20, DY18]. misbehavior [SGMT22]. misconfigurations [MdCdAdS24]. Missing [Yu19]. mission [SAN24]. mission-critical [SAN24]. Mitigated [ABEAA10]. mitigation [RSS24]. ML [RKS^{+}25, SK24]. MMOG [Wu08]. Mobile [AJS08, CL11a, Cur02, EKAO08, HAA05, KBL21, LH12, LL03, Lic06, OE12, PS12, RS04, SN12, SD03, WZ05, WLC05, XYSX10, ZJH08, dMRF^{+}14, Ano22b, Bay18, BSG19, CZL^{+}17, Che20, Dud19, KKM17, Li20b, MJ19b, RMG19, TS15, TW19, VR19, VNR19, Yan18]. Mobile-Edge [KBL21]. Mobilenet [ZFH24]. Mobility [RWL19, KKK22]. MObility-support [KKK22]. Modalities [Oh10]. modality [GSJ24]. Mode [AAA02, EHA14]. Model [AAAZ00, ABMKO14, BSB10, CGL12, HYLS12, HB09, HB01, INyN11, IIDI01, KS07, LYY04, LL11b, LB04a, LYR09, MMCM06, OSKC02, OW11, PSP19, Rad04, RS04, SCCB10, SP14, SdSNL06, Smi03, SWLH08, VT01, WLC10, XY02, YYZ08, Yan18, AJP18, AS24a, AA24, Ano22a, Bai21, BS18, BM17, CM20, CXJ24, CDR24, CMG20, DZ18, DRF18, EMM24, EBA19, EMN17, FT20, GGK^{+}24, GCMS20, HKS19, HY18, JCI18, JXG25, KC20, KKK22, KV17, KK23b, LLB17, LZL^{+}17, LHZ18, MJ18, NMK23, PN20, PCD25, QD18, RV19, RST23, Rui19, SSS21a, SA24, Sha20, SSS20, SBA^{+}22, SK24, Tia18, WG17, XZCL17, YLZ^{+}15, Yan19, ZLF19, ZFH24, ZY19a]. Model-Based [BSB10]. Modeling [BRR17, CZZ18, HDS17, ZWC24, BS24, BZT23, JMMS18, KB22, MZZC18, Oka19]. Modelling [AKK10, BDB14, FJZ06, GA08, HB09, Juh06, KLB06, Mag05, MB10, SEB06, Wu08, XXR09]. Models [CCG01, RW07, Sat08, dMM03, AA18, AS20, BH21, CU19, CRAV24, DND25, GH20, JLA20, JSM19, MPS^{+}23, RKS^{+}25, SV23, SGMT22, TS19]. modes [KSB19]. modifications [NM16]. Modified [LC07, LLH08, MD14, CPJ17, GMS19, HT19, MJ19b, PS16, WhP20]. Modular [MBS01]. module [XM23]. molecular [Man15]. Moment [CS22, SG05]. Moment-to-Moment [SG05]. MOMENTAP [SG05]. Momentum [OS08]. Monitoring [BD11, GP11, HLW04, MKH02, TAD03, ZWLH14, Ano21a, Ano21c, Ano21d, EBA19, FCW^{+}20, HT19, PV18, RV19, RH23, ZP19, ZSWF20, ZY19b]. Monte [TLC18]. monthly [SG16]. MOODLE [MP17]. Morphological [DSV12, DS19b]. Morse [YCYL04]. Motion [CDB13, GG05, MC07, AUHH^{+}18, BKK17, Dua23, DRF18, ZDW^{+}21]. Motor [AR05, PS16]. MOUT [ZTSL08]. move [SSS21b]. movement [KH19, Wan19]. Movie [CS22]. moving [HY18, TLC18]. MPEG [CJNY^{+}07, FJK01, GG05]. MPEG-2 [CJNY^{+}07, GG05]. MPEG-4 [CJNY^{+}07]. MPI [PLK^{+}04, Tsu07]. MPI-I [Tsu07]. MPI-I/O [Tsu07]. MPLS [ACGP10]. MPPT [PS16]. MPS [WLZZ12]. MQTT [AFLdS25, DT20]. MR [QDQ24]. MRI [BRB22, DSBR19, HHU24]. MSVM [KJ22]. Mtens [ACGP10]. MTTF [CZZ^{+}23]. MuDeLA [AYI23]. Multi [AKAJ23, Amm07, CL07, DSV12, Dua23, FMI11, GP23, GTY14, HL11, HYS11, HB09, HWL07, JS08, JYYJ08, KJ22, KMGS11, KMCV09, LS10, LQLX11, MSH21, MBP23, MFK^{+}15, MGK^{+}00, PC11, Rui19, SMT19, SHDY10, TLC18, TT08, fWfH19, Yin18a, YHK^{+}10a, AJJ18, AYI23, AH18, BS25, BSG19, BS18, DB20, GP17, IB23, KS21, KSD23, KH19, Li20a, MPS^{+}19, Men18, QD18, RS19, RHJ24, Ren18, SSMM16, SK18, SN20, VS24, WL18, WhP20]. multi-angle [QD18]. Multi-Camera [KMGS11]. Multi-Channel [KJ22]. Multi-Class [LQLX11]. Multi-classification [Yin18a, Li20a]. multi-cloud [BSG19]. multi-cloudlet [RS19]. multi-constraint [SSMM16]. Multi-criteria [SMT19]. Multi-Domain [TT08]. multi-factor [SK18]. Multi-feature [Rui19, Ren18]. multi-floor [MPS^{+}19]. Multi-Formalism [HB09]. Multi-FPGA [SHDY10]. Multi-frame [TLC18]. Multi-Grain [HWL07]. Multi-Hop [Amm07, HL11, SN20]. multi-label [VS24]. Multi-layered [MFK^{+}15]. Multi-Level [PC11, AYI23, GP17, KH19]. Multi-Objective [GTY14, JS08, GP23, BS25, RHJ24]. Multi-Paradigm [HWL07]. Multi-Path [JYYJ08, fWfH19, KS21]. Multi-Pitch [FMI11]. Multi-QoS [CL07]. Multi-Receiver [KMCV09]. Multi-scale [Dua23, MBP23]. multi-scenarios [AH18]. Multi-Secret [HYS11]. multi-server [AJJ18]. Multi-Signatures [YHK^{+}10a]. Multi-source [MSH21, Men18]. multi-step [BS18]. Multi-Structure [DSV12]. multi-SVMs [WhP20]. Multi-Threaded [MGK^{+}00]. multi-threading [WL18]. multi-time [Dua23]. multi-user [KSD23]. multi-viewpoints [DB20]. Multi-wavelet [AKAJ23]. Multiagent [VGPB09]. Multibody [FJZ06]. Multicast [Beg06, EKAO08, HCSC06, HM07, NS10a, NS10b, ÖZ11, PCS08].

Multicast-Favourable [ÖZ11].

Multicasting [AJS08, AK06]. multiclass [KJ19]. Multicomputers [GG04]. multicriteria [MP18]. Multidatabases [RS04]. Multidimensional [HY00, SCCB10, KBD21, PRR20, Tia18]. multifactor [GAS23]. multilabel [MP18]. multilabel-multicriteria [MP18]. multilayer [DMB23, EHA16]. Multilevel [Dai19, WCC06, Di18]. multilingual [KB23]. Multimedia [ACA07, AAA04, Fur07, HB01, KGGJ01, LLRW07, MBS01, OSKC02, VBM02, SKS19, ZL18]. Multimodal [AR19, SV23, CZZ^{+}23]. multiobjective [ABB19]. Multipath [HL10, LH12, WTLW07]. Multiplayer [FRP07]. Multiple [AJAk00, CL10, LH13, Li19, MK11a, MS01, Nya10, PBS22, WWL^{+}14, WXYN13, ZDG^{+}23, ABK24, GB20, HCC16, JSM19, MBS^{+}24, NN17, PW19, WG19]. Multiple-Layered [WWL^{+}14]. multiplicative [QDQ24]. Multipliers [LC07]. Multipoint [FR00, SVM^{+}00]. Multiprocessor [DTDE06, LH13]. Multiprocessors [Che07, CJNY^{+}07, LS03]. Multiresolution [ZC07]. Multispectral [SX12]. multitenant [MT24]. Multiversion [XYSX10]. Municipal [WXYN13]. mutant [GSP19]. mutation [Man15]. mutual [DS19a, KPMR20, MKA^{+}18]. MUX [GMD16]. MVC [AP01]. MVC-Based [AP01]. Mykil [HM07]. Myrinet [MPDK13]. Myrinet-Connected [MPDK13].

Naive [JWC^{+}09, JCW10, AEDS07].

Naive-Bayes [JWC^{+}09]. NAM [JMMS18]. named [KSS19, YNA23]. Nanoparticles [PKW^{+}11]. nanoscale [OIB17]. nanotube [OIB17]. NARMA [BKK17]. NARMA-L2 [BKK17]. National [dSTE05]. Natural [Sam10, SKN18]. naturally [BS24].

Navigation [AR05, FSL17, JSM19, RJS18]. NCDC [GH20]. Near [SVM^{+}00, SLQR20, BS17]. Near- [SVM^{+}00]. near-fuzzy [BS17]. Nearest [JWC^{+}09, AAH25, JMMS18, MMS20]. neighbor [AAH25]. Neighborhood [Che20, BRB22, KBL21]. Neighbour [JWC^{+}09, KMCV09, MMS20]. Neighbourhood [SS12]. NEM [Mag05]. Nemo [HC11]. Nets [LH00]. Network [AAA02, BY03, Bha00, BMM^{+}08, CBL06, DK05, EKAO08, HL10, HC11, INyN11, KYL11, KB10, LH12, LVK03, LJL^{+}19, LE02, Mag05, MK11a, OL04, PC11, PTLH08, Sam10, SSP07, Sar13, SKP09, Sha10, VG14, XY02, Yan19, ZXF07, ZJH08, AJK23, AASL23, Ano21a, Ano21i, Ano22a, BS18, BN24, CSFS17, Che20, DU19, Di18, Dua23, FSNK22, Fu25, GBBZ19, HFG22, HZ19, Hou18, Hua18a, ISM25, JSYT21, JC19, KV17, KH19, LLB17, LC20, MSG18, Mal24a, MWC^{+}17, MBN^{+}24, Men18, ND18, OOG^{+}18, Oka19, PBM16, PCD25, Pei18, PC19, RPZW17, Rah19, RMG19, RV19, SG16, SN20, SS20d, SP21, SCS19, SPKS24, WYZ^{+}19, Wu18, Yan17, Yan18, YNA23, Yin18b, YPY20, YxLmLx17, Zha17a, ZY19a, ZY19b, Zhe19]. network-on-chip [GBBZ19]. Network-Routing [BY03]. networking [KSS19]. Networks [ACGP10, AJS08, AKK10, Bab11, Beg06, CHC01, CEJ12, CLBP09, GT10, GHRK11, HHGG11, JR12, Kar12, KNS07, KGGJ01, LLR05, LXM^{+}12, LM02, LS03, LYY^{+}08, MS01, MB09, MKMA07, MMHJ10, NS10b, OS08, SAD09, SEB06, SS04, SBS12, SN12, Sha10, SS19b, SD03, SRJS08, TT08, WZ05, WTLW07, WLC05, ZJR08, AFLdS25, AASL23, AB24b, Ano21j, BA16, BYBM21, BB23a, BP17, CZL^{+}17, EHA16, EBA19, Gen19, IM19, JNS18, KS21, KPMR20, LSDJ23, MMS18, MG19, MSD24, MSK24, Ogu18, OMA24, PC22, PR18, RK24, RHJ24, SS18, SC19, SN20, SBA^{+}22, SN23, TKY23, TS15, TG24, TK24, VP17, VR19, VNR19, WC24, fWfH19, YPY20, ZX24, udQ22]. Networks* [KMCV09]. Neural [KB10, LXM^{+}12, LE02, LM02, NC10, OS08, Sam10, ZXF07, AJK23, AB24b, Ano21i, BB23a, BN24, EHA16, FSNK22, Fu25, JSYT21, KH19, LC20, LSDJ23, Mal24a, MMS18, MBN^{+}24, MSD24, OOG^{+}18, PBM16, RV19, SCS19, WC24, WYZ^{+}19, XM23, udQ22]. neuro [KC20]. neuro-fuzzy [KC20]. Neuron [ZXF07]. Neutral [Sha11]. news [APS^{+}21, ECBR20]. Newsfeeds [FJK01]. Newsgroup [AWA20]. Next [TVV^{+}02, CZL^{+}17]. Next-Generation [TVV^{+}02]. No [RABB08]. No-Reference [RABB08]. NoC [JS08]. NOCs [JYYJ08]. Node [SG05, VR19, DWT23, Yan19]. nodes [Ano22a, TYBwY17, WGE17, YMRLJ17, ZY19a]. Nodule [LSMI11]. Noise [LWH12, RABB08, Wan11, ARAAMD24, PN20, QDQ24, WLHZ18, Zhe19, ZZ20]. Noisy [Kat07, MR19b, PBM16]. NOMA [RVS24]. NOMA-MEC [RVS24]. Non [SBS12, SR19, WLL10, BJ21, CSB19, HFG22, PR17, SKS19, SAN24]. Non-Classical [SBS12]. non-factoid [BJ21]. non-functional [SAN24]. non-linear [SKS19]. non-local [HFG22]. Non-Parametric [WLL10]. non-recursive [PR17]. Non-small-cell [SR19]. non-volatile [CSB19]. Nonblocking [LYY^{+}08]. Nonlinear [EHA14, FJZ06]. nonparametric [HY18]. Normalization [RGNMPM12]. normalized [DS19a]. noSQL [ACM20]. Note [Ano05, HYS11]. noticeable [ZZ20]. Novel [Abd11, AEDS07, GA11, HC11, IEZA10, KM09a, LYY^{+}08, MK11b, SSM12, SD03, TS08, AS24a, Bad19, DMB23, GSP19, IB23, JXLZ15, KBL21, LLB17, MMS20, MJ19b, NMK23, NNDP22, SSP19, SCM18, SPKS24, TN24, YWDC24, ZEZ17]. NSCT [SX12]. Nt [MKH02]. Nt-Based [MKH02]. nucleus [SR20]. null [Don18]. numerical [BO18].

O [Tsu07, VT01]. OA [Yu19]. OALO [PN20]. Object [AEHSES08, AAA04, CHL01, DK05, FA06, FWQ00, HYT^{+}05, Lee01, LLR05, LV06, MDY14, OL04, PG04, Pon04, Pon05, SCCB10, Sat08, SEH05, TN04, VNBA03, Wer09, ZMY^{+}00, CM20, PBM16, RPS18, SCOD17, YF18]. Object-Oriented [AEHSES08, CHL01, DK05, FA06, FWQ00, HYT^{+}05, Lee01, OL04, Pon04, SCCB10, SEH05, SCOD17, YF18]. Objective [GTY14, JS08, Lap05, BS25, GP23, RHJ24]. Objects [HWS09, Wan06, CZZ18, JPJ^{+}19]. Obsolescence [FRP07]. Obsolescence-Based [FRP07]. obstacle [RvW20]. occurrence [HNH19]. OCDM [AFbMSA10]. OCL [RW07]. OCL-Specifications [RW07]. OCR [DJQS04, KB09]. OCT [DK23]. OFDM [PYMD09]. Off [MMPM21]. Off-chain [MMPM21]. Offering [RCJJ06, SP14]. office [Yu19]. Offline [HEZ05]. offloading [KBL21]. often [Pat18]. older [YB18]. on-arrival [NN21]. On-Chip [LHK03]. One [HWS09, SZT18]. Online [CRAV24, Di18, FRP07, LLC11, Tia18, APS^{+}21, Bai21, CBXh18, RV19, SAM23]. Only [ZKS05]. ONoC [ZJR08]. onto [DDL08]. Ontologies [MJ18, WGE17]. ontology [AESBE17, BS19, BHZGG22, DS19a, DB20, Fan18, JPJ^{+}19, KKM17, MZZC18]. ontology-driven [BS19]. Open [DK05, HE20, PKD18, PKM18]. Operating [DHSR05, Lap05]. Operation [DDL08, LWG17]. Operations [Loo04, DS19b, HL24]. Operative [Ano00b]. Opinion [MMH19, OMA24]. opinions [Xia18]. opportunistic [SN20]. opportunities [NBS21]. Optical [OS08, SYJ11, ZJR08, GCMS20]. Optimal [Bab11, BY03, CMG20, MJ21, PBS22, SVM^{+}00, Sar13, VR08, Al 18, ABB19, RK24, RR17, SS19a, TYBwY17]. Optimistic [FRP07, XYSX10]. Optimization [CL10, GMS19, GTY14, He18, JNS18, KZ07, LQLX11, PR19, SML^{+}13, YWZJ06, AKAJ24, ASM18, AASA22, AD18, Ano21g, AZK23, BS25, BH21, Che20, FMP^{+}19, GP23, HLW^{+}18, HC17, Hua18b, IG23, INR24, JS16, KK23b, LJL^{+}19, MBP23, MR23, Mal24a, MJ21, RRRR17, RHJ24, SZT18, SK17, SPC19, SSS21b, SR23a, SPKS24, TS15, VNR19, WWHW18, Wu18, XZ18, Yan18, Ye17, Yin18b, YZ19a, yZ19b, ZLHW18]. Optimization-based [GMS19, IG23, WWHW18, Wu18]. Optimized [YTW^{+}06, BBL23, SR23b]. Optimizer [CS22]. Optimizing [BO12, BA23, BB18, Bha00, RH23, TBS00, KD19, SK24]. Optimum [CGH11, KC19]. Optional [Rus06, AJJ18]. Oracles [XLQ09]. ORB [OA01]. OrBAC [BM18]. orchestration [DE17, OOG^{+}18]. order [GS22, MR19b]. Organisms [NZT^{+}20]. Organizational [KL03]. Organized [NS10b]. organizing [PBM16]. Orientation [DSV12]. Orientation-Driven [DSV12]. Oriented [AEHSES08, CHL01, DK05, FA06, FWQ00, HYT^{+}05, Lee01, OL04, PG04, Pon04, SCCB10, SEH05, BRR17, CMG20, LZL^{+}17, ST16, SCOD17, VP19b, YF18]. OSRAD [FM18]. Osteoarthritis [CDE20]. OTIS [ASA05]. OTIS-Cube [ASA05]. out-of-band [SCM18]. outlier [CGC17, SR23b, SR23a, ZJZ15]. outlook [AAB19]. Overhead [AP04b, Bai10, LHK03]. Overlap [LL11a]. overloaded [KPSK23]. overview [KK23a, MSYG18, RPZW17]. OVSF [SS10].

P2P [KB09, Lin11, SMSI23, SAS07].

Package [WMB02]. Packet

[AAED06, BLDD06, MMS18, NN17]. paddy [ZDG^{+}23]. pads [BO18]. Page [AEDS07, DSP^{+}18]. Pages [Ani14, Kim01]. Pairing [LXH07, XGC12]. Pairing-Based [LXH07, XGC12]. Pairings [LXH08]. Pairs [ANAKS18]. Pairwise [Wer09]. Pakistan [SJZ^{+}18]. Palmtop [HB01]. Panchromatic [SX12]. pancreatic [PBS19]. pandemic [Pat20, Pat22a]. Panorama [ZW08]. Papers [Ano00a, Ano00b, Ano02a, YM01]. PAPR [PYMD09]. Paradigm [HWL07, VCP00, KS21]. Parallel [AH02, AAAZ00, CDB13, CJNY^{+}07, Dar09, EGL05, Haq06, KC10, LL14, LC07, Loo04, MM00, MRU04, NPY07, OL00, OL04, PKW^{+}11, PS06, Rad04, Tsu07, WR00, ZFV04, LBYB17, PBM16]. Paralleling [ABEAA10]. Parallelism [HWL07, LS03, YWZJ06]. Parallelization [MPDK13]. Parallelized [SNS18]. parameter [SPKS24]. Parameters [Bab11, VR08, DRF18, ND18, RRRR17]. Parametric [WLL10, AA18, BO18]. parametric-based [AA18]. parasite [SCS19]. parents [BBL23]. parents-children [BBL23]. Parma [AJS08]. partial [KPSK23, YZ21]. Partially [CLX10, DJQS04, KK18, PW19]. Particle [GTY14, LQLX11, FMP^{+}19, JS16, MJ21, SPC19, TLC18, XZ18, Ye17]. Partition [Loo04]. Partition/Sorting [Loo04]. Partitioning [VTS06, YP05]. Party [XCCL13]. pass [RMG19]. pass-codes [RMG19]. Passing [Juh06]. Passive [BHB09, Lee07, MMCM06, RvW20]. Password [XCCL13, Dud19]. Password-Based [XCCL13]. patch [LZXB24]. Path [JYYJ08, JOYB13, KNS07, LS10, TBS00, TT08, GMS18, HZ19, KS21, RK24, fWfH19, YYM17]. pathological [PSP19]. Paths [JOYB13]. pathway [MZZC18]. Patient [OE12, SLSS24, BM18]. Pattern [LYR09, MD14, BRB22, CBXh18, DD16, SS24, VR19]. Patterns [AEHSES08, SEH05, TN04, ABK24, SJTl25]. payload [IG23]. payload-efficient [IG23]. payment [Pat23]. PC [GG05, LGD^{+}08a, LCZD02, ZFV04]. PC-based [LGD^{+}08a, LCZD02]. PCA [DA16, PGK19, SX12]. PCM [HL11]. PCNet [YZ21]. Peak [ZLHW18]. peaks [Bad19]. peculiar [BM17]. pedestrian [FSL17]. Peer [Lin11, PKDP12, Wu08, SS18]. Peer-To-Peer [Lin11, PKDP12, Wu08, SS18]. penetration [AFLdS25, MP17]. People [AR05]. Perceived [NvV10]. perception [LL18, ZJZ15]. perceptron [EHA16]. Perceptual [RGNMPM12, Fu25]. Performance [AA12, AAA02, Ali24, AKK10, BP17, CTLT04, Din11, EGL05, GP11, HAA05, HL10, JV12, Kao10, LB04a, LE01, MR12, MKH02, MMHJ10, PR18, QDQ24, Rad04, RK15, SSP07, VP17, Vee05, YW08, ZNLA08, AA24, CSB19, HWL^{+}18, KK23b, LLB17, OB15, PS16, SLSS24, WL18]. performance-aware [KK23b]. Performance-Driven [Kao10]. permutation [MB19]. permutation-based [MB19]. Personal [CHC01]. personality [BB23b]. Personalized [LZ18, DSP^{+}18]. Personalizing [Lee07]. Persons [YCYL04]. perspective [Pei18, SK21]. Pervasive [ZKNL10]. Petri [LH00]. Phase [Kat07, LD13, JS19, PC19]. Phish [LTS25]. phishing [EMM24, LTS25]. phone [WT15]. Phonetic [SOC07]. photos [LGSS18]. phrase [SC25]. phrases [SC25]. Physical [ARAAMD24, Bab11, Oka19]. Physical-layer [ARAAMD24]. physiological [ND18]. Pillar [AWA20]. pipeline [EBA19]. Pipelined [MB10]. Pipelines [ABEAA10, AJAk00]. Pitch [FMI11]. Pixel [SS12, ABB19, EHA16]. Placement [AAAZ00]. plagiarism [Kar19]. Plan [IIDI01]. Planning [Sha10, CPKT18, RK24]. plant [Hua18b, NMK23, PS19]. Platform [LL03, LGD^{+}08a, MPDK13, AMM^{+}17, Fu25, MKN^{+}18, NN21, Zhu17]. PM [BK23]. PM-ECC [BK23]. point [SMT^{+}23]. Points [AB09, Don18]. points-to [Don18]. pointwise [DS19a]. Policers [SR06]. Policy [BLDD06, CL11b]. pollination [AM16]. Polynomial [YWZJ06]. Pool [MB09]. pools [LCZA17]. Population [YLY07, AJJ18]. Porous [WLZZ12]. Position [MS17]. positioned [BM17]. Positioning [Yin18b, MPS^{+}19, MKN^{+}18, Wu18]. positive [ZJG17]. Possibilistic [KMR13]. postgraduate [HWL^{+}18]. potential [Don18, SN20]. Power [AJS08, HB01, KS14, KHVW04, MK11a, MS17, PC11, SS04, SHLA17, SHDY10, MJ21, SD17, SAM23, YWH19, Zhu17]. Power-Aware [PC11, SS04]. powered [SMT^{+}23]. Ppctsi [WMB02]. PPFM [MS17]. Practical [Chi08, RPZW17]. Practice [LL07]. practices [Dud19]. PRAM [BB08]. Pre [FA06, MB09, KB23, MPS^{+}23, SK17, SAM23]. Pre-Defined [FA06]. Pre-Distribution [MB09]. pre-processing [SK17]. pre-trained [KB23, MPS^{+}23, SAM23]. precise [Ren18]. Precision [MKN^{+}18, TYBwY17]. Predict [GS11, AMM^{+}17]. Predictable [YP05]. Predicting [YB18, ABK24, BB23b]. Prediction [ABEAA10, AD18, AAA^{+}11, Che04, KC20, RRRR17, WVK07, WXYN13, ZLF19, AS20, AA24, Ano21j, BS25, BH21, CMG20, DZ18, EhMZ24, FD23, GH20, GKMR19, Gen19, ISM25, JN17, JSM19, KH19, MSG18, Mal24a, Man15, RN17, SHE^{+}23, SR19, SS20b, SV23, SR23a, Tia18, VJTG24, YLZ^{+}15]. Predictive [CU19, DP17, NZT^{+}20, TKY23, XZCL17]. predictive-based [TKY23]. predictor [SSP19]. Preface [Ano18, DF09, YW08]. prefer [Pat23]. Preliminary [DGR09, MMHJ10]. preparing [KBD21]. Preprocessing [DY18, ACM20, KJ19, MP18]. prescriptive [MP18]. prescriptive-based [MP18].

Present [PJ19]. Presentation [FJK01]. preservation [ASM18, KR19a]. preserved [VV19]. Preserving [BM18, RABB08, BD24, BSG19, Ogu18, Ogu19, SA24, ZJZ15]. presume [BM17]. preventing [Ano21h, CZC20]. Prevention [SdSNL06]. previously [PKD18]. Pricing [FM06]. primary [ARAAMD24]. Principal [SML^{+}13, SC25, DS19b, GGMM23]. principles [AJP18]. Printed [AAD06]. prior [LZXB24]. Priorities [ANB09]. Prioritization [SSM12]. Prioritizing [RST23, NA19]. Privacy [BD24, ASM18, APZ17, BM18, Bad19, BSG19, BS24, OHZ^{+}23, Ogu18, Ogu19, SA24, VV19]. privacy-centric [OHZ^{+}23]. Privacy-preserving [BD24, Ogu18, Ogu19]. private [SSS21a]. Proactive [LL07, DE17, LXZ08, KKK22]. Proactively [BS20]. Probabilistic [SSO05, BJBA24, HNB24, JXG25]. Probabilities [Man17]. Probability [SP18, IQ22]. probability-based [IQ22]. Probability-weighted [SP18]. Problem [AAA04, AD00, Dar09, SWL07, TZ07, AH18]. Problems [WF06, AI18, LHZ18, SSMM16]. Process [LZW07, MZZC18, VHL11, VBM02, DU19, GSJ24, OW11, RRRR17]. Processing [AP04a, DS10, KHVW04, LCHR06, LVK03, LXM^{+}12, Nya10, OL04, Vee05, Ano21g, Ano21b, Ano21f, DWT23, HZ20, LGL20, SK17, WZW20, YZ19a]. Processor [CTLT04, MKMA07, PR08, SKP09]. Processors [YWZJ06, WL18]. Product [BD11, GB20]. production [CPKT18]. products [Lou19]. professional [HWL^{+}18]. profile [BA23]. Profit [SWL07]. prognostic [BTZA18]. Program [Shi13]. Programming [HYT^{+}05, KC10, SIT11, SRCXLC10, YTW^{+}06]. Programs [Haq06, MGK^{+}00, WHBS01, MMPM21, SCOD17]. progress [SJTl25]. progressive [HS24]. Project [FO09, VHL11, Hua18b]. Prolog [MGK^{+}00, Rus06]. PROMO [KKK22]. proneness [RR17]. propagation [WZTH18, WWHW18]. Properties [ITK08, KB10, PW02]. property [Don18]. Proportional [SLH08]. Proposal [KJ03]. prospects [SJTl25]. protecting [HRMA17]. Protection [TT08, APZ17, AS24b, Bad19, BD24, DNP24]. protein [AD18]. Protocol [AAA02, AL06, Beg06, DLDB04, HL11, HHGG11, MMCM06, PT14, PCS08, Pop05, RS08, RCJJ06, SN12, SG05, VG14, XYSX10, XCCL13, ZLC06, BYBM21, DT20, JXLZ15, JC19, KS21, KPMR20, MSK24, PCK21, PNAB18, RK15, YPY20, Yut18]. Protocols [Amm07, ASA04, BCP09, KYL11, LB04b, DG19, HDS17, MG19, SC19, WQX^{+}15]. Prototyping [FR00]. protract [YPY20]. Provable [XCCL13, Ogu18, Ogu19]. Provably [XLQ09]. provenance [AWO^{+}18]. Providing [RMF19]. Provision [Dua07]. Provisioning [HJR08]. Proxy [AL06, CBL06]. Prune [LZW07]. pruning [LSDJ23]. pseudo [INR24]. pseudorandom [PR17]. PSO [BS18, BBL23, KC20, SSMM16]. PSO-ACO [SSMM16]. PSO-based [KC20]. PSO-K2PC [BBL23]. psychoactive [ABK24]. psychology [FT20]. psychology-inspired [FT20]. PTS [PYMD09]. Public [Chi08, LXH07, MR10, XGC12, dSTE05, SS15, Xia18]. Published [Lap05]. Pulse [LXM^{+}12]. Pupil [DS10]. Pure [PBM16]. PV [SD17]. pyramid [BRB22].

Q [INR24, SN23]. QoE [KKKV19]. QOS [LS10, Alm11, Beg06, CL07, Cur02, FM06, HJR08, NS10b, SN12, WTLW07, ZKNL10, ZL18]. QR [MFK^{+}15]. Qrobot [LCZD02]. QSD [HNB24]. Quadrant [ZNLA08]. Quadtree [TCC^{+}14, CDE20]. Qualitative [GA08]. Quality [BD11, Dua07, Oh10, SCCB10, SSM12, BYBM21, DD16, KB22, Sen24, SKS19, SHLA17, XZCL17, KKKV19].

Quality-of-Experience [KKKV19]. quantifying [SCOD17]. Quantitative [Lin17, PGK19]. Quantization [LS00, JXG25, TTW15]. quantization-based [JXG25]. Quantized [CTW12]. Quantizer [CKC09, TN24]. Quantum [KYPS06, GMD16, PBM16, KK18]. Quasi [Alm11]. Quasi-Synchronous [Alm11]. Queries [Bha00, LCHR06]. Query [KPP02, OL00, OL04, Rus06, SHS^{+}02, BHZGG22, INR24, KKSR22, LZ18]. Quest [MM24]. question [BJ21, GK23, PN22, SHI24, SJTl25]. questions [APZ17, BHZGG22]. queueing [AJJ18]. Quorum [PT14].

R [ABT09, JMMS18]. R-Tree

[JMMS18, ABT09]. Radial [LE02]. Radio [KMCV09, ARAAMD24, DU19, EhMZ24]. Radiographs [Nom12]. Radiology [SHS^{+}02]. radiomic [SR19]. Radon [RGNMPM12]. rail [Ano21i, LC20]. rain [HFG22]. Rainfall [ZXF07, SG16]. Random [LL11b, RDEDAE09, XLQ09, AR19, EK23, HS24, LLL^{+}15, RR17, ZZ20]. random-grid [HS24]. random-valued [ZZ20]. randomized [HMP20]. Range [EJ05, WZ05, CT15]. ranging [TYBwY17]. rank [CJW^{+}23, MSH21, MSG18, SHI24]. Ranking [JR12]. Rapid [LLH08, WMB02]. Rapidly [LLC11]. Rapsody [OSKC02]. Rate [CTW12, DS11, VR19]. Rate-Distortion [CTW12]. Rater [VNBA03]. Rates [PBA10]. rating [SMT19]. rationality [Wan19]. Ray [Nom12, BM17, CDE20]. Rayleigh [Bab11, GT09]. Re [AK06, BBTH18, Har03]. re-composition [BBTH18]. Re-Keying [AK06]. RE-USE [Har03]. reactive [AH18]. Readability [DJQS04]. Readers [DJQS04]. Reading [PT14, ZKS05]. readings [KR19a]. Real [AYV10, Ano22b, Bay18, BA03, BA07, CO00, DHSR05, LCHR06, Lap05, Li20b, LH00, MWC^{+}17, OA01, Pon04, RPS18, SSO05, SMF12, SR06, VR08, WLC10, YMTbB15, YP05, dMRF^{+}14, Aa20, CGC17, HDS17, HRMA17, KPSK23, KPMR20, OB15, RK24, SLQR20, ZLF19]. Real-Scenario [SR06]. Real-Time [AYV10, BA03, CO00, DHSR05, LCHR06, Lap05, LH00, OA01, Pon04, SSO05, SMF12, WLC10, YP05, Ano22b, Bay18, Li20b, RPS18, YMTbB15, Aa20, CGC17, HDS17, HRMA17, KPSK23, KPMR20, OB15, RK24, SLQR20]. Real-world [MWC^{+}17]. Reality [LZ07, dMRF^{+}14, dMM03]. Realization [MMM13]. Reasoning [LM02, JPJ^{+}19]. Receive [SEB06]. Receiver [KMCV09, SN11]. Recognition [AO09, HEZ05, KMGS11, LLR05, MD14, NC10, SML^{+}13, SS05, SOC07, Ano21i, Ano22b, AUHH^{+}18, CBXh18, DZ23, Dua23, ECBR20, GT19, GP17, HCC16, Li20b, LC20, LLL^{+}15, MMS18, Men18, PS19, SLSS24, SMAM25, TG24, ZDG^{+}23]. recognize [Ano21b, WZW20]. Recognized [CTLT04]. Recognizing [AAD06]. Recommendation [Kim01, KM09c, MS03, TW19, ANAKS18, CXJ24, DSP^{+}18, DZ18, Li19, QD18]. Recommendations [CS22]. recommender [BB23a, JLA20, SMT19]. Reconfigurable [CBL06, DDL08, MR12]. Reconfiguration [HYT^{+}05]. Reconstruction [IQ22, KGEL06, WLZZ12, AA18]. record [AESBE17, KR18]. records [OHZ^{+}23, TS19]. Recovery [AEHSES08, PN24]. Rectangles [CGH11]. rectifying [Gho17]. Recurrent [Kat07, KH19]. Recursive [CGH11, PR17]. Reduce [Bad19, SJMO02, Ano22a, KK23b, ZY19a]. Reduced [Fur07]. Reducing [AP04b, RR17]. Reduction [BW07, CS22, KAKF11, Lam13, PYMD09, SRCXLC10, WC24]. redundancies [LSDJ23]. refactoring [NBS21, SSS20]. Reference [RABB08, PGK19]. Refined [LHJ07]. Region [KR19b, NB24, Yut18, ZEZ17]. Regions [AAY02, NA19, PW19]. Registration [Wer09, BJ17, KB16]. Regression [HL10, LL11b, Shi12a, WXYN13, YTW^{+}06, RST23]. regular [AI18, PV18]. regularized [ZJZ15]. Regulating [EHA14]. Rehearsed [AO09]. reinforcement [DP17, RST23, SS18]. Related [CS09, DK23, LCZA17]. relation [BR17, Lin17, SMT19]. relational [MYA16]. Relationship [DSD^{+}13, JPJ^{+}19]. Relationships [Lin11]. relays [SN20]. release [SSS21a]. released [PKD18]. ReLeC [RHJ24]. ReLeC-based [RHJ24]. relevance [INR24]. relevant [Lin17]. Reliability [PKD18, RCJJ06, SAN24, VNBA03, XXR09, BRR17, CZL^{+}17, Oka19, Yan17, Zha17a]. Reliable [AJS08, AMJ12, Bai10, MSK24, VR19]. reliefF [AR19]. remnant [RWS17]. Remote [GSV08, HLW04, KHVW04, Tsu07, Ano21a, Ano21b, Ano21d, DS19b, ENER23, WZW20, ZP19, ZSWF20, ZY19b]. Removal [RABB08, HFG22, ZZ20]. removing [KSB19]. rename [NBS21]. Rendering [AC06, Day17]. rendezvous [VNR19]. Repairing [GTY14]. repeated [CPJ17]. repeating [FR18]. Replacement [AL06]. Replicated [PT14]. Replication [AH02, LV06, RG23]. report [PSP19]. Reporting [KLRH13]. repositories [HE20]. Repository [SGD07, VHL11]. Representation [MD14, AUHH^{+}18, Kar19, MSH21, PS19]. reproduction [CT15]. Reputation [SP14]. Request [SJMO02]. Requests [AAA04, NBS21]. Requirement [SSP07, SAN24]. Requirements [BL01, SSM12, BS19, SKM20, WDW17]. rerouting [TKY23]. ResDenseNet [HHU24]. Research [BSB10, CGC17, HWH03, Hua18a, LL07, LTY^{+}24, RCNM03, Ren18, YZ19a, ZWLH14, Zha17a, Zhu17, Ano21g, CPKT18, GMS18, HSR20, Hou18, KKXB19, PJ19, Yao17, YNA23]. Reservation [CGC01, HHGG11, SBS12, WLC05]. reserve [SHE^{+}23]. Residual [TCC^{+}14, Dua23, NMK23, XM23]. resiliency [SS24]. Resilient [LD13]. Resnik [DS19a]. resolution [GGB23, HS24, KLT15, KKS18, PK19, SSP19]. Resolver [PKDP12]. Resonance [ZXT^{+}06]. Resource [CO00, CGC01, CL07, CL11a, CL11b, Dua07, SWL07, WLC05, ZYS13, AASA22, AZK23, DB20, HDS17, KBL21, LJL^{+}19, NY17, RVS24, yZ19b]. resources [AKB^{+}19, DE17, KKM17, KS17, SZT18]. response [RS19]. Restorable [KNS07]. restricted [Ali24]. Results [JSHB10, BH21]. Retina [LCHR06]. retinex [CT15]. retinopathy [MBN^{+}24]. Retracted [CZC20, FCW^{+}20, Gen19, HZ20, Li20b, LC20, WZW20, YZ19a, ZSWF20, ZY19a, ZY19b, ZB20, Ano19, Ano21j, Ano21g, Ano21a, Ano21c, Ano21i, Ano21b, Ano21d, Ano21h, Ano21f, Ano21e]. Retraction [Ano19, Ano21e, Ano22a, Ano22b, Ano21j, Ano21g, Ano21a, Ano21c, Ano21i, Ano21b, Ano21d, Ano21h, Ano21f]. Retransmission [PBA10]. retrial [AJJ18]. Retrievability [SP14]. Retrieval [JS14, Lee07, PW02, Pon05, SHS^{+}02, Wan06, BRB22, HNH19, INR24, Kar19, KR18, KKM17, NX18, Rui19, VP19a, VP19b, WG17, WG19, Yan18]. Reusability [Lic06]. Reusable [SV03, UU03]. Reversible [APSK14, PSP19, GMD16, HPC18]. review [AEEO23, BS20, DG18, DM20, EhMZ24, GR19, KK24, MG19, MdCdAdS24, NPN^{+}23, NY17, RO24, SC19, SAM23, ST16, SP21, TAUK25, udQ22]. Revisited [ABT09]. Revitalizing [MM24]. revocation [CG23]. RF [Gho17, MMCM06]. RFID [AGM23, MKA^{+}18, ZKS05]. rigid [CZZ18]. Ring [JJY^{+}10, OHZ^{+}23]. Risk [LSMI11, LZL^{+}17]. risks [Ano21h, CZC20]. road [Di18, RvW20]. robin [Dai19]. Robinson [CDM^{+}21]. Robot [KK13, LCZD02, BKK17]. Robotic [RCJJ06, BTAB21]. Robots [Ano00b, ZYS13]. Robust [JCI18, KMGS11, PN24, SOC07, Wer09, FM18, GGMM23, GP17, IG23, RV19, SLSS24]. robustness [AEH17, GMD^{+}24]. role [Ano21h, CZC20]. Rood [PC13]. room [ZDW^{+}21]. rotation [VP19a]. Rough [SRCXLC10, Mal24b, Man17]. round [Dai19]. round-robin [Dai19]. Route [ZW08, VR19]. Routed [ZJR08]. Routers [GH06, ZYM03]. Routing [AAAZ00, ASA05, BY03, Beg06, JYYJ08, LS10, LYY^{+}08, NS10a, NS10b, RS08, SVM^{+}00, SN12, WZ05, WTLW07, YNA23, ZJR08, DP17, EBA19, GBBZ19, KS21, KV17, MG19, MSK24, ND18, PNAB18, RK15, Sen24, TS15, VNR19, fWfH19, Yut18]. RSA [YHK^{+}10b]. RST [LD13]. RTS4MTS [MT24]. RTTSMCE [RS19]. rule [BKXK19, BD24, RN17]. rules [Hua18a]. Run [GP11]. Run-Time [GP11]. Runtime [TOG^{+}05].

S [DG18]. S-boxes [DG18]. Sa [KC10].

Sa-C [KC10]. SaaS [MT24]. Safe

[DHSR05, WHBS01]. safety [NPN^{+}23].

Saisense [MK11b]. saliency

[EMN17, Lou19, RPAU17]. salp [SR23a]. SAMA [Aa20]. sampling [LLL^{+}15, YxLmLx17]. Sat [Dar09]. Satellite [KGGJ01, GR19]. satisfaction [TA23]. satisfaction-based [TA23]. SBCDetector [GS22]. SBMT [HCSC06]. Scalability [MBS01]. Scalable [HR04, Kao10, LB04a, MPDK13, MK11b]. Scale [KM09a, Dua23, HFK19, MBP23, Pat22b, VP19a, WZTH18]. Scaling [JWC^{+}09]. Scan [Bad19, WMB02]. Scanning [BSB10]. scavenging [Gho17]. Scenario [BL01, SR06, MPĆ15]. Scenario-Based [BL01]. scenarios [AH18, DA16, ND18, ZDG^{+}23]. schedule [Li17]. Scheduled [FA06]. Scheduler [NvV10]. Schedulers [ZYM03]. Scheduling [AAA04, AJAk00, CL07, GT10, GHRK11, HH07, HC05, HR04, JH00, Lam13, MKMA07, RHR06, SNS18, SJMO02, VT01, Ano21j, Dai19, Gen19, GGB23, Hua18b, KPSK23, KKK22, KK23b, MJ21, NN17, NN21, PC22, RS19, RG23, SZT18, Zha17b]. Schema [KJ03, Lee01, MDY14]. Schema-Free [MDY14]. Scheme [AFbMSA10, ABR10, AK06, Bai10, CTLT04, CCLC10, CCZC13, CGY^{+}13, CL10, DS11, HYS11, HYLS12, JYYJ08, LXH07, LXH08, LXZ08, MB09, NS10a, NS10b, SD03, Tan11, TS08, WCC06, XGC12, YHK^{+}10b, ZB08, CKC15, CJG18, GSJ24, HKE20, HPC18, HA19, IAT20, Ogu18, Ogu19, RRA^{+}23, SK18, SCM18, TKY23, IM19]. Schemes [BJ07, CHC01, SS10, MB19, TK24]. Schunck [CDB13]. SCI [Mal24a]. scientific [KRMS19]. scoring [HKE20]. Scour [AAA^{+}11]. scrambling [GSJ24]. screening [GCMS20, WYZ^{+}19]. Script [KM06, SS05]. sea [AKAJ24]. Seamless [ZW08]. Search [DNP19, DB04, DSD^{+}13, JSHB10, LL14, OL00, PC13, PK07, Aa20, AH18, AAA17, BH21, Fan18, HZW^{+}19, JXG25, KBL21, MBP23, SK17, SS18, XF18, XM23, NZT^{+}20]. Searcher [SGD07]. Searcher-Browser [SGD07]. Searching [CKC09, SAS07]. second [GS22]. second-order [GS22]. secondary [ARAAMD24, OB15]. Secret [Bai10, CC06, HYS11, KK18, NK24, DD16, HMP20, HS24]. Secure [AP04b, AK06, BK23, HYLS12, HM07, LH12, PSP19, Pop05, PBS22, TS11, WBV04, XLQ09, YWZ^{+}06, YHK^{+}10a, YHK^{+}10b, BSG19, JC19, KR19a, MMPM21, RRA^{+}23, SA24, SCOD17, SMT^{+}23, WQX^{+}15]. Secured [CJG18, MSK24]. Securing [CEJ12, HL24, JCI18]. Securities [WWL^{+}14]. Security [BCP09, DG19, GAS23, LB04b, MP17, PS12, XGC12, XCCL13, ARAAMD24, AWO^{+}18, Ano21h, APZ17, CZC20, HA20, KC19, Li20a, Mal24a, MdCdAdS24, MFK^{+}15, NN17, Ogu18, Ogu19, RPZW17, RSS24, RO24, TK24, VLK19]. security-aware [NN17]. Segmentation [CS09, JKRS02, KMR13, Nom12, Rus06, SL03, SS05, AL19, AB24b, DSBR19, KR19b, LZXB24, MBP23, NB24, SR20, XZ18, ZEZ17]. seismogeology [LCZA17]. seismogeology-related [LCZA17]. Selecting [Lap05]. Selection [Amm07, Bab11, BLDD06, JOYB13, KYL11, PR08, PR19, SKM20, VR08, ABB19, AR19, Di18, FSNK22, HZW^{+}19, JS19, KSS19, LLL^{+}15, Mal24b, MT24, MV18, SMSI23, SS19a, SHE^{+}23, SV23, TD23, VR19, VS24, WhP20, Zhe19]. selector [IG23]. Self [LH12, LXH07, LYY^{+}08, NS10b, NB24, OSKC02, SKP09, XGC12, Hao17, PBM16, PN24]. Self-Authentication [LH12]. Self-Certified [LXH07, XGC12]. Self-Configurable [SKP09]. self-embedding [PN24]. Self-Organized [NS10b]. Self-Routing [LYY^{+}08]. self-sensing [Hao17]. Self-sparse [NB24]. Self/Collaborative [OSKC02]. Self/Collaborative-Learning [OSKC02]. Semantic [AM14, KM09c, OL04, PW02, Sha20, Shi13, AESBE17, AA18, DSP^{+}18, KKM17, LZXB24, MJ18, NNDP22, NA19, PJ19, WG17]. semantic-based [KKM17]. semantics [SHP25]. Semi [CLTY17, GA11, HB01, JSHB10, WWHW18]. Semi-Batch [HB01]. Semi-Statistical [GA11]. Semi-Structured [JSHB10]. Semi-supervised [CLTY17, WWHW18]. Semiconductor [DJQS04]. SemTraClus [NA19]. Sensing [CCZC13, Ano21b, DS19b, EhMZ24, ENER23, Hao17, PR18, WZW20]. sensitive [BD24, VNR19]. Sensitivity [Loo04]. Sensor [Bab11, CEJ12, CGL12, HHGG11, KYL11, Kar12, MB09, SAD09, SS04, VG14, Ano22a, BS18, EBA19, IM19, JNS18, KS21, KV17, KPMR20, MMS18, MG19, MV18, NCGA24, PC19, RPZW17, RK24, SC19, SS20d, TK24, VNR19, fWfH19, Wu18, Yin18b, YPY20, ZX24, ZY19a]. sensors [AS19, CBXh18, DZ23, SMAM25, WT15]. Sentence [BB10, HKE20]. sentiment [AS20, AAA17, AR19, KKXB19, KD19, MMH18]. separate [KSB19]. Separation [CX03, PKW^{+}11]. Sequence [Nya10, Hou18, PR17, PW19, WDW17]. Sequences [AJAk00, GA11, SSP19]. sequency [IG23]. Series [KLB06, BN24]. Server [ACA07, BW07, CO00, Che04, DWT23, GM01, MBS01, Shi12a, ZC07, AJJ18]. Server-Side [Shi12a]. Service [BJ07, CEJ12, Dua07, GP11, KZ07, KM09c, MS01, ZYS13, AJJ18, BRR17, BYBM21, Fan18, He18, JYN19, MS19, Mal24b, NN17, Sen24, Sha20, ST16, WG17]. service-oriented [BRR17]. service-oriented-architecture-based [ST16]. Services [BK01, CGC01, CHY00, FR00, Lee07, Sha10, SWL07, WL07, AMM^{+}17, Bar18, BBTH18, BTAB21, CZL^{+}17, JYN19, KKKV19, DB04]. Sessions [CS09]. Set [Rus06, SRCXLC10, AL19, BS17, WZTH18, WhP20, Xia18]. Set-Of-Mappings [Rus06]. Setrequest [WBV04]. sets [MR23, Mal24b]. Several [LE01, WCC06]. shallows [BS24]. Shape [SLZZ10]. Share [PBS22, DD16]. Shared [Che07, CJNY^{+}07, MBS01, TT08]. Shared-Memory [Che07, CJNY^{+}07]. Sharing [Bai10, BBAR08, FR00, HYS11, DD16, HMP20, HS24, OHZ^{+}23, KK18]. Sheep [ABMKO14]. shifting [TA23]. ship [HC17]. Shop [WF06]. Shops [HC05]. Short [KYPS06, SGZL20, Yin18a]. Shots [MC07]. Shouted [Sha11]. Shows [JKRS02]. Shuffled [MR19b]. Sib [YYD10]. Side [LS00, Shi12a]. Side-Match [LS00]. Signal [AS09, CX03, BK24, Kao17].

Signature [CGY^{+}13, HYLS12, XLQ09, YHK^{+}10b, OHZ^{+}23]. Signatures [CLX10, FL04, YHK^{+}10a]. Signcryption [LXH07, LXH08, LXZ08, PBS22, XGC12]. significant [Mal24a]. similar [HNH19]. Similarity [HCH12, BR17, CLTY17, GS22, LSDJ23, QD18]. similarity-induced [LSDJ23]. Simple [CC06, SN11, XCCL13, MP18]. Simplex [MPDK13]. Simulated [AAA04, SWLH08, YYD10]. Simulation [ANB09, BK12, HP04, INyN11, SC10, ZFV04, BZT23, CZZ18, CPKT18, HC17, LBYB17, Yao17]. Simulator [ACGP10, Che07, MMHJ10]. Simulators [GT09, dMM03]. Simultaneous [SLH08, WL18]. Single [DLL05, TTW15, TBS00, WLC10, MM24]. Single-Address-Space [DLL05]. Single-Hop [TBS00]. singular [GMD^{+}24, GGMM23]. sink [VNR19]. situation [LZL^{+}17, MJ19a]. situational [JPJ^{+}19]. Size [BJ07, BA07, MB09, OHZ^{+}23]. Sized [HWS09]. skeletal [AA18]. Sketcher [CH10]. Ski [AAA^{+}11]. Ski-Jump [AAA^{+}11]. skin [GGK^{+}24, SK24]. SKIP [KJ22]. SLA [AEA19]. Slackness [AJAk00]. Slackness-Based [AJAk00]. Slc [CS03]. Sliding [EHA14]. slipping [BO18]. SLTC [VP17]. SLTC-D2D [VP17]. Small [AC06, Juh06, SR19]. Smart [Al 18, AKB^{+}19, Hua18b, dSTE05, Ano21f, DG18, GT19, HZ20, KR19a, MMPM21, MMS18, SK18, SMT^{+}23, TA23, WT15, YWH19, ZP19, YLZ^{+}15]. smartphone [DZ23]. Smooth [KGEL06, CJW^{+}23]. smoothed [Dai19]. SMP [HWL07]. SMP-Cluster [HWL07]. SMRT [IG23]. SNMP [WBV04]. SOA [LZ18]. Social [LZL^{+}17, RWL19, OMA24, PCD25, RMG19]. Society [BK01]. SoCs [WPJZ08]. Soft [AAA^{+}11, Jin10, SKS19]. Software [AA12, AJAk00, CB05, FWQ00, GS11, Har03, KHVW04, LH00, LYR09, DF09, MRU04, PW02, RO24, SSM12, SH10, WMB02, YWZ^{+}06, Bar18, BS25, FD23, Fu25, HE20, LCZA17, PKD18, RR17, RST23, RS17, SKM20, SSS20, Sop23, SP21, TKY23, VP17, YF18, ZL18]. software-defined [SP21, TKY23, VP17]. sold [RWS17]. Solid [WXYN13]. Solution [Lam13, YTW^{+}06, He18]. solutions [AI18, NPN^{+}23, udQ22]. solve [SSMM16]. Solver [EA05, SHDY10]. Solving [AD00, Dar09, GG04, AH18, LHZ18, LBYB17]. some [Pat23, RCNM03]. Sorting [Loo04]. Sound [KO00, Don18]. Source [Kar19, HE20, MSH21, Men18, PKD18]. sources [WG19]. SP [MP18, SBA^{+}22]. SP-DDPT [MP18]. SP-TRUST [SBA^{+}22]. Space [BA07, DLL05, DDL08, KC10, Lam13, Ano21c, FCW^{+}20, ZX24]. spacecraft [RS17]. Spam [DMB23, SLQR20, VLK19]. Spams [AS20]. span [RK15]. Spare [KNS07]. SPARQL [BHZGG22]. Sparse [EA05, MD14, NB24, PS19]. Spatial [DR15, LCHR06, SSG06, YP05, HFG22, He18, ISM25, Lin17, NX18, WLHZ18]. Spatial-contextual [DR15]. spatial-temporal [ISM25]. Spatio [PC13, SMF12, NA19]. Spatio-Temporal [PC13, SMF12, NA19]. Speaker [Sha11, SS19a]. Special [Ano00b, Ano02a, DF09, YW08]. specific [AAA17, BS19]. Specification [BCP09, TN04]. Specifications [RW07]. Spectral [LQLX11, MR19a, SPKS24]. Spectrum [AS09, Wan11, DU19, EhMZ24, SPKS24]. Speculation [JOYB13]. Speculative [JOYB13, Pon05]. Speech [KAKF11, SOC07, GS19, SS19a]. Speed [AYV10, SEB06, ZYM03, JXG25, RRRR17, SBA^{+}22, YMTbB15]. Speeding [WhP20]. SPHMMS [Sha11]. SPIHT [PN24]. Spillways [AAA^{+}11]. Splitting [Wan11, TKY23]. Stability [LLH08, RV19]. Stack [NK24]. stacking [BB23b]. stadium [Li17]. staging [HHU24]. stand [SD17]. stand-alone [SD17]. Standard [HYLS12, MPDK13, AEH17]. Star [JR12]. start [BB23a]. State [BSB10, LS00, MS19, SM10, SMAM25, WVK07, MMPM21, PBM16, SJTl25, SCOD17, VR19, WDW17]. State-Based [SM10]. State-of-the-art [SMAM25]. state-tracking [MMPM21]. Stateless [SdSNL06]. Statement [Ano21j, Ano21g, Ano21a, Ano21c, Ano21i, Ano21b, Ano21d, Ano21h, Ano21f, Ano21e, Ano22a, Ano22b, PJ19]. states [AD18, KK18]. Static [AEHSES08]. Statistical [GA11, KLB06, TD23, AL19, AB24a, CPKT18]. Steel [LM02, RRRR17]. steerable [BRB22]. steganographic [CJG18]. Steganography [MAL10, NK24, PSP19, WWL^{+}14, ABB19, FR18, KR19a, MS17]. stego [DD16, HPC18]. stego-image [HPC18]. Steiner [HCSC06, SVM^{+}00]. step [BFB^{+}20, BS18]. stereo [CGI18]. Stochastic [SSO05]. stock [KH19]. Storage [BMM^{+}08, MMHJ10, VT01, BSG19, BK23, CSB19, RWS17, SS15, YLZ^{+}15]. Storages [SP14]. stores [MYA16]. stories [ZWC24]. Storing [ACM20]. Strategies [Vee05]. Strategy [AEDS07, GHRK11, LH12, OL00, ANAKS18, LLL^{+}15, RG23, SAN24, SGZL20, YxLmLx17]. Stream [GYQC14, Dai19, SP18]. Streaming [SC10, ZW08, KKKV19, PRR20, ZLF19]. Streams [SMF12, CGC17, YWDC24]. Street [BY03]. strength [GKMR19]. Stress [BK12, JN17]. Strictly [LYY^{+}08]. Structural [AM14, AAD06, CS03, Kar19]. Structure [AFbMSA10, AAED06, DSV12, LB04b, BBL23, Di18, HFG22, Man15, Yao17]. Structured [JSHB10]. structures [EK23]. Student [TS11, KKSR22, Tia18]. Students [Kar03]. Study [ASA04, DJQS04, DGR09, DHSR05, GT09, KO00, Loo04, MU09, PBA10, SC10, TAD03, BO18, EHA16, GH20, GMD16, HFK19, LCZA17, MWC^{+}17, MP17, OMA24, SJZ^{+}18, VJTG24]. Studying [FD23]. Subdivision [KGEL06, LHJ07]. SUBSCRIBE [DT20]. Subsections [CS09]. subsequent [BB23b]. Subspace [SOC07, VV19, XM23]. Subspace-based [VV19]. substance [ABK24, BB23b]. substances [ABK24]. substitution [DD16]. Subsystems [VT01]. successive [RVS24]. Suggestion [PK07]. suitable [ACM20]. summarization [HKE20, IB23, MLM^{+}24]. Summation [RDB07]. super [HS24, KLT15, Pei18, SD17]. super-capacitor [SD17]. super-resolution [HS24]. supervised [CLTY17, LZXB24, SS20c, SGZL20, WWHW18]. supply [KSD23, RO24]. Support [ACGP10, DAC03, FMI11, HWL07, KHVW04, LZW07, OSKC02, PF10, SML^{+}13, SSG06, ZXT^{+}06, FSNK22, He18, KKK22, MP18, WYZ^{+}19, ZA18]. Supported [FJK01]. Supporting [CLLC10, FO09, KL03, Lic06, SGD07, VBM02, YP05, dSTE05]. Surface [BK12, KGEL06, ZMY^{+}00]. Surfaces [BD11, MU09]. Surgical [RJS18]. surround [CT15]. surveillance [TS19]. Survey [DK23, KYPS06, LE01, SS15, AESBE17, AWO^{+}18, BKXK19, BD24, BB23a, BJ21, MV18, SMAM25, SK21, YNA23]. suspected [ZJG17]. suspension [WL18]. Suspicious [KLRH13]. sustainable [HSR20, RK24, YWH19]. SVD [JLA20]. SVM [CXJ24, DZ18, JS19, Li20a, NMK23, NC10, QD18, ZDW^{+}21]. SVMs [WhP20]. Swarm [GTY14, LQLX11, MSYG18, DU19, FMP^{+}19, JS16, MJ21, SPC19, SR23a, XZ18, Ye17]. swarms [SN23]. Switch [BLDD06, MS01, OOG^{+}18]. Switched [TS08]. Switches [Dua07, SYJ11, ZYM03]. Switches/Routers [ZYM03]. Switching [HR04, LYY^{+}08, OW11]. Sybil [BMSP24].

Symbiotic [NZT^{+}20]. Symbolic [KM06]. Symmetric [RDB07, SYJ11, ZNLA08, DNP19, HNH19]. Synchronization [DTDE06, FRP07, AS19]. Synchronous [Alm11]. Syntax [SHP25]. Synthesis [AI18, BO12, HH07, KS14, PC11, RDB07, Sop23]. synthetic [BS24]. System [AYV10, AH02, AFbMSA10, AA06, AR05, AEDA03, CO00, CT13, CLLC10, CJNY^{+}07, CL10, DLL05, FA06, FH05, GG10, Hag03, HEZ05, HYT^{+}05, Hsi10, HM07, HLW04, IAD10, JH00, Jin10, KO00, LZ07, LCZD02, MLT09, PF10, PKDP12, PR08, PBS22, SS05, SM10, SSG06, SSM12, Shi12b, SAS07, TBS00, TS11, Tsu07, VGPB09, Wan06, WL18, XXR09, YHL04, YWZ^{+}06, ZFV04, ZXF07, ZJH08, ZKNL10, ASGW20, AJJ18, ANAKS18, ARAAMD24, Al 18, ACM20, Ano21d, Ano22b, AEA19, Bai21, BB18, DNP24, FSNK22, GT19, GP17, Hao17, HT19, He18, HS24, HA20, HCC16, HWL^{+}18, JCI18, KC20, KRMS19, KKA20, Li20b, MZZC18, MPS^{+}19, MJ19b, NNDP22, PN22, RVS24, RS17, SSMM16, SR20, SJZ^{+}18, SMT^{+}23, SPKS24, Yu19, ZP19, ZSWF20, ZL18, Zhu17]. System-On-Chip [CJNY^{+}07]. Systematic [GR19, MdCdAdS24, SGZL20, BS20, RO24, TAUK25]. systematized [MG19]. Systems [AL06, Alm11, AMJ12, BBAR08, Bha00, CB05, DHSR05, EA05, FJZ06, Har03, INyN11, KAKF11, KM09a, KM09b, KM06, Lam13, Lap05, LH13, LB04a, LV06, Pon04, PYMD09, RCJJ06, SS10, SSO05, SJMO02, SS19b, VTS06, VBM02, YA06, AYI23, BJBA24, BB23a, BJ21, CSB19, EhMZ24, JLA20, KPSK23, KB22, MSYG18, NPN^{+}23, PKD18, SD17, SAN24, SMT19, SD23, YLZ^{+}15, YWH19]. Systolic [ITK08, MM00, MMM13].

Table [AAED06]. tables [MYA16]. Tabu [LL14]. Tackling [KB23]. Tag [Shi12b]. Tag-Based [Shi12b]. Tags [ZKS05]. Tailed [LGD08b]. Taiwanese [Hsi10]. Talk [JKRS02]. Talking [Sha11]. tamper [PN24]. tangent [ZX24]. Target [BS18, SC19, Ano21b, HY18, LL18, WZW20]. Task [AH02, CS09, Lam13, RG23, SNS18, GGB23, JGM25, KKK22, NN21, RS19, SHI24]. Task-Related [CS09]. Tasks [FA06, AZK23, SZT18]. taxonomy [RSS24, WQX^{+}15]. TCP [KGGJ01, PBA10]. TCP/IP [KGGJ01]. tea [NMK23]. Teacher [OSKC02]. teaching [SPC19]. Technical [KZ07, SSM12]. Technique [AYV10, AAED06, IEZA10, AS24b, BMSP24, DMB23, FMP^{+}19, HZ19, HA20, KSB19, KV17, KR19a, LCZA17, MP18, MJ21, PR17, PC19, SSMM16, Sen24, SHE^{+}23]. Techniques [AH02, AAD06, Ano00b, AAA^{+}11, CCG01, CH10, CT13, DTDE06, JS14, KAKF11, KS07, SL03, YA06, AFLdS25, AA24, AZK23, BMS24, BD24, CRAV24, DG18, EhMZ24, GR19, NY17, RSS24, RKS^{+}25, Rui19, ST16, SK21, SLQR20, TAUK25, VJTG24]. Technological [Pat18]. Technologies [JSZ09, Ano21f, Ano21e, HZ20, PJ19, ZB20]. Technology [JSZ09, Lin11, YWZ^{+}06, AJK23, Ano21h, Che20, CZC20, DY18, MZZC18, Pat18, Pat22a, RJS18, RDYGE18, YLZ^{+}15, Yao17, Ye17, YWH19]. Telecommunications [LLH08]. telemedicine [FM18]. Telemetry [LE01]. Telerobotic [LCZD02]. temperature [CZZ^{+}23, PS16]. Template [GP17]. Temporal [PC13, SMF12, Sat08, GK23, ISM25, Lin17, NX18, NA19]. term [AWA20]. term-cluster [AWA20]. Terminal [GSV08, HB01, Yan18]. Terminals [Chi08]. Termination [TCC^{+}14, JXG25]. terms [AAA17]. Tertiary [VT01]. TESLDAR [MSK24]. Test [AB09, EKAO08, HH07, LWG17, LH00, BH21, MT24, MP17]. Test-Bed [EKAO08]. Testable [RDB07]. Testing [SM10, Shi12a, AFLdS25, Fu25, Gho17, MP17, RR17, RST23]. Text [ABR10, KAKF11, SHS^{+}02, AEEO23, BA23, FR18, HKE20, Hua18a, MLM^{+}24, MMS20, SC25, SGZL20, WG19, Yin18a]. Text-Based [SHS^{+}02, BA23]. Text-To-Speech [KAKF11]. texts [LLL^{+}15]. texture [DR15, EMN17, HNH19, RPS18, SKN18, VP19a]. their [AS20, LBYB17, NvV10]. Theoretic [CX03, MAL10]. Theory [AKAJ23, AKAJ24, CSFS17, PNAB18, YMRLJ17]. Thermal [MSK24]. Thermal-energy-efficient-secured-link [MSK24]. things [ASM21, AZK23, BJBA24, BRR17, CMG20, Ano21g, Dar20, HL24, KC19, Ogu19, PR19, Ye17, YZ19a]. Thread [CJNY^{+}07, WL18]. Thread-Parallel [CJNY^{+}07]. Threaded [MGK^{+}00]. threading [WL18]. threat [JPJ^{+}19, RSS24]. Three [CKC09, CHL01, XCCL13, ZWLH14]. Three-Dimensional [ZWLH14]. Three-Party [XCCL13]. threefold [AAH25, AS24b]. threefold-ensemble [AAH25]. Threshold [HYS11, HYLS12, LXZ08, LSC03, ND18, RDB07, Tan11, XLQ09, YHK^{+}10b, ZZ20]. Threshold-based [ND18]. Thresholding [WCC06, PGK19]. Throughput [GT10, RDEDAE09, Tsu07]. Thyroid [LSMI11]. Ti [SYJ11]. Timbre [SS19a]. Time [AYV10, BCP09, BN24, BA03, BA07, CO00, DHSR05, GP11, KLB06, KM09b, LCHR06, Lap05, LLH08, LH00, OA01, Pon04, SSO05, SMF12, VG14, WLC10, YP05, ASM21, Aa20, Ano22b, Bay18, BB18, CGC17, Dua23, HDS17, HRMA17, KPSK23, KS17, KPMR20, Li20b, OB15, PV18, RS19, RK24, RPS18, SLQR20, VJTG24, YMTbB15, ZLF19]. time-bound [ASM21]. Time-Constrained [KM09b]. Time-Dependent [BCP09]. Timely [Alm11]. Times [SSO05]. Timetabling [RHR06]. Token [SR06, LZXB24]. token-based [LZXB24].

TokenLink [SMT^{+}23]. Tolerance [HY00, JJY^{+}10, ZJR08, KRMS19, RMF19]. Tolerant [AH02, ASA05, CBL06, KM09b, WL07]. tomography [GCMS20]. tone [CT15]. Tool [MKH02, RHR06, KKS18]. Tools [DSD^{+}13, Kom03, TAD03]. topic [AWA20, AS20, SGZL20]. topologies [SD17]. Topology [EHA14, Mag05, SWLH08, LSDJ23]. TOPSIS [SKM20, VS24]. TOPSIS-ACO [VS24]. Torus [AKK10, JYYJ08]. Torus-Based [JYYJ08]. tournament [HZW^{+}19]. Trace [Che07, Lou19]. Trace-Driven [Che07]. traces [AMM^{+}17]. Track [LC20, Ano21i]. Tracking [EJ05, AGM23, BS18, DRF18, Hu17, MMPM21, MWC^{+}17]. trading [ANAKS18]. Traffic [ACGP10, Dua07, Gen19, KGGJ01, PC22, SBS12, TKY23, AASL23, Ano21j, Ano21b, Di18, Hao17, HT19, HZ19, ISM25, PN20, WZW20]. Traffic-aware [PC22]. trained [KB23, MPS^{+}23, SAM23]. Training [OSKC02, OS08, HWL^{+}18]. traits [BB23b]. trajectories [AUHH^{+}18, NA19]. trajectory [NX18]. Transaction [RS04]. Transactions [AP04b]. Transfer [AAA02, ENER23, HHU24, JSYT21, SK24]. Transform [AKAJ23, MM00, MMM13, RGNMPM12, Ano21g, BRB22, DNP24, GS19, GSJ24, Guo19, SHLA17, YZ19a, NPY07]. Transform-Based [RGNMPM12, DNP24]. Transformation [KC10]. Transformations [KS07, HNB24]. transformative [DND25]. transformer [BN24, CZZ^{+}23, CRAV24, DND25, LZXB24, ZDG^{+}23, ZFH24]. transformer-based [CRAV24, DND25]. transformers [MBS^{+}24, MLM^{+}24]. Transforming [RW07]. transistors [OIB17]. transit [Ano21i, LC20]. Transition [SG05]. translation [MYA16]. Translator [OE12]. Transmission [ABEAA10, BK12, WZ05, ZKS05, Ano21j, Bay18, FT20, Gen19, ND18, VR19]. Transmission-Only [ZKS05]. Transmissions [Kar12]. Transmitter [SN11]. Transmitter/Receiver [SN11]. Transport [KGGJ01, RCJJ06]. Transportation [CHY00]. Trapezoidal [AAY02]. travel [AMM^{+}17]. treasure [MJ19a]. Tree [ABT09, AAD06, HCSC06, LB04b, SP18, JMMS18, SHP25]. treed [OW11]. Trees [LL11b, SVM^{+}00]. trend [Wan19]. trending [BTZA18]. trends [JJS18]. Triangulation [ZMY^{+}00]. Triple [PC19]. Troubles [Shi12b]. Trust [LYY04, SP14, YMRLJ17, DSP^{+}18, FT20, HL24, LZL^{+}17, SBA^{+}22, VR19, SBA^{+}22]. trust-based [DSP^{+}18]. trust-rate [VR19]. Tucker [CJW^{+}23]. Tumour [KMR13]. tuned [BKK17, CRAV24, MPS^{+}23, MBS^{+}24, MLM^{+}24]. Tuning [AYV10]. Turbo [MR12]. Tutoring [AEDA03, Smi03, TAD03]. TWE [SGZL20]. tweets [AAM^{+}18, ACM20]. Twilight [ZTSL08]. twins [HL24]. Twitter [AS20, PCD25, SLQR20]. Two [GG04, JOYB13, KB22, KS07, SYZS10, CM20, DG18, ECBR20, HC05, JS19, YB18]. Two-Dimensional [SYZS10]. two-level [ECBR20]. Two-Path [JOYB13]. two-phase [JS19]. two-way [CM20]. Type [YHK^{+}10a, BS17, JC19, KK18]. type-2 [BS17]. typical [PV18].

UAV [SN23]. UAVs [JPJ^{+}19]. Ubidata [ZJH08]. ulcer [BJ17]. Ultrasound [LSMI11]. UML [KS07]. unbalanced [YxLmLx17]. Unbounded [HWS09]. uncertain [MR23]. uncertainty [KB22]. underground [MKN^{+}18]. underlay [ARAAMD24]. Understanding [TK24]. Underwater [HHGG11, LGD^{+}08a, AS19, PK20]. UNet [ENER23]. Unified [YHL04]. uniformed [JXG25]. Unit [APSK14, BO12, SMAM25]. Units [GSV08, QD18]. Universal [CLX10].

Unleashing [SAM23]. unlocking [TD23]. Unsafety [ASA05]. unstructured [SS18]. Unsupervised [YLY07, BMS24, DR15]. Unveiling [DND25, BS24]. Unwrapping [Kat07]. Update [WVK07]. Updates [Lee01]. Upon [TOG^{+}05]. Urban [PK19, PN20]. URLGBM [EMM24]. Usage [LYR09]. USB [RWS17]. Use [AO09, LH13, RW07, BB23b, DS19b, PK19, Har03]. Use-Case [LH13]. used [YMTbB15]. User [GS13, TA23, AS24a, APZ17, AMM^{+}17, BA23, CG23, DY18, GAS23, JCI18, KSD23, NX18, SK18, TK24, ZWC24]. user-generated [AMM^{+}17]. Users [ANB09, ARAAMD24]. Using [AEHSES08, AP04a, AAAZ00, ABEAA10, ASA05, AS09, AAD06, AZK23, ABMKO14, AAA^{+}11, BW07, CH10, CKC09, CGC01, CC06, DK05, DG09, DB04, EJ05, FMI11, GG05, GA11, HH07, HL10, HHGG11, HR04, JR12, JSHB10, JS08, Jin10, KAKF11, KC10, LXH07, LS00, LGD08b, LYR09, MM00, Man15, MC07, NK24, Nom11, OS08, PW02, PK07, Pop05, RDEDAE09, Sha11, SR06, SEH05, SV03, Smi03, SPW02, TBS00, Tsu07, VHL11, WWL^{+}14, Wan11, WCC06, YCYL04, YWZ^{+}06, dMRF^{+}14, dSTE05, AB24a, AKAJ24, AESBE17, AWA20, AA18, AGM23, Al 18, AJK23, AAA17, AA24, AD18, AR19, AB24b, BM18, BFB^{+}20, BMS24, BS18, BK23, BMSP24, BJ17, BN24, BS17, BRB22, BBL23, BB23b, CM20, CZZ18, CRAV24, CSB19, CPJ17, CJG18, CGI18, DMB23, DNP24, DU19, DG18, DS19a]. using [DS19b, DZ23, DVT20, EMN17, FR18, GMS19, GS19, GMD^{+}24, GGMM23, GP17, GCMS20, GAS23, HHU24, HPC18, INR24, IQ22, IB23, JNS18, JCI18, JS19, JGM25, JN17, JS14, JSM19, JMMS18, JSYT21, KKXB19, KC19, KC20, KKR25, KJ22, KKKV19, KMR13, KH19, KR19a, KR19b, LL18, MR23, MT24, MR19b, MKN^{+}18, MBS^{+}24, MLM^{+}24, MJ21, MMS20, Nom12, OB15, OOG^{+}18, PC13, PBM16, PR17, PGK19, PN22, PV18, PK19, PN20, PR18, PN24, RN17, RGNMPM12, RR17, RPS18, RK15, RST23, RMF19, Rui19, SKM20, SK17, SP18, Sen24, SHE^{+}23, SS18, SR19, SS20c, SAAAOH10, SN23, SSS21b, SHP25, VLK19, VJTG24, VNR19, WC24, WLZZ12, XGC12, Yan17, YB18, ZLHW18, ZDW^{+}21]. Utility [CL10, VV19]. Utilization [Ani14, SEB06]. utilize [KS17]. utilizing [LBYB17].

V [KJ03]. V3 [ZFH24]. V3-transformer [ZFH24]. Vaidman’s [KK18]. Vaidman’s-type [KK18]. validating [MS19]. Value [Hsi10, SH10, GMD^{+}24, GGMM23]. Value-Driven [SH10]. valued [LWG17, ZZ20]. VANET [BMSP24, SGMT22]. Variable [BJ07, Bai21, CLLC10, RCJJ06, KBL21, PS16]. Variance [CS22, LWH12]. Variants [JV12]. variation [Lin17, LBYB17, ZZ20]. variational [IQ22, Mal24a, TN24]. varicose [BJ17]. varied [SLSS24]. Various [SR06, VJTG24]. varying [KB16]. VCC [AFbMSA10]. Vector [CKC09, FMI11, LZW07, LS00, SML^{+}13, ZXT^{+}06, TTW15, WYZ^{+}19, ZA18]. Vectorized [EA05]. Vectors [ASA05, AAA17]. Vehicle [WLC10, Ano21b, HT19, Hu17, MJ21, WZW20, YxLmLx17]. vehicle-to-grid [MJ21]. Vehicles [FT20, LGD^{+}08a, Sen24]. vehicular [BYBM21, Ogu18, PN20, SS20d, SBA^{+}22]. ventricle [Jij15]. Verifiable [HYS11, WHBS01]. Verification [ASA04, BCP09, MB10, GAS23, HDS17, HA20]. Verifier [CLX10]. version [KKS18, PKD18, PBM16]. Vertical [AEDS07]. VFS [Din11]. Via [HWH03, ARAAMD24, AAA17, CJW^{+}23, PGK19, ZZ20]. Video [AUHH^{+}18, CO00, CCG01, Cur02, GT19, GM01, Kao10, PC13, SC10, SMF12, Sha10, DWT23, GGMM23, RPS18, TLC18, Yan18]. View [KMGS11, LCHR06, ZC07, Rui19]. view-based [Rui19]. View-Dependent [ZC07]. viewpoints [DB20]. Virtual [GS13, KL03, LZ07, TZ07, TBS00, Tsu07, WZ05, ZTSL08, dMM03, AJP18, AA24, MPĆ15]. Virtual-Reality [dMM03]. visible [RDYGE18]. vision [Hu17, LL18, RvW20, RJS18, ZDW^{+}21]. vision-based [ZDW^{+}21]. Visual [JKRS02, KO00, SIT11, Bay18, HMP20, IAT20, Lou19, ZJZ15]. Visualization [PKW^{+}11, Shi13, YWZ^{+}06, dMRF^{+}14, Hou18, KBD21]. VLSI [MMM13]. VM [JMMS18, MV18]. VOD [SJMO02]. Voice [Wan11]. Voiceover [SR06]. volatile [CSB19]. Volatility [OW11]. Voltage [BK12, EHA14, RV19]. Voltages [ABEAA10]. Volume [BK12]. Voluntary [VCP00]. Voting [AZP03, FSNK22]. VR [RJS18]. vs [Pat20]. Vulnerability [CSFS17, BS25].

Wading [GYQC14]. WAP [RS04, YHL04]. WAP-Enabled [RS04, YHL04]. war [MBP23]. warehouse [KR18]. Warehousing [CHL01]. Waste [WXYN13]. water [EBA19, GMS19]. Waterfall [VHL11]. Watermark [Jin10, IG23]. Watermarking [BDB14, CCZC13, LSC03, LD13, DNP24, GSJ24, IG23, PN24]. Watershed [SL03]. waterways [CSFS17]. wave [GMS19]. Wavelength [ZJR08]. Wavelet [Abd11, EGL05, LWH12, PC13, SS12, AKAJ23, Ano21g, GS19, GP17, Guo19, KR19a, YZ19a]. Wavelet-Based [Abd11, LWH12, PC13, SS12]. Wavelets [LHJ07]. way [CM20]. Ways [HWH03]. WBE [SV03]. WBSGA [RHR06]. WCDMA [SS10]. WDM [CHC01]. WDM-Based [CHC01]. Weakly [AYV10, LZXB24]. weapon [SD23]. wearable [SCM18]. weather [GH20]. Web [DB04, BBTH18, DSP^{+}18, DS19a, Fan18, JYN19, JSM19, MS19, MdCdAdS24, PJ19, VLK19, WG17, AL06, AR05, Aa20, AEDS07, Ani14, BW07, CCG01, Che04, DP03, DG09, FH05, JYN19, JSZ09, Kar03, Kim01, KM09c, LTS25, PK07, Pon05, RHR06, SIT11, Shi12a, SWL07, TAD03, UU03, VBM02, WPJZ08]. Web-Based [AR05, DP03, FH05, Kar03, RHR06, TAD03, UU03]. Web-Graph-Based [WPJZ08]. Webchair [AR05]. WebCODS [SPW02]. website [AS24a, VLK19]. website-based [AS24a]. websites [LTS25, MWC^{+}17]. Weight [Mal24b, RABB08, Sam10]. Weighted [LL11a, CPJ17, CJG18, HHU24, KPSK23, PS19, SP18, TS15, ZZ20]. weighting [AWA20]. whale [ASM18, ABB19]. Wheel [PT14]. Wheelchair [AR05]. whispering [SS19a]. whitening [XM23]. Width [AAAZ00]. WiFi [MKN^{+}18]. Window [DDL08, SN23]. Windows [MKH02]. Wire [KM09b, RRRR17]. Wireless [AAA02, AA06, AR05, Amm07, BJ07, Bab11, CEJ12, EKAO08, JR12, KYL11, Kar12, KM09a, MB09, SS04, SBS12, SN11, SRJS08, VG14, WTLW07, AS19, Ano21j, Ano21a, BS18, Che20, EBA19, Gen19, IM19, JNS18, KS21, KV17, KPMR20, MG19, MSK24, ND18, PC19, RPZW17, RK24, SC19, SN20, SS20d, SPKS24, TK24, VNR19, fWfH19, YPY20, ZX24, ZY19b]. Without [NPY07, XLQ09, SMT19]. WLAN [YA06]. Word [BB10]. work [PR18]. Workbenches [DGR09]. workflow [KRMS19]. working [WhP20]. Workload [Che04, SS20a]. workspaces [LBYB17]. Workstations [LVK03]. world [MWC^{+}17, TG24]. Wormhole [GH06]. wrapping [PV18]. write [OB15]. write-backs [OB15]. writeprint [LLL^{+}15]. WSN [Hao17, XZCL17, YMRLJ17]. WSNs [HRMA17, PNAB18]. WWW [Smi03].

X [MV18, BM17, CDE20, KM09b, Nom12]. X-By-Wire [KM09b]. X-IoT [MV18].

X-Ray [Nom12, BM17, CDE20]. XML [DK05, KJ03, PMR^{+}03, SGD07]. XML-Based [DK05]. XMLrepository.Org [KJ03]. XOR [PC11].

YOLOv5 [SD23]. YOLOv5-based [SD23]. Yoruba [OE12].

Zero [CTW12, HWS09, Kar12]. Zero-Quantized [CTW12]. Zeta [Kar12]. Zone [CT13, RK15].

References

Al-Ali:2006:WDE

[AA06] A. R. Al-Ali. Wireless distributed energy billing system. International Journal of Computers and Applications, 28 (4):388–393, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441825.

Abraham:2012:MIP

[AA12] Blanca T. Abraham and Jose L. Aguilar. Middleware for improving performance in a component-based software architecture. International Journal of Computers and Applications, 34(1): 25–28, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2012.
1.202-3047.

Agathos:2018:RSM

[AA18] Alexander Agathos and Philip Azariadis. 3D reconstruction of skeletal mesh models and human foot biomodel generation using semantic parametric-based deformation. International Journal of Computers and Applications, 42(2):127–140, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1443316.

Al-akashi:2020:SRT

[Aa20] Falah Al-akashi. SAMA: a real-time Web search architecture. International Journal of Computers and Applications, 44(7):633–640, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1859245.

Alubaidan:2024:PMI

[AA24] Haya A. Alubaidan and Sumayh S. Aljameel. A prediction model for improving virtual machine live migration performance in cloud computing using artificial intelligence techniques. International Journal of Computers and Applications, 46 (12):1069–1087, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2409988.

Aderounmu:2002:PCD

[AAA02] G. A. Aderounmu, E. R. Adagunodo, and A. D. Akinde. Performance comparison of data-link control protocol for wireless asynchronous transfer mode network. International Journal of Computers and Applications, 24(3): 144–152, 2002. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2002.11441674.

Al-Anzi:2004:HSA

[AAA04] F. S. Al-Anzi and A. Allahverdi. A hybrid simulated annealing heuristic for multimedia object requests scheduling problem. International Journal of Computers and Applications, 26 (4):1–6, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2004.11441745.

Ayoubloo:2011:PSD

[AAA^{+}11] Mohammad K. Ayoubloo, Hazi Md. Azamathulla, Zulfequar Ahmad, Aminuddin Ab. Ghani, Javad Mahjoobi, and Amin Rasekh. Prediction of scour depth in downstream of ski-jump spillways using soft computing techniques. International Journal of Computers and Applications, 33 (1):92–97, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2011.
1.202-3078.

Alqasemi:2017:CAD

[AAA17] Fahd Alqasemi, Amira Abdelwahab, and Hatem Abdelkader. Constructing automatic domain-specific sentiment lexicon using KNN search via terms discrimination vectors. International Journal of Computers and Applications, 41 (2):129–139, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2017.1409477.

Al-Anzi:2000:IPC

[AAAZ00] F. S. Al-Anzi and K. M. Al-Zamel. Improved parallel cell placement algorithm using bounded edgewidth routing model. International Journal of Computers and Applications, 22 (1):8–12, 2000. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2000.11441593.

Aurisch:2019:OAM

[AAB19] Renae Aurisch, Mohiuddin Ahmed, and Abu Barkat. An outlook at agile methodologies for the independent games developer. International Journal of Computers and Applications, 43 (8):812–818, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/1206212X.2019.1621463.

Amin:2006:SDR

[AAD06] A. Amin and N. Al-Darwish.

Structural description to recognizing hand-printed Arabic characters using decision tree learning techniques. International Journal of Computers and Applications, 28(2): 129–134, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441796.

Ali:2006:IPF

[AAED06] H. Arafat Ali, M. F. Ared, and A. I. El-Desouky. An IP packet forwarding technique based on a new structure of lookup table. International Journal of Computers and Applications, 28(2): 112–121, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441794.

Abdalla:2025:TEN

[AAH25] Hassan I. Abdalla, Aneela Altaf, and Ali A. Hamzah. A threefold-ensemble k-nearest neighbor algorithm. International Journal of Computers and Applications, 47 (1):70–83, 2025. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2446896.

Abdullah:2018:EEA

[AAM^{+}18] Manal Abdullah, Muna AlMasawa, Ibtihal Makki, Maha Alsolmi, and Samar Mahrous. Emotions extraction from Arabic tweets. International Journal of Computers and Applications, 42 (7):661–675, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1482395.

Akhtar:2022:EHA

[AASA22] Md Mobin Akhtar, Danish Ahamad, Abdallah Saleh Ali Shatat, and Alameen Eltoum M. Abdalrahman. Enhanced heuristic algorithmbased energy-aware resource optimization for cooperative IoT. International Journal of Computers and Applications, 44(10):959–970, 2022. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2022.2028999.

Anande:2023:GAN

[AASL23] Tertsegha J. Anande, Sami Al-Saadi, and Mark S. Leeson. Generative adversarial networks for network traffic feature generation. International Journal of Computers and Applications, 45 (4):297–305, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2191072.

Al-Ayyoub:2002:DDA

[AAY02] A. E. Al-Ayyoub and A. Yazici.

Data dependence analysis on trapezoidal regions. International Journal of Computers and Applications, 24 (1):1–7, 2002. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2002.11441653.

Aranha:2009:EMT

[AB09] E. Aranha and P. Borba. Estimating manual test execution effort and capacity based on execution points. International Journal of Computers and Applications, 31(3): 167–172, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2009.11441938.

Abdu:2024:MLC

[AB24a] Nail Adeeb Ali Abdu and Khaled Omer Basulaim. Machine learning in concept drift detection using statistical measures. International Journal of Computers and Applications, 46 (5):281–291, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2289706.

Annavarapu:2024:FGS

[AB24b] Ambika Annavarapu and Surekha Borra. Figure-ground segmentation based medical image denoising using deep convolutional neural networks. International Journal of Computers and Applications, 46 (12):1179–1205, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2420870.

Ambika:2019:EBS

[ABB19] Ambika, Rajkumar L. Biradar, and Vishwanath Burkpalli. Encryption-based steganography of images by multiobjective whale optimal pixel selection. International Journal of Computers and Applications, 44 (12):1140–1149, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1692442.

Abdou:2011:FWB

[Abd11]  Mohamed  Abd-ElRahman

Abdou. A fuzzy waveletbased approach to a novel image denoising engine. International Journal of Computers and Applications, 33(4): 335–340, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/

10.2316/Journal.202.2011.
4.202-3084.

Al-Badi:2010:PVM

[ABEAA10] A. H. Al-Badi, K. Ellithy, and S. Al-Alawi. Prediction of voltages on mitigated pipelines paralleling electric transmission lines using ANN. International Journal of Computers and Applications, 32 (1):15–22, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2010.11441956.

Abada:2024:HML

[ABK24] Wafia Abada, Abdelkrim Bouramoul, and Moustafa Sadek Kahil. How machine learning algorithms can examine patterns in multiple substance addictions: predicting the effects and interactions of psychoactive substances. International Journal of Computers and Applications, 46 (11):1045–1055, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2409398.

Ayati:2014:HDU

[ABMKO14] Seyedeh B. Ayati, Kaddour Bouazza-Marouf, David Kerr, and Michael O’Toole. Haematoma detection using EIT in a sheep model. International Journal of Computers and Applications, 36 (3):87–92, 2014. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2014.
                                        3.202-3834.

Al-Bahadili:2010:BLT

[ABR10] H. Al-Bahadili and A. Rababa’a. A bit-level text compression scheme based on the HCDC algorithm. International Journal of Computers and Applications, 32(3): 355–361, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2010.
                                        3.202-2914.

Al-Badarneh:2009:LRT

[ABT09] A. Al-Badarneh and M. Tawil. Linear R-tree revisited. International Journal of Computers and Applications, 31 (2):74–83, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2009.11441927.

Amoroso:2006:FBF

[AC06] A. Amoroso and G. Casciola. Fast bivariate function rendering for small devices. International Journal of Computers and Applications, 28(4): 321–328, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2006.11441817.

Ahmed:2007:ACA

[ACA07] D. T. Ahmed, N. M. M. K.

Chowdhury, and M. M. Akbar. Admission control algorithm for multimedia server: A hybrid approach. International Journal of Computers and Applications, 29(4): 414–419, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441873.

Adami:2010:MNS

[ACGP10] D. Adami, C. Callegari, S. Giordano, and M. Pagano. Mtens: a new simulator for the design of MPLS networks with traffic engineering support. International Journal of Computers and Applications, 32 (1):110–118, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2010.11441967.

Amghar:2020:SPA

[ACM20] Souad Amghar, Safae Cherdal, and Salma Mouline. Storing, preprocessing and analyzing tweets: finding the suitable noSQL system. International Journal of Computers and Applications, 44 (6):586–595, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1846946.

Arcelli:2000:DAP

[AD00] F. Arcelli and M. De Santo. Distributed agents for problem solving. International Journal of Computers and Applications, 22(1):38–45, 2000. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

                                        /www.tandfonline.com/doi/
                                        full/10.1080/1206212X.2000.
                                        11441597.

Anbarasi:2018:PPF

[AD18] M. Anbarasi and M. A. Saleem Durai. Prediction of protein folding kinetics states using hybrid brainstorm optimization. International Journal of Computers and Applications, 42(7):635–643, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1479348.

Al-Debagy:2021:DBM

[ADM21] Omar Al-Debagy and Péter Martinek. Dependenciesbased microservices decomposition method. International Journal of Computers and Applications, 44 (9):814–821, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2021.1915444.

Awad:2019:ASM

[AEA19] Wael A. Awad and Noha E. El-Attar. Adaptive SLA mechanism based on fuzzy system for dynamic cloud environment. International Journal of Computers and Applications, 44(1):12–22, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1683956.

Ali:2003:EIA

[AEDA03] H. A. Ali, A. I. El-Dousky, and A. A. A. Ali. Exploitation of intelligent agent for building expandable and flexible computerized tutoring system. International Journal of Computers and Applications, 25 (2):119–129, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2003.11441693.

Ali:2007:NSV

[AEDS07] H. A. Ali, A. I. El-Desouky, and A. I. Saleh. A novel strategy for a vertical Web page classifier based on continuous learning Naïve Bayes algorithm. International Journal of Computers and Applications, 29(3): 259–277, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441856.

Azumah:2023:CTC

[AEEO23] Sylvia W. Azumah, Nelly Elsayed, Zag ElSayed, and Murat Ozer. Cyberbullying in text content detection: an analytical review. International Journal of Computers and Applications, 45 (9):579–586, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2256048.

Ahmed:2017:IRD

[AEH17] Kareem Ahmed and Ibrahim El-Henawy. Increasing robustness of Data Encryption Standard by integrating DNA cryptography. International Journal of Computers and Applications, 39(2): 91–105, 2017. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2017.1289690.

Abd-El-Hafiz:2008:ROO

[AEHSES08] S. K. Abd-El-Hafiz, D. M. Shawky, and A. L. El-Sedeek. Recovery of object-oriented design patterns using static and dynamic. International Journal of Computers and Applications, 30(3): 220–233, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2008.11441902.

Adel:2017:DEH

[AESBE17] Ebtsam Adel, Shaker El-Sappagh, Sherif Barakat, and Mohammed Elmogy. Distributed electronic health record based on semantic interoperability using fuzzy ontology: a survey. International Journal of Computers and Applications, 40 (4):223–241, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1418237.

Ahmed:2010:OSN

[AFbMSA10] Hassan Y. Ahmed, Ibrahima Faye, Naufal bin M. Saad, and Syed A. Aljunid. OCDM a system: New detection scheme and encoder–decoder structure based on fibre Bragg gratings (FBGs) for VCC code. International Journal of Computers and Applications, 32 (4):461–468, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2010.
4.202-2881.

Alencar:2025:ATA

[AFLdS25] Reno C. Alencar, Bruno J. T. Fernandes, Paulo H. E. S. Lima, and Carlo M. R. da Silva. AI techniques for automated penetration testing in MQTT networks: a literature investigation. International Journal of Computers and Applications, 47 (1):106–121, 2025. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2443504.

Ahmed:2023:IBC

[AGM23] Nadia Ahmed, Sadik Kamel Gharghan, and Ammar Hussein Mutlag. IoT-based child tracking using RFID and GPS. International Journal of Computers and Applications, 45(5):367–378, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2215077.

Aguilar:2002:FTM

[AH02] J. Aguilar and M. Hernandez. A fault-tolerant mechanism for distributed/parallel system based on task replication techniques. International Journal of Computers and Applications, 24(3): 129–135, 2002. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2002.11441672.

Aldouri:2018:HRS

[AH18] Thekra Aldouri and Mhand Hifi. A hybrid reactive search for solving the max-min knapsack problem with multiscenarios. International Journal of Computers and Applications, 40(1):1–13, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.

tandfonline.com/doi/full/
10.1080/1206212X.2016.1207426.

Ade-Ibijola:2018:SRE

[AI18] Abejide Ade-Ibijola. Synthesis of regular expression problems and solutions. International Journal of Computers and Applications, 42 (8):748–764, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1482398.

Al-Jarah:2000:SBC

[AJAk00] O. M. Al-Jarah and T. A.

Al-khdour. Slackness-based co-scheduling hardware and software pipelines with multiple latency sequences. International Journal of Computers and Applications, 22 (3):121–128, 2000. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2000.11441610.

Ahuja:2018:FPM

[AJJ18] Anjali Ahuja, Anamika Jain, and Madhu Jain. Finite population multi-server retrial queueing system with an optional service and balking. International Journal of Computers and Applications, 41(1):54–61, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/1206212X.2018.1505023.

Aljabri:2023:CNN

[AJK23] Ahmed Aljabri, Farah Jemili, and Ouajdi Korbaa. Convolutional neural network for intrusion detection using blockchain technology. International Journal of Computers and Applications, 46 (2):67–77, 2023. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2284443.

Al-Jarrah:2018:CVA

[AJP18] Ahmad Al-Jarrah and Enrico Pontelli. The collaborative virtual affinity group model: principles, design, implementation, and evaluation. International Journal of Computers and Applications, 42(5):485–513, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1486557.

Al-Jarrah:2008:PPA

[AJS08] O. M. Al-Jarrah and B. H. Sababha. Parma: a power aware reliable multicasting algorithm for mobile ad hoc networks. International Journal of Computers and Applications, 30(3): 244–250, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.

tandfonline.com/doi/full/
10.1080/1206212X.2008.11441904.

Asem:2006:CEK

[AK06] Y. M. Asem and A. Kara. A computationally efficient keyhiding based group re-keying scheme for secure multicasting. International Journal of Computers and Applications, 28(1):65–73, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441790.

Abdul-Kareem:2023:IEA

[AKAJ23] Ali Akram Abdul-Kareem and Waleed Ameen Mahmoud Al-Jawher. Image encryption algorithm based on Arnold transform and chaos theory in the multi-wavelet domain. International Journal of Computers and Applications, 45(4):306–322, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2196902.

Abdul-Kareem:2024:IEA

[AKAJ24] Ali Akram Abdul-Kareem and Waleed Ameen Mahmoud Al-Jawher. An image encryption algorithm using hybrid sea lion optimization and chaos theory in the Hartley domain. International Journal of Computers and Applications, 46(5):324–337, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2313300.

Almutawakel:2019:SFA

[AKB^{+}19] Abdallah Almutawakel, Okba Kazar, Mouadh Bali, Houcine Belouaar, and Abdelbasset Barkat. Smart and fuzzy approach based on CSP for cloud resources allocation. International Journal of Computers and Applications, 44 (2):117–129, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1701241.

Alzeidi:2010:PMT

[AKK10] N. Alzeidi, M. O. Khaoua, and A. Khonsari. Performance modelling of torus interconnection networks with deep buffers. International Journal of Computers and Applications, 32(1):1–8, 2010. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2010.11441954.

AlJanabi:2018:SSC

[Al 18] Samaher Al Janabi. Smart system to create an optimal higher education environment using IDA and IOTs. International Journal of Computers and Applications, 42 (3):244–259, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1512460.

Aguilar:2006:CRP

[AL06] J. Aguilar and E. L. Leiss. A coherence-replacement protocol for Web proxy cache systems. International Journal of Computers and Applications, 28(1):12–18, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441783.

Abdelkader:2019:ELS

[AL19] Birane Abdelkader and Hamami Latifa. An efficient level set method based on global statistical information for image segmentation. International Journal of Computers and Applications, 44 (1):48–56, 2019. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1690797.

Ali:2024:PGR

[Ali24] Shahzad Ali. Performance of geographically restricted flooding for emergency alert applications. International Journal of Computers and Applications, 46(6):390–396, 2024. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (electronic). URL

http://www.tandfonline.
com/doi/full/10.1080/1206212X.
2024.2328488.

Almeida:2011:TDQ

[Alm11]  Carlos Almeida.  Timely

and dependable QoS adaptation in quasi-synchronous systems. International Journal of Computers and Applications, 33(3):179–188, 2011. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2011.
                                        3.202-2172.

Alenezi:2014:EEN

[AM14] Mamdouh Alenezi and Kenneth Magel. Empirical evaluation of a new coupling metric: Combining structural and semantic coupling. International Journal of Computers and Applications, 36(1): 34–44, 2014. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2014.
                                        1.202-3902.

Agarwal:2016:EFP

[AM16] Parul Agarwal and Shikha Mehta. Enhanced flower pollination algorithm on data clustering. International Journal of Computers and Applications, 38(2–3):144–155, 2016. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        full/10.1080/1206212X.2016.
                                        1224401.

Awasthi:2012:ECC

[AMJ12] Lalit K. Awasthi, Manoj Misra, and Ramesh C. Joshi. An efficient coordinated checkpointing approach for distributed computing systems with reliable channels. International Journal of Computers and Applications, 34 (1):1–10, 2012. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2012.
1.202-2118.

Ammari:2007:DSP

[Amm07] H. M. Ammari. Discovery and selection protocols for multihop wireless Internet access. International Journal of Computers and Applications, 29 (1):25–32, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441829.

Arain:2017:ITI

[AMM^{+}17] Qasim Ali Arain, Hina Memon, Imran Memon, Muhammad Hammad Memon, Riaz Ahmed Shaikh, and Farman Ali Mangi. Intelligent travel information platform based on location base services to predict user travel behavior from usergenerated GPS traces. International Journal of Computers and Applications, 39(3): 155–168, 2017. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2017.1309222.

Al-Naymat:2018:PTS

[ANAKS18] Ghazi Al-Naymat, Mouhammd Al-Kasassbeh, and Zyad Sober. Pairs trading strategy: a recommendation system. International Journal of Computers and Applications, 42(8):787–797, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1493074.

Al-Naamany:2009:DSF

[ANB09] A. Al-Naamany and H. Bourdoucen. Design and simulation of a fuzzy logic bandwidth controller for users classification and priorities allocations. International Journal of Computers and Applications, 31 (1):23–29, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2009.11441920.

Anirban:2014:EBU

[Ani14] Kundu Anirban. Efficient bandwidth utilization for downloading Web pages. International Journal of Computers and Applications, 36 (1):1–6, 2014. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/

10.2316/Journal.202.2014.
1.202-3443.

Anonymous:2000:CPI

[Ano00a] Anonymous. Call for papers — IMSA 2000. International Journal of Computers and Applications, 22(1): 48–49, 2000. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2000.11441599.

Anonymous:2000:CPS

[Ano00b] Anonymous. Call for papers special issue on computational intelligence techniques in co-operative robots. International Journal of Computers and Applications, 22 (3):185, 2000. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2000.11441613.

Anonymous:2002:CPS

[Ano02a] Anonymous. Call for papers for special issue on cluster/grid computing. International Journal of Computers and Applications, 24(3): 161–162, 2002. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2002.11441677.

Anonymous:2002:E

[Ano02b] Anonymous. Erratum. International Journal of Computers and Applications, 24 (1):39, 2002. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2002.11441658.

Anonymous:2005:EN

[Ano05] Anonymous. Editors’ note. International Journal of Computers and Applications, 27 (1):??, 2005. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441755.

Anonymous:2017:C

[Ano17a] Anonymous. Corrigendum. International Journal of Computers and Applications, 39 (1):57, 2017. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2016.1251122.

Anonymous:2017:EB

[Ano17b] Anonymous. Editorial board. International Journal of Computers and Applications, 39 (4):(ei), 2017. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1389680.

Anonymous:2018:P

[Ano18] Anonymous. Preface. International Journal of Computers and Applications, 40(3): 1, 2018. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (electronic). URL

http://www.tandfonline.
com/doi/full/10.1080/1206212X.
2018.1460446.

Anonymous:2019:RAR

[Ano19] Anonymous. RETRACTED ARTICLE: Retraction. International Journal of Computers and Applications, 42 (5):530, 2019. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1572585.

Anonymous:2021:RASc

[Ano21a] Anonymous. RETRACTED ARTICLE: Statement of retraction: a wireless network remote monitoring method driven by artificial intelligence. International Journal of Computers and Applications, 44(6):598, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2021.1994762.

Anonymous:2021:RASf

[Ano21b] Anonymous. RETRACTED ARTICLE: Statement of retraction: based on deep learning in traffic remote sensing image processing to recognize target vehicle. International Journal of Computers and Applications, 44 (6):601, 2021. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2021.1994767.

Anonymous:2021:RASd

[Ano21c] Anonymous. RETRACTED ARTICLE: Statement of retraction: collaborative correlation space big data clustering algorithm for abnormal flow monitoring. International Journal of Computers and Applications, 44 (6):599, 2021. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2021.1994763.

Anonymous:2021:RASg

[Ano21d] Anonymous. RETRACTED ARTICLE: Statement of retraction: development of computer-based agricultural remote intelligent information monitoring system. International Journal of Computers and Applications, 44 (6):602, 2021. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2021.1994770.

Anonymous:2021:RASj

[Ano21e] Anonymous. RETRACTED ARTICLE: Statement of retraction: Key technologies of cloud computing-based IoT data mining. International Journal of Computers and Applications, 44 (6):605, 2021. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2021.1994774.

Anonymous:2021:RASi

[Ano21f] Anonymous. RETRACTED ARTICLE: Statement of retraction: key technologies of massive concurrent data processing in smart city based on cloud computing. International Journal of Computers and Applications, 44 (6):604, 2021. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2021.1994773.

Anonymous:2021:RASb

[Ano21g] Anonymous. RETRACTED ARTICLE: Statement of retraction: research on digital image wavelet transform filtering optimization processing method based on DSP Internet of Things. International Journal of Computers and Applications, 44 (6):597, 2021. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2021.1994761.

Anonymous:2021:RASh

[Ano21h] Anonymous. RETRACTED ARTICLE: Statement of retraction: the role of computer security management in preventing financial technology risks. International Journal of Computers and Applications, 44(6):603, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2021.1994771.

Anonymous:2021:RASe

[Ano21i] Anonymous. RETRACTED ARTICLE: Statement of retraction: track recognition algorithm based on neural network for rail transit. International Journal of Computers and Applications, 44(6):600, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2021.1994765.

Anonymous:2021:RASa

[Ano21j] Anonymous. RETRACTED ARTICLE: Statement of retraction: traffic prediction and transmission scheduling of artificial intelligencedriven cognitive wireless networks. International Journal of Computers and Applications, 44(6):596, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2021.1994760.

Anonymous:2022:SRI

[Ano22a] Anonymous. Statement of retraction: An intelligent model to reduce the energy consumption of sensor network nodes. International Journal of Computers and Applications, 44(10):1003, 2022. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2022.2034230.

Anonymous:2022:SRR

[Ano22b] Anonymous. Statement of retraction: Real-time image recognition algorithm and system design of Android mobile devices. International Journal of Computers and Applications, 44(10):1002, 2022. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2022.2034225.

Abada:2009:ERC

[AO09] G. O. Abada and E. A.

Onibere. The effect of rehearsed computer use on Icon recognition. International Journal of Computers and Applications, 31(1): 9–15, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2009.11441918.

Althammer:2001:DIM

[AP01] E. Althammer and W. Pree.

Design and implementation of an MVC-based architecture for e-commerce applications.

International Journal of Computers and Applications, 23 (3):173–185, 2001. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2001.11441649.

Adam:2004:GIP

[AP04a] G. K. Adam and V. G. Pappas. Geographical image processing using GisTool. International Journal of Computers and Applications, 26 (2):1–6, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2004.11441728.

Aljifri:2004:ROS

[AP04b] H. Aljifri and A. P. Pons. Reducing overhead of secure ecommerce transactions. International Journal of Computers and Applications, 26 (4):1–6, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2004.11441746.

Abhishek:2021:IAM

[APS^{+}21] Kumar Abhishek, Vaibhav Pratihar, Shishir Kumar Shandilya, Sanju Tiwari, Vinay Kumar Ranjan, and Sudhakar Tripathi. An intelligent approach for mining knowledge graphs of online news. International Journal of Computers and Applications, 44(9):838–846, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2021.1957551.

Arunachalam:2014:DIR

[APSK14] Kamaraj Arunachalam, Marichamy Perumalsamy, C. Kalyana Sundaram, and J. Senthil Kumar. Design and implementation of a reversible logic based 8-bit arithmetic and logic unit. International Journal of Computers and Applications, 36(2):49–55, 2014. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2014.
2.202-3832.

Anvari:2017:GSQ

[APZ17] Armin Anvari, Lei Pan, and Xi Zheng. Generating security questions for better protection of user privacy. International Journal of Computers and Applications, 42 (4):329–350, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1395132.

Al-Rousan:2005:WWB

[AR05] M. Al-Rousan. Webchair: a Web-based wireless navigation wheelchair system for people with motor disabilities. International Journal of Computers and Applications, 27(4):

274–284, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441781.

Angadi:2019:MSA

[AR19] Sujay Angadi and Venkata Siva Reddy. Multimodal sentiment analysis using reliefF feature selection and random forest classifier. International Journal of Computers and Applications, 43 (9):931–939, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1658054.

Al-Rjoob:2024:PLS

[ARAAMD24] Ayat M. Al-Rjoob, Ahmad A. Ababnah, Mamoun F. Al-Mistarihi, and Khalid A. Darabkh. Physical-layer security for primary users in 5G underlay cognitive radio system via artificial-noise-aided by secondary users. International Journal of Computers and Applications, 46 (9):702–714, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2380666.

Al-Shoshan:2009:ASD

[AS09] A. I. Al-Shoshan. Audio signal discrimination using evolutionary spectrum. International Journal of Computers and Applications, 31 (2):69–73, 2009. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2009.11441926.

Anand:2019:CIU

[AS19] J. V. Anand and P. Sivanesan.

Certain investigations of underwater wireless sensors synchronization and funneling effect. International Journal of Computers and Applications, 43(5):423–430, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1573947.

Ahsan:2020:SCT

[AS20] Mohammad Ahsan and T. P.

Sharma. Spams classification and their diffusibility prediction on Twitter through sentiment and topic models. International Journal of Computers and Applications, 44(4):365–375, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1758430.

Alhasyim:2024:NHC

[AS24a] Abdul Ghofar Alhasyim and Nabila Husna Shabrina. A novel hybrid CNN-GRU model for website-based user interface code generator. International Journal of Computers and Applications, 46 (12):1129–1142, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2416955.

Anwar:2024:TCI

[AS24b] Shamama Anwar and Anshuman Srivastava. A threefold chaotic image encryption technique for efficacious image data protection. International Journal of Computers and Applications, 46 (11):974–982, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2401606.

Azizi:2004:CSV

[ASA04] M. Azizi, X. Song, and E.- M. Aboulhamid. A case study on the verification of cache coherence protocols. International Journal of Computers and Applications, 26 (1):1–10, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2004.11441722.

Al-Sadi:2005:NFT

[ASA05] J. Al-Sadi and A. M. Awwad. A new fault-tolerant routing algorithm for OTIS-cube using unsafety vectors. International Journal of Computers and Applications, 27(4):

244–251, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441777.

Abdellatief:2020:ACC

[ASGW20] Majdi Abdellatief, Abu Bakar Md Sultan, Abdul Azim Abdul Ghani, and Abubaker Wahaballa. Assessing the changeability of componentbased system design: a controlled experiment. International Journal of Computers and Applications, 44 (6):513–520, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1822597.

Adhirai:2018:CIC

[ASM18] S. Adhirai, Paramjit Singh, and Rajendra Prasad Mahapatra. Circular interpolation and chronological-whale optimization based privacy preservation in cloud. International Journal of Computers and Applications, 43 (3):292–304, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1560668.

AboDoma:2021:ATB

[ASM21] Noran AboDoma, Eman Shaaban, and Ahmad Mostafa. Adaptive time-bound access control for internet of things in fog computing architecture. International Journal of Computers and Applications, 44(8):779–790, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2021.1935653.

Arif:2018:VRD

[AUHH^{+}18] Sheeraz Arif, Tehseen Ul-Hassan, Fida Hussain, Jing Wang, and Zesong Fei. Video representation by dense trajectories motion map applied to human activity recognition. International Journal of Computers and Applications, 42(5):474–484, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1486001.

Adinugroho:2020:NTE

[AWA20] Sigit Adinugroho, Randy C. Wihandika, and Putra P. Adikara. Newsgroup topic extraction using term-cluster weighting and Pillar Kmeans clustering. International Journal of Computers and Applications, 44 (4):357–364, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2020.1757246.

Ametepe:2018:DPC

[AWO^{+}18] Wolali Ametepe, Changda Wang, Selasi Kwame Ocansey, Xiaowei Li, and Fida Hussain. Data provenance collection and security in a distributed environment: a survey. International Journal of Computers and Applications, 43(1):11–25, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1501937.

Al-Yaseen:2023:MML

[AYI23] Wathiq Laftah Al-Yaseen and Ali Kadhum Idrees. MuDeLA: multi-level deep learning approach for intrusion detection systems. International Journal of Computers and Applications, 45 (12):755–763, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2275084.

Agrawal:2010:SFT

[AYV10] S. Agrawal, R. S. Yadav, and R. Vijay. A speed fine tuning technique for system energy minimization of weakly hard real-time system. International Journal of Computers and Applications, 32(2): 197–205, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2010.11441975.

Arvaneh:2023:UFC

[AZK23] Fatemeh Arvaneh, Faraneh Zarafshan, and Abbas Karimi. Using fog computing (FC) and optimization techniques for tasks migration and resource allocation in the internet of things (IoT). International Journal of Computers and Applications, 46 (2):113–121, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2287257.

Azadmanesh:2003:EVA

[AZP03] A. Azadmanesh, L. Zhou, and D. Peak. Egophobic voting algorithms. International Journal of Computers and Applications, 25(4): 236–246, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2003.11441711.

Bruda:2003:RTC

[BA03] S. D. Bruda and S. G. Akl. Real-time computation: a formal definition and its applications. International Journal of Computers and Applications, 25(4): 247–257, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2003.11441712.

Bruda:2007:SML

[BA07] S. D. Bruda and S. G. Akl.

Size matters: Logarithmic space is real time. International Journal of Computers and Applications, 29(4): 327–336, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441863.

Banati:2016:DCC

[BA16] Hema Banati and Nidhi Arora.

Detecting communities in complex networks — a discrete hybrid evolutionary approach. International Journal of Computers and Applications, 38 (1):29–40, 2016. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2016.1210280.

Benkhedda:2023:OUP

[BA23] Youcef Benkhedda and Faical Azouaou. Optimizing user profile matching: a textbased approach. International Journal of Computers and Applications, 45 (5):403–412, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2218244.

Babu:2011:SOP

[Bab11] Anchare V. Babu. On selection of optimal physical layer parameters for wireless sensor networks over Rayleigh fading channels. International Journal of Computers and Applications, 33(1):15–21, 2011. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

/www.tandfonline.com/doi/
full/10.2316/Journal.202.
2011.1.202-2823.

Badri:2019:NMS

[Bad19]  SatyaJaswanth Badri.  A

novel map-scan-reduce based density peaks clustering and privacy protection approach for large datasets. International Journal of Computers and Applications, 43 (7):663–673, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1624314.

Bai:2010:RKI

[Bai10] L. Bai. A reliable (K, N) image secret sharing scheme with low information overhead. International Journal of Computers and Applications, 32 (1):9–14, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2010.11441955.

Bai:2021:VIA

[Bai21]  Zhongxian Bai.  Variable

incremental adaptive learning model based on knowledge graph and its application in online learning system. International Journal of Computers and Applications, 44(7):650–658, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2021.1878419.

Bardsiri:2018:NCF

[Bar18] Amid Khatibi Bardsiri. A new combinatorial framework for software services development effort estimation. International Journal of Computers and Applications, 40 (1):14–24, 2018. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1395103.

Bayi:2018:RTV

[Bay18] Xiao Bayi. Real-time visual transmission mechanism of graphics diversity based on mobile 3D graphics matching algorithm. International Journal of Computers and Applications, 43 (4):340–345, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1550165.

Bahig:2008:MP

[BB08] Hazem M. Bahig and Hatem M.

Bahig. Merging on PRAM. International Journal of Computers and Applications, 30 (1):51–55, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.

                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2008.11441877.

Bhattacharya:2010:DAS

[BB10] S. Bhattacharya and A. Basu. Design of a word to sentence generator for augmentative. International Journal of Computers and Applications, 32 (1):73–83, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2010.11441962.

Bhakar:2018:OLT

[BB18] Suman Bhakar and Devershi Pallavi Bhatt. Optimizing latency time of the AR system through glyph detection. International Journal of Computers and Applications, 41(1):14–30, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1495863.

Berisha:2023:ACS

[BB23a] Fjolla Berisha and Eliot Bytyçi. Addressing cold start in recommender systems with neural networks: a literature survey. International Journal of Computers and Applications, 45 (7–8):485–496, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2237766.

Bouhadja:2023:BCH

[BB23b] Amina Bouhadja and Abdelkrim Bouramoul. Beyond cocaine and heroin use: a stacking ensemblebased framework for predicting the likelihood of subsequent substance use disorder using demographics and personality traits. International Journal of Computers and Applications, 45 (11):722–733, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2273011.

Benlamri:2008:ACL

[BBAR08] R. Benlamri, H. Barada, and A. Al-Raqabani. Analysis of coordinated load sharing for large distributed systems. International Journal of Computers and Applications, 30 (2):151–162, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441894.

Bouazizi:2023:PKB

[BBL23] Samar Bouazizi, Emna Benmohamed, and Hela Ltifi. PSO-K2PC: Bayesian structure learning using optimized K2 algorithm for parentschildren detection. International Journal of Computers and Applications, 45 (9):553–563, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2250143.

Bentlemsan:2018:HRC

[BBTH18] Khadidja Bentlemsan, Djamal Bennouar, Dalila Tamzalit, and Khaled Walid Hidouci. A hybrid re-composition based on components and web services. International Journal of Computers and Applications, 42(5):449–462, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1482820.

Benerecetti:2009:ESV

[BCP09] M. Benerecetti, N. Cuomo, and A. Peron. An environment for the specification and verification of time-dependent security protocols. International Journal of Computers and Applications, 31(3): 183–192, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2009.11441940.

Bhandari:2011:AES

[BD11] Smriti H. Bhandari and Sunil M. Deshpande. Analysis of engineered surfaces for product quality monitoring. International Journal of Computers and Applications, 33 (4):284–292, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.

tandfonline.com/doi/full/
10.2316/Journal.202.2011.
4.202-2670.

Barua:2024:PPA

[BD24] Eirene Barua and Mala Dutta.

Privacy-preserving association rule mining: a survey of techniques for sensitive rule identification and enhanced data protection. International Journal of Computers and Applications, 46 (4):252–265, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2307086.

Belkacem:2014:DCM

[BDB14] Samia Belkacem, Zohir Dibi, and Ahmed Bouridane. DCT coefficients modelling for image watermarking. International Journal of Computers and Applications, 36(4): 155–163, 2014. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2014.
4.202-4017.

Beghdad:2006:FBQ

[Beg06]  R. Beghdad.  A flows-

based QoS routing protocol for multicast communication networks. International Journal of Computers and Applications, 28(1): 59–64, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.

                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2006.11441789.

Bagui:2020:ISC

[BFB^{+}20] Sikha Bagui, Xingang Fang, Subhash Bagui, Jeremy Wyatt, Patrick Houghton, Joe Nguyen, John Schneider, and Tyler Guthrie. An improved step counting algorithm using classification and double autocorrelation. International Journal of Computers and Applications, 44 (3):250–259, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2020.1726006.

Belete:2021:GSH

[BH21] Daniel Mesafint Belete and Manjaiah D. Huchaiah. Grid search in hyperparameter optimization of machine learning models for prediction of HIV/AIDS test results. International Journal of Computers and Applications, 44 (9):875–886, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2021.1974663.

Bhasker:2000:OQD

[Bha00]  B. Bhasker.  Optimizing

queries in distributed database systems on a broadcast local area network. International Journal of Computers and Applications, 22(3):

166–173, 2000. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2000.11441620.

Beauchemin:2009:PAD

[BHB09] S. S. Beauchemin, H. O.

Hamshari, and M. A. Bauer. Passive atmospheric diffusion with Gaussian fragmentation. International Journal of Computers and Applications, 31 (2):97–108, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2009.11441930.

Benhocine:2022:TAS

[BHZGG22] Karim Benhocine, Adel Hansali, Leila Zemmouchi-Ghomari, and Abdessamed Reda Ghomari. Towards an automatic SPARQL query generation from ontology competency questions. International Journal of Computers and Applications, 44 (10):971–980, 2022. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2022.2031722.

Babu:2007:SDS

[BJ07] A. V. Babu and L. Jacob. Service differentiation schemes in IEEE 802.11 wireless LANs with variable frame size. International Journal of Computers and Applications, 29(2): 187–195, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2007.11441847.

Bhavani:2017:IRV

[BJ17] R. R. Bhavani and G. Wiselin Jiji. Image registration for varicose ulcer classification using KNN classifier. International Journal of Computers and Applications, 40(2):88–97, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2017.1395108.

Breja:2021:SNF

[BJ21] Manvi Breja and Sanjay Kumar Jain. A survey on nonfactoid question answering systems. International Journal of Computers and Applications, 44(9):830–837, 2021. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        full/10.1080/1206212X.2021.
                                        1949117.

Bali:2024:IDS

[BJBA24] Nadia Bali, Zied Jaoua, Olfa Bzeouich, and Imed Abbassi. Intrusion detection systems for the internet of things: a probabilistic anomaly detection approach. International Journal of Computers and Applications, 46 (11):933–944, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2400148.

Bouras:2001:BIS

[BK01] C. Bouras and V. Kapoulas.

Basic Internet services environment for the education society. International Journal of Computers and Applications, 23(1):35–44, 2001. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2001.11441632.

Basappa:2012:SSV

[BK12] Subba Reddy Basappa and Udaya Kumar. Simulation of surface and volume stress for high voltage transmission insulators. International Journal of Computers and Applications, 34(1):29–35, 2012. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

/www.tandfonline.com/doi/
full/10.2316/Journal.202.
2012.1.202-3077.

Bhagat:2023:SMD

[BK23] Shinzeer C. K. Avinash Bhagat and Ajay Shriram Kushwaha. Secure medical data storage in DPOS-hyper ledger fabric block chain using PM-ECC and L2-DWT. International Journal of Computers and Applications, 45 (7–8):516–522, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/1206212X.2023.2243676.

Babu:2024:AFC

[BK24] Nirmal Varghese Babu and E. Grace Mary Kanaga. Advanced functional connectivity analysis with integrated EEG signal enhancement and GTW-EEG-GAN for emotion detection. International Journal of Computers and Applications, 46 (12):1152–1165, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2417847.

Bhatia:2017:CGT

[BKK17] Vishank Bhatia, V. Kalaichelvi, and R. Karthikeyan. Comparison of GA tuned fuzzy logic and NARMA-L2 controllers for motion control in 5-DOF robot. International Journal of Computers and Applications, 39(2): 69–78, 2017. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1281038.

Badhon:2019:SAR

[BKXK19] Bodrunnessa Badhon, Mir Md. Jahangir Kabir, Shuxiang Xu, and Monika Kabir. A survey on association rule mining based on evolutionary algorithms. International Journal of Computers and Applications, 43 (8):775–785, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1612993.

Babin:2001:AFM

[BL01] G. Babin and F. Lustman.

Application of formal methods to scenario-based requirements engineering. International Journal of Computers and Applications, 23 (3):141–151, 2001. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2001.11441645.

Bilami:2006:CSP

[BLDD06] A. Bilami, M. Lalam, M. Daoui, and B. Djamah. A cell selection policy for an inputbuffered packet switch. International Journal of Computers and Applications, 28(3): 234–242, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441808.

Buhari:2017:CFP

[BM17] M. P. A. Buhari and S. K.

Mohideen. A combination of fuzzy positioned dental X-ray analysis model to presume the peculiar images. International Journal of Computers and Applications, 42(1):17–22, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2017.1396422.

Babrahem:2018:PCP

[BM18] Afnan Salem Babrahem and Muhammad Mostafa Monowar. Preserving confidentiality and privacy of the patient’s EHR using the OrBAC and AES in cloud environment*. International Journal of Computers and Applications, 43 (1):50–61, 2018. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1505025.

Brothers:2008:MEI

[BMM^{+}08] T. Brothers, N. Mandagere, S. Muknahallipatna, J. C. Hamann, and H. Johnson. Microsoft exchange implementation on a distributed storage area network. International Journal of Computers and Applications, 30(3): 251–264, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2008.11441905.

Bairy:2024:EHD

[BMS24] Maithri Bairy, Balachandra Muniyal, and Nisha P. Shetty. Enhancing healthcare data integrity: fraud detection using unsupervised learning techniques. International Journal of Computers and Applications, 46 (11):1006–1019, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2408262.

Bhanja:2024:DSD

[BMSP24] Urmila Bhanja, Ankit Majhi, Smrutirekha Sahu, and Debapriya Parida. Detection of Sybil & DDoS attacks in VANET using intelligent technique. International Journal of Computers and Applications, 46 (10):811–829, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2389476.

Bhogade:2024:TSF

[BN24] Vaibhav Bhogade and B. Nithya.

Time series forecasting using transformer neural network. International Journal of Computers and Applications, 46(10):880–888, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2396321.

Bassil:2012:OFU

[BO12] Layale Bassil and Iyad Ouaiss.

Optimizing functional unit binding during high-level synthesis. International Journal of Computers and Applications, 34(1):58–65, 2012. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

                                        /www.tandfonline.com/doi/
                                        full/10.2316/Journal.202.
                                        2012.1.202-3223.

Belhocine:2018:NPS

[BO18] Ali Belhocine and Wan Zaidi Wan Omar. A numerical parametric study of mechanical behavior of dry contacts slipping on the disc-pads interface. International Journal of Computers and Applications, 40(1):42–60, 2018. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2017.1395105.

Budati:2017:PAH

[BP17] Anil Kumar Budati and Trinatha Rao Polipalli. Performance analysis of HFDI computing algorithm in intelligent networks. International Journal of Computers and Applications, 41 (4):255–261, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2017.1396424.

Beg:2017:FSM

[BR17] Ismat Beg and Tabasam Rashid. A fuzzy similarity measure based on equivalence relation with application in cluster analysis. International Journal of Computers and Applications, 39(3):

148–154, 2017. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1309220.

Biswas:2022:MCI

[BRB22] Ranjit Biswas, Sudipta Roy, and Abhijit Biswas. MRI and CT image indexing and retrieval using steerable pyramid transform and local neighborhood difference pattern. International Journal of Computers and Applications, 44 (11):1005–1014, 2022. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2022.2092937.

Behera:2017:MAR

[BRR17] Ranjit Kumar Behera, K. Hemant Kumar Reddy, and Diptendu Sinha Roy. Modeling and assessing reliability of service-oriented internet of things. International Journal of Computers and Applications, 41 (3):195–206, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1420589.

Biswas:2017:MIF

[BS17] Biswajit Biswas and Biplab Kanti Sen. Medical image fusion using type-2 fuzzy and near-fuzzy set approach.

International Journal of Computers and Applications, 42 (4):399–414, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1402588.

Bhagat:2018:TTW

[BS18] Dhiren P. Bhagat and Himanshukumar Soni. Target tracking in a wireless sensor network using a multistep KF-PSO model. International Journal of Computers and Applications, 43 (5):401–412, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1564172.

Banerjee:2019:DSR

[BS19] Shreya Banerjee and Anirban Sarkar. Domain-specific requirements analysis framework: ontology-driven approach. International Journal of Computers and Applications, 44(1):23–47, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1687177.

Bharti:2020:PMC

[BS20] Sarveshwar Bharti and Hardeep Singh. Proactively managing clones inside an IDE: a systematic literature review. International Journal of Computers and Applications, 44(3):230–249, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1720952.

Bhat:2024:SDU

[BS24] K. A. Bhat and S. A. Sofi.

From shallows to depths: unveiling hybrid synthetic data modeling for enhanced learning with privacy considerations in naturally imbalanced datasets. International Journal of Computers and Applications, 46 (12):1088–1103, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2409989.

Bassi:2025:EAM

[BS25] Deepali Bassi and Hardeep Singh. The empirical analysis of multi-objective hyperparameter optimization in software vulnerability prediction. International Journal of Computers and Applications, 47(2):197–215, 2025. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2025.2452849.

Bhattacharya:2010:MBD

[BSB10] S. Bhattacharya, D. Samanta, and A. Basu. Model-based design of scanning input communication aids: State of the art and research issues. International Journal of Computers and Applications, 32(3): 290–296, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2010.
                                        3.202-2597.

Bedi:2019:ESP

[BSG19] Rajeev Kumar Bedi, Jaswinder Singh, and Sunil Kumar Gupta. An efficient and secure privacy preserving multicloud storage framework for mobile devices. International Journal of Computers and Applications, 43 (5):472–482, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1572847.

Bouziane:2021:TAC

[BTAB21] Radhia Bouziane, Labib Sadek Terrissa, Soheyb Ayad, and Jean-François Brethé. Towards an architecture for cloud robotic services. International Journal of Computers and Applications, 44 (9):863–874, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2021.1964790.

Bouzidi:2018:ECP

[BTZA18] Zohra Bouzidi, Labib Sadek Terrissa, Noureddine Zerhouni, and Soheyb Ayad. An efficient cloud prognostic approach for aircraft engines fleet trending. International Journal of Computers and Applications, 42 (5):514–529, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1488024.

Bai:2007:WDA

[BW07] Y. W. Bai and Y. C. Wu. Web delay analysis and reduction by using load balancing of a DNS-based Web server cluster. International Journal of Computers and Applications, 29(1):79–88, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441835.

Bataineh:2003:OMS

[BY03] S. Bataineh and A. Younis.

An optimal Manhattan street network-routing algorithm. International Journal of Computers and Applications, 25 (2):146–153, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2003.11441696.

Bennai:2021:FAM

[BYBM21] Yani-Athmane Bennai, Samira Yessad, and Louiza Bouallouche-Medjkoune. A flexible and adaptive medium access control protocol for improving quality of service in vehicular ad-hoc networks. International Journal of Computers and Applications, 44 (10):929–938, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2021.1988234.

Bouhouch:2023:DEC

[BZT23] Laila Bouhouch, Mostapha Zbakh, and Claude Tadonki. DFMCloudsim: an extension of Cloudsim for modeling and simulation of data fragments migration over distributed data centers. International Journal of Computers and Applications, 46 (1):1–20, 2023. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2277554.

Chalasani:2005:SAC

[CB05] S. Chalasani and R. V. Boppana. Software architectures for e-commerce computing systems with external hosting. International Journal of Computers and Applications, 27 (3):190–198, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441769.

Crocker:2006:FTR

[CBL06] M. Crocker, J. Baranski, and G. Y. Lazarou. Fault-tolerant reconfigurable Ethernet-based IP network proxy. International Journal of Computers and Applications, 28(3): 270–277, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441812.

Cheng-Bing:2018:ASO

[CBXh18] Li Cheng-Bing and Mao Xihao. Array sensors online pattern recognition based on FCM and ANFIS. International Journal of Computers and Applications, 43 (4):352–359, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1550167.

Chuang:2006:USF

[CC06] J.-C. Chuang and C.-C.

Chang. Using a simple and fast image compression algorithm to hide secret information. International Journal of Computers and Applications, 28(4): 329–333, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441818.

Chambel:2001:HWM

[CCG01] T. Chambel, N. Correia, and N. Guimarães. Hypervideo on the Web: Models and techniques for video integration. International Journal of Computers and Applications, 23 (2):90–98, 2001. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2001.11441637.

Chen:2010:BCD

[CCLC10] J. Chen, T. S. Chen, C. N. Lin, and C. Y. Cheng. A bitrate controlled data hiding scheme for Jpeg2000. International Journal of Computers and Applications, 32(2): 238–241, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2010.11441981.

Chen:2013:WSB

[CCZC13] Guoming Chen, Qiang Chen, Dong Zhang, and Yiqun Chen. A watermarking scheme based on compressive sensing and Bregman iteration. International Journal of Computers and Applications, 35(4): 173–180, 2013. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2013.
4.202-3844.

Charif:2013:PIH

[CDB13] Fella Charif, Noureddine Djedi, and Abderrazak Benchabane. On parallel implementation of Horn and Schunck motion estimation method. International Journal of Computers and Applications, 35 (2):79–85, 2013. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2013.
2.202-3506.

Chan:2020:ODA

[CDE20] Sophal Chan, Kwankamon Dittakan, and Subhieh El Salhi. Osteoarthritis detection by applying quadtree analysis to human joint knee X-ray imagery. International Journal of Computers and Applications, 44(6):571–578, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1838145.

Cirillo:2021:ECB

[CDM^{+}21] Albenzio Cirillo, Vito Dalena, Antonio Mauro, Francesco Mogavero, Diego Pennino, Maurizio Pizzonia, Andrea Vitaletti, and Marco Zecchini. Empowering citizens by a blockchain-based Robinson list. International Journal of Computers and Applications, 44(10):920–928, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2021.1986245.

Cherukuri:2024:FLB

[CDR24] Raghavendra Varma Cherukuri, Golagani Lavanya Devi, and Neelapu Ramesh. Federated learning with blockchainbased model aggregation and incentives. International Journal of Computers and Applications, 46 (6):407–417, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2328492.

Choi:2012:SWS

[CEJ12] Seonho Choi, Hyeonsang Eom, and Edward Jung. Securing wireless sensor networks against broadcast service attacks. International Journal of Computers and Applications, 34(3):185–191, 2012. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

/www.tandfonline.com/doi/
full/10.2316/Journal.202.
2012.3.202-3335.

Chawla:2023:CBE

[CG23] Shobha Chawla and Neha Gupta. A cloud based enhanced CPABE framework for efficient user and attributelevel revocation. International Journal of Computers and Applications, 45 (7–8):523–533, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2250149.

Chen:2001:RRC

[CGC01] Jiann-Liang Chen, Jian-Hong Gong, and Han-Chieh Chao. Resource reservation for cellular data services using fuzzy matrices. International Journal of Computers and Applications, 23(1): 51–59, 2001. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2001.11441634.

Chen:2017:RRT

[CGC17] Liangchen Chen, Shu Gao, and Xiufeng Cao. Research on real-time outlier detection over big data streams. International Journal of Computers and Applications, 42 (1):93–101, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1397388.

Cui:2011:RAO

[CGH11] Yaodong Cui, Tianlong Gu, and Wei Hu. Recursive algorithms for the optimum cutting of equal rectangles. International Journal of Computers and Applications, 33 (2):103–107, 2011. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2011.
2.202-2422.

Chowdhury:2018:EDI

[CGI18] Mozammel Chowdhury, Junbin Gao, and Rafiqul Islam. Extracting depth information from stereo images using a fast correlation matching algorithm. International Journal of Computers and Applications, 42 (8):798–803, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1494895.

Chunlin:2012:SDI

[CGL12] Li Chunlin, Chen Gang, and Li Layuan. Sensor data integration with the grid: Model and algorithm. International Journal of Computers and Applications, 34(4): 235–240, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2012.
                                        4.202-3209.

Cheng:2013:NIB

[CGY^{+}13] Xiangguo Cheng, Lifeng Guo, Jia Yu, Huiran Ma, and Yuexiu Wu. A new identity-based group signature scheme. International Journal of Computers and Applications, 35(1):1–5, 2013. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/

10.2316/Journal.202.2013.
1.202-3136.

Chandesa:2010:ASU

[CH10] Tissa Chandesa and Michael Hartley. Automated sketcher using edge detection techniques. International Journal of Computers and Applications, 32(4):404–411, 2010. ISSN 1206-212X (print), 1925-7074 (electronic). URL

https://www.tandfonline.
com/doi/full/10.2316/Journal.
202.2010.4.202-2160.

Chao:2001:CAS

[CHC01] Han-Chieh Chao, Jenn-Yuh Hong, and Jiann-Liang Chen. Channel assignment schemes for WDM-based personal communications networks. International Journal of Computers and Applications, 23 (1):30–34, 2001. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2001.11441631.

Cheong:2004:WSW

[Che04] C. W. Cheong. Web server workload prediction: Fuzzy Markovian approach. International Journal of Computers and Applications, 26 (2):1–6, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2004.11441730.

Chen:2007:NTD

[Che07] S.-N. Chen. A new tracedriven shared-memory multiprocessors machine simulator. International Journal of Computers and Applications, 29 (3):239–244, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2007.11441853.

Chen:2020:NOI

[Che20] Li Chen. Neighborhood optimization of intelligent wireless mobile network based on big data technology. International Journal of Computers and Applications, 43 (4):391–400, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2020.1723258.

Chien:2008:EPA

[Chi08] H. Y. Chien. Efficient and practical approach to authenticating public terminals. International Journal of Computers and Applications, 30 (4):319–324, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2008.11441911.

Chen:2001:TMA

[CHL01] Wei-Chou Chen, Tzung-Pei Hong, and Wen-Yang Lin. Three maintenance algorithms for compressed object-oriented data warehousing. International Journal of Computers and Applications, 23(1): 68–75, 2001. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2001.11441624.

Chong:2000:FDF

[CHY00] L. S. K. Chong, S. C. Hui, and C. K. Yeo. Faxportal: Distributed FAX transportation for Internet FAX services. International Journal of Computers and Applications, 22 (2):78–88, 2000. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2000.11441604.

Chowdhuri:2018:SSS

[CJG18] Partha Chowdhuri, Biswapati Jana, and Debasis Giri. Secured steganographic scheme for highly compressed color image using weighted matrix through DCT. International Journal of Computers and Applications, 43 (1):38–49, 2018. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1505024.

Chouliaras:2007:TPM

[CJNY^{+}07] V. A. Chouliaras, T. R. Jacobs, J. L. Núñez-Yanez, K. Manolopoulos, K. Nakos, and D. Reisis. Threadparallel MPEG-2 and MPEG-4 encoders for shared-memory system-on-chip multiprocessors. International Journal of Computers and Applications, 29(4):353–361, 2007. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

                                        /www.tandfonline.com/doi/
                                        full/10.1080/1206212X.2007.
                                        11441866.

Cai:2023:IIS

[CJW^{+}23] Jing Cai, Jiawei Jiang, Yibing Wang, Jianwei Zheng, and Honghui Xu. Image inpainting via Smooth Tucker decomposition and low-rank Hankel constraint. International Journal of Computers and Applications, 45 (6):421–432, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2219836.

Chang:2009:TIC

[CKC09] Chin-Chen Chang, Ching-Lin Kuo, and Chang-Chu Chen. Three improved codebook searching algorithms for image compression using vector quantizer. International Journal of Computers and Applications, 31(1): 16–22, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2009.11441919.

Chain:2015:EKA

[CKC15] Kai Chain, Wen-Chung Kuo, and Kuei-Hu Chang. Enhancement key agreement scheme based on chaotic maps. International Journal of Computers and Applications, 37 (2):67–72, 2015. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2015.1088211.

Chunlin:2007:MQG

[CL07] L. Chunlin and L. Layuan.

A multi-QoS guaranteed dynamic grid resource scheduling algorithm. International Journal of Computers and Applications, 29(3): 245–252, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441854.

Chunlin:2010:MGC

[CL10] L. Chunlin and L. Layuan.

Multiple granularity control scheme for system utility optimization in grid environments. International Journal of Computers and Applications, 32 (3):282–289, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2010.
3.202-2588.

Chunlin:2011:DRA

[CL11a] Li Chunlin and Li Layuan.

Device resource allocation in context-aware mobile grid. International Journal of Computers and Applications, 33 (1):57–63, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2011.
                                        1.202-3000.

Chunlin:2011:MBR

[CL11b] Li Chunlin and Li Layuan. A market-based resource allocation policy in ad hoc grid. International Journal of Computers and Applications, 33 (3):252–257, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2011.
                                        3.202-3134.

Crocker:2009:BDM

[CLBP09] M. Crocker, G. Y. Lazarou, J. Baca, and J. Picone. A bandwidth determination method for IPv6-based networks. International Journal of Computers and Applications, 31(2):109–118, 2009. ISSN 1206-212X (print), 1925-7074 (electronic). URL

                                        https://www.tandfonline.
                                        com/doi/full/10.1080/1206212X.
                                        2009.11441931.

Chou:2010:SVD

[CLLC10] C. Y. Chou, C. J. Lin, J. L. Lu, and T. W. Chan. Supporting variable dyad learning activities in an instructional game system. International Journal of Computers and Applications, 32(3): 267–274, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2010.
3.202-2534.

Chen:2017:SSC

[CLTY17] Xia Chen, Chang Lu, Qiaoyu Tan, and Guoxian Yu. Semisupervised classification based on clustering adjusted similarity. International Journal of Computers and Applications, 39(4):210–219, 2017. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

/www.tandfonline.com/doi/
full/10.1080/1206212X.2017.
1329262.

Cao:2010:UDV

[CLX10] Tianjie Cao, Dongdai Lin, and Rui Xue. Universal designated-verifier partially blind signatures for e-commerce. International Journal of Computers and Applications, 32(4):399–403, 2010. ISSN 1206-212X (print), 1925-7074 (electronic). URL

https://www.tandfonline.
com/doi/full/10.2316/Journal.
202.2010.4.202-1992.

Chebbout:2020:HCM

[CM20] Samira Chebbout and Hayet Farida Merouani. A hybrid codebook model for object categorization using two-way clustering based codebook generation method. International Journal of Computers and Applications, 44 (2):178–186, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2020.1712775.

Chowdhury:2020:OFE

[CMG20] Subrata Chowdhury, P. Mayilvahanan, and Ramya Govindaraj. Optimal feature extraction and classificationoriented medical insurance prediction model: machine learning integrated with the internet of things. International Journal of Computers and Applications, 44 (3):278–290, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2020.1733307.

Cha:2000:RTR

[CO00] Hojung Cha and Jaehak Oh. A real-time resource management system for video server. International Journal of Computers and Applications, 22 (1):13–22, 2000. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2000.11441594.

Chowdhuri:2017:IDH

[CPJ17] Partha Chowdhuri, Pabitra Pal, and Biswapati Jana. Improved data hiding capacity through repeated embedding using modified weighted matrix for color image. International Journal of Computers and Applications, 41 (3):218–232, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1422587.

Chin:2018:ACS

[CPKT18] Jeng Feng Chin, Joshua Prakash, Shahrul Kamaruddin, and Melissa Chea Ling Tan. Automating computer simulation and statistical analysis in production planning and control research. International Journal of Computers and Applications, 40 (1):25–41, 2018. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1395104.

Chinivar:2024:OAD

[CRAV24] Sneha Chinivar, M. S. Roopa, J. S. Arunalatha, and K. R. Venugopal. Online aggression detection using ensemble techniques on finetuned transformer-based language models. International Journal of Computers and Applications, 46 (8):567–579, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2380649.

Choi:2003:FAS

[CS03] S. J. Choi and W. Scacchi.

Formal analysis of the structural correctness of Slc descriptions. International Journal of Computers and Applications, 25(2):91–97, 2003. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2003.11441688.

Coman:2009:ASD

[CS09] I. D. Coman and A. Sillitti. Automated segmentation of development sessions into task-related subsections. International Journal of Computers and Applications, 31 (3):159–166, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2009.11441937.

Chetana:2022:CAC

[CS22] V. Lakshmi Chetana and Hari Seetha. CF-AMVRGO: Collaborative filtering based adaptive moment variance reduction gradient optimizer for movie recommendations. International Journal of Computers and Applications, 44 (11):1015–1023, 2022. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2022.2097769.

Choi:2019:ISP

[CSB19] Hyunkyoung Choi, Yong-Hyeon Shin, and Hyokyung Bahn. Improving storage performance of high-performance computing systems by using the non-volatile buffer cache. International Journal of Computers and Applications, 43 (9):949–953, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1662168.

Changhai:2017:VAI

[CSFS17] Huang Changhai, Hu Shenping, Kong Fancun, and Xuan Shaoyong. Vulnerability analysis of inland waterways network base on complex network theory. International Journal of Computers and Applications, 42(1):67–75, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1397385.

Chang:2013:IEA

[CT13] Chia-Ying Chang and Shen-Chuan Tai. Image enhancement algorithm based on zone system and dualexposure techniques. International Journal of Computers and Applications, 35(4): 162–172, 2013. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2013.
4.202-3821.

Chen:2015:CCS

[CT15] Zih-Siou Chen and Shen-Chuan Tai. Corrected centersurround retinex: application to tone reproduction for high dynamic range images. International Journal of Computers and Applications, 37 (1):37–51, 2015. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2015.1076604.

Chen:2004:EIR

[CTLT04] C.-Y. Chen, J. Tang, D.- L. Lee, and J.-F. Tu. Exploiting instruct recognized scheme to improve processor performance. International Journal of Computers and Applications, 26(2): 1–9, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2004.11441729.

Chen:2012:GRD

[CTW12] Bo-Jhih Chen, Shen-Chuan Tai, and Yung-Gi Wu. A Gaussian rate-distortion method for detecting zero-quantized blocks. International Journal of Computers and Applications, 34(3):200–212, 2012. ISSN 1206-212X (print), 1925-7074 (electronic). URL

                                        https://www.tandfonline.
                                        com/doi/full/10.2316/Journal.
                                        202.2012.3.202-3417.

Chaya:2019:PAE

[CU19] Jagtap D. Chaya and Rani N. Usha. Predictive analysis by ensemble classifier with machine learning models. International Journal of Computers and Applications, 45(1):19–26, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1675019.

Curcio:2002:MVQ

[Cur02] I. D. D. Curcio. Mobile video QoS metrics. International Journal of Computers and Applications, 24 (2):41–51, 2002. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2002.11441659.

Chen:2003:CCK

[CW03] W. Chen and B. Wasson. Coordinating collaborative knowledge building. International Journal of Computers and Applications, 25(1):1–10, 2003. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2003.11441678.

Cheung:2003:NIT

[CX03] Y.-M. Cheung and L. Xu.

A new information-theoretic based ICA algorithm for blind signal separation. International Journal of Computers and Applications, 25(2): 106–110, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.

tandfonline.com/doi/full/
10.1080/1206212X.2003.11441691.

Chen:2024:ASM

[CXJ24] Lan Chen, Rui Xiong, and Yifan Ji. Application of SVM model based on collaborative filtering hybrid algorithm in e-commerce recommendation. International Journal of Computers and Applications, 46(5):292–300, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2309809.

Chen:2020:RAR

[CZC20] Caixia Chen, Sheng Zhou, and Qingqing Chang. Retracted article: The role of computer security management in preventing financial technology risks. International Journal of Computers and Applications, 45(2):187–195, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1080/1206212X.2020.1738664.

Chen:2017:IRC

[CZL^{+}17] Hao Chen, Dongmei Zhao, Kan Liu, Wen Wu, and Li Wang. Improving reliability of cloud computing services for next generation mobile communication networks. International Journal of Computers and Applications, 40(3):1–7, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1413626.

Chen:2018:MDO

[CZZ18] Wanglong Chen, Lifeng Zhu, and Xiaorui Zhang. Modeling deformable objects using local rigid body simulation. International Journal of Computers and Applications, 42(5):439–448, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1479621.

Cao:2023:MMT

[CZZ^{+}23] Yang Cao, Junhai Zhai, Wei Zhang, Xuesong Zhou, and Feng Zhang. MTTF: a multimodal transformer for temperature forecasting. International Journal of Computers and Applications, 46 (2):122–135, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2289708.

Dagher:2016:DPS

[DA16] Issam Dagher and Rima Antoun. Different PCA scenarios for email filtering. International Journal of Computers and Applications, 38 (1):41–54, 2016. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.

                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2016.1218237.

Dicheva:2003:CCA

[DAC03] D. Dicheva, L. Aroyo, and A. Cristea. Cooperative courseware-authoring support. International Journal of Computers and Applications, 25(3):179–187, 2003. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

                                        /www.tandfonline.com/doi/
                                        full/10.1080/1206212X.2003.
                                        11441701.

Dai:2019:MCD

[Dai19] Xiuzhen Dai. Multilevel CLOS data stream exchange algorithm based on smoothed round-robin fair scheduling. International Journal of Computers and Applications, 44 (8):721–728, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1710361.

Darehmiraki:2009:PAS

[Dar09] M. Darehmiraki. A parallel algorithm for solving sat problem based on DNA computing. International Journal of Computers and Applications, 31 (2):128–131, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2009.11441933.

Darwish:2020:CAD

[Dar20] R. R. Darwish. A congestionaware decision-driven architecture for information-centric Internet-of-Things applications. International Journal of Computers and Applications, 44(4):324–337, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1738088.

Dayal:2017:IAF

[Day17]  Abhinav Dayal.  Improv-

ing adaptive frameless rendering. International Journal of Computers and Applications, 40(2):110–120, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1397341.

Ding:2004:GGM

[DB04] C. H. Ding and R. Buyya.

Guided Google: a meta search engine and its implementation using the Google Distributed Web Services. International Journal of Computers and Applications, 26 (3):1–7, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2004.11441740.

Djama:2020:IMV

[DB20] Ouahiba Djama and Zizette Boufaida. Instantiation of the multi-viewpoints ontology from a resource. International Journal of Computers and Applications, 44 (2):154–165, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1711615.

Das:2016:ACB

[DD16] Sujit Kumar Das and Bibhas Chandra Dhara. Arithmetic coding based secret image sharing with inverted pattern LSB substitution for share hiding to achieve better quality stego images. International Journal of Computers and Applications, 38 (2–3):134–143, 2016. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2016.1224400.

Dong:2008:DSE

[DDL08] Y. Dong, Y. Dou, and M. Liu. A design space exploration algorithm in compiling window operation onto reconfigurable hardware. International Journal of Computers and Applications, 30(1): 36–43, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441881.

Darwish:2017:GPO

[DE17] R. R. Darwish and Abdullah Elewi. A green proactive orchestration architecture for cloud resources. International Journal of Computers and Applications, 41 (2):112–128, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1408983.

Lucia:2009:PSI

[DF09] A. De Lucia and F. Ferrucci. Preface: Special issue on software engineering. International Journal of Computers and Applications, 31(3): 211–212, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2009.11441943.

DiMartino:2009:EWA

[DG09] S. Di Martino and C. Gravino.

Estimating Web application development effort using COSMIC-FFP method. International Journal of Computers and Applications, 31(3): 153–158, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2009.11441936.

Dey:2018:SRT

[DG18] Sankhanil Dey and Ranjan Ghosh. A smart review and two new techniques using 4-bit Boolean functions for cryptanalysis of 4bit crypto S-boxes. International Journal of Computers and Applications, 43 (3):199–217, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1504459.

Dharminder:2019:SAA

[DG19] Dharminder Dharminder and Pratik Gupta. Security analysis and application of Chebyshev chaotic map in the authentication protocols. International Journal of Computers and Applications, 43 (10):1095–1103, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1682238.

Deufemia:2009:AGC

[DGR09] V. Deufemia, C. Gravino, and M. Risi. Automatically generated case workbenches: A preliminary case study. International Journal of Computers and Applications, 31(3): 173–182, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2009.11441939.

Deville:2005:SCE

[DHSR05] D. Deville, Y. Hodique, and I. Simplot-Ryl. Safe collaboration in extensible operating systems: A study on real-time extensions. International Journal of Computers and Applications, 27(1): 20–26, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441752.

Di:2018:OTF

[Di18] Zhang Di. Online traffic flow assignment method based on structure matching selection of multilevel road network. International Journal of Computers and Applications, 43(2):171–175, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1536393.

Ding:2011:IMV

[Din11] Ing-Jr Ding. Improvement of map-VFS adaptation performance by fuzzy control. International Journal of Computers and Applications, 33(2): 116–123, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2011.
2.202-2733.

Desrochers:2004:AGP

[DJQS04] D. Desrochers, Y. Jin, Z. Qu, and A. Saengdeejing. Algorithms to generate partially damaged characters and readability study for OCR readers in semiconductor manufacturing. International Journal of Computers and Applications, 26(4):1–6, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2004.11441742.

Dada:2005:XBO

[DK05] J. O. Dada and H.-D. Kochs. XML-based open electricity market information exchange network using objectoriented methods. International Journal of Computers and Applications, 27(3): 153–160, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2005.11441772.

Deepti:2023:SAR

[DK23] S. Renu Deepti and P. C. Karthik. Survey on agerelated macular degeneration detection in OCT image. International Journal of Computers and Applications, 46 (2):101–112, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2286032.

Daoui:2004:HCC

[DLDB04] M. Daoui, M. Lalam, B. Djamah, and A. Bilami. Hardware cache coherence protocol. International Journal of Computers and Applications, 26 (4):1–13, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.

tandfonline.com/doi/full/
10.1080/1206212X.2004.11441749.

Dini:2005:DFD

[DLL05] G. Dini, G. Lettieri, and L. Lopriore. Design framework for a distributed, singleaddress-space system. International Journal of Computers and Applications, 27(2): 108–118, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441763.

Dubey:2020:CRA

[DM20] Ajay Kumar Dubey and Vimal Mishra. Crowd review and attribute-based credit computation for an access control mechanism in cloud data centers. International Journal of Computers and Applications, 45(2):212–219, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1746500.

Das:2023:SED

[DMB23] Subhajit Das, Sourav Mandal, and Rohini Basak. Spam email detection using a novel multilayer classification-based decision technique. International Journal of Computers and Applications, 45 (9):587–599, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2258328.

deMoraes:2003:HMM

[dMM03] R. M. de Moraes and L. S. Machado. Hidden Markov models for learning evaluation in virtual-reality simulators. International Journal of Computers and Applications, 25 (3):212–215, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2003.11441705.

deMacedo:2014:UEA

[dMRF^{+}14] Daniel V. de Macedo, Maria Andréia F. Rodrigues, João J. V. P. Furtado, Elizabeth S. Furtado, and Daniel A. Chagas. Using and evaluating augmented reality for mobile data visualization in real estate classified Ads. International Journal of Computers and Applications, 36 (1):7–14, 2014. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2014.
                                        1.202-3737.

Dehkordi:2025:UAT

[DND25] Sajad Balali Dehkordi, Saeed Nasri, and Sina Dami. Unveiling anomalies: transformative insights from transformerbased autoencoder models. International Journal of Computers and Applications, 47 (1):29–44, 2025. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2441147.

Das:2019:FCC

[DNP19] Poulami Das, Sudip Kumar Naskar, and Sankar Narayan Patra. Fast converging cuckoo search algorithm to design symmetric FIR filters. International Journal of Computers and Applications, 43 (6):547–565, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1588512.

Dash:2024:ECP

[DNP24] Alina Dash, Kshiramani Naik, and Priyanka Priyadarshini. Enhancing copyright protection with Hadamard transformbased blind digital watermarking using the cycling chaos system. International Journal of Computers and Applications, 46(7):467–480, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2336329.

Dong:2018:DAP

[Don18]  Yukun Dong.  Detecting

all potential null dereferences based on points-to property sound analysis. International Journal of Computers and Applications, 42 (2):187–195, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1476805.

Devedzic:2003:WDC

[DP03] V. Devedzic and V. Pocajt. What does current Web-based education lack? International Journal of Computers and Applications, 25(1): 65–71, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2003.11441686.

Desai:2017:ARB

[DP17] Rahul Desai and B. P. Patil. Adaptive routing based on predictive reinforcement learning. International Journal of Computers and Applications, 40(2):73–81, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2017.1395106.

David:2015:SCT

[DR15] Beulah David and Dorai Rangasamy. Spatial-contextual texture and edge analysis approach for unsupervised change detection of faces in counterfeit images. International Journal of Computers and Applications, 37(3–4): 143–159, 2015. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.

tandfonline.com/doi/full/
10.1080/1206212X.2016.1188555.

Duo:2018:ABT

[DRF18] Zhang Duo and Guo Rui-Fang. An adaptive Bayesian tracking algorithm for Internet of motion parameters based on explicit formulation and integral contour model. International Journal of Computers and Applications, 43 (4):320–325, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1539435.

Dey:2010:EAP

[DS10] S. Dey and D. Samanta. An efficient and accurate pupil detection method for iris biometric processing. International Journal of Computers and Applications, 32(2): 141–148, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2010.11441970.

Domnic:2011:FLB

[DS11] Sandanam Domnic and Karuppanagounder Somasundaram. Fast and low bit-rate AMBTC image compression scheme. International Journal of Computers and Applications, 33 (2):108–115, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/

10.2316/Journal.202.2011.
2.202-2528.

Dhanith:2019:OLB

[DS19a] P. R. Joe Dhanith and B. Surendiran. An ontology learning based approach for focused web crawling using combined normalized pointwise mutual information and resnik algorithm. International Journal of Computers and Applications, 44 (12):1123–1129, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1684023.

Dharani:2019:LUL

[DS19b] M. Dharani and G. Sreenivasulu. Land use and land cover change detection by using principal component analysis and morphological operations in remote sensing applications. International Journal of Computers and Applications, 43(5):462–471, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1578068.

Debakla:2019:CBF

[DSBR19] Mohammed Debakla, Mohammed Salem, Rochdi Bachir Bouiadjra, and Mohammed Rebbah. CMA-ES based fuzzy clustering approach for MRI images segmentation. International Journal of Computers and Applications, 45(1):1–7, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1662984.

Ding:2013:FFI

[DSD^{+}13] Li Ding, Dana Steil, Brandon Dixon, Nicholas A. Kraft, David B. Brown, and Allen Parrish. First: Framework to integrate relationship search tools. International Journal of Computers and Applications, 35(3):114–124, 2013. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

/www.tandfonline.com/doi/
full/10.2316/Journal.202.
2013.3.202-3609.

Deepak:2018:HST

[DSP^{+}18] Gerard Deepak, B. N. Shwetha, C. N. Pushpa, J. Thriveni, and K. R. Venugopal. A hybridized semantic trustbased framework for personalized web page recommendation. International Journal of Computers and Applications, 42(8):729–739, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1480472.

dosSantos:2005:SNP

[dSTE05] A. L. M. dos Santos, M. E.

Torrey, and A. El Sheshai. Supporting national public key infrastructures using smart cards. International Journal of Computers and Applications, 27(1): 35–40, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441754.

Deva:2012:FOD

[DSV12] Jemi Florinabel Deva, Ebenezer Juliet Selwyn, and Sadasivam Velayuthum. Fast orientationdriven multi-structure morphological inpainting. International Journal of Computers and Applications, 34(2): 127–134, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2012.
2.202-3343.

Dikii:2020:DAE

[DT20] Dmitrii Dikii and Aleksey Tikhomirov. Detection of DoS attacks exploiting SUB-SCRIBE messages of the MQTT protocol. International Journal of Computers and Applications, 44 (6):579–585, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1846945.

DeMara:2006:CCM

[DTDE06] R. F. DeMara, Y. Tseng, K. Drake, and A. Ejnioui. Capability classes of multiprocessor synchronization techniques. International Journal of Computers and Applications, 28(4):342–349, 2006. ISSN 1206-212X (print), 1925-7074 (electronic). URL

https://www.tandfonline.
com/doi/full/10.1080/1206212X.
2006.11441820.

Devi:2019:IPS

[DU19] M. Kalpana Devi and K. Umamaheswari. Intelligent process of spectrum handoff for dynamic spectrum access in cognitive radio network using swarm intelligence. International Journal of Computers and Applications, 44 (12):1186–1192, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1704483.

Duan:2007:RAQ

[Dua07] Q. Duan. Resource allocation for quality of service provision in buffered crossbar switches with traffic aggregation. International Journal of Computers and Applications, 29(3): 283–290, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441858.

Duan:2023:MSR

[Dua23] Fang Duan. Multi-scale residual aggregation feature network based on multi-time division for motion behavior recognition. International Journal of Computers and Applications, 45 (6):452–459, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2232169.

Dudheria:2019:APP

[Dud19] Rishabh Dudheria. Assessing password practices of mobile apps. International Journal of Computers and Applications, 44(1):64–82, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1696446.

Dwivedi:2020:DAD

[DVT20] Shubhra Dwivedi, Manu Vardhan, and Sarsij Tripathi. Defense against distributed DoS attack detection by using intelligent evolutionary algorithm. International Journal of Computers and Applications, 44 (3):219–229, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2020.1720951.

Du:2023:SNV

[DWT23] Yuanhan Du, Ling Wang, and Yebo Tao. Server node video processing based on feature depth analysis algorithm. International Journal of Computers and Applications, 46(1):58–65, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2283648.

Dagao:2018:PTC

[DY18] Duan Dagao and Gao Yang.

Preprocessing technology of consuming big data based on user interest with Internet of Location Mining. International Journal of Computers and Applications, 43 (2):147–152, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1534372.

Dianhui:2018:PMM

[DZ18] Mao Dianhui and Song Zihao.

A prediction model of microblog affective hotspots based on SVM collaborative filtering recommendation model. International Journal of Computers and Applications, 43 (2):176–180, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1536394.

Djemili:2023:EDL

[DZ23] Rafik Djemili and Merouane Zamouche. An efficient deep learning-based approach for human activity recognition using smartphone inertial sensors. International Journal of Computers and Applications, 45(4):323–336, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2198785.

El-Awad:2005:VCG

[EA05] M. M. El-Awad. A vectorized conjugate-gradient solver for sparse systems of algebraic equations. International Journal of Computers and Applications, 27(4): 260–265, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2005.11441779.

Elleuchi:2019:EER

[EBA19] Manel Elleuchi, Manel Boujeleben, and Mohamed Abid. Energy-efficient routing model for water pipeline monitoring based on wireless sensor networks. International Journal of Computers and Applications, 44(1):3–11, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1682239.

Elsayed:2020:TLD

[ECBR20] Hoda Ahmed Galal Elsayed, Soumaya Chaffar, Samir Brahim Belhaouari, and Hafsa Raissouli. A two-level deep learning approach for emotion recognition in Arabic news headlines. International Journal of Computers and Applications, 44(7):604–613, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1851501.

El-Ghazawi:2005:PWD

[EGL05] T. A. El-Ghazawi and J. Le Moigne. Performance of the wavelet decomposition on massively parallel architectures. International Journal of Computers and Applications, 27(2):72–81, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441760.

Emar:2014:CBS

[EHA14] Walid Emar, Sofyan Hayajneh, and Musbah J. Aqel. A computer based sliding mode controller topology for nonlinear voltage regulating chopper. International Journal of Computers and Applications, 36(3):110–114, 2014. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

/www.tandfonline.com/doi/
full/10.2316/Journal.202.
2014.3.202-3950.

El-Henawy:2016:ACB

[EHA16] Ibrahim El-Henawy and Kareem Ahmed. Accelerating convergence of backpropagation for multilayer perceptron neural networks: a case study on character bit-mapped pixel image to ASCII conversion. International Journal of Computers and Applications, 38 (1):9–18, 2016. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2016.1188560.

El-haryqy:2024:RDL

[EhMZ24] Noureddine El-haryqy, Zhour Madini, and Younes Zouine. A review of deep learning techniques for enhancing spectrum sensing and prediction in cognitive radio systems: approaches, datasets, and challenges. International Journal of Computers and Applications, 46 (12):1104–1128, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2414042.

Edwards:2005:LTU

[EJ05] G. Edwards and R. Jayne. Location tracking using differential range measurements. International Journal of Computers and Applications, 27 (3):199–205, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441768.

Etem:2023:FIE

[EK23] Taha Etem and Turgay Kaya.

Fast image encryption algorithm with random structures. International Journal of Computers and Applications, 45(10):626–637, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2260617.

Eu:2008:MHF

[EKAO08] Y. C. Eu, S. Khatun, B. M.

Ali, and M. Othman. Multicast, hierarchical and fast handover in mobile IPv6 wireless network: a testbed experience. International Journal of Computers and Applications, 30(3): 207–212, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441900.

Elkouay:2024:GBP

[EMM24] Abdelali Elkouay, Najem Moussa, and Abdallah Madani. Graph-based phishing detection: URLGBM model driven by machine learning. International Journal of Computers and Applications, 46 (7):481–495, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2342710.

Esmaili:2017:NMF

[EMN17] Somayeh Saraf Esmaili, Keivan Maghooli, and Ali Motie Nasrabadi. A new model for face detection in cluttered backgrounds using saliency map and C_{2} texture features. International Journal of Computers and Applications, 40(4):214–222, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2017.1399721.

ElAsri:2023:EBE

[ENER23] Smail Ait El Asri, Ismail Negabi, Samir El Adib, and Naoufal Raissouni. Enhancing building extraction from remote sensing images through UNet and transfer learning. International Journal of Computers and Applications, 45 (5):413–419, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2219117.

El-Qawasmeh:2007:IEB

[EQSMB07] E. El-Qawasmeh, M. Strauss, M. Mack, and S. Berkovich. Increasing the efficiency of bitcounting. International Journal of Computers and Applications, 29(1):51–58, 2007. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

                                        /www.tandfonline.com/doi/

full/10.1080/1206212X.2007.
11441832.

Fergany:2006:EAD

[FA06] T. Fergany and Sarhan Amany. Efficient allocation of distributed objectoriented tasks to a pre-defined scheduled system. International Journal of Computers and Applications, 28(1): 35–42, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441786.

Fan:2018:KWS

[Fan18] Yang Fan. Key web search algorithm based on service ontology. International Journal of Computers and Applications, 43 (2):165–170, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1535392.

Fu:2020:RAC

[FCW^{+}20] Ting Fu, Hong Chen, Fei Wu, YunDong Su, and Li Zhuang. Retracted article: Collaborative correlation space big data clustering algorithm for abnormal flow monitoring. International Journal of Computers and Applications, 45 (2):136–143, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.

tandfonline.com/doi/abs/
10.1080/1206212X.2020.1727659.

Feyzi:2023:SED

[FD23] Farid Feyzi and Arman Daneshdoost. Studying the effectiveness of deep active learning in software defect prediction. International Journal of Computers and Applications, 45 (7–8):534–552, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2252117.

Fong:2005:WBS

[FH05] A. C. M. Fong and S. C. Hui. A Web-based system for B2B e-commerce. International Journal of Computers and Applications, 27 (4):209–217, 2005. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441773.

Falkemeier:2001:ISA

[FJK01] G. Falkemeier, G. R. Joubert, and O. Kao. Internet supported analysis and presentation of MPEG compressed newsfeeds. International Journal of Computers and Applications, 23(2): 129–136, 2001. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2001.11441641.

Fox:2006:MCO

[FJZ06] B. Fox, L. S. Jennings, and A. Y. Zomaya. On the modelling and computation of nonlinear multibody cable systems. International Journal of Computers and Applications, 28(2): 154–161, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441799.

Fan:2004:DBS

[FL04] C. I. Fan and C. L. Lei. Divisible blind signatures based on hash chains. International Journal of Computers and Applications, 26 (1):1–9, 2004. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2004.11441718.

Faizullah:2006:MPQ

[FM06] S. Faizullah and I. Marsic.

Measuring and pricing QoS. International Journal of Computers and Applications, 28 (3):189–209, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441803.

Fredj:2018:FRO

[FM18] Amira Hadj Fredj and Jihene Malek. A fast and robust OS-RAD filter for telemedicine applications. International Journal of Computers and Applications, 43(1):70–79, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1512235.

Fujieda:2011:AMP

[FMI11] Masaru Fujieda, Takahiro Murakami, and Yoshihisa Ishida. An approach to multi-pitch estimation using a support vector machine. International Journal of Computers and Applications, 33(3): 202–210, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2011.
3.202-2818.

Ferdush:2019:EIE

[FMP^{+}19] Jannatul Ferdush, Greetashree Mondol, Amrita Pritom Prapti, Mahbuba Begum, Mohammad Nowsin Amin Sheikh, and Syed Md. Galib. An enhanced image encryption technique combining genetic algorithm and particle swarm optimization with chaotic function. International Journal of Computers and Applications, 43 (9):960–967, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1662170.

Fasano:2009:SPM

[FO09] F. Fasano and R. Oliveto.

Supporting project management with fine-grained artefact managementin Adams. International Journal of Computers and Applications, 31 (3):145–152, 2009. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2009.11441935.

Farimani:2000:MAS

[FR00] M. Farimani and J. A. Robinson. Multipoint activity sharing services for fast prototyping of groupware applications. International Journal of Computers and Applications, 22 (1):23–28, 2000. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2000.11441595.

Fateh:2018:EBH

[FR18] Mansoor Fateh and Mohsen Rezvani. An email-based high capacity text steganography using repeating characters. International Journal of Computers and Applications, 43(3):226–232, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1517713.

Ferretti:2007:OOB

[FRP07] S. Ferretti, M. Roccetti, and C. E. Palazzi. An optimistic obsolescence-based approach to event synchronization for massively multiplayer online games. International Journal of Computers and Applications, 29 (1):33–43, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441830.

Fei:2017:MHE

[FSL17] Cheng-Yu Fei, Zhong Su, and Qing Li. A method of heading estimation for pedestrian navigation based on information compression MDL criteria. International Journal of Computers and Applications, 42(4):392–398, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1397383.

Fathima:2022:MVE

[FSNK22] M. Dhilsath Fathima, S. Justin Samuel, R. Natchadalingam, and V. Vijeya Kaveri. Majority voting ensembled feature selection and customized deep neural network for the enhanced clinical decision support system. International Journal of Computers and Applications, 44 (10):991–1001, 2022. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/1206212X.2022.2069643.

Fabi:2020:PIT

[FT20] A. K. Fabi and Sabu M.

Thampi. A psychologyinspired trust model for emergency message transmission on the Internet of Vehicles (IoV). International Journal of Computers and Applications, 44 (5):480–490, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1814557.

Fu:2025:APG

[Fu25] Xiaobao Fu. Application of perceptual graph neural network intelligent algorithm in software testing platform. International Journal of Computers and Applications, 47 (2):216–225, 2025. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2025.2452882.

Furini:2007:MRE

[Fur07] M. Furini. MCDL: a reduced but extensible multimedia contents description language. International Journal of Computers and Applications, 29 (2):204–210, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441849.

Wang:2019:MPD

[fWfH19] Fei fei Wang and Hai feng Hu. Multi-path data fusion method based on routing algorithm for wireless sensor networks. International Journal of Computers and Applications, 43 (9):916–923, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1652786.

Fernandez:2000:CFO

[FWQ00] E. B. Fernandez, Jie Wu, and Haifeng Qian. A combined functional and objectoriented approach to software design. International Journal of Computers and Applications, 22(2):51–61, 2000. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2000.11441601.

Garani:2008:QMD

[GA08] G. Garani and G. K. Adam. Qualitative modelling at the design of concrete manufacturing machinery. International Journal of Computers and Applications, 30(4): 325–330, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2008.11441912.

Gupta:2011:NAC

[GA11] Ashutosh Gupta and Suneeta Agarwal. A novel approach for compressing DNA sequences using semi-statistical compressor. International Journal of Computers and Applications, 33(3):245–251, 2011. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

/www.tandfonline.com/doi/
full/10.2316/Journal.202.
2011.3.202-3114.

Gupta:2023:SEC

[GAS23] Megha Gupta, Laxmi Ahuja, and Ashish Seth. Security enhancement in a cloud environment using a hybrid chaotic algorithm with multifactor verification for user authentication. International Journal of Computers and Applications, 45(11):680–696, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2267839.

Guendouz:2020:AAM

[GB20] Amina Guendouz and Djamal Bennouar. AMPL: aspect multiple product lines. International Journal of Computers and Applications, 44 (4):305–315, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1735761.

Gaffour:2019:NCA

[GBBZ19] K. Gaffour, M. K. Benhaoua, A. H. Benyamina, and H. E. Zahaf. A new congestion-aware routing algorithm in network-on-chip: 2D and 3D comparison. International Journal of Computers and Applications, 45(1):27–35, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1679529.

Gupta:2020:MSE

[GCMS20] Sanchay Gupta, Siddharth Chandra, N. Maheswari, and M. Sivagami. A model for screening eye diseases using optical coherence tomography images. International Journal of Computers and Applications, 44 (4):396–400, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1759857.

Geng:2019:RAT

[Gen19] Bin Geng. Retracted article: Traffic prediction and transmission scheduling of artificial intelligence-driven cognitive wireless networks. International Journal of Computers and Applications, 45 (1):96–104, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1080/1206212X.2019.1706812.

Garzon:2004:SEM

[GG04] E. M. Garzón and I. García.

Solving eigenproblems on multicomputers: Two different approaches. International Journal of Computers and Applications, 26(4): 1–10, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2004.11441748.

Grosu:2005:MME

[GG05] D. Grosu and H. Gâlmeanu.

MPEG-2 motion estimation using a PC cluster. International Journal of Computers and Applications, 27(2): 51–56, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441757.

Ghumbre:2010:ISH

[GG10] Shashikant U. Ghumbre and Ashok A. Ghatol. An intelligent system for hepatitis B disease diagnosis. International Journal of Computers and Applications, 32(4): 455–460, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2010.
4.202-2874.

Goyal:2023:ITS

[GGB23] Abhishek Goyal, Rakesh Garg, and Komal Kumar Bhatia. An improved task scheduling algorithm for conflict resolution in cloud environment. International Journal of Computers and Applications, 46(4):218–226, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2301266.

Gautam:2024:FIF

[GGK^{+}24] Yuvika Gautam, Piyush Gupta, Deepika Kumar, Bhavya Kalra, Abhinav Kumar, and Jude D. Hemanth. FusionEXNet: an interpretable fused deep learning model for skin cancer detection. International Journal of Computers and Applications, 46 (9):743–753, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2385923.

Gowda:2023:BIV

[GGMM23] Vishruth B. Gowda, M. T. Gopalakrishna, J. Megha, and Shilpa Mohankumar. Background initialization in video data using singular value decomposition and robust principal component analysis. International Journal of Computers and Applications, 45 (9):600–609, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2258329.

Ge:2006:BEW

[GH06] J. Ge and R. He. Buffer efficiency in wormhole routers. International Journal of Computers and Applications, 28 (4):314–320, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441816.

Gad:2020:CSP

[GH20] Ibrahim Gad and Doreswamy Hosahalli. A comparative study of prediction and classification models on NCDC weather data. International Journal of Computers and Applications, 44 (5):414–425, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1766769.

Ghosh:2017:DTR

[Gho17] Saswati Ghosh. Design and testing of rectifying antenna for RF energy scavenging in GSM 900 band. International Journal of Computers and Applications, 39(1): 36–44, 2017. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2016.1259801.

Guesmi:2011:DIG

[GHRK11] Tarek Guesmi, Salem Hasnaoui, Houria Rezig, and Ouajdi Korbaa. Design and implementation of a global scheduling strategy for data distribution over CAN-based networks. International Journal of Computers and Applications, 33(4):271–283, 2011. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

                                        /www.tandfonline.com/doi/
                                        full/10.2316/Journal.202.
                                        2011.4.202-2637.

Gupta:2023:FBG

[GK23] Tanu Gupta and Ela Kumar. Fusion of Bi-GRU and temporal CNN for biomedical question classification. International Journal of Computers and Applications, 45 (6):460–470, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2235458.

Gaidhane:2019:EAC

[GKMR19] Vilas H. Gaidhane, Nand Kumar, Ravi Kant Mittal, and J. Rajevenceltha. An efficient approach for cement strength prediction. International Journal of Computers and Applications, 45 (1):8–18, 2019. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/1206212X.2019.1673288.

Goularte:2001:MAV

[GM01] R. Goularte and E. D. S.

Moreira. Managing access to a video server through a metadata base. International Journal of Computers and Applications, 23(1): 25–29, 2001. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2001.11441630.

Ghosh:2016:SDD

[GMD16] Mili Ghosh, Debarka Mukhopadhyay, and Paramartha Dutta. A study on 2 dimensional 2 dot 1 electron quantum dot cellular automata based reversible 2:1 MUX design: an energy analytical approach. International Journal of Computers and Applications, 38 (2–3):82–95, 2016. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2016.1218239.

Ghosh:2024:BIA

[GMD^{+}24] Bibek Ranjan Ghosh, Jyotsna Kumar Mandal, Swarnadeep Das, Ayantanu Laha, Rajarshi Saha, and Siddhartha Banerjee. Blind image authentication using singular value decomposition with enhanced robustness by augmenting Hamming code. International Journal of Computers and Applications, 46 (11):1020–1044, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2408748.

Gao:2018:REL

[GMS18] Xuexin Gao, Yongmin Mu, and Meie Shen. The research of endless loop detection method based on the basic path. International Journal of Computers and Applications, 42(8):740–747, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1481182.

Gangappa:2019:CLC

[GMS19] Malige Gangappa, C. Kiran Mai, and P. Sammulal. Classification of land cover images using modified water wave optimizationbased hybrid classifier. International Journal of Computers and Applications, 43 (10):1054–1064, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1660835.

Gnanamani:2011:DDC

[GP11] Maria Kalavathy Gnanamani and Seethalakshmi Pandian. Designing dynamic composite media service and monitoring its run-time performance. International Journal of Computers and Applications, 33 (3):229–237, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2011.
3.202-2980.

Goyani:2017:TMM

[GP17] Mahesh Goyani and Narendra Patel. Template matching and machine learningbased robust facial expression recognition system using multi-level Haar wavelet. International Journal of Computers and Applications, 42 (4):360–371, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1395134.

Geeta:2023:MOC

[GP23] Koppula Geeta and V. Kamakshi Prasad. Multiobjective cloud load-balancing with hybrid optimization. International Journal of Computers and Applications, 45 (10):611–625, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2260616.

Gavade:2019:SAS

[GR19] Anil B. Gavade and Vijay S. Rajpurohit. Systematic analysis of satellite imagebased land cover classification techniques: literature review and challenges. International Journal of Computers and Applications, 43 (6):514–523, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1573946.

Goel:2011:EAM

[GS11] Bindu Goel and Yogesh Singh. An empirical analysis of metrics to predict the software defect fix-effort. International Journal of Computers and Applications, 33(2): 124–131, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2011.
                                        2.202-2749.

Ghosh:2013:AUE

[GS13] Soumalya Ghosh and Debasis Samanta. Analysis on user errors in virtual keyboards. International Journal of Computers and Applications, 35(4):145–151, 2013. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

                                        /www.tandfonline.com/doi/
                                        full/10.2316/Journal.202.
                                        2013.4.202-3673.

Garg:2019:HAS

[GS19] Anil Garg and O. P. Sahu. A hybrid approach for speech enhancement using Bionic wavelet transform and Butterworth filter. International Journal of Computers and Applications, 42 (7):686–696, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1614293.

Garg:2022:SHA

[GS22] Ritu Garg and R. K. Singh.

SBCDetector: a hybrid approach to detect second-order similarity or change. International Journal of Computers and Applications, 45 (3):238–247, 2022. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2022.2149117.

Gorbal:2024:IWS

[GSJ24] Mangalagowri Gorbal, Ramesh D.

Shelke, and Manuj Joshi. An image watermarking scheme: combining the transform domain and deep learning modality with an improved scrambling process. International Journal of Computers and Applications, 46 (5):310–323, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2313301.

Gupta:2019:NAM

[GSP19] Neha Gupta, Arun Sharma, and Manoj Kumar Pachariya. A novel approach for mutant diversity-based fault localization: DAM-FL. International Journal of Computers and Applications, 43 (8):795–804, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1617504.

Gupta:2008:RTU

[GSV08] R. P. Gupta, S. C. Srivastava, and R. K. Varma. Remote terminal units for distribution automation: Development and commissioning experience. International Journal of Computers and Applications, 30 (2):80–91, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2008.11441885.

Gupta:2009:CSD

[GT09] R. Gupta and A. Trivedi. A comparative study of different Rayleigh fading channel simulators. International Journal of Computers and Applications, 31(4): 222–229, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2009.11441945.

Ganeshkumar:2010:BTF

[GT10] Perumal Ganeshkumar and Kuppusamy Thyagarajah. Balancing throughput and fairness for concurrent flows based on per flow scheduling in ad hoc networks. International Journal of Computers and Applications, 32(4): 447–454, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2010.
4.202-2856.

Gautam:2019:VAB

[GT19] K. S. Gautam and Senthil Kumar Thangavel. Video analytics-based facial emotion recognition system for smart buildings. International Journal of Computers and Applications, 43 (9):858–867, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1642438.

Girsang:2014:MOP

[GTY14] Abba Suganda Girsang, Chun-Wei Tsai, and Chu-Sing Yang. Multi-objective particle swarm optimization for repairing inconsistent comparison matrices. International Journal of Computers and Applications, 36(3): 101–109, 2014. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2014.
3.202-3940.

Guo:2019:AFW

[Guo19] Chunhua Guo. The application of fractional wavelet transform in image enhancement. International Journal of Computers and Applications, 43(7):684–690, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1626573.

Gao:2014:WAS

[GYQC14] Shang Gao, Hualong Yu, Ling Qiu, and Cungen Cao. The wading across stream algorithm. International Journal of Computers and Applications, 36(4):127–132, 2014. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

                                        /www.tandfonline.com/doi/
                                        full/10.2316/Journal.202.
                                        2014.4.202-3916.

Hussein:2019:EEC

[HA19] Haval I. Hussein and Wafaa M. Abduallah. An efficient ElGamal cryptosystem scheme. International Journal of Computers and Applications, 43 (10):1088–1094, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1678799.

Hossain:2020:ICD

[HA20] Md. Alamgir Hossain and Md. Abdullah Al Hasan.

Improving cloud data security through hybrid verification technique based on biometrics and encryption system. International Journal of Computers and Applications, 44(5):455–464, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1809177.

Hamza:2005:EMA

[HAA05] M. A. Hamza, I. F. Awad, and G. M. B. Awad. Enhancing mobile agent communication performance. International Journal of Computers and Applications, 27 (2):45–50, 2005. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441756.

Haga:2003:DHC

[Hag03]  H. Haga.  Development

of hypermedia computermediated learning system and its evaluation. International Journal of Computers and Applications, 25(2): 111–118, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2003.11441692.

Hao:2017:DWT

[Hao17] Zhang Hao. Design of WSN traffic forecasting system with delayed self-sensing. International Journal of Computers and Applications, 42 (1):30–35, 2017. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2017.1396427.

Haque:2006:CDD

[Haq06]  W. Haque.  Concurrent

deadlock detection in parallel programs. International Journal of Computers and Applications, 28(1): 19–25, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2006.11441784.

Harrer:2003:SEM

[Har03] A. Harrer. Software engineering methods for RE-USE of components and design in educational systems. International Journal of Computers and Applications, 25(1): 17–23, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2003.11441681.

Hu:2001:PEL

[HB01] J. R. Hu and Y. W. Bai. Power efficiency and latency for a semi-batch power management model in a palmtop multimedia terminal. International Journal of Computers and Applications, 23 (1):10–16, 2001. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2001.11441628.

Hardebolle:2009:MFM

[HB09] C. Hardebolle and F. Boulanger.

Multi-formalism modelling and model execution. International Journal of Computers and Applications, 31(3): 193–203, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2009.11441941.

Hong:2005:FGS

[HC05] T.-P. Hong and T.-N. Chuang.

Fuzzy Gupta scheduling for flow shops with more than two machines. International Journal of Computers and Applications, 27(3): 169–177, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441765.

Huang:2011:NHM

[HC11] Hui-Min Huang and Jia-Lin Chang. Novel handoff method for the integrated 3G and Nemo network. International Journal of Computers and Applications, 33(3): 238–244, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2011.
3.202-3034.

Hu:2017:ESL

[HC17] Min Hu and Wei Cai. Evacuation simulation and layout optimization of cruise ship based on cellular automata. International Journal of Computers and Applications, 42(1):36–44, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2017.1396428.

Hsiao:2016:HFR

[HCC16] Ju-Yuan Hsiao, Shu-Ju Chuang, and Po-Yueh Chen. A hybrid face recognition system based on multiple facial features. International Journal of Computers and Applications, 38(1):1–8, 2016. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2016.1188553.

Hung:2012:IIB

[HCH12] Kuo-Ming Hung, Yen-Liang Chen, and Ching-Tang Hsieh. Image inpainting based on geometric similarity. International Journal of Computers and Applications, 34(1): 11–18, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2012.
                                        1.202-3020.

Hsu:2006:SSB

[HCSC06] W. H. Hsu, J. Chen, S.- T. Sheu, and C.-F. Chao. SBMT–Steiner backup multicast tree. International Journal of Computers and Applications, 28(2):87–98, 2006. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441792.

Hazra:2017:MAV

[HDS17] Rumpa Hazra, Shouvik Dey, and Jayashree Singha. Modeling, analysis and verification of real-time resource access control protocols: a formal approach. International Journal of Computers and Applications, 40(2):63–72, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1367530.

He:2018:OLD

[He18] Taibo He. Optimization of the large data spatial information decision support service system dependent on the solution of the information entropy equation. International Journal of Computers and Applications, 42 (2):179–186, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1475336.

Hamdy:2020:DMO

[HE20] Abeer Hamdy and Gloria Ezzat. Deep mining of open source software bug repositories. International Journal of Computers and Applications, 44(7):614–622, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2020.1855705.

Haraty:2005:AHO

[HEZ05] R. A. Haraty and H. M. El-Zabadani. Abjad Hawwaz: An offline Arabic handwriting recognition system. International Journal of Computers and Applications, 27(3): 178–189, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2005.11441767.

Han:2022:NEE

[HFG22] Yandong Han, Lijuan Feng, and Jian Gao. A new end-to-end framework based on nonlocal network structure and spatial attention mechanism for image rain removal. International Journal of Computers and Applications, 44 (11):1083–1091, 2022. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2022.2112325.

Hossain:2019:SIL

[HFK19] Md. Alamgir Hossain, Jannatul Ferdush, and Marjia Khatun. A study and implementation of largescale log-determinant computation to cloud. International Journal of Computers and Applications, 43 (10):1020–1028, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1648632.

Harmanani:2007:CST

[HH07] H. M. Harmanani and A. M. K.

Hajar. Concurrent BIST synthesis and test scheduling using genetic algorithms. International Journal of Computers and Applications, 29(2): 132–142, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441841.

Hong:2011:HHM

[HHGG11] Lu Hong, Feng Hong, Zhongwen Guo, and Ying Guo. HCR: a hybrid MAC protocol for underwater sensor networks using channel reservation. International Journal of Computers and Applications, 33(2):154–159, 2011. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

/www.tandfonline.com/doi/
full/10.2316/Journal.202.
2011.2.202-3014.

Hasan:2024:ERA

[HHU24] Md Rabiul Hasan, A. B. M. Aowlad Hossain, and Shah Muhammad Azmat Ullah. Ensemble ResDenseNet: Alzheimer’s disease staging from brain MRI using deep weighted ensemble transfer learning. International Journal of Computers and Applications, 46 (7):539–554, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2380648.

Hota:2008:ABM

[HJR08] C. Hota, S. Jha, and G. Raghurama. Adaptive bandwidth management and QoS provisioning in IPVPNs. International Journal of Computers and Applications, 30(2): 142–150, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2008.11441893.

Hosseinabadi:2020:INI

[HKE20] Saeed Hosseinabadi, Manoochehr Kelarestaghi, and Farshad Eshghi. ISSE: a new iterative sentence scoring and extraction scheme for automatic text summarization. International Journal of Computers and Applications, 44 (6):535–540, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1829844.

Haidri:2019:CBD

[HKS19] Raza Abbas Haidri, Chittaranjan Padmanabh Katti, and Prem Chandra Saxena. Capacity based deadline aware dynamic load balancing (CP-DALB) model in cloud computing environment. International Journal of Computers and Applications, 43 (10):987–1001, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1640932.

Hiranvanichchakorn:2010:URA

[HL10] P. Hiranvanichchakorn and S. Lertvorratham. Using regression analysis for improving multipath ad hoc network performance. International Journal of Computers and Applications, 32(2): 206–214, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2010.11441976.

Hamid:2011:CBP

[HL11] Yaser M. A. Hamid and Daya K. Lobiyal. COMPOWbased PCM protocol for multi-hop manet. International Journal of Computers and Applications, 33(2): 160–166, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2011.
                                        2.202-2590.

Huang:2024:SFI

[HL24] Chun Huang and Shudan Liu. Securing the future of industrial operations: a blockchainenhanced trust mechanism for digital twins in the industrial Internet of Things. International Journal of Computers and Applications, 46 (5):338–347, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2318821.

Hui:2004:ERI

[HLW04] S. C. Hui, M. K. H. Leung, and F. Wang. Eleview: Remote intelligent elevator monitoring system. International Journal of Computers and Applications, 26(2):1–8, 2004. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2004.11441727.

Han:2018:IFF

[HLW^{+}18] Xuming Han, Qiaoming Liu, Limin Wang, Huimin Lu, Liyuan Zhou, and Junhua Wang. An improved fruit fly optimization algorithm based on knowledge memory. International Journal of Computers and Applications, 42(6):558–568, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1479349.

Huang:2007:MSH

[HM07] J.-H. Huang and S. Mishra.

Mykil: a secure and highly available key management system for large group multicast. International Journal of Computers and Applications, 29 (3):300–308, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441860.

Holla:2020:GBR

[HMP20] Raviraja Holla, Nikhil C.

Mhala, and Alwyn R. Pais. GPGPU-based randomized visual secret sharing (GRVSS) for grayscale and colour images. International Journal of Computers and Applications, 44(6):552–560, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1830246.

Harkati:2024:LPI

[HNB24] Oussama Harkati, Lemnouar Noui, and Assia Beloucif. A lossless probabilistic image encryption algorithm based on QSD decomposition and matrix transformations. International Journal of Computers and Applications, 46 (10):868–879, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2392168.

Hatibaruah:2019:ETD

[HNH19] Rakcinpha Hatibaruah, Vijay Kumar Nath, and Deepika Hazarika. An effective texture descriptor for retrieval of biomedical and face images based on co-occurrence of similar center-symmetric local binary edges. International Journal of Computers and Applications, 43 (6):589–600, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1590953.

Hou:2018:CSD

[Hou18] Yan Hou. Chaos sequence data classification mining research under network visualization analysis. International Journal of Computers and Applications, 42 (2):150–156, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1475332.

Hong:2004:SAA

[HP04] B. Hong and V. K. Prasanna. Simulation of adaptive applications in heterogeneous computing environments. International Journal of Computers and Applications, 26 (3):1–8, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2004.11441737.

Hsiao:2018:ARD

[HPC18] Ju-Yuan Hsiao, Ai-Chieh Pan, and Po-Yueh Chen. An adaptive reversible data hiding scheme using dual stegoimage. International Journal of Computers and Applications, 43(3):282–291, 2018. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1080/1206212X.2018.
1557782.

Hung:2004:SSC

[HR04] J. T. Hung and T. G. Robertazzi. Scalable scheduling for clusters and grids using cut through switching. International Journal of Computers and Applications, 26 (3):1–10, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2004.11441736.

Huang:2017:EAP

[HRMA17] Xu Huang, Raul Fernandez Rojas, Allan C. Madoc, and Dua’a Ahmad. Evidentiary assessment for protecting WSNs from internal attacks in realtime. International Journal of Computers and Applications, 39(1):1–8, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2016.1249718.

Holla:2024:ARG

[HS24] M. Raviraja Holla and D. Suma. Adaptive random-grid and progressive color secret sharing with CNN super-resolution on a many-core system. International Journal of Computers and Applications, 46 (10):830–839, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2389477.

Hsieh:2010:VEE

[Hsi10] K. L. Hsieh. Value expansion expert system (CVE-ES) in Taiwanese leisure farming. International Journal of Computers and Applications, 32 (3):275–281, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2010.
                                        3.202-2578.

Haldorai:2020:AIS

[HSR20] Anandakumar Haldorai, Muhammad Sharif, and Sri Devi Ravana. Artificial intelligence for sustainable internet research. International Journal of Computers and Applications, 44 (6):501–502, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1737776.

Haritha:2019:MDL

[HT19] H. Haritha and Senthil Kumar Thangavel. A modified deep learning architecture for vehicle detection in traffic monitoring system. International Journal of Computers and Applications, 43 (9):968–977, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1662171.

Hu:2017:LVT

[Hu17] Mingyuan Hu. Logistics vehicle tracking method based on intelligent vision. International Journal of Computers and Applications, 41 (4):276–282, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1397346.

Huang:2018:RFC

[Hua18a] Hongben Huang. Research on feature classification method of network text data based on association rules. International Journal of Computers and Applications, 42 (2):157–163, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1475333.

Huang:2018:SPP

[Hua18b]  Junjun Huang.  Smart

plant project scheduling optimization based on hidden Markov algorithm. International Journal of Computers and Applications, 43 (4):346–351, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1550166.

Henke:2003:LIN

[HWH03] K. Henke, H. D. Wuttke, and S. Hellbach. Laboratory via Internet — new ways in education and research. International Journal of Computers and Applications, 25 (3):157–163, 2003. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2003.11441698.

Hu:2007:LSM

[HWL07] C. Hu, J. Wang, and J. Li. Language support for multiparadigm and multi-grain parallelism on SMP-cluster. International Journal of Computers and Applications, 29 (2):196–203, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.

tandfonline.com/doi/full/
10.1080/1206212X.2007.11441848.

Huang:2018:IHP

[HWL^{+}18] Wenzhun Huang, Peng Wang, Lintao Lv, Liping Wang, and Harry Haoxiang Wang. An inventive high-performance computing electronic information system for professional postgraduate training. International Journal of Computers and Applications, 42 (4):422–428, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1468590.

Ho:2009:UZO

[HWS09] K. I.-J. Ho, J. Wu, and J. Sum.

On unbounded zero–one knapsack with discrete-sized objects. International Journal of Computers and Applications, 31(2):84–89, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2009.11441928.

Huang:2000:MFT

[HY00] J. M. Huang and T. C. Yang.

Multidimensional fault tolerance in cube-connected cycles architecture. International Journal of Computers and Applications, 22(3): 140–150, 2000. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2000.11441617.

Hongli:2018:IMT

[HY18] Li Hongli and Ma Yaofeng. Internet of moving target detection method based on nonparametric background model. International Journal of Computers and Applications, 43 (2):193–198, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1537096.

Hao:2012:FST

[HYLS12] Rong Hao, Jia Yu, Jing Li, and Zhiling Song. Forward secure threshold signature scheme in the standard model. International Journal of Computers and Applications, 34(2): 98–104, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2012.
                                        2.202-3135.

Hao:2011:NTV

[HYS11] Rong Hao, Jia Yu, and Zhiling Song. A note on a threshold verifiable multisecret sharing scheme. International Journal of Computers and Applications, 33(4): 330–334, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2011.
                                        4.202-3074.

Horie:2005:DOO

[HYT^{+}05] T. Horie, K. Yamasaki, T. Tsuji, T. Hochin, and K. Higuchi. A distributed object-oriented programming system with dynamic reconfiguration capability. International Journal of Computers and Applications, 27(2): 63–71, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441759.

He:2019:AIT

[HZ19] Shaofu He and Xiaochang Zheng. Accurate identification technique of access path information for hierarchical geographic traffic network under big data construction. International Journal of Computers and Applications, 43(9):881–888, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1645417.

He:2020:RAK

[HZ20] Yongqiang He and Xuyang Zhang. Retracted article: Key technologies of massive concurrent data processing in smart city based on cloud computing. International Journal of Computers and Applications, 45 (2):204–211, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1080/1206212X.2020.1743040.

Han:2019:FAB

[HZW^{+}19] Xuming Han, Li Zheng, Limin Wang, Hong Zheng, and Xin Wang. Fireworks algorithm based on dynamic search and tournament selection. International Journal of Computers and Applications, 43(6):577–588, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1590034.

Ismail:2010:EAE

[IAD10] I. A. Ismail, M. Amin, and H. Diab. An efficient adaptive ergodic matrix and chaotic system for image encryption. International Journal of Computers and Applications, 32 (3):381–388, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2010.
                                        3.202-2330.

Ibrahim:2020:ECV

[IAT20] Dyala R. Ibrahim, Rosni Abdullah, and Je Sen Teh. An enhanced color visual cryptography scheme based on the binary dragonfly algorithm. International Journal of Computers and Applications, 44 (7):623–632, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1859244.

Iyyappan:2023:NMD

[IB23] K. Sakkaravarthy Iyyappan and S. R. Balasundaram. A novel multi document summarization with documentelements augmentation for learning materials using concept based ILP and clustering methods. International Journal of Computers and Applications, 46(2):78–89, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2284446.

Ismail:2010:NTD

[IEZA10] I. A. Ismail, S. F. El-Zoghdy, and A. A. Abdo. A novel technique for datahiding. International Journal of Computers and Applications, 32 (1):32–37, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2010.11441958.

Ikbal:2023:BSK

[IG23] Febina Ikbal and Rajamma Gopikakumari. Block SMRT and knapsack optimizationbased sequency selector for robust, imperceptible, and payload-efficient color image watermarking for binary watermark. International Journal of Computers and Applications, 45 (3):269–283, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2175435.

Iwata:2001:AMD

[IIDI01] Kazunori Iwata, Nobuhiro Ito, Xiaoyong Du, and Naohiro Ishii. Agent model for dynamically changing plan in π-calculus. International Journal of Computers and Applications, 23(3): 166–172, 2001. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2001.11441648.

Iabbassen:2019:CHE

[IM19] Dalila Iabbassen and Samira Moussaoui. Cooperative and hybrid energy efficient schemE for wireless sensor networks. International Journal of Computers and Applications, 43 (8):750–763, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1609648.

Inje:2024:DRU

[INR24] Bhushan Inje, Kapil Kumar Nagwanshi, and Radhakrishna Rambola. Document retrieval using clustering-based Aquila hash-Q optimization with query expansion based on pseudo relevance feedback. International Journal of Computers and Applications, 46(7):496–507, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2342715.

Ikeda:2011:CMC

[INyN11] Ritsuya Ikeda, Kensuke Narita, and Shin ya Nishizaki. Cooperation of model checking and network simulation for cost analyses of distributed systems. International Journal of Computers and Applications, 33(4):323–329, 2011. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

/www.tandfonline.com/doi/
full/10.2316/Journal.202.
2011.4.202-3068.

Iqbal:2022:RPB

[IQ22] Touseef Iqbal and Shaima Qureshi. Reconstruction probability-based anomaly detection using variational auto-encoders. International Journal of Computers and Applications, 45 (3):231–237, 2022. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2022.2143026.

Izadi:2025:KDS

[ISM25] Mohammad Izadi, Mehran Safayani, and Abdolreza Mirzaei. Knowledge distillation on spatial-temporal graph convolutional network for traffic prediction. International Journal of Computers and Applications, 47 (1):45–56, 2025. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2442699.

Ishihara:2008:CSA

[ITK08] M. Ishihara, M. Tanaka, and K. Kuriyama. A cubic systolic array and its properties. International Journal of Computers and Applications, 30(3): 173–182, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2008.11441896.

Jyothi:2019:SCB

[JC19] K. Krishna Jyothi and Shilpa Chaudhari. A secure clusterbased authentication and key management protocol for machine-type communication in the LTE network. International Journal of Computers and Applications, 44 (12):1150–1160, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1693000.

Jahan:2018:RUA

[JCI18] Sharmin Jahan, Mozammel Chowdhury, and Rafiqul Islam. Robust user authentication model for securing electronic healthcare system using fingerprint biometrics. International Journal of Computers and Applications, 41 (3):233–242, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1437651.

Jiang:2010:INB

[JCW10] L. Jiang, Z. Cai, and D. Wang. Improving naive Bayes for classification. International Journal of Computers and Applications, 32(3): 328–332, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2010.
3.202-2747.

Jin:2018:DAE

[JCYZ18] Chunhua Jin, Guanhua Chen, Changhui Yu, and Jianyang Zhao. Deniable authenticated encryption for e-mail applications. International Journal of Computers and Applications, 42(5):429–438, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1478193.

Jain:2025:ITC

[JGM25] Tarun Jain, Dinesh Gopalani, and Yogesh Kumar Meena. Informative task classification with concatenated embeddings using deep learning on crisisMMD. International Journal of Computers and Applications, 47 (2):123–140, 2025. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2447066.

Jin:2000:ELA

[JH00] Hai Jin and Kai Hwang. Efficient LRU algorithm for cache scheduling in a disk array system. International Journal of Computers and Applications, 22(3): 134–139, 2000. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2000.11441616.

Jiji:2015:AFL

[Jij15] G. Wiselin Jiji. Analysis of functionality of left ventricle. International Journal of Computers and Applications, 37(3– 4):168–180, 2015. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2016.1188552.

Jin:2010:ADW

[Jin10]  C. Jin.  Adaptive digi-

tal watermark system using soft computation. International Journal of Computers and Applications, 32(3): 341–346, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/

10.2316/Journal.202.2010.
3.202-2846.

Joshi:2018:RTC

[JJS18] Sandeep Joshi, Monika Jain, and Rahul Saxena. Recent trends in communication and computational intelligence. International Journal of Computers and Applications, 41(1):1, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1520783.

Jha:2010:IAB

[JJY^{+}10] S. K. Jha, P. K. Jana, R. Yadav, B. Sinha, and S. Srivastava. Improved algorithms for balanced ring formation for fault tolerance in a 2D mesh. International Journal of Computers and Applications, 32 (2):232–237, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2010.11441980.

Javed:2002:VCB

[JKRS02] O. Javed, S. Khan, Z. Rasheed, and M. Shah. Visual contentbased segmentation of talk and game shows. International Journal of Computers and Applications, 24(2): 77–82, 2002. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2002.11441663.

Jallouli:2020:WCI

[JLA20] Maryam Jallouli, Sonia Lajmi, and Ikram Amous. When contextual information meets recommender systems: extended SVD++ models. International Journal of Computers and Applications, 44 (4):349–356, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2020.1752971.

Jiyani:2018:NNA

[JMMS18] Ankita Jiyani, Mehul Mahrishi, Yogesh Meena, and Girdhari Singh. NAM: a nearest acquaintance modeling approach for VM allocation using R-Tree. International Journal of Computers and Applications, 43(3):218–225, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1514726.

Jayaprabha:2017:ESP

[JN17] D. Jayaprabha and K. Nirmala. Efficiency stress prediction in BPO industries using hybrid k-means and artificial bee colony algorithm. International Journal of Computers and Applications, 42(1):9–16, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/1206212X.2017.1396416.

Jafari:2018:OEC

[JNS18] Hasan Jafari, Mousa Nazari, and Shahaboddin Shamshirband. Optimization of energy consumption in wireless sensor networks using density-based clustering algorithm. International Journal of Computers and Applications, 43(1):1–10, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1497117.

Jutori:2013:DSS

[JOYB13] Hiroyoshi Jutori, Kanemitsu Ootsu, Takashi Yokota, and Takanobu Baba. Dynamic selection of speculative paths in two-path limited speculation method. International Journal of Computers and Applications, 35(2):86–95, 2013. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

/www.tandfonline.com/doi/
full/10.2316/Journal.202.
2013.2.202-3835.

Jeon:2019:RBU

[JPJ^{+}19] Myung-Joong Jeon, Hyun-Kyu Park, Batselem Jagvaral, Hyung-Sik Yoon, Yun-Geun Kim, and Young-Tack Park. Relationship between UAVs and ambient objects with threat situational awareness through grid map-based ontology reasoning. International Journal of Computers and Applications, 44 (2):101–116, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1698694.

Janakiraman:2012:DSE

[JR12] Thenkarai N. Janakiraman and John Janet L. Rani. Double star embedded clustering algorithm for wireless Ad Hoc networks using ranking. International Journal of Computers and Applications, 34(2):71–80, 2012. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2012.
2.202-2614.

Jena:2008:AMM

[JS08] R. K. Jena and G. K. Sharma. Application mapping of mesh based-NoC using multi-objective genetic algorithm. International Journal of Computers and Applications, 30(1): 17–22, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441878.

Jiji:2014:RDU

[JS14] Ganasigamony W. Jiji and Peter Raj J. D. Savariraj. Retrieval in dermatology using intelligent techniques. International Journal of Computers and Applications, 36(3): 115–124, 2014. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2014.
3.202-3952.

Jana:2016:IPS

[JS16] Nanda Dulal Jana and Jaya Sil. Interleaving of particle swarm optimization and differential evolution algorithm for global optimization. International Journal of Computers and Applications, 38 (2–3):116–133, 2016. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2016.1218242.

Jain:2019:TPH

[JS19] Divya Jain and Vijendra Singh. A two-phase hybrid approach using feature selection and adaptive SVM for chronic disease classification. International Journal of Computers and Applications, 43(6):524–536, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1577534.

Jay:2010:SSI

[JSHB10] G. T. Jay, R. K. Smith, M. Hudnall, and B. Bonds. Semi-structured information:

An architecture improving search results using domain guidance1. International Journal of Computers and Applications, 32(1): 47–55, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2010.11441960.

Jindal:2019:EWN

[JSM19] Honey Jindal, Neetu Sardana, and Raghav Mehta. Efficient web navigation prediction using hybrid models based on multiple evidence combinations. International Journal of Computers and Applications, 42(7):715–728, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1680011.

Joshi:2021:DBA

[JSYT21] Bhavya Joshi, Akhilesh Kumar Sharma, Narendra Singh Yadav, and Shamik Tiwari. DNN based approach to classify Covid’19 using convolutional neural network and transfer learning. International Journal of Computers and Applications, 44 (10):907–919, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2021.1983289.

Jou:2009:AWT

[JSZ09] M. Jou, J.-K. Shiau, and H.-W. Zhang. Application of Web technologies in automation technology education. International Journal of Computers and Applications, 31(4):215–221, 2009. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2009.11441944.

Juhasz:2006:MAM

[Juh06] S Juhasz. Modelling asynchronous message passing in small cluster environments. International Journal of Computers and Applications, 28 (1):43–49, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441787.

Jeyakumar:2012:DED

[JV12] Gurusamy Jeyakumar and Chinthamani Nathan S. Velayutham. Differential evolution and dynamic differential evolution variants — an empirical comparative performance analysis. International Journal of Computers and Applications, 34(2): 135–144, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2012.
2.202-3412.

Jiang:2009:SAN

[JWC^{+}09] L. Jiang, D. Wang, Z. Cai, S. Jiang, and X. Yan. Scaling up the accuracy of Knearest-neighbour classifiers: a naive-Bayes hybrid. International Journal of Computers and Applications, 31 (1):36–43, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2009.11441922.

Jiang:2025:UAE

[JXG25] Jianfeng Jiang, Shen Xu, and Ying Gao. A uniformed adaptive early termination model through probabilistic feature to speed up quantization-based search. International Journal of Computers and Applications, 47 (1):17–28, 2025. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2431229.

Jin:2015:NCD

[JXLZ15] Chunhua Jin, Chunxiang Xu, Fagen Li, and Xiaojun Zhang. A novel certificateless deniable authentication protocol. International Journal of Computers and Applications, 37(3– 4):181–192, 2015. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2016.1188564.

Jin:2001:FAI

[JYHS01] Kai Jin, Chai Kiat Yeo, Siu Cheung Hui, and Ing Yann Soon. A FAX adapter for Internet FAX-to-FAX communication. International Journal of Computers and Applications, 23(1): 60–67, 2001. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2001.11441635.

Jalal:2019:WSD

[JYN19] Sunita Jalal, Dharmendra Kumar Yadav, and Chetan Singh Negi. Web service discovery with incorporation of web services clustering. International Journal of Computers and Applications, 45 (1):51–62, 2019. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1698131.

Jiao:2008:MPR

[JYYJ08] Y. Jiao, Y. Yang, M. Yang, and Y. Jiang. A multipath routing scheme for torus-based NOCs. International Journal of Computers and Applications, 30(1): 9–16, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441874.

Khasawneh:2011:UCT

[KAKF11] Natheer Khasawneh, Maisa M. Al-Khudair, and Mohammad Fraiwan. On using classification techniques for corpus reduction in Arabic textto-speech systems. International Journal of Computers and Applications, 33(4): 347–354, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2011.
4.202-3188.

Kao:2010:PDM

[Kao10] Chi-Chou Kao. Performancedriven methods for fine granularity scalable video coding. International Journal of Computers and Applications, 32 (4):412–419, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2010.
4.202-2394.

Kao:2017:IFE

[Kao17] Chi-Chou Kao. Improved feature extraction and classification methods for electroencephalographic signal based brain–computer interfaces. International Journal of Computers and Applications, 39 (4):189–197, 2017. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1313490.

Karweit:2003:WBL

[Kar03] M. Karweit. A Web-based laboratory for distance-learning engineering students. International Journal of Computers and Applications, 25(3): 164–169, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2003.11441699.

Karakehayov:2012:ZZE

[Kar12] Zdravko Karakehayov. Zeta:

Zero energy transmissions algebra for wireless sensor networks. International Journal of Computers and Applications, 34(1):19–24, 2012. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

/www.tandfonline.com/doi/
full/10.2316/Journal.202.
2012.1.202-3040.

Karnalim:2019:SCP

[Kar19] Oscar Karnalim. Source code plagiarism detection with lowlevel structural representation and information retrieval. International Journal of Computers and Applications, 43 (6):566–576, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1589944.

Kattoush:2007:DRN

[Kat07] A. H. Kattoush. An Ndimensional recurrent noisy phase unwrapping algorithm.

International Journal of Computers and Applications, 29 (4):402–407, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441871.

Khemakhem:2009:PGA

[KB09] M. Khemakhem and A. Belghith. A P2P grid architecture for distributed Arabic OCR based on the DTW algorithm. International Journal of Computers and Applications, 31 (1):44–49, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2009.11441923.

Kumar:2010:ANN

[KB10] J. Kumar and A. Bansal. Application of neural network for estimating properties of diesel–biodiesel blends. International Journal of Computers and Applications, 32 (1):99–103, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2010.11441965.

Kashyap:2016:DDR

[KB16] Manish Kashyap and Mahua Bhattacharya. Detectors and descriptors for registration of illumination varying, globally distorted images. International Journal of Computers and Applications, 38 (2–3):75–81, 2016. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2016.1218238.

Kchaou:2022:TIQ

[KB22] Amina Kchaou and Sonda Ammar Bouhamed. Two image quality assessment methods based on evidential modeling and uncertainty: application to automatic iris identification systems. International Journal of Computers and Applications, 45 (3):254–268, 2022. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2022.2162671.

Kanfoud:2023:TMH

[KB23] Mohamed Raouf Kanfoud and Abdelkrim Bouramoul. Tackling the multilingual and heterogeneous documents with the pre-trained language identifiers. International Journal of Computers and Applications, 45 (5):391–402, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2218236.

Kahil:2021:GGA

[KBD21] Moustafa Sadek Kahil, Abdelkrim Bouramoul, and Makhlouf Derdour. Greedy-BigVis — a greedy approach for preparing large datasets to multidimensional visualization. International Journal of Computers and Applications, 44(8):760–769, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2021.1920670.

Kaci:2021:NVN

[KBL21] Mohamed Younes Kaci, Malika Bessedik, and Amina Lammari. A novel variable neighborhood search for the offloading and resource allocation in mobile-edge computing. International Journal of Computers and Applications, 44(9):897–905, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2021.1977895.

Kuo:2010:DDD

[KC10] J. L. Kuo and K. L. Chao. Data-dependency decomposition of colour space transformation for microprocessor by using parallel Sa-C programming. International Journal of Computers and Applications, 32(3):333–340, 2010. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

                                        /www.tandfonline.com/doi/
                                        full/10.2316/Journal.202.
                                        2010.3.202-2829.

Kalyani:2019:EAE

[KC19] G. Kalyani and Shilpa Chaudhari. An efficient approach for enhancing security in Internet of Things using the optimum authentication key. International Journal of Computers and Applications, 42 (3):306–314, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1619277.

Kaur:2020:PCD

[KC20] Sandeep Kaur and Kuljit Kaur Chahal. Prediction of Chikungunya disease using PSObased adaptive neuro-fuzzy inference system model. International Journal of Computers and Applications, 44 (7):641–649, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1870196.

Krishna:2019:MOD

[KD19] Muddada Murali Krishna and G. Lavanya Devi. Method of optimizing the dimensional features in sentiment analysis. International Journal of Computers and Applications, 43(7):643–652, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1609190.

Karam:2006:TFS

[KGEL06] H. Hussein Karam, F. F. M.

Ghaleb, and Y. M. Abd El-Latif. Towards fast and smooth subdivision surface reconstruction. International Journal of Computers and Applications, 28(2): 170–176, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2006.11441801.

Kota:2001:MSN

[KGGJ01] S. Kota, M. Goyal, R. Goyal, and R. Jain. Multimedia satellite networks and TCP/ IP traffic transport. International Journal of Computers and Applications, 23(2): 115–128, 2001. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2001.11441640.

Kumar:2019:BCM

[KH19] Krishna Kumar and Md. Tanwir Uddin Haider. Blended computation of machine learning with the recurrent neural network for intra-day stock market movement prediction using a multi-level classifier. International Journal of Computers and Applications, 43(8):733–749, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1593614.

Kaefer:2004:ASP

[KHVW04] G. Kaefer, J. Haid, K. Voit, and R. Weiss. Architectural software power estimation support for power aware remote processing. International Journal of Computers and Applications, 26 (2):1–6, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2004.11441733.

Kim:2001:HFW

[Kim01]  J. Kim.  Hybrid filtering

of Web pages for a recommendation agent. International Journal of Computers and Applications, 23(2): 99–105, 2001. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2001.11441638.

Klever:2003:PXS

[KJ03] N. Klever and M. Jeckle. Proposal for XML Schema V 1.1

as a basis for XMLrepository.Org.

International Journal of Computers and Applications, 25 (1):11–16, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2003.11441680.

Kumar:2019:PMC

[KJ19] M. R. Pavan Kumar and Prabhu Jayagopal. A preprocessing method combined with an ensemble framework for the multiclass imbalanced data classification. International Journal of Computers and Applications, 44 (12):1178–1185, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1700335.

Kesav:2022:MCC

[KJ22] Nivea Kesav and M. G. Jibukumar. Multi-channel CNN based image classification using SKIP connection and MSVM. International Journal of Computers and Applications, 44 (10):981–990, 2022. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2022.2047443.

Keshavdas:2013:FMH

[KK13] Shanker Keshavdas and Geert-Jan M. Kruijff. Functional mapping for human–robot collaborative exploration. International Journal of Computers and Applications, 35(3): 125–135, 2013. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2013.
                                        3.202-3897.

Kaur:2018:AEP

[KK18] Hargeet Kaur and Atul Kumar. Analyzing the efficiency of partially entangled states in vaidman’s-type games and its applications in Quantum Secret Sharing. International Journal of Computers and Applications, 41(1):2–13, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1465666.

Kinger:2023:DBB

[KK23a] Shakti Kinger and Vrushali Kulkarni. Demystifying the black box: an overview of explainability methods in machine learning. International Journal of Computers and Applications, 46 (2):90–100, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2285533.

Kumar:2023:HMM

[KK23b] Vishal Kumar and Sumit Kushwaha. Hybrid metaheuristic model based performanceaware optimization for map reduce scheduling. International Journal of Computers and Applications, 45 (12):776–788, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2277553.

Kinger:2024:REA

[KK24] Shakti Kinger and Vrushali Kulkarni. A review of explainable AI in medical imaging: implications and applications. International Journal of Computers and Applications, 46(11):983–997, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2404082.

Kumari:2020:EEB

[KKA20] Adesh Kumari, Vinod Kumar, and M. Yahya Abbasi. EAAF: ECC-based anonymous authentication framework for cloud-medical system. International Journal of Computers and Applications, 44(5):491–500, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2020.1815334.

Kaur:2022:PPM

[KKK22] Navjeet Kaur, Ashok Kumar, and Rajesh Kumar. PROMO: PROactive MObility-support model for task scheduling in fog computing. International Journal of Computers and Applications, 44 (11):1092–1101, 2022. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2022.2117763.

Kesavan:2019:IAH

[KKKV19] Selvaraj Kesavan, E. Saravana Kumar, Abhishek Kumar, and K. Vengatesan. An investigation on adaptive HTTP media streaming Quality-of-Experience (QoE) and agility using cloud media services. International Journal of Computers and Applications, 43 (5):431–444, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1575034.

Krishnan:2017:SBO

[KKM17] Kalyani Krishnan, Reshmy Krishnan, and Ayyakannu Muthumari. A semantic-based ontology mapping — information retrieval for mobile learning resources. International Journal of Computers and Applications, 39(3): 169–178, 2017. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1309223.

Kaushik:2025:MIM

[KKR25] Neha Kaushik, Harish Kumar, and Vinay Raj. Maintainability improvement of microservices based applications using micro frontend. International Journal of Computers and Applications, 47 (2):188–196, 2025. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2025.2450250.

Kaur:2018:EEA

[KKS18] Inderdeep Kaur, Parminder Kaur, and Hardeep Singh. An empirical evaluation of an advanced version control tool in conflict detection and resolution. International Journal of Computers and Applications, 41(1):78–87, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1512223.

Kumar:2022:ACS

[KKSR22] Naveen Kumar, Hare Krishna, Shashi Shubham, and Prabhu Padarbind Rout. Automated categorization of student’s query. International Journal of Computers and Applications, 44 (11):1056–1064, 2022. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2022.2105899.

Kabir:2019:ERS

[KKXB19] Monika Kabir, Mir Md. Jahangir Kabir, Shuxiang Xu, and Bodrunnessa Badhon. An empirical research on sentiment analysis using machine learning approaches. International Journal of Computers and Applications, 43 (10):1011–1019, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1643584.

Kuljis:2003:SOL

[KL03] J. Kuljis and D. Y. Lees.

Supporting organizational elearning with a distributed, virtual, collaborative learning environment. International Journal of Computers and Applications, 25(1): 42–49, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2003.11441684.

Kannoth:2006:SCM

[KLB06] V. Kannoth, B. Suk Lee, and J. Buzas. Statistical cost-modelling of financial time series functions. International Journal of Computers and Applications, 28 (3):181–188, 2006. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441802.

Khan:2013:BAS

[KLRH13] Nida S. Khan, Asma S. Larik, Quratulain Rajput, and Sajjad Haider. A Bayesian approach for suspicious financial activity reporting. International Journal of Computers and Applications, 35(4): 181–187, 2013. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2013.
4.202-3864.

Kao:2015:IED

[KLT15] Chi-Chou Kao, Yen-Tai Lai, and Chao-Feng Tseng. Improved edge-directed super resolution. International Journal of Computers and Applications, 37(3–4):160– 167, 2015. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2016.1188558.

Koutsikas:2006:NSL

[KM06] C. Koutsikas and N. Malevris. A new script language applicable to symbolic execution systems. International Journal of Computers and Applications, 28(1):1–11, 2006. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2006.11441782.

Khilar:2009:NHC

[KM09a] P. M. Khilar and S. Mahapatra. A novel hierarchical clustering approach for diagnosing large scale wireless adhoc systems. International Journal of Computers and Applications, 31(4): 260–267, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2009.11441949.

Khilar:2009:TCF

[KM09b] P. M. Khilar and S. Mahapatra. Time-constrained fault tolerant X-by-wire systems. International Journal of Computers and Applications, 31 (4):230–238, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2009.11441946.

Kumar:2009:TFC

[KM09c] S. Kumar and R. B. Mishra.

Towards a framework for classification and recommendation of semantic Web service composition approaches. International Journal of Computers and Applications, 31(4): 274–281, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2009.11441951.

Krishnamurthy:2009:NDM

[KMCV09] S. Krishnamurthy, N. Mittal, R. Chandrasekaran, and S. Venkatesan. Neighbour discovery in multi-receiver cognitive radio networks*. International Journal of Computers and Applications, 31 (1):50–57, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2009.11441924.

Kisku:2011:RMC

[KMGS11] Dakshina R. Kisku, Hunny Mehrotra, Phalguni Gupta, and Jamuna K. Sing. Robust multi-camera view face recognition. International Journal of Computers and Applications, 33(3):211–219, 2011. ISSN 1206-212X (print), 1925-7074 (electronic). URL

                                        https://www.tandfonline.
                                        com/doi/full/10.2316/Journal.
                                        202.2011.3.202-2922.

Kumar:2013:ASL

[KMR13] Sikamony S. Kumar, Rama S. Moni, and Jayapathy Rajeesh. Automatic segmentation of liver tumour using a possibilistic alternative fuzzy C-means clustering. International Journal of Computers and Applications, 35(1):6–12, 2013. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

                                        /www.tandfonline.com/doi/
                                        full/10.2316/Journal.202.
                                        2013.1.202-3246.

Kennington:2007:DAS

[KNS07] J. L. Kennington, V. S. S. Nair, and G. Spiride. A decomposition approach for spare capacity assignment for path restorable mesh networks. International Journal of Computers and Applications, 29(2):170–179, 2007. ISSN 1206-212X (print), 1925-7074 (electronic). URL

                                        https://www.tandfonline.
                                        com/doi/full/10.1080/1206212X.
                                        2007.11441845.

Kobayashi:2000:BSS

[KO00] M. Kobayashi and M. Ohta. Basic study of sound localization for an acoustic visual aid system. International Journal of Computers and Applications, 22(3):

129–133, 2000. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2000.11441615.

Kommers:2003:ELT

[Kom03] P. Kommers. Experiential learning through constructivist learning tools. International Journal of Computers and Applications, 25(1): 72–83, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2003.11441687.

Kumar:2020:CIM

[KPMR20] Devender Kumar, Sai Kishore Pachigolla, Shubham Singh Manhas, and Karan Rawat. Cryptanalysis and improvement of mutual authentication protocol for real-time data access in industrial wireless sensor networks. International Journal of Computers and Applications, 44 (6):521–534, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1825159.

Keramopoulos:2002:GGQ

[KPP02] E. Keramopoulos, P. Pouyioutas, and T. Ptohos. The goql graphical query language. International Journal of Computers and Applications, 24 (3):122–128, 2002. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2002.11441671.

Kallehbasti:2023:EWP

[KPSK23] Mohammad Mehdi Pourhashem Kallehbasti, Jamshid Pirgazi, Ali Ghanbari Sorkhi, and Ali Kermani. An efficient weighted partial MaxSAT encoding for scheduling in overloaded real-time systems. International Journal of Computers and Applications, 46 (3):175–183, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2297478.

Kishore:2018:EAL

[KR18] C. B. David Joel Kishore and T. Bhaskara Reddy. An efficient approach for land record classification and information retrieval in data warehouse. International Journal of Computers and Applications, 43(1):80–89, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1514290.

Kumar:2019:CBS

[KR19a] Ashish Kumar and N. S. Raghava. Chaos-based steganography technique to secure information and integrity preservation of smart grid readings using wavelet. International Journal of Computers and Applications, 44 (1):57–63, 2019. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1692511.

Kumar:2019:RIB

[KR19b] B. P. Santosh Kumar and K. Venkata Ramanaiah. Region of interest-based adaptive segmentation for image compression using hybrid Jaya– Lion mathematical approach. International Journal of Computers and Applications, 43 (10):1035–1046, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1651987.

Khaldi:2019:FTS

[KRMS19] Miloud Khaldi, Mohammed Rebbah, Boudjelal Meftah, and Omar Smail. Fault tolerance for a scientific workflow system in a cloud computing environment. International Journal of Computers and Applications, 42 (7):705–714, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1647651.

Kuzniarz:2007:TTU

[KS07] L. Kuzniarz and M. Staron.

Two techniques for UML model transformations. International Journal of Computers and Applications, 29 (1):10–17, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441827.

Kadiyala:2014:CGA

[KS14] Sai P. Kadiyala and Debasis Samanta. Clock-gating approach to low power Domino circuit synthesis. International Journal of Computers and Applications, 36(4): 140–147, 2014. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2014.
4.202-3938.

Kumar:2017:DLB

[KS17] Mohit Kumar and S. C. Sharma. Dynamic load balancing algorithm to minimize the makespan time and utilize the resources effectively in cloud environment. International Journal of Computers and Applications, 42 (1):108–117, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1404823.

Kanouni:2021:NPM

[KS21] Lakhdar Kanouni and Fouzi Semchedine. A new paradigm for multi-path routing protocol for data delivery in wireless sensor networks. International Journal of Computers and Applications, 44 (10):939–952, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2021.2014044.

Kaur:2019:ADT

[KSB19] Amanpreet Kaur, Jagroop Singh Sidhu, and Jaskarn Singh Bhullar. Adaptive deblocking technique based on separate modes for removing compression effects in JPEG coded images. International Journal of Computers and Applications, 43(6):501–513, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1567045.

Karumanchi:2023:EIB

[KSD23] Mani Deep Karumanchi, J. I.

Sheeba, and S. Pradeep Devaneyan. An efficient integrity based multi-user blockchain framework for heterogeneous supply chain management applications. International Journal of Computers and Applications, 45 (4):337–351, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2199966.

Kumar:2019:FSI

[KSS19] Naveen Kumar, Ashutosh Kumar Singh, and Shashank Srivastava. Feature selection for interest flooding attack in named data networking. International Journal of Computers and Applications, 43(6):537–546, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1583820.

Kurkovsky:2007:HFD

[Kur07] A. Kurkovsky. Hierarchy of formal descriptions for architectural analysis and design of complex CS. International Journal of Computers and Applications, 29(4): 346–352, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2007.11441865.

Krishnamoorthy:2017:EAR

[KV17] A. Krishnamoorthy and V. Vijayarajan. Energy aware routing technique based on Markov model in wireless sensor network. International Journal of Computers and Applications, 42 (1):23–29, 2017. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2017.1396423.

Kao:2011:LEC

[KYL11] Chi-Chou Kao, Chia-Nan Yeh, and Yen-Tai Lai. Low-energy cluster head selection for clustering communication protocols in wireless sensor network. International Journal of Computers and Applications, 33 (1):9–14, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2011.
1.202-2801.

Kanamori:2006:SSQ

[KYPS06] Y. Kanamori, S.-M. Yoo, W. D. Pan, and F. T. Sheldon. A short survey on quantum computers. International Journal of Computers and Applications, 28(3): 227–233, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441807.

Klostermann:2007:CBO

[KZ07] T. Klostermann and E. O. K.

Zhan. Control based optimization of collaborative technical customer service. International Journal of Computers and Applications, 29(1): 96–102, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441837.

Lam:2013:SSR

[Lam13] Yuet M. Lam. Solution space reduction in task scheduling for heterogeneous computing systems. International Journal of Computers and Applications, 35(1):29–35, 2013. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

/www.tandfonline.com/doi/
full/10.2316/Journal.202.
2013.1.202-3441.

Laplante:2005:COA

[Lap05] P. A. Laplante. Criteria and an objective approach to selecting commercial real-time operating systems based on published information. International Journal of Computers and Applications, 27 (2):82–96, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441761.

Lien:2004:EEP

[LB04a] C.-H. Lien and Y.-W. Bai. Energy efficiency and performance of a controlled model for scalable computing systems. International Journal of Computers and Applications, 26(4):1–9, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2004.11441744.

Liu:2004:CTD

[LB04b] D. X. Liu and Y. C. Bai. Cross-tree: a data structure for analysis of security protocols. International Journal of Computers and Applications, 26(1):1–7, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2004.11441724.

Lu:2017:AES

[LBYB17] Yi Lu, Cuicui Bi, Nijia Ye, and Hu Bo. Autoestablishing simulation parallel manipulators with linear legs and auto-solving their workspaces by utilizing CAD variation geometry. International Journal of Computers and Applications, 39(4): 220–233, 2017. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1309221.

Lee:2007:LCB

[LC07] C.-Y. Lee and C. W. Chiou.

Low-complexity bit-parallel multipliers for a class of GF(2^{m}) based on modified Booth’s algorithm. International Journal of Computers and Applications, 29(4): 337–345, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441864.

Li:2020:RAT

[LC20] Pengcheng Li and HuiQin Cheng. Retracted article: Track recognition algorithm based on neural network for rail transit. International Journal of Computers and Applications, 45 (2):144–150, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1080/1206212X.2020.1730070.

Lam:2006:PLD

[LCHR06] K. Y. Lam, E. Chan, D. Hung, and K. Ramamritham. Processing of location-dependent continuous queries on realtime spatial data: the view from retina. International Journal of Computers and Applications, 28(2): 135–143, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2006.11441797.

Liu:2017:SSP

[LCZA17] ShuRen Liu, ChangNing Cai, QiWei Zhu, and N. Arunkumar. A study of software pools for seismogeology-related software based on the Docker technique. International Journal of Computers and Applications, 42(1):45–51, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2017.1396429.

Loffler:2002:TDD

[LCZD02] M. S. Loffler, N. P. Costescu, E. Zergeroglu, and D. M. Dawson. Telerobotic decontamination and decommissioning with Qrobot, a PC-based robot control system. International Journal of Computers and Applications, 24(3): 112–121, 2002. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2002.11441670.

Louchene:2013:WMR

[LD13] Ahmed Louchene and Ammar Dahmani. Watermarking method resilient to RST and compression based on DWT, LPM and phase correlation. International Journal of Computers and Applications, 35 (1):36–43, 2013. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2013.
1.202-3503.

Logeswaran:2001:PSS

[LE01] R. Logeswaran and C. Eswaran.

Performance survey of several lossless compression algorithms for telemetry applications. International Journal of Computers and Applications, 23(1):1–9, 2001. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2001.11441623.

Logeswaran:2002:RBN

[LE02] R. Logeswaran and C. Eswaran.

Radial basis neural network for lossless data compression. International Journal of Computers and Applications, 24 (1):14–19, 2002. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2002.11441655.

Lee:2001:DUI

[Lee01]  Ming-Chi Lee.  Dynamic

updates of inheritance hierarchy for object-oriented database schema. International Journal of Computers and Applications, 23(1): 17–24, 2001. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2001.11441629.

Lee:2007:PII

[Lee07]  W.-P. Lee.  Personaliz-

ing Internet information services: Passive filtering and active retrieval. International Journal of Computers and Applications, 29(2): 124–131, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2007.11441840.

Liu:2008:PBC

[LGD^{+}08a] W. Liu, L. Gao, Y. Ding, J. Xu, and Y. Ji. PCbased CAN-bus experimental platform for autonomous underwater vehicles. International Journal of Computers and Applications, 30(3): 213–219, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.

tandfonline.com/doi/full/
10.1080/1206212X.2008.11441901.

Liu:2008:UHT

[LGD08b] W. J. Liu, J. Gong, and W. Ding. Using heavytailed feature to estimate flow length distributions. International Journal of Computers and Applications, 30(3): 201–206, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441899.

Li:2020:IFC

[LGL20] Wei Li, William W. Guo, and Michael Li. The impact factors on the competence of big data processing. International Journal of Computers and Applications, 44(3):201–218, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1719623.

Liang:2018:CCE

[LGSS18] Hong Liang, Yuan Gao, Yunlei Sun, and Xiao Sun. CEP: calories estimation from food photos. International Journal of Computers and Applications, 42(6):569–577, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1486558.

Lin:2000:GRT

[LH00] J. C. Lin and I. Ho. Generating real-time software test cases by time Petri nets. International Journal of Computers and Applications, 22(3): 151–158, 2000. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2000.11441618.

Lertvorratham:2012:ISM

[LH12] Supachote Lertvorratham and Pipat Hiranvanichakorn. Integrating secure multipath mobile ad hoc network with selfauthentication strategy. International Journal of Computers and Applications, 34(3): 174–184, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2012.
                                        3.202-3245.

Li:2013:MSA

[LH13] Da Li and Yibin Hou. Multiprocessor systems auto-design for multiple use-case applications on FPGA. International Journal of Computers and Applications, 35(3): 108–113, 2013. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2013.
                                        3.202-3529.

Linsen:2007:WAR

[LHJ07] L. Linsen, B. Hamann, and K. I. Joy. Wavelets for adaptively refined ^{3} \sqrt{2}- subdivision meshes. International Journal of Computers and Applications, 29(3): 223–231, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441851.

Lee:2003:AMM

[LHK03] J.-S. Lee, W.-K. Hong, and S.-D. Kim. Adaptive methods to minimize decompression overhead for compressed on-chip caches. International Journal of Computers and Applications, 25(2): 98–105, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2003.11441690.

Liu:2018:CAM

[LHZ18] Zhongliang Liu, Lili Huang, and Xiaoliang Zong. Clustering analysis method and implementation in the MapReduce model for solving dataintensive problems. International Journal of Computers and Applications, 42 (6):533–543, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1477321.

Li:2017:ESC

[Li17] Hongtao Li. Estimation of stadium construction schedule based on big data analysis. International Journal of Computers and Applications, 41(4):268–275, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2017.1397345.

Li:2019:MCF

[Li19] Ya-Wen Li. Multiple collaborative filtering recommendation algorithms for electronic commerce information. International Journal of Computers and Applications, 43 (9):903–909, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1649813.

Li:2020:ISA

[Li20a]  Jingfu Li.  IOT security

analysis of BDT-SVM multiclassification algorithm. International Journal of Computers and Applications, 45 (2):170–179, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2020.1734313.

Li:2020:RAR

[Li20b] Juan Li. Retracted article: Real-time image recognition algorithm and system design of Android mobile devices. International Journal of Computers and Applications, 45(2):161–169, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1080/1206212X.2020.1733791.

Licea:2006:SRF

[Lic06] G. Licea. Supporting reusability in fixed and mobile groupware applications. International Journal of Computers and Applications, 28(2): 99–111, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441793.

Lin:2011:ECR

[Lin11]  Jyhjong Lin.  Enhancing

customer relationships with customer knowledge management and P2P (peer-to-peer) technology. International Journal of Computers and Applications, 33(4): 303–315, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2011.
4.202-3036.

Lin:2017:QAS

[Lin17] Wang Lin. Quantitative analysis of spatial–temporal variation and relation of LST and relevant ecological elements. International Journal of Computers and Applications, 42(1):57–66, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2017.1397384.

Liu:2019:NRO

[LJL^{+}19] Yong Liu, Jiabao Jiang, Yun Liu, Yong Zhang, and Qilin Wu. Network resource optimization configuration in edge computing environment. International Journal of Computers and Applications, 45(1):88–95, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1706811.

Li:2003:HAP

[LL03] C. Li and L. Li. A hybrid agent platform composed of cooperating intelligent agents and mobile agents. International Journal of Computers and Applications, 25 (4):217–224, 2003. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2003.11441707.

Li:2007:CRP

[LL07] Y. Li and Z. Lan. Current research and practice in proactive fault management. International Journal of Computers and Applications, 29(4): 408–413, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2007.11441872.

Li:2011:CWH

[LL11a] Chaoqun Li and Hongwei Li. Correlation weighted heterogeneous Euclidean-overlap metric. International Journal of Computers and Applications, 33(4):341–346, 2011. ISSN 1206-212X (print), 1925-7074 (electronic). URL

https://www.tandfonline.
com/doi/full/10.2316/Journal.
202.2011.4.202-3179.

Li:2011:LRM

[LL11b] Chaoqun Li and Hongwei Li. Learning random model trees for regression. International Journal of Computers and Applications, 33(3): 258–265, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2011.
3.202-3162.

Lam:2014:MCB

[LL14] Yuet M. Lam and Wayne Luk. A many-core based parallel tabu search. International Journal of Computers and Applications, 36(1): 15–22, 2014. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2014.
1.202-3808.

Li:2018:ITP

[LL18] Xu Li and Lan Ling. Internet of target perception event driven based on machine vision using a fast extraction algorithm. International Journal of Computers and Applications, 43 (2):186–192, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1537095.

Li:2017:NMI

[LLB17] Guodong Li, Qian Liu, and Shuo Bai. A novel model to improve network performance. International Journal of Computers and Applications, 40(2):82–87, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1395107.

Lee:2011:CBF

[LLC11] Wei-Po Lee, Li-Jen Liu, and Jeng-An Chiou. A component-based framework for rapidly developing online board games. International Journal of Computers and Applications, 33(4): 293–302, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2011.
4.202-2813.

Li:2008:RCT

[LLH08] M. Li, C.-S. Liao, and J.- K. Hwang. Rapid calculation of time deviation and modified Allan deviation for characterizing telecommunications clock stability. International Journal of Computers and Applications, 30 (2):92–96, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441886.

Liu:2015:DRS

[LLL^{+}15] Zhi Liu, Sanya Liu, Lin Liu, Meng Wang, Jianwen Sun, and Xian Peng. A discriminative random sampling strategy with individual-author feature selection for writeprint recognition of Chinese texts. International Journal of Computers and Applications, 37(3– 4):94–101, 2015. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2016.1160643.

Liang:2005:FGO

[LLR05] J. M. Liang, J. Q. Liang, and Q. L. Ren. A framework for generic object recognition with Bayesian networks. International Journal of Computers and Applications, 27 (3):123–138, 2005. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441764.

Lian:2007:MDE

[LLRW07] S. Lian, Z. Liu, Z. Ren, and Z. Wang. Multimedia data encryption in blockbased codecs. International Journal of Computers and Applications, 29(1): 18–24, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2007.11441828.

Lotfy:2002:ANN

[LM02] E. A. Lotfy and A. S. Mohamed. Applying neural networks in case-based reasoning adaptation for cost assessment of steel buildings. International Journal of Computers and Applications, 24 (1):28–38, 2002. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2002.11441657.

Loo:2004:SAD

[Loo04] A. Loo. Sensitivity analysis of database operations: a case study of parallel partition/sorting algorithm. International Journal of Computers and Applications, 26 (4):1–8, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2004.11441743.

Lou:2019:DTE

[Lou19]  Yuesheng Lou.  Detection

of trace elements in metallurgical products under complex background based on visual saliency. International Journal of Computers and Applications, 43 (7):706–714, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1634328.

Liu:2011:MCS

[LQLX11] Li-Feng Liu, Yan-Yun Qu, Cui-Hua Li, and Yuan Xie. Multi-class spectral clustering based on particle swarm optimization. International Journal of Computers and Applications, 33(1):64–69, 2011. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

/www.tandfonline.com/doi/
full/10.2316/Journal.202.
2011.1.202-3006.

Lin:2000:ICU

[LS00] Shinfeng D. Lin and Shih-Chieh Shie. Image coding using adaptive classified side-match finite-state vector quantization. International Journal of Computers and Applications, 22(3): 174–180, 2000. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2000.11441621.

Lowenthal:2003:HAI

[LS03] D. K. Lowenthal and R. Subramanian. Hyfi: Architectureindependent parallelism on networks of multiprocessors. International Journal of Computers and Applications, 25 (4):272–282, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2003.11441715.

Leela:2010:DMC

[LS10] R. Leela and S. Selvakumar. Dynamic multi constraint multi path QOS routing algorithm (DMCMPRA). International Journal of Computers and Applications, 32 (3):347–354, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2010.
                                        3.202-2886.

Lin:2003:DBI

[LSC03] S. D. Lin, S. C. Shie, and C. F. Chen. A DCT-based image watermarking with threshold embedding. International Journal of Computers and Applications, 25(2): 130–135, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2003.11441694.

Liu:2023:ESI

[LSDJ23] J. Liu, H. J. Shao, X. Deng, and Y. T. Jiang. Exploiting similarity-induced redundancies in correlation topology for channel pruning in deep convolutional neural networks. International Journal of Computers and Applications, 45 (5):379–390, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/1206212X.2023.2218061.

Legakis:2011:CBN

[LSMI11] Ioannis Legakis, Michalis A.

Savelonas, Dimitris Maroulis, and Dimitris K. Iakovidis. Computer-based nodule malignancy risk assessment in thyroid ultrasound images. International Journal of Computers and Applications, 33 (1):29–35, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2011.
1.202-2955.

Lohar:2025:FPD

[LTS25] Rohit Lohar, Meenakshi Tripathi, and Hossein Shirazi. Free-Phish: detecting phishing websites hosted on free Web hosting domains. International Journal of Computers and Applications, 47 (2):162–168, 2025. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2448835.

Liu:2024:RFI

[LTY^{+}24] HanFan Liu, Pei Tian, HaiFang Yao, Cong Shen, and FengYang Hu. Research on fast-DTW and isolation forestbased lightweight gesture authentication. International Journal of Computers and Applications, 46 (7):455–466, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2335717.

Lin:2006:DOA

[LV06] W. J. Lin and B. Veeravalli. A dynamic object allocation and replication algorithm for distributed systems with centralized control. International Journal of Computers and Applications, 28(1): 26–34, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441785.

Li:2003:DIP

[LVK03] X L. Li, B. Veeravalli, and C. C. Ko. Distributed image processing on a network of workstations. International Journal of Computers and Applications, 25(2): 136–145, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2003.11441695.

Li:2017:TIG

[LWG17] Feng Li, Yawen Wang, and Yunzhan Gong. Test input generation of bitwise operation based on four-valued logic. International Journal of Computers and Applications, 40(3):1–8, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1413624.

Li:2012:IWB

[LWH12] Tianyi Li, Minghui Wang, and Zujian Huang. An improved wavelet-based approach for estimating the variance of noise in images. International Journal of Computers and Applications, 34(4): 229–234, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2012.
4.202-3195.

Li:2007:PBS

[LXH07] F. Li, X. Xin, and Y. Hu.

A pairing-based signcryption scheme using self-certified public keys. International Journal of Computers and Applications, 29(3):278–282, 2007. ISSN 1206-212X (print), 1925-7074 (electronic). URL

https://www.tandfonline.
com/doi/full/10.1080/1206212X.
2007.11441857.

Li:2008:ECB

[LXH08] F. Li, X. Xin, and Y. Hu. Efficient certificate-based signcryption scheme from bilinear pairings. International Journal of Computers and Applications, 30(2): 129–133, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441891.

Liu:2012:BFI

[LXM^{+}12] Qing Liu, Lu-Ping Xu, Yi-De Ma, Yong Wang, and Qiang Xie. Bilateral filtering for image processing based on pulse coupled neural networks. International Journal of Computers and Applications, 34(4):217–222, 2012. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

                                        /www.tandfonline.com/doi/
                                        full/10.2316/Journal.202.
                                        2012.4.202-3120.

Li:2008:IPT

[LXZ08] F. Li, C. Xu, and S. Zhou. Improvement of a proactive threshold signcryption scheme. International Journal of Computers and Applications, 30(4):345–347, 2008. ISSN 1206-212X (print), 1925-7074 (electronic). URL

                                        https://www.tandfonline.
                                        com/doi/full/10.1080/1206212X.
                                        2008.11441915.

Liu:2009:SLU

[LYR09] C. Liu, E. Ye, and D. J. Richardson. Software library usage pattern extraction using a software model checker. International Journal of Computers and Applications, 31 (4):247–259, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2009.11441948.

Li:2004:EFM

[LYY04] B. Li, W. Yao, and J. You.

An effort to formal model of trust management in grids. International Journal of Computers and Applications, 26(3):1–6, 2004. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2004.11441735.

Lu:2008:NDS

[LYY^{+}08] E. Lu, M. Yang, B. Yang, X. Feng, and S. Q. Zheng. A novel design of self-routing strictly nonblocking switching networks. International Journal of Computers and Applications, 30(1): 44–50, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441876.

Liu:2007:DNV

[LZ07] G. H. Liu and S. Zhang.

Development of a new virtual reality system. International Journal of Computers and Applications, 29 (4):321–326, 2007. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441862.

Liu:2018:PFD

[LZ18] Siyang Liu and Cheng Zhang.

Personalized file data query matching method based on SOA. International Journal of Computers and Applications, 43(2):159–164, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1534374.

Li:2017:SBD

[LZL^{+}17] Limiao Li, Zhihong Zhang, Zhixiong Liu, Hui Ye, and Keke He. Social big data oriented trust situation risk model based on individual behavior. International Journal of Computers and Applications, 42 (2):119–126, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1309209.

Li:2007:PSV

[LZW07] Y. G. Li, W. D. Zhang, and G. L. Wang. Prune support vector machines by an iterative process. International Journal of Computers and Applications, 29(2): 164–169, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441844.

Li:2024:APF

[LZXB24] Linjuan Li, Haoxue Zhang, Gang Xie, and Yanhong Bai. Adjustable patch and feature prior token-based transformer for weakly supervised semantic segmentation. International Journal of Computers and Applications, 46 (7):445–454, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2333122.

Magoni:2005:NTA

[Mag05] D. Magoni. Network topology analysis and Internet modelling with NEM. International Journal of Computers and Applications, 27(4): 252–259, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441778.

Moskowitz:2010:ITE

[MAL10] I. S. Moskowitz, F. Ahmed, and P. A. Lafferty. Information theoretic effects of JPEG compression on image steganography. International Journal of Computers and Applications, 32(3): 318–327, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2010.
3.202-2736.

Mallick:2024:ECS

[Mal24a] Bhawna Mallick. Enhancing cyber security: a comprehensive approach to the classification and prediction of significant cyber incidents (SCI) through data mining and variational neural network with fox optimization algorithm. International Journal of Computers and Applications, 46 (8):621–627, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2380652.

Malouche:2024:WAC

[Mal24b] Hela Malouche. Weight assignment in cloud service selection based on FAHP and rough sets. International Journal of Computers and Applications, 46(9):754–768, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2387658.

Mansour:2015:UAM

[Man15] Romany F. Mansour. Using adaptive mutation to accelerate the convergence of immune algorithms for prediction of 3D molecular structure. International Journal of Computers and Applications, 37(3– 4):127–133, 2015. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2016.1188568.

Mani:2017:PDR

[Man17] A. Mani. Probabilities, dependence and rough membership functions. International Journal of Computers and Applications, 39(1): 17–35, 2017. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2016.1259800.

Mittal:2009:MKP

[MB09] N. Mittal and T. R. Belagodu.

On maximum key pool size for a key pre-distribution scheme in wireless sensor networks. International Journal of Computers and Applications, 31 (1):30–35, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2009.11441921.

Merniz:2010:MVP

[MB10] S. Merniz and M. Benmohammed. Modelling and verification of pipelined micro-architectures: Functional approach. International Journal of Computers and Applications, 32(1): 84–92, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2010.11441963.

Mekhaznia:2019:NAA

[MB19] Tahar Mekhaznia and Akram Bennour. New approach for attack of permutation-based image encryption schemes. International Journal of Computers and Applications, 43 (7):697–705, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1632003.

Mecili:2024:EED

[MBN^{+}24] Oualid Mecili, Hadj Barkat, Farid Nouioua, Abdelouahab Attia, and Samir Akhrouf. An efficient explainable deep neural network classifier for diabetic retinopathy detection. International Journal of Computers and Applications, 46(9):795–810, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2389342.

Madhukar:2023:MSC

[MBP23] B. N. Madhukar, S. H. Bharathi, and Ashwin M. Polnaya. Multi-scale convolution based breast cancer image segmentation with attention mechanism in conjunction with war search optimization. International Journal of Computers and Applications, 45(5):353–366, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2212945.

Maierhofer:2001:ASA

[MBS01] M. Maierhofer, C. Bailey, and R. Sotudeh. “Affordable scalability”: An analysis of modular multimedia server architectures with shared local memory. International Journal of Computers and Applications, 23(2):79–89, 2001. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2001.11441636.

Mudigonda:2024:CME

[MBS^{+}24] Krishna Siva Prasad Mudigonda, Karthik Bulusu, Yashmitha Sri, Adithi Damera, and Vignesh Kode. Capturing multiple emotions from conversational data using fine tuned transformers. International Journal of Computers and Applications, 46 (12):1166–1178, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2420189.

Mittal:2007:DES

[MC07] A. Mittal and L.-F. Cheong.

Detecting establishment shots using camera motion. International Journal of Computers and Applications, 29(3): 232–238, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441852.

Mawloud:2014:MLB

[MD14] Guermoui Mawloud and Melaab Djame. Modified local binary pattern for human face recognition based on sparse representation. International Journal of Computers and Applications, 36(2):64–71, 2014. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

                                        /www.tandfonline.com/doi/
                                        full/10.2316/Journal.202.
                                        2014.2.202-3908.

Martins:2024:SLR

[MdCdAdS24] Samuel Luna Martins, Felipe Mendes da Cruz, Rogério Pontes de Araújo, and Carlo Marcelo Revoredo da Silva. Systematic literature review on security misconfigurations in web applications. International Journal of Computers and Applications, 46(10):840–852, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2390977.

Ma:2014:IOM

[MDY14] Kun Ma, Fusen Dong, and Bo Yang. Incremental object matching approach of schemafree data with Mapreduce. International Journal of Computers and Applications, 36 (2):72–77, 2014. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2014.
                                        2.202-3912.

Meng:2018:CNA

[Men18]  Xiangfei Meng.  Complex

network anomaly recognition integrating multi-source and large data. International Journal of Computers and Applications, 42 (2):174–178, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1475335.

Meruga:2015:MLC

[MFK^{+}15] Jeevan M. Meruga, Carly Fountain, Jon Kellar, Grant Crawford, Aravind Baride, P. Stanley May, William Cross, and Randy Hoover. Multi-layered covert QR codes for increased capacity and security. International Journal of Computers and Applications, 37(1):17–27, 2015. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

/www.tandfonline.com/doi/
full/10.1080/1206212X.2015.
1061254.

Maratha:2019:CSR

[MG19] Priti Maratha and Kapil Gupta. A comprehensive and systematized review of energy-efficient routing protocols in wireless sensor networks. International Journal of Computers and Applications, 44(1):83–100, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1697513.

Mohamed:2000:DDM

[MGK^{+}00] A. S. Mohamed, A. Galal, I. Khalil, K. Sobh, and M. Selim. Dispo: Distributed multi-threaded execution of Prolog programs. International Journal of Computers and Applications, 22(2): 100–108, 2000. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2000.11441606.

Menassel:2019:IBI

[MGT19] Rafik Menassel, Idriss Gaba, and Khalil Titi. Introducing BAT inspired algorithm to improve fractal image compression. International Journal of Computers and Applications, 42(7):697–704, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1638631.

Mishra:2018:OSM

[MJ18] Sanju Mishra and Sarika Jain. Ontologies as a semantic model in IoT. International Journal of Computers and Applications, 42 (3):233–243, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1504461.

Mehla:2019:DEK

[MJ19a] Sonia Mehla and Sarika Jain. Development and evaluation of knowledge treasure for emergency situation awareness. International Journal of Computers and Applications, 43(5):483–493, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1574950.

Muthalagu:2019:NMK

[MJ19b] Raja Muthalagu and Subeen Jain. A novel modified KA-SUMI block cipher for global system for mobile communications. International Journal of Computers and Applications, 43(8):805–811, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1619988.

Mulla:2021:OSV

[MJ21] Arkan Mulla and H. T. Jadhav. Optimal scheduling of vehicle-to-grid power exchange using particle swarm optimization technique. International Journal of Computers and Applications, 44 (7):687–704, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2021.1903707.

Manoharan:2011:PAM

[MK11a] Arun Manoharan and Ammasi Krishnan. Power analysis of multiple hashing Bloom filter architecture for network applications. International Journal of Computers and Applications, 33(4):

316–322, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2011.
                                        4.202-3052.

Meetoo:2011:SNS

[MK11b] Anuja Meetoo and Kavi K. Khedo. Saisense: a novel scalable, adaptive and intelligent context-aware architecture. International Journal of Computers and Applications, 33(3):189–201, 2011. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

                                        /www.tandfonline.com/doi/
                                        full/10.2316/Journal.202.
                                        2011.3.202-2772.

Mohsin:2018:ERM

[MKA^{+}18] Syed Muhammad Mohsin, Imran Ali Khan, Syed Muhammad Abrar Akber, Shahaboddin Shamshirband, and Anthony T. Chronopoulos. Exploring the RFID mutual authentication domain. International Journal of Computers and Applications, 43 (2):127–141, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1533614.

Muknahallipatna:2002:MTP

[MKH02] S. Muknahallipatna, A. Kadkol, and J. Hamann. Monitoring tool for performance measurements of Windows Nt-based lans. International Journal of Computers and Applications, 24(3): 136–143, 2002. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2002.11441673.

Mohammad:2007:PAJ

[MKMA07] S. B. Mohammad, M. O.

Khaoua, L. M. Mackenzie, and I. Ababneh. Processor allocation and job scheduling on 3D mesh interconnection networks. International Journal of Computers and Applications, 29(3): 309–317, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441861.

Mohapatra:2018:PLP

[MKN^{+}18] Ambarish G. Mohapatra, Bright Keswani, Shivani Nanda, Abhishek Ray, Ashish Khanna, Deepak Gupta, and Poonam Keswani. Precision local positioning mechanism in underground mining using IoT-enabled WiFi platform. International Journal of Computers and Applications, 42(3):266–277, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1551178.

Mudigonda:2024:ETS

[MLM^{+}24] Krishna Siva Prasad Mudigonda, Nikitha Lingineni, Y. Manisai, Manoj Pennada, Mallesh Gadde, and Revanth Sai Aluri. Extractive text summarization on medical insights using fine-tuned transformers. International Journal of Computers and Applications, 46 (11):957–973, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2401081.

Marwala:2009:IHC

[MLT09] T. Marwala, M. Lagazio, and T. Tettey. An integrated human–computer system for controlling interstate disputes. International Journal of Computers and Applications, 31 (4):239–246, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2009.11441947.

Maharana:2000:PAS

[MM00] G. S. Maharana and P. K. Meher. Parallel algorithms and systolic architectures for 1- and 2-D interpolation using discrete Hartley transform. International Journal of Computers and Applications, 22(1):1–7, 2000. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2000.11441592.

Manna:2024:RSB

[MM24] Supriya Manna and Krishna Siva Prasad Mudigonda. Revitalizing the single batch environment: a ‘quest’ to achieve fairness and efficiency. International Journal of Computers and Applications, 46 (8):651–665, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2380660.

Mickle:2006:CMP

[MMCM06] M. H. Mickle, M. Mi, J. T.

Cain, and T. Minor. A circuit model for passive RF autonomous devices with protocol considerations. International Journal of Computers and Applications, 28(3): 243–250, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441809.

Malloy:2001:EII

[MMH01] B. A. Malloy, J. D. McGregor, and S. R. Hughes. Exploiting iostreams to integrate a GUI into a command driven application. International Journal of Computers and Applications, 23(3): 152–158, 2001. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2001.11441646.

Moussa:2018:GLB

[MMH18] Mohammed Elsaid Moussa, Ensaf Hussein Mohamed, and Mohamed Hassan Haggag. A generic lexicon-based framework for sentiment analysis. International Journal of Computers and Applications, 42(5):463–473, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1483813.

Moussa:2019:OMH

[MMH19] Mohammed Elsaid Moussa, Ensaf Hussein Mohamed, and Mohamed Hassan Haggag. Opinion mining: a hybrid framework based on lexicon and machine learning approaches. International Journal of Computers and Applications, 43 (8):786–794, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1615250.

Muknahallipatna:2010:LFS

[MMHJ10] S. Muknahallipatna, J. Miles, J. C. Hamann, and H. L. Johnson. Large fabric storage area networks: Fabric simulator development and preliminary performance analysis. International Journal of Computers and Applications, 32(2): 167–180, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2010.11441973.

Maharana:2013:ESA

[MMM13] Gouri S. Maharana, Pramod K.

Meher, and Basant K. Mohanty. Efficient systolic architecture for VLSI realization of 2-D Hartley-like transform. International Journal of Computers and Applications, 35 (1):22–28, 2013. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2013.
1.202-3379.

Mallaki:2021:CMS

[MMPM21] Mahdi Mallaki, Babak Majidi, Amirhossein Peyvandi, and Ali Movaghar. Offchain management and statetracking of smart programs on blockchain for secure and efficient decentralized computation. International Journal of Computers and Applications, 44(9):822–829, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2021.1948170.

Mansouri:2018:MNN

[MMS18] Amin Mansouri, Babak Majidi, and Abdolah Shamisa. Metaheuristic neural networks for anomaly recognition in industrial sensor networks with packet latency and jitter for smart infrastructures. International Journal of Computers and Applications, 43 (3):257–266, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1533613.

Mustafi:2020:NAT

[MMS20] D. Mustafi, A. Mustafi, and G. Sahoo. A novel approach to text clustering using genetic algorithm based on the nearest neighbour heuristic. International Journal of Computers and Applications, 44(3):291–303, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1735035.

Mudiyanselage:2017:STM

[MP17] Akalanka Karunarathne Mudiyanselage and Lei Pan. Security test MOODLE: a penetration testing case study. International Journal of Computers and Applications, 42 (4):372–382, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1396413.

Mehfooza:2018:SDS

[MP18] M. Mehfooza and V. Pattabiraman. SP-DDPT: a simple prescriptive-based domain data preprocessing technique to support multilabelmulticriteria learning with expert information. International Journal of Computers and Applications, 43 (4):333–339, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1547475.

Mihajlovic:2015:ISC

[MPĆ15] Željka Mihajlović, Siniša Popović, and Krešimir Ćosić. Interactive scenario control in virtual environments. International Journal of Computers and Applications, 37 (2):53–59, 2015. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2015.1079955.

Mamalis:2013:HSP

[MPDK13] Basilis Mamalis, Grammati Pantziou, Georgios Dimitropoulos, and Dimitris Kremmydas. Highly scalable parallelization of standard simplex method on a Myrinetconnected cluster platform. International Journal of Computers and Applications, 35 (4):152–161, 2013. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2013.
4.202-3691.

Malar:2019:IMF

[MPS^{+}19] A. Christy Jeba Malar, M. Deva Priya, J. Sengathir, N. Kiruthiga, R. Anitha, and T. Sangeetha. An intelligent multi-floor indoor positioning system for cloudbased environment. International Journal of Computers and Applications, 44 (12):1170–1177, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1696447.

Majhi:2023:MIC

[MPS^{+}23] Santosh Kumar Majhi, Abhipsa Panda, Suresh Kumar Srichandan, Usha Desai, and Biswaranjan Acharya. Malware image classification: comparative analysis of a fine-tuned CNN and pre-trained models. International Journal of Computers and Applications, 45 (11):709–721, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2270804.

Muniyal:2010:EMM

[MR10] Balachandra Muniyal and Prema K. Venkat Reddy. An efficient method to merge hierarchical public key infrastructures. International Journal of Computers and Applications, 32(4):442–446, 2010. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

/www.tandfonline.com/doi/
full/10.2316/Journal.202.
2010.4.202-2855.

Mathana:2012:RAH

[MR12] Joseph Michael M. Mathana and Parthasarathy Rangarajan. Reconfigurable architecture for high performance turbo decoder. International Journal of Computers and Applications, 34(3): 166–173, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2012.
3.202-3222.

Maheswari:2019:KSC

[MR19a] K. Maheswari and M. Ramakrishnan. Kernelized spectral clustering based conditional MapReduce function with big data. International Journal of Computers and Applications, 43 (7):601–611, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1587892.

Mohan:2019:EDF

[MR19b] G. S. S. S. S. V. Krishna Mohan and Yarravarapu Srinivasa Rao. An efficient design of fractional order differentiator using hybrid shuffled frog leaping algorithm for handling noisy electrocardiograms. International Journal of Computers and Applications, 43 (5):494–500, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1573948.

Malipatil:2023:DIF

[MR23] Sridevi Malipatil and T. Hanumantha Reddy. Discovery of interesting frequent item sets in an uncertain database using ant colony optimization. International Journal of Computers and Applications, 45(11):673–679, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2263689.

Montejano:2004:SCM

[MRU04] G. Montejano, D. Riesco, and R. Uzal. Software configuration management in a parallel development environment. International Journal of Computers and Applications, 26 (2):1–10, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2004.11441732.

Marshall:2001:ICS

[MS01] A. Marshall and S. Sezer. The influence of cumulative switch delay in multiple service class networks. International Journal of Computers and Applications, 23(2): 106–114, 2001. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2001.11441639.

Manouselis:2003:ABL

[MS03] N. Manouselis and D. Sampson. Agent-based e-learning course recommendation: Matching learner characteristics with content attributes. International Journal of Computers and Applications, 25(1): 50–64, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2003.11441685.

Mohammed:2016:ECA

[MS16] Ahmed Abdulateef Mohammed and Atul Sajjanhar. Experimental comparison of approaches for feature extraction of facial attributes. International Journal of Computers and Applications, 38 (4):187–198, 2016. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2016.1207427.

Mukherjee:2017:EPP

[MS17] Srilekha Mukherjee and Goutam Sanyal. Enhanced position power first mapping (PPFM) based image steganography. International Journal of Computers and Applications, 39 (2):59–68, 2017. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2016.1273624.

Maheswari:2019:SBA

[MS19] S. Maheswari and Justus Selwyn. State behavior analysis in validating the web service framework. International Journal of Computers and Applications, 43 (7):632–642, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1602331.

Meybodi:2024:DBG

[MSD24] Mohsen Alambardar Meybodi, Mahdi Safari, and Ensieh Davoodijam. Domination based graph neural networks. International Journal of Computers and Applications, 46 (11):998–1005, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2404087.

Mallick:2018:EBF

[MSG18] Partho Mallick, Priyanka Seth, and Anupam Ghosh. Entropy-based fuzzy hybrid framework for gene prediction network — an application to identify and rank the biomarkers for human lung adenocarcinoma. International Journal of Computers and Applications, 41(1):62–77, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1508865.

Madadi:2021:MSD

[MSH21] Yeganeh Madadi, Vahid Seydi, and Reshad Hosseini. Multisource domain adaptationbased low-rank representation and correlation alignment. International Journal of Computers and Applications, 44(7):670–677, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2021.1885786.

Mohanty:2024:TEE

[MSK24] Rasmita Kumari Mohanty, Satya Prakash Sahoo, and Manas Ranjan Kabat. Thermalenergy-efficient-secured-link reliable and delay-aware routing protocol (TESLDAR) for wireless body area networks. International Journal of Computers and Applications, 46 (9):715–727, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2380675.

Mishra:2018:SIA

[MSYG18] Sanju Mishra, Rafid Sagban, Ali Yakoob, and Niketa Gandhi. Swarm intelligence in anomaly detection systems: an overview. International Journal of Computers and Applications, 43 (2):109–118, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1521895.

Mangwani:2024:ETC

[MT24] Poonam Mangwani and Vrinda Tokekar. An efficient test case selection framework for multitenant SaaS applications using RTS4MTS. International Journal of Computers and Applications, 46 (12):1143–1151, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2417374.

Monterde:2009:CSB

[MU09] J. Monterde and H. Ugail. A comparative study between biharmonic Bézier surfaces and biharmonic extremal surfaces. International Journal of Computers and Applications, 31 (2):90–96, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2009.11441929.

Mekala:2018:SEE

[MV18] Mahammad Shareef Mekala and P. Viswanathan. A survey: energy-efficient sensor and VM selection approaches in green computing for x-IoT applications. International Journal of Computers and Applications, 42 (3):290–305, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1558511.

Mansoori:2017:RWI

[MWC^{+}17] Masood Mansoori, Ian Welch, Kim-Kwang Raymond Choo, Roy A. Maxion, and Seyed Ebrahim Hashemi. Real-world IP and network tracking measurement study of malicious websites with HAZOP. International Journal of Computers and Applications, 39(2): 106–121, 2017. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1283910.

Ma:2016:ADT

[MYA16] Kun Ma, Bo Yang, and Ajith Abraham. Asynchronous data translation framework for converting relational tables to document stores. International Journal of Computers and Applications, 38(1): 19–28, 2016. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2016.1188563.

Ma:2018:POT

[MZZC18] Jian Ma, Runtong Zhang, Xiaomin Zhu, and Runqi Cao.

Process ontology technology in modeling clinical pathway information system. International Journal of Computers and Applications, 42 (6):550–557, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1478195.

Nishad:2019:SAC

[NA19] A. Nishad and Sajimon Abraham. SemTraClus: an algorithm for clustering and prioritizing semantic regions of spatio-temporal trajectories. International Journal of Computers and Applications, 43(8):841–850, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1655853.

Nayak:2024:SSF

[NB24] Tunirani Nayak and Nilamani Bhoi. Self-sparse fuzzy clustering with automatic region merging approach for image segmentation. International Journal of Computers and Applications, 46 (9):728–742, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2380975.

Nyamawe:2021:IRR

[NBS21] Ally S. Nyamawe, Khadidja Bakhti, and Sulis Sandiwarno. Identifying rename refactoring opportunities based on feature requests. International Journal of Computers and Applications, 44 (8):770–778, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2021.1922151.

Nemmour:2010:HDR

[NC10] H. Nemmour and Y. Chibani.

Handwritten digit recognition based on a neural– SVM combination. International Journal of Computers and Applications, 32(1): 104–109, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2010.11441966.

Nicho:2024:ECC

[NCGA24] Mathew Nicho, Brian Cusack, Shini Girija, and Nalin Arachchilage. Evaluating the cost of classifier discrimination choices for IoT sensor attack detection. International Journal of Computers and Applications, 46 (11):945–956, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2401069.

Navya:2018:TBE

[ND18] V. Navya and P. Deepalakshmi. Threshold-based energyefficient routing for transmission of critical physiological parameters in a wireless body area network under emergency scenarios. International Journal of Computers and Applications, 43 (4):367–376, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1554334.

Nahar:2024:DSE

[NK24] Mahbubun Nahar and A. H. M. Kamal. DNA steganography: Embedding the secret messages using glass stack method and detecting errors made by different attacks. International Journal of Computers and Applications, 46(7):508–525, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2347456.

Nandi:2016:EAF

[NM16] Utpal Nandi and Jyotsna Kumar Mandal. Efficiency of adaptive fractal image compression with archetype classification and its modifications. International Journal of Computers and Applications, 38(2–3):156–163, 2016. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/1206212X.2016.1237130.

Nath:2023:NRL

[NMK23] Manabendra Nath, Pinaki Mitra, and Deepak Kumar. A novel residual learning-based deep learning model integrated with attention mechanism and SVM for identifying tea plant diseases. International Journal of Computers and Applications, 45 (6):471–484, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2235750.

Nandakumar:2017:ABI

[NN17] R. Nandakumar and K. Nirmala. Anonymity-based intra– inter and multiple layer service dependent security-aware packet scheduling algorithm (AIIMLSDSPS). International Journal of Computers and Applications, 42 (4):383–391, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1396421.

Nath:2021:GAB

[NN21] Rintu Nath and A. Nagaraju.

Genetic algorithm based onarrival task scheduling on distributed computing platform. International Journal of Computers and Applications, 44(9):887–896, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2021.1974751.

Nguyen:2022:BNS

[NNDP22] Minh Duc Nguyen, Cuong H. Nguyen-Dinh, and Le Anh Phuong. BOCA: A novel semantic blockchain-based authentication system of educational certificates. International Journal of Computers and Applications, 44 (11):1074–1082, 2022. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2022.2111509.

Nomir:2011:IBH

[Nom11] Omaima Nomir. Iris biometrics: Human identification using force field. International Journal of Computers and Applications, 33 (1):41–48, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2011.
                                        1.202-2985.

Nomir:2012:HIN

[Nom12] Omaima Nomir. Human identification: a new X-ray dental radiographs segmentation algorithm using Graphcut. International Journal of Computers and Applications, 34 (2):81–89, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2012.
2.202-3044.

Nair:2023:RRD

[NPN^{+}23] Akhil Nair, Varad Patil, Rohan Nair, Adithi Shetty, and Mimi Cherian. A review on recent driver safety systems and its emerging solutions. International Journal of Computers and Applications, 46(3):137–151, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2293348.

Namneh:2007:PID

[NPY07] R. A. Namneh, W. D. Pan, and S.-M. Yoo. Parallel implementations of 1-D Fast Fourier Transform without interprocessor communication. International Journal of Computers and Applications, 29(2): 180–186, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441846.

Nargunam:2010:CBM

[NS10a] A. S. Nargunam and M. P.

Sebastian. Cluster-based MANET multicast routing scheme. International Journal of Computers and Applications, 32(1):38–46, 2010. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

                                        /www.tandfonline.com/doi/
                                        full/10.1080/1206212X.2010.
                                        11441959.

Nargunam:2010:SOQ

[NS10b] A. S. Nargunam and M. P. Sebastian. Self-organized QoS aware multicast routing scheme for ad hoc networks. International Journal of Computers and Applications, 32 (1):23–31, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2010.11441957.

Ngubiri:2010:CFM

[NvV10] J. Ngubiri and M. van Vliet. Characteristics of fairness metrics and their effect on perceived scheduler effectiveness. International Journal of Computers and Applications, 32 (2):188–196, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2010.11441974.

Nguyen:2018:IBU

[NX18] Van Luong Nguyen and Ying Xia. IoT based user trajectory retrieval and discovery algorithm based on spatial– temporal correlation analysis. International Journal of Computers and Applications, 42(3):282–289, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/1206212X.2018.1553272.

Nzanywayingoma:2017:ERM

[NY17] Frederic Nzanywayingoma and Yang Yang. Efficient resource management techniques in cloud computing environment: a review and discussion. International Journal of Computers and Applications, 41 (3):165–182, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1416558.

Nyamapfene:2010:HAM

[Nya10] A. Nyamapfene. A hypermap approach to multiple sequence processing. International Journal of Computers and Applications, 32(2): 160–166, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2010.11441972.

Noureddine:2020:NPM

[NZT^{+}20] Samia Noureddine, Baarir Zineeddine, Abida Toumi, Abir Betka, and Aïcha-Nabila Benharkat. A new predictive medical approach based on data mining and Symbiotic Organisms Search algorithm. International Journal of Computers and Applications, 44 (5):465–479, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2020.1809825.

Oquendo:2001:DCO

[OA01] L. S. Oquendo and A. Attoui. Deterministic CORBA ORB (DORB) for distributed real-time applications. International Journal of Computers and Applications, 23(3): 186–190, 2001. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2001.11441650.

OFarrell:2015:ERT

[OB15] John O’Farrell and Sanjeev Baskiyar. Enhanced realtime performance using a secondary bus for cache writebacks. International Journal of Computers and Applications, 37(1):1–9, 2015. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2015.1061157.

Oladosu:2012:YEL

[OE12] John B. Oladosu and Justice O. Emuoyibofarhe. A Yoruba–English language translator for doctor–patient mobile chat application. International Journal of Computers and Applications, 34 (3):149–156, 2012. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/

10.2316/Journal.202.2012.
3.202-3079.

Ogundoyin:2018:ALC

[Ogu18] Sunday Oyinlola Ogundoyin.

An autonomous lightweight conditional privacy-preserving authentication scheme with provable security for vehicular ad-hoc networks. International Journal of Computers and Applications, 42 (2):196–211, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1477320.

Ogundoyin:2019:LPP

[Ogu19] Sunday Oyinlola Ogundoyin.

A lightweight privacy-preserving data aggregation scheme with provable security for Internet-of-Things. International Journal of Computers and Applications, 42 (3):315–327, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1674008.

Oh:2010:JJI

[Oh10]  Tick H. Oh.  JPEG2000

and JPEG: Image quality comparison of compressed medical modalities. International Journal of Computers and Applications, 32 (4):393–398, 2010. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2010.
                                        4.202-1503.

Odoom:2023:BAS

[OHZ^{+}23] Justice Odoom, Xiaofang Huang, Zuhong Zhou, Samuel Danso, Benedicta Nana Esi Nyarko, Jinan Zheng, and Yanjie Xiang. Blockchainassisted sharing of electronic health records: a feasible privacy-centric constant-size ring signature framework. International Journal of Computers and Applications, 45 (9):564–578, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2252238.

Obite:2017:CNF

[OIB17] Felix Obite, Geoffrey Ijeomah, and Joseph Stephen Bassi. Carbon nanotube field effect transistors: toward future nanoscale electronics. International Journal of Computers and Applications, 41 (2):149–164, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2017.1415111.

Okafor:2019:DRM

[Oka19] K. C. Okafor. Dynamic reliability modeling of cyberphysical edge computing network. International Journal of Computers and Applications, 43(7):612–622, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1600830.

Oh:2000:PED

[OL00] Sang H. Oh and Won S. Lee.

Parallel evaluation of a deductive query with a breadthfirst search strategy. International Journal of Computers and Applications, 22(2): 89–99, 2000. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2000.11441605.

Oh:2004:DQP

[OL04] S. H. Oh and W. S. Lee.

Deductive query processing with an object-oriented semantic network in a massively parallel environment. International Journal of Computers and Applications, 26 (2):1–11, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2004.11441731.

Oueslati:2024:CSS

[OMA24] Wided Oueslati, Siwar Mejri, and Jalel Akaichi. A comprehensive study on social networks analysis and mining to detect opinion leaders. International Journal of Computers and Applications, 46(8):641–650, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2380655.

Okafor:2018:TCD

[OOG^{+}18] K. C. Okafor, G. C. Ononiwu, Sam Goundar, V. C. Chijindu, and C. C. Udeze. Towards complex dynamic fog network orchestration using embedded neural switch. International Journal of Computers and Applications, 43 (2):91–108, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1517440.

Otair:2008:ETN

[OS08] M. A. Otair and W. A. Salameh. Efficient training of neural networks using optical backpropagation with momentum factor. International Journal of Computers and Applications, 30 (3):167–172, 2008. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441895.

Okamoto:2002:RSC

[OSKC02] T. Okamoto, K. Seki, M. Kayama, and A. I. Cristea. Rapsody: a self/collaborativelearning multimedia based teacher training distance support model. International Journal of Computers and Applications, 24(2): 52–57, 2002. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2002.11441660.

Ou:2011:FVS

[OW11] Phichhang Ou and Hengshan Wang. Forecasting volatility switching ARCH by treed Gaussian process with jumps to the limiting linear model. International Journal of Computers and Applications, 33 (4):355–361, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2011.
4.202-3260.

Osterberg:2011:MFM

[ÖZ11] Patrik Österberg and Tingting Zhang. Multicast-favourable max-min fairness — the definition and how to comply. International Journal of Computers and Applications, 33(1):1–8, 2011. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2011.
1.202-2567.

Pathan:2018:TAI

[Pat18] Al-Sakib Khan Pathan. Technological advancements and innovations are often detrimental for concerned technology companies. International Journal of Computers and Applications, 40 (4):189–191, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1515412.

Pathan:2020:AIV

[Pat20] Al-Sakib Khan Pathan. Access to information vs blocking of information during COVID-19 pandemic: a governance dilemma in the era of crowdsourcing based on ICT. International Journal of Computers and Applications, 42(6):531–532, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2020.1767396.

Pathan:2022:CPI

[Pat22a] Al-Sakib Khan Pathan. COVID-19 pandemic, information and communications technology (ICT), and the Digital Divide. International Journal of Computers and Applications, 44(1):1–2, 2022. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2022.2020022.

Pathan:2022:SCC

[Pat22b] Al-Sakib Khan Pathan. On the scale of cyberspace and cybersecurity. International Journal of Computers and Applications, 44(9):805–806, 2022. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (electronic). URL

http://www.tandfonline.
com/doi/full/10.1080/1206212X.
2022.2125185.

Pathan:2023:WDS

[Pat23] Al-Sakib Khan Pathan. Why do some IT freelancers in certain countries prefer digital payment with cryptocurrency despite it being illegal? International Journal of Computers and Applications, 45(4):285–287, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2193778.

Pathan:2024:ECF

[Pat24] Al-Sakib Khan Pathan. Editorin-chief’s foreword. International Journal of Computers and Applications, 46 (7):443–444, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2314926.

Pentikousis:2010:CSA

[PBA10] Kostas Pentikousis, Hussein Badr, and Asha Andrade. A comparative study of aggregate TCP retransmission rates. International Journal of Computers and Applications, 32(4):435–441, 2010. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

                                        /www.tandfonline.com/doi/
                                        full/10.2316/Journal.202.
                                        2010.4.202-2660.

Pal:2016:PCO

[PBM16] Pankaj Pal, Siddhartha Bhattacharyya, and Ashish Mani. Pure color object extraction from a noisy state using quantum version parallel self organizing neural network. International Journal of Computers and Applications, 38 (2–3):164–186, 2016. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2016.1237164.

Pahari:2019:EBB

[PBS19] Purbanka Pahari, Piyali Basak, and Anasua Sarkar. Ensemble based biomarker identification on pancreatic ductal adenocarcinoma gene expressions. International Journal of Computers and Applications, 43(9):851–857, 2019. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (electronic). URL

                                        http://www.tandfonline.
                                        com/doi/full/10.1080/1206212X.
                                        2019.1636524.

Prabhu:2022:DMS

[PBS22] D. Prabhu, S. Vijay Bhanu, and S. Suthir. Design of multiple share creation with optimal signcryption based secure biometric authentication system for cloud environment. International Journal of Computers and Applications, 44 (11):1047–1055, 2022. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2022.2103890.

Parikh:2011:PAM

[PC11] Ritesh Parikh and Santanu Chattopadhyay. Power-aware multi-level AND-XOR network synthesis. International Journal of Computers and Applications, 33(1): 22–28, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2011.
1.202-2938.

Padmanabhan:2013:WBV

[PC13] Sriraman A. Padmanabhan and Soundarrajan Chandramathi. A wavelet-based video compression using adaptive rood search with spatiotemporal correlation and LSK. International Journal of Computers and Applications, 35 (1):13–21, 2013. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2013.
1.202-3349.

Pooja:2019:TPH

[PC19] Pooja and R. K. Chauhan.

Triple phase hybrid cryptography technique in a wireless sensor network. International Journal of Computers and Applications, 44 (2):148–153, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1710342.

Pradhan:2022:TAA

[PC22] Nilam Pradhan and Bharat S. Chaudhari. Traffic-aware autonomous scheduling for 6TiSCH networks. International Journal of Computers and Applications, 44 (11):1039–1046, 2022. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2022.2103889.

Paul:2025:DGL

[PCD25] Amit Paul, Narayan Changder, and Animesh Dutta. DSF-GAM: a location inference model in social network Twitter. International Journal of Computers and Applications, 47(2):141–161, 2025. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2448497.

Perumalla:2021:BCB

[PCK21] Subhadra Perumalla, Santanu Chatterjee, and A. P. Siva Kumar. Block chain-based access control protocol in Internet of Drones. International Journal of Computers and Applications, 44 (9):847–862, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2021.1961366.

Papazis:2008:NAL

[PCS08] K. Papazis, N. K. Chilamkurti, and B. Soh. A new adaptive layered multicast protocol. International Journal of Computers and Applications, 30(2): 73–79, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441884.

Pei:2018:CDM

[Pei18] Li Pei. Community discovery method based on complex network of data fusion based on the super network perspective. International Journal of Computers and Applications, 43(4):383–390, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1554836.

Pan:2010:EIS

[PF10] T. Pan and K. Fang. An effective information support system for medical management: Indicator based intelligence system. International Journal of Computers and Applications, 32(1):

119–124, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2010.11441968.

Pathan:2018:FEC

[PF18] Al-Sakib Khan Pathan and Zubair Md Fadlullah. Foreword from the Editors-in-Chief. International Journal of Computers and Applications, 40(2):61–62, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1446619.

Panigrahi:2004:AFO

[PG04] P. K. Panigrahi and A. Goswami. Algebra for fuzzy object oriented database language. International Journal of Computers and Applications, 26 (1):1–9, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2004.11441721.

Panigrahi:2019:QED

[PGK19] Susant Kumar Panigrahi, Supratim Gupta, and S. Vamsee Krishna. Quantitative evaluation of different thresholding methods using automatic reference image creation via PCA. International Journal of Computers and Applications, 43 (7):653–662, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1615175.

Patel:2019:PFS

[PJ19] Archana Patel and Sarika Jain. Present and future of semantic web technologies: a research statement. International Journal of Computers and Applications, 43 (5):413–422, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1570666.

Park:2007:WSU

[PK07] Y. Park and B. S. Kim.

Web search using dynamic keyword suggestion. International Journal of Computers and Applications, 29 (1):1–9, 2007. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441826.

Parveen:2019:ULU

[PK19] Rubina Parveen and Subhash Kulkarni. Urban land use dynamics and change implication by using low-resolution LISS III imagery. International Journal of Computers and Applications, 43 (10):1029–1034, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/1206212X.2019.1649042.

Prasath:2020:EAB

[PK20] R. Prasath and T. Kumanan. Enhanced artificial bee colony approach for the enhancement and classification of underwater images. International Journal of Computers and Applications, 44 (5):433–443, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1784591.

Pachauri:2018:RAO

[PKD18] Bhoopendra Pachauri, Ajay Kumar, and Joydip Dhar. Reliability analysis of open source software systems considering the effect of previously released version. International Journal of Computers and Applications, 41 (1):31–38, 2018. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1497575.

Papadakis:2012:ALR

[PKDP12] Nikos Papadakis, Pavlos Kefalas, Antonis Delidakis, and Kostas Papadakis. An approach of a logic resolver in a distributed peer-to-peer system. International Journal of Computers and Applications, 34(1):42–47, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2012.
1.202-3092.

Parveen:2018:AED

[PKM18] Rubina Parveen, Subhash Kulkarni, and V. D. Mytri. Automated extraction and discrimination of open land areas from IRS-1C LISS III imagery. International Journal of Computers and Applications, 42(7):676–685, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1558937.

Phongpensri:2011:PAV

[PKW^{+}11] Chantana Phongpensri (Chantrapornchai), Saran Keinprapai, Opas Wongtaveesap, Kanok Hournkumnuard, and Sergei Gorlatch. Parallel algorithm and visualization of high gradient magnetic separation of nanoparticles. International Journal of Computers and Applications, 33 (1):70–82, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2011.
1.202-3018.

Park:2004:DID

[PLK^{+}04] K.-L. Park, H.-J. Lee, O.- Y. Kwon, S.-Y. Park, H.-W. Park, and S.-D. Kim. Design and implementation of a dynamic communication MPI library for the grid. International Journal of Computers and Applications, 26 (3):1–8, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2004.11441738.

Palau:2003:XAA

[PMR^{+}03] C. E. Palau, V. Manso, J. M. Raga, R. Romero, J. C. Guerri, and M. Esteve. An XML approach for assessment in education. International Journal of Computers and Applications, 25(1): 24–37, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2003.11441682.

Patil:2020:EAN

[PN20] Vilas K. Patil and P. P. Nagrale. Effective analysis of noise levels due to vehicular traffic in urban area using deep learning with OALO model. International Journal of Computers and Applications, 44 (6):561–570, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2020.1835269.

Paramasivam:2022:EQG

[PN22] Aarthi Paramasivam and S. Jaya Nirmala. Exploring question generation in medical intelligent system using entailment. International Journal of Computers and Applications, 45(3):248–253, 2022. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2022.2161147.

Priyadarshini:2024:RIT

[PN24] Priyanka Priyadarshini and Kshiramani Naik. Robust image tamper detection and recovery with selfembedding watermarking using SPIHT and LDPC. International Journal of Computers and Applications, 46 (8):580–603, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2380650.

Premananda:2018:EEC

[PNAB18] B. S. Premananda, Nupur K.

Neti, K. P. Ashwin, and Shreya V. Basu. An energyefficient, coalition game theory based hierarchical routing protocol for WSNs. International Journal of Computers and Applications, 43 (3):246–256, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1529726.

Pons:2004:OOL

[Pon04] A. P. Pons. An object-oriented language for real-time systems. International Journal of Computers and Applications, 26(1):1–7, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2004.11441720.

Pons:2005:EWO

[Pon05]  A. P. Pons.  Enhance-

ment of Web object speculative retrieval. International Journal of Computers and Applications, 27(3): 139–146, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2005.11441766.

Popescu:2005:SKA

[Pop05]  C. Popescu.  A secure

key agreement protocol using elliptic curves. International Journal of Computers and Applications, 27(3): 147–152, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2005.11441770.

Parimala:2008:PSE

[PR08] N. Parimala and S. R. N. Reddy. Processor selection for embedded system design. International Journal of Computers and Applications, 30 (4):348–353, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2008.11441916.

Pandian:2017:AAN

[PR17] K. K. Soundra Pandian and K. C. Ray. An algorithm and architecture for nonrecursive pseudorandom sequence generation using sequence folding technique. International Journal of Computers and Applications, 39 (1):45–56, 2017. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2016.1262165.

Prasad:2018:PBD

[PR18] K. Venkata Vara Prasad and P. Trinatha Rao. Performance of blind detection frame work using energy detection approach for local sensing in intelligent networks. International Journal of Computers and Applications, 43 (4):326–332, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1546659.

Poluru:2019:IFF

[PR19] Ravi Kumar Poluru and Lokesh Kumar R. An Improved Fruit Fly Optimization (IFFOA) based cluster head selection algorithm for Internet of Things. International Journal of Computers and Applications, 43 (7):623–631, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1600831.

Prasad:2020:EFC

[PRR20] Kalli Srinivasa Nageswara Prasad, Annaluri Sreenivasa Rao, and Attili Venkata Ramana. Ensemble framework for concept-drift detection in multidimensional streaming data. International Journal of Computers and Applications, 44(12):1193–1200, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2020.1711617.

Prasad:2006:PID

[PS06] M. V. N. K. Prasad and K. K. Shukla. Parallel implementation of domain decomposition based image compression algorithm. International Journal of Computers and Applications, 28(4): 334–341, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2006.11441819.

Prem:2012:CIS

[PS12] Monickaraj Vigilson Prem and Sankaranarayanan Swamynathan. Code and itinerary security for mobile agents. International Journal of Computers and Applications, 34(4): 260–266, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.

tandfonline.com/doi/full/
10.2316/Journal.202.2012.
4.202-3601.

Pakkiraiah:2016:NMM

[PS16] B. Pakkiraiah and G. Durga Sukumar. A new modified MPPT controller for improved performance of an asynchronous motor drive under variable irradiance and variable temperature. International Journal of Computers and Applications, 38 (2–3):61–74, 2016. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2016.1188586.

Pankaja:2019:HAC

[PS19] K. Pankaja and V. Suma.

A hybrid approach combining CUR matrix decomposition and weighted kernel sparse representation for plant leaf recognition. International Journal of Computers and Applications, 43 (8):830–840, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1632031.

Patel:2019:CBR

[PSP19] Sunil Kumar Patel, Chandran Saravanan, and Vikash Kumar Patel. Cloud-based reversible dynamic secure steganography model for embedding pathological report in medical images. International Journal of Computers and Applications, 43(10):1002–1010, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1641273.

Pandey:2014:WQP

[PT14] Parul Pandey and Maheshwari Tripathi. The wheel quorum protocol: An efficient approach for reading replicated data. International Journal of Computers and Applications, 36(2):78–83, 2014. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2014.
                                        2.202-3946.

Peng:2008:NEA

[PTLH08] J. Peng, C.-J. Tang, C. Li, and J.-J. Hu. A new evolutionary algorithm based on chromosome hierarchy network. International Journal of Computers and Applications, 30(3): 183–191, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2008.11441897.

Parthasarathy:2018:TIA

[PV18] P. Parthasarathy and S. Vivekanandan. A typical IoT architecturebased regular monitoring of arthritis disease using time wrapping algorithm. International Journal of Computers and Applications, 42 (3):222–232, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1457471.

Park:2002:SCR

[PW02] Y. Park and L. Wu. Software component retrieval by composition using semantic properties. International Journal of Computers and Applications, 24(1): 8–13, 2002. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2002.11441654.

Perera:2019:HAI

[PW19] Gamage Kokila Kasuni Perera and Champi Thusangi Wannige. A hybrid algorithm for identifying partially conserved regions in multiple sequence alignment. International Journal of Computers and Applications, 43 (10):979–986, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1628468.

Pradabpet:2009:NPR

[PYMD09] C. Pradabpet, S. Yoshizawa, Y. Miyanaga, and K. Dejhan. New PAPR reduction in OFDM systems by hybrid algorithm of PTS and APPR methods. International Journal of Computers and Applications, 31(2): 119–127, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2009.11441932.

Qianqian:2018:SRI

[QD18] Zhu Qianqian and Wang Dongliang. A SVM recommendation IoT model based on similarity evaluation and collaborative filtering of multiangle knowledge units. International Journal of Computers and Applications, 42 (3):278–281, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1553271.

Qadir:2024:PCD

[QDQ24] Imran Qadir, V. Devendran, and Fasel Qadir. Performance comparison of decisionbased median filtering for brain MR images with high multiplicative noise. International Journal of Computers and Applications, 46 (9):687–701, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2380665.

Roomi:2008:NRE

[RABB08] S. M. Mansoor Roomi, V. Abhaikumar, S. Baskar, and N. S. Balaji. A no-reference edge preserving genetic weight adaptation filter for impulse noise removal. International Journal of Computers and Applications, 30(4): 338–344, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441914.

Radaideh:2004:DPM

[Rad04] M. A. Radaideh. A distributed and parallel model for high-performance indexing of database content. International Journal of Computers and Applications, 26 (4):1–6, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2004.11441747.

Rahman:2019:HCE

[Rah19] M. M. Hafizur Rahman.

HMMN: a cost-effective derivative of midimew-connected mesh network. International Journal of Computers and Applications, 43 (8):727–732, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1591690.

Roy:2006:NTL

[RCJJ06] D. Roy, A. Chatterjee, N. Jasapara, and N. Jadhav. A new transport layer protocol offering variable reliability in bidirectional communication: Application for robotic systems.

International Journal of Computers and Applications, 28 (3):251–258, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2006.11441810.

Ramos:2003:MLE

[RCNM03] F. Ramos, A. Conde, L. Neves, and A. Moreira. Management of e-learning environments: some issues and research clues. International Journal of Computers and Applications, 25 (1):38–41, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2003.11441683.

Rahaman:2007:TDD

[RDB07] H. Rahaman, D. K. Das, and B. B. Bhattacharya. Testable design of digital summation threshold logic array for synthesis of symmetric functions. International Journal of Computers and Applications, 29 (2):115–123, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2007.11441839.

Rizk:2009:FTE

[RDEDAE09] M. R. M. Rizk, M. I. Dessouky, S. A. El-Dolil, and M. Abd-Elnaby. Fairness and throughput enhancement-based random access using fuzzy controlled backoff interval. International Journal of Computers and Applications, 31 (1):58–64, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2009.11441925.

Ru-Da:2018:FAI

[RDYGE18] Jia Ru-Da, Chen Yong-Gao, and Zhang Yin E. Fusion algorithm of infrared and visible images based on frame difference detection technology and area feature. International Journal of Computers and Applications, 42 (7):655–660, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1522722.

Ren:2018:RAP

[Ren18]  Dan Ren.  Research and

analysis on precise matching method for multi-feature of fuzzy digital image. International Journal of Computers and Applications, 42 (2):141–149, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1475331.

Rambabu:2023:TSD

[RG23] D. Rambabu and A. Govardhan. Task scheduling and data replication in cloud with improved correlation strategy. International Journal of Computers and Applications, 45(11):697–708, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2267840.

Ramirez-Gutierrez:2012:IRT

[RGNMPM12] Kelsey A. Ramirez-Gutierrez, Mariko Nakano-Miyatake, and Hector M. Perez-Meana. Improvement of Radon transformbased perceptual hashing using image normalization. International Journal of Computers and Applications, 34 (4):249–259, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2012.
4.202-3530.

Regilan:2023:OEM

[RH23] S. Regilan and L. K. Hema. Optimizing environmental monitoring in IoT: integrating DBSCAN with genetic algorithms for enhanced clustering. International Journal of Computers and Applications, 46(1):21–31, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2277966.

Regilan:2024:RBC

[RHJ24] S. Regilan, L. K. Hema, and J. Jenitha. ReLeC-based clustering and multi-objective optimization for efficient energy optimization in IoT networks.

International Journal of Computers and Applications, 46 (7):526–538, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2380646.

Radaideh:2006:WWB

[RHR06] M. Radaideh, S. Horani, and M. Raseen. WBSGA: a Web-based tool for course timetabling and scheduling. International Journal of Computers and Applications, 28 (1):74–83, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441791.

Rong:2018:SNT

[RJS18] Fan Rong, Zhang Juan, and Zhao ShuoFeng. Surgical navigation technology based on computer vision and VR towards IoT. International Journal of Computers and Applications, 43 (2):142–146, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1534371.

Ravi:2015:PAE

[RK15] G. Ravi and K. R. Kashwan. Performance analysis of an energy aware zone routing protocol using span. International Journal of Computers and Applications, 37 (1):10–16, 2015. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2015.1061158.

Ranjan:2024:TOP

[RK24] Rajeev Ranjan and Prabhat Kumar. Towards optimal path planning for sustainable data collection in real-time wireless sensor networks. International Journal of Computers and Applications, 46 (5):301–309, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2311590.

Reddy:2025:ETE

[RKS^{+}25] Y. C. A. Padmanabha Reddy, Sai Sathwik Kosuru, Nageswara Rao Sirisalla, G. N. Vivekananda, and Venkata Akanksha Perala. Empirical techniques for effort estimation in designing effective ML models. International Journal of Computers and Applications, 47(2):226–234, 2025. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2025.2454509.

Rezaeipanah:2019:PNA

[RMF19] Amin Rezaeipanah, Musa Mojarad, and Ahad Fakhari. Providing a new approach to increase fault tolerance in cloud computing using fuzzy logic. International Journal of Computers and Applications, 44(2):139–147, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1709288.

Rajasekaran:2019:AMP

[RMG19] Rajkumar Rajasekaran, Jolly Masih, and K. Govinda. An analysis of mobile pass-codes in case of criminal investigations through social network data. International Journal of Computers and Applications, 43(9):954–959, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1662169.

Ramasamy:2017:DPD

[RN17] S. Ramasamy and K. Nirmala. Disease prediction in data mining using association rule mining and keyword based clustering algorithms. International Journal of Computers and Applications, 42(1):1–8, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1396415.

Reichert:2024:SSC

[RO24] Beatriz M. Reichert and Rafael R. Obelheiro. Software supply chain security: a systematic literature review. International Journal of Computers and Applications, 46(10):853–867, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2390978.

Rahman:2017:FFA

[RPAU17] Ziaur Rahman, Yi-Fei Pu, Muhammad Aamir, and Farhan Ullah. A framework for fast automatic image cropping based on deep saliency map detection and Gaussian filter. International Journal of Computers and Applications, 41(3):207–217, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1422358.

Rao:2018:RTV

[RPS18] N. Venkatesvara Rao, D. Venkatavara Prasad, and M. Sugumaran. Real-time video object detection and classification using hybrid texture feature extraction. International Journal of Computers and Applications, 43(2):119–126, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1525929.

Radhappa:2017:POS

[RPZW17] Harish Radhappa, Lei Pan, James Xi Zheng, and Sheng Wen. Practical overview of security issues in wireless sensor network applications. International Journal of Computers and Applications, 40(4):202–213, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1398214.

Rao:2017:RIF

[RR17] K. Koteswara Rao and G. S.

V. P. Raju. Reducing interactive fault proneness in software application using genetic algorithm based optimal directed random testing. International Journal of Computers and Applications, 41(4):296–305, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1417769.

Ramakrishna:2023:SAI

[RRA^{+}23] Ch. Jnana Ramakrishna, D. Bharath Kalyan Reddy, P. P. Amritha, K. V. Lakshmy, and Vasily Sachnev. A secure authenticated image encryption scheme based on elliptic curve cryptography. International Journal of Computers and Applications, 46(3):184–193, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2299575.

Rao:2017:POP

[RRRR17] S. Sivanaga Malleswara Rao, K. Venkata Rao, K. Hemachandra Reddy, and Ch. V. S. Parameswara Rao. Prediction and optimization of process parameters in wire cut electric discharge machining for high-speed steel (HSS). International Journal of Computers and Applications, 39 (3):140–147, 2017. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2017.1309219.

Radhamani:2004:WET

[RS04] G. Radhamani and M. U. Siddiqi. A WAP-enabled transaction model for mobile multidatabases. International Journal of Computers and Applications, 26 (1):1–6, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2004.11441726.

Raj:2008:HIE

[RS08] A. A. Raj and M. Suganthi. Hardware implementation of an efficient Internet protocol routing filter design. International Journal of Computers and Applications, 30(2): 124–128, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2008.11441890.

Ramachandran:2016:NAG

[RS16] V. Ramachandran and C. Sekar.

(1, N)-arithmetic graphs. International Journal of Computers and Applications, 38 (1):55–59, 2016. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2016.1218240.

Russell:2017:CCS

[RS17] Matthew Russell and Jeremy Straub. Characterization of command software for an autonomous attitude determination and control system for spacecraft. International Journal of Computers and Applications, 39(4): 198–209, 2017. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1329261.

Ramasubbareddy:2019:RRT

[RS19] Somula Ramasubbareddy and R. Sasikala. RTTSMCE: a response time aware task scheduling in multi-cloudlet environment. International Journal of Computers and Applications, 43(7):691–696, 2019. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (electronic). URL

http://www.tandfonline.
com/doi/full/10.1080/1206212X.
2019.1629098.

Rani:2024:CCS

[RSS24] Pinki Rani, Sukhdip Singh, and Karanbir Singh. Cloud computing security: a taxonomy, threat detection and mitigation techniques. International Journal of Computers and Applications, 46 (5):348–361, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2319937.

Rawat:2023:PSR

[RST23] Neelam Rawat, Vikas Somani, and Arun Kr. Tripathi. Prioritizing software regression testing using reinforcement learning and hidden Markov model. International Journal of Computers and Applications, 45 (12):748–754, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2273585.

Ruiz:2019:MFM

[Rui19] Conrado Ruiz Jr. Multifeature 3D model retrieval using view-based techniques. International Journal of Computers and Applications, 43 (8):819–829, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1625529.

Russomanno:2006:PQI

[Rus06] D. J. Russomanno. Prolog query interface: a set-of-mappings approach with optional segmentation. International Journal of Computers and Applications, 28(2): 144–153, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441798.

Rao:2019:RNN

[RV19] A. Nageswara Rao and P. Vijayapriya. A robust neural network model for monitoring online voltage stability. International Journal of Computers and Applications, 44 (12):1103–1112, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1666224.

Rana:2024:EER

[RVS24] Sandeep Singh Rana, Gaurav Verma, and O. P. Sahu. Energy-efficient resource allocation for NOMA-MEC system under imperfect successive interference cancellation. International Journal of Computers and Applications, 46 (10):911–920, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2397014.

Rateke:2020:PVR

[RvW20] Thiago Rateke and Aldo von Wangenheim. Passive vision road obstacle detection: a literature mapping. International Journal of Computers and Applications, 44 (4):376–395, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2020.1758877.

Roussev:2007:TUC

[RW07] B. Roussev and J. Wu. Transforming use case models to class models and OCL-specifications. International Journal of Computers and Applications, 29(1): 59–69, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2007.11441833.

Ranganathan:2019:EIS

[RWL19] G. Ranganathan, Hui-Ming Wee, and Pavel Lafata. Editorial on IoT in social, mobility, analytics & cloud [ISMAC]. International Journal of Computers and Applications, 41 (4):243–244, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1600776.

Robins:2017:IRD

[RWS17] Nikki Robins, Patricia A. H. Williams, and Krishnun Sansurooah. An investigation into remnant data on USB storage devices sold in Australia creating alarming concerns. International Journal of Computers and Applications, 39 (2):79–90, 2017. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1289689.

Sanober:2024:SPP

[SA24] Adla Sanober and Shamama Anwar. A secure and privacy preserving model for healthcare applications based on blockchain-layered architecture. International Journal of Computers and Applications, 46(12):1206–1218, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2422427.

Sleit:2010:ECM

[SAAAOH10] A. Sleit, S. Al-Adaileh, N. Al-Omari, and H. Hurani. Extending the cluster map algorithm using automated cluster identifier. International Journal of Computers and Applications, 32(2): 222–225, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2010.11441978.

Saad:2009:AEE

[SAD09] E. M. Saad, M. H. Awadalla, and R. R. Darwish. Adaptive and energy-efficient clustering architecture for dynamic sensor networks. International Journal of Computers and Applications, 31(4): 282–289, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2009.11441952.

Sammouda:2010:DDW

[Sam10] R. Sammouda. Data-dependent weight initialization in the Hopfield neural network classifier: Application to natural colour images. International Journal of Computers and Applications, 32(2): 242–249, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2010.11441982.

Shetty:2023:UPC

[SAM23] Amrithkala M. Shetty, Mohammed Fadhel Aljunid, and D. H. Manjaiah. Unleashing the power of 2D CNN with attention and pretrained embeddings for enhanced online review analysis. International Journal of Computers and Applications, 46(1):46–57, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2283647.

Shirvani:2024:RNF

[SAN24] Mirsaeid Hosseini Shirvani, Amir Akbarifar, and Amir Salar Nazokkar. Reliability nonfunctional requirement evaluation in mission-critical systems with an architectural strategy for future systems. International Journal of Computers and Applications, 46 (4):227–251, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2304385.

Sarker:2013:OBN

[Sar13] Goutam Sarker. An optimal backpropagation network for face identification and localization. International Journal of Computers and Applications, 35(2):63–69, 2013. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2013.
2.202-3388.

Sygkouna:2007:SIP

[SAS07] I. Sygkouna, M. Anagnostou, and E. Sykas. Searching for information in a P2P system. International Journal of Computers and Applications, 29 (4):394–401, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441870.

Sato:2008:COM

[Sat08] H. Sato. Construction of object migration behaviour models from temporal constraints. International Journal of Computers and Applications, 30 (3):265–275, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441906.

Souissi:2022:STT

[SBA^{+}22] Ilhem Souissi, Nadia Ben Azzouna, Rihab Abidi, Tahar Berradia, and Lamjed Ben Said. SP-TRUST: a trust management model for speed trust in vehicular networks. International Journal of Computers and Applications, 44 (11):1065–1073, 2022. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2022.2108562.

Samanta:2012:ACW

[SBS12] Raj K. Samanta, Partha Bhattacharjee, and Goutam Sanyal. Analysis of cellular wireless networks with non-classical traffic and channel reservation. International Journal of Computers and Applications, 34(3): 157–165, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2012.
3.202-3182.

Sankar:2010:SSV

[SC10] Padmanabhan Sankar and Chinnagounder Chellamuthu. Study and simulation of video streaming. International Journal of Computers and Applications, 32(4): 428–434, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2010.
4.202-2658.

Sharma:2019:TCC

[SC19] Anamika Sharma and Siddhartha Chauhan. Target coverage computation protocols in wireless sensor networks: a comprehensive review. International Journal of Computers and Applications, 43 (10):1065–1087, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1663382.

Small:2025:PPM

[SC25] Ellie Small and Javier Cabrera. Principal phrase mining: an automated method for extracting meaningful phrases from text. International Journal of Computers and Applications, 47(1):84–92, 2025. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2448494.

Sarkar:2010:QEC

[SCCB10] A. Sarkar, S. Choudhury, N. Chaki, and S. Bhattacharya. Quality evaluation of conceptual level object-oriented multidimensional data model. International Journal of Computers and Applications, 32(3): 362–371, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2010.
                                        3.202-2970.

Singh:2018:NBB

[SCM18] Manmeet Mahinderjit Singh, Ke Wan Ching, and Asrulnizam Abd Manaf. A novel out-of-band biometrics authentication scheme for wearable devices. International Journal of Computers and Applications, 42 (6):589–601, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1547347.

Stewart:2017:FGF

[SCOD17] Arran Stewart, Rachel Cardell-Oliver, and Rowan Davies. A fine-grained framework for quantifying secure management of state in objectoriented programs. International Journal of Computers and Applications, 39(1): 9–16, 2017. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2016.1253911.

Suriya:2019:EDC

[SCS19] M. Suriya, V. Chandran, and M. G. Sumithra. Enhanced deep convolutional neural network for malarial parasite classification. International Journal of Computers and Applications, 44 (12):1113–1122, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1672277.

Sivaradje:2003:NDG

[SD03] G. Sivaradje and P. Dananjayan. A novel dynamic guard channel scheme for channel allocation in cellular mobile networks. International Journal of Computers and Applications, 25(4): 225–235, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2003.11441710.

Sehil:2017:CAP

[SD17] K. Sehil and M. Darwish.

Critical analysis of power conversion topologies for stand-alone PV systems with super-capacitor. International Journal of Computers and Applications, 39 (4):179–188, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1309224.

Sumi:2023:YBW

[SD23] Lucy Sumi and Shouvik Dey.

YOLOv5-based weapon detection systems with data augmentation. International Journal of Computers and Applications, 45 (4):288–296, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2182966.

Singh:2006:SMP

[SdSNL06] A. Singh, A. L. Mora dos Santos, O. Nordstrom, and C. Lu. Stateless model for the prevention of malicious communication channels. International Journal of Computers and Applications, 28(3): 285–297, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2006.11441804.

Salah:2006:MAR

[SEB06] K. Salah and K. El-Badawi. On modelling and analysis of receive livelock and CPU utilization in highspeed networks. International Journal of Computers and Applications, 28(2): 162–169, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2006.11441800.

Shatnawi:2005:COO

[SEH05] R. A. Shatnawi, L. H. Etzkorn, and W. E. Hughes, Jr. Comparing object-oriented languages using design patterns.

International Journal of Computers and Applications, 27 (4):228–236, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441775.

Senouci:2024:ERQ

[Sen24] Oussama Senouci. Enhancing routing and quality of service using AI-driven technique for Internet of Vehicles contexts. International Journal of Computers and Applications, 46(8):628–640, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2380653.

Sharma:2005:MMN

[SG05] A. K. Sharma and A. Goel.

Moment-to-moment node transition awareness protocol (MOMENTAP). International Journal of Computers and Applications, 27(1): 1–9, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441750.

Sharma:2016:BNM

[SG16] Ashutosh Sharma and Manish Kumar Goyal. Bayesian network for monthly rainfall forecast: a comparison of K2 and MCMC algorithm. International Journal of Computers and Applications, 38(4):

199–206, 2016. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2016.1237131.

Steele:2007:XRS

[SGD07] R. Steele, W. Gardner, and T. S. Dillon. XML repository searcher-browser supporting fine-grained access control. International Journal of Computers and Applications, 29 (1):44–50, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441831.

Sultana:2022:EML

[SGMT22] Rukhsar Sultana, Jyoti Grover, Jitesh Meghwal, and Meenakshi Tripathi. Exploiting machine learning and deep learning models for misbehavior detection in VANET. International Journal of Computers and Applications, 44 (11):1024–1038, 2022. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2022.2099122.

Sun:2020:SFS

[SGZL20] Baoshan Sun, Mengying Ge, Peng Zhao, and Chunqing Li. Systematic framework for short text classification based on improved TWE and supervised MCFS topic merging strategy. International Journal of Computers and Applications, 44 (5):401–413, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1761597.

Sneed:2010:VDS

[SH10] H. Sneed and S. Huang. Valuedriven software maintenance. International Journal of Computers and Applications, 32 (2):215–221, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2010.11441977.

Sharafeddine:2010:NPV

[Sha10] S. Sharafeddine. On network planning for video services over IP networks. International Journal of Computers and Applications, 32(3): 297–308, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2010.
3.202-2671.

Shahin:2011:SIE

[Sha11] Ismail Shahin. Speaker identification in each of the neutral and shouted talking environments based on gender-dependent approach using SPHMMS. International Journal of Computers and Applications, 33(1): 83–91, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.

                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2011.
                                        1.202-3019.

Shang:2020:SSC

[Sha20] Kun Shang. Semantic service composition model based on cloud computing. International Journal of Computers and Applications, 44 (7):597–603, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2020.1738089.

Sudarsanam:2010:PEL

[SHDY10] A. Sudarsanam, T. Hauser, A. Dasu, and S. Young. A power efficient linear equation solver on a multi-FPGA accelerator. International Journal of Computers and Applications, 32(1): 56–72, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2010.11441961.

Shaheen:2023:MFP

[SHE^{+}23] Amal Shaheen, Mustafa Hammad, Wael Elmedany, Riadh Ksantini, and Saeed Sharif. Machine failure prediction using joint reserve intelligence with feature selection technique. International Journal of Computers and Applications, 45 (10):638–646, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2260619.

Shimomura:2012:ASS

[Shi12a] Takao Shimomura. Automated server-side regression testing for Web applications. International Journal of Computers and Applications, 34 (2):119–126, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2012.
2.202-3225.

Shimomura:2012:CTB

[Shi12b] Takao Shimomura. Command tag-based automatic diagnosis for computer system troubles. International Journal of Computers and Applications, 34 (3):192–199, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2012.
3.202-3408.

Shimomura:2013:SPV

[Shi13] Takao Shimomura. Semantic program visualization with attachable display classes. International Journal of Computers and Applications, 35 (2):70–78, 2013. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2013.
2.202-3463.

Sajid:2024:FEL

[SHI24] Nafis Sajid, Md. Rashidul Hasan, and Muhammad Ibrahim. Feature engineering in learningto-rank for community question answering task. International Journal of Computers and Applications, 46 (8):555–566, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2380647.

Shen:2017:PQD

[SHLA17] Yue Shen, Fida Hussain, Hui Liu, and Destaw Addis. Power quality disturbances classification based on curvelet transform. International Journal of Computers and Applications, 40(4):192–201, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1398213.

Swilam:2025:ICS

[SHP25] Zeina Swilam, Abeer Hamdy, and Andreas Pester. Improving code semantics learning using enhanced Abstract Syntax Tree. International Journal of Computers and Applications, 47(1):57–69, 2025. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2443506.

Shyu:2002:HQA

[SHS^{+}02] C.-R. Shyu, J. Harnsomburana, Y. Sethi, R. Singh, and J. C. Reid. A hybrid query approach for radiology education: Integrating content-based medical image retrieval and text-based information retrieval. International Journal of Computers and Applications, 24(2): 83–92, 2002. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2002.11441664.

Shimomura:2011:FWC

[SIT11] Takao Shimomura, Kenji Ikeda, and Muneo Takahashi. Functional Web component generation for visual Web application programming. International Journal of Computers and Applications, 33(2): 167–174, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2011.
2.202-2730.

Singh:2017:CDB

[SJ17] Laxman Singh and Zainul Abdin Jaffery. Computerized detection of breast cancer in digital mammograms. International Journal of Computers and Applications, 40 (2):98–109, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2017.1395131.

Shin:2002:GRS

[SJMO02] G. S. Shin, S. I. Jin, P. S. Mah, and C. S. Oho. Greedy request scheduling to reduce initial latency for VOD systems. International Journal of Computers and Applications, 24(2):93–99, 2002. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2002.11441665.

Sharma:2025:CSK

[SJTl25] Sumit Sharma, Sarika Jain, Mukesh Kumar Tiwari, and Sohan lal. Classifying the state of knowledge-based question answering: patterns, progress, and prospects. International Journal of Computers and Applications, 47 (1):93–105, 2025. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2426512.

Shahzad:2018:EFA

[SJZ^{+}18] Khuram Shahzad, Zeng Jianqiu, Muhammad Azam Zia, Aliya Shaheen, and Taiba Sardar. Essential factors for adopting hospital information system: a case study from Pakistan. International Journal of Computers and Applications, 43(1):26–37, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1504460.

Satish:2017:FBD

[SK17] K. V. Rama Satish and N. P.

Kavya. A framework for big data pre-processing and search optimization using HMGA-ACO: a hierarchical optimization approach. International Journal of Computers and Applications, 41 (3):183–194, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1417768.

Sharma:2018:AMF

[SK18] Geeta Sharma and Sheetal Kalra. Advanced multifactor user authentication scheme for e-governance applications in smart cities. International Journal of Computers and Applications, 41 (4):312–327, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1445352.

Singh:2021:SID

[SK21] Geeta Singh and Neelu Khare.

A survey of intrusion detection from the perspective of intrusion datasets and machine learning techniques. International Journal of Computers and Applications, 44 (7):659–669, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2021.1885150.

Szabo:2024:MLA

[SK24] Aron Benedek Szabo and Jozsef Katona. A machine learning approach for skin lesion classification on iOS: implementing and optimizing a convolutional transfer learning model with Create ML. International Journal of Computers and Applications, 46(8):666–685, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2380664.

Sadiq:2020:SSR

[SKM20] Mohd. Sadiq, Shadab Khan, and Chaudhary Wali Mohammad. Selection of software requirements using TOPSIS under fuzzy environment. International Journal of Computers and Applications, 44 (6):503–512, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1820689.

Suruliandi:2018:EER

[SKN18] A. Suruliandi, J. C. Kavitha, and D. Nagarajan. An empirical evaluation of recent texture features for the classification of natural images. International Journal of Computers and Applications, 42 (2):164–173, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1475334.

Satheesh:2009:SCE

[SKP09] A. Satheesh, S. Krishnaveni, and S. Ponkarthick. Selfconfigurable environment for the Intel ixp2400 network processor. International Journal of Computers and Applications, 31(4):268–273, 2009. ISSN 1206-212X (print), 1925-7074 (electronic). URL

https://www.tandfonline.
com/doi/full/10.1080/1206212X.
2009.11441950.

Shahila:2019:SCB

[SKS19] D. Ferlin Deva Shahila, S. H.

Krishnaveni, and Valantina Stephen. Soft computingbased non-linear discriminate classifier for multimedia image quality enhancement. International Journal of Computers and Applications, 43 (7):674–683, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1625152.

Salman:2003:ISE

[SL03] N. H. Salman and C.-Q.

Liu. Image segmentation and edge detection based on watershed techniques. International Journal of Computers and Applications, 25(4): 258–263, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2003.11441713.

Selvaraj:2008:SPB

[SLH08] M. Selvaraj, G. Y. Lazarou, and R. Hu. Simultaneous and proportional bandwidth, delay, and loss differentiation. International Journal of Computers and Applications, 30 (2):105–116, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2008.11441888.

Sun:2020:NRT

[SLQR20] Nan Sun, Guanjun Lin, Junyang Qiu, and Paul Rimba. Near real-time Twitter spam detection with machine learning techniques. International Journal of Computers and Applications, 44 (4):338–348, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2020.1751387.

Sain:2024:PAR

[SLSS24] Manoj Kumar Sain, Rabul Laskar, Joyeeta Singha, and Sandeep Saini. Patient activity recognition for robust performance under varied illumination and cluttered background in indoor environment. International Journal of Computers and Applications, 46 (12):1219–1245, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2427287.

Sajjanhar:2010:CBS

[SLZZ10] A. Sajjanhar, G. Lu, D. S.

Zhang, and W. Zhou. Connectivitybased shape descriptors. International Journal of Computers and Applications, 32 (1):93–98, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2010.11441964.

Sarma:2010:SBA

[SM10] M. Sarma and R. Mall. Statebased approach to system testing. International Journal of Computers and Applications, 32(2):181–187, 2010. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

/www.tandfonline.com/doi/
full/10.1080/1206212X.2010.
11441983.

Sepahvand:2025:SAH

[SMAM25] Majid Sepahvand, Maytham N.

Meqdad, and Fardin Abdali-Mohammadi. State-of-the-art in human activity recognition based on inertial measurement unit sensors: survey and applications. International Journal of Computers and Applications, 47(1):1–16, 2025. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2426501.

Sarhan:2012:STE

[SMF12] Amany Sarhan, Rasha Mahmoud, and Mohamed Faheem. Spatio-temporal error concealment algorithms for encoded video streams for realtime applications. International Journal of Computers and Applications, 34(2): 105–118, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2012.
                                        2.202-3140.

Smid:2003:ITM

[Smi03] J. Smid. Interactive tutoring model using information cycling on the WWW. International Journal of Computers and Applications, 25 (1):84–90, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2003.11441679.

Sahak:2013:OPC

[SML^{+}13] Rohilah Sahak, Wahidah Mansor, Khuan Y. Lee, Azlee Zabidi, and Ahmad I. M. Yassin. Optimization of principal component analysis and support vector machine for the recognition of infant cry with asphyxia. International Journal of Computers and Applications, 35(3):99–107, 2013. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2013.
2.202-3456.

Saman:2023:FSP

[SMSI23] Muhammad Sam’an, Muhammad Munsarif, Safuan, and Yahya Nur Ifriza. Feature selection in P2P lending based on hybrid genetic algorithm with machine learning. International Journal of Computers and Applications, 45 (12):764–775, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2276553.

Sitkrongwong:2019:MCR

[SMT19] Padipat Sitkrongwong, Saranya Maneeroj, and Atsuhiro Takasu. Multi-criteria rating conversion without relation loss for recommender systems. International Journal of Computers and Applications, 44 (2):130–138, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1701798.

Suhagiya:2023:TSL

[SMT^{+}23] Hartik Suhagiya, Hrithik Mistry, Aryan Trivedi, Ramchandra Mangrulkar, and Pallavi Chavan. TokenLink: a secure loyalty point exchange system powered by smart contracts. International Journal of Computers and Applications, 46 (3):152–166, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2293592.

Sone:2011:SFB

[SN11] Michael E. Sone and Ndeh N. Ningo. A simple FPGA-based wireless transmitter/receiver convolutional cryptosystem. International Journal of Computers and Applications, 33 (2):137–143, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2011.
                                        2.202-2905.

Santhoshbaboo:2012:QBB

[SN12] Sethuraman Santhoshbaboo and Balakrishnan Narasimhan. A QoS backbone based minimum delay routing protocol for mobile Ad Hoc networks. International Journal of Computers and Applications, 34 (1):36–41, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2012.
                                        1.202-3088.

Singh:2020:HBO

[SN20] Amit Singh and Aitha Nagaraju. Heuristic-based opportunistic network coding at potential relays in multihop wireless networks. International Journal of Computers and Applications, 45 (2):124–135, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1724679.

Subash:2023:DAC

[SN23] Neethu Subash and B. Nithya.

Dynamic adaptation of contention window boundaries using deep Q networks in UAV swarms. International Journal of Computers and Applications, 46 (3):167–174, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2296720.

Sheikh:2018:PDT

[SNS18] Sophiya Sheikh, Aitha Nagaraju, and Mohammad Shahid. A parallelized dynamic task scheduling for batch of task in a computational grid. International Journal of Computers and Applications, 41 (1):39–53, 2018. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1505018.

Selouani:2007:IPK

[SOC07] S. A. Selouani, D. O’Shaughnessy, and J. Caelen. Incorporating phonetic knowledge into an evolutionary subspace approach for robust speech recognition. International Journal of Computers and Applications, 29(2): 143–154, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2007.11441842.

Sophatsathit:2023:BLS

[Sop23] Peraphon Sophatsathit. A biological-like synthesis framework for software engineering environments. International Journal of Computers and Applications, 46 (4):208–217, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2301183.

Shyamala:2014:TRM

[SP14] Coimbatore K. Shyamala and Tattamangalam R. Padmanabhan. A trust–reputation model offering data retrievability and correctness in distributed storages. International Journal of Computers and Applications, 36(2): 56–63, 2014. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/

10.2316/Journal.202.2014.
2.202-3841.

Sayed:2018:PWE

[SP18] Shabina Sayed and Rakesh Poonia. Probability-weighted ensemble classifier using holoentropyenabled decision tree for data stream classification. International Journal of Computers and Applications, 43 (3):267–281, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1542766.

Srivastava:2021:LBS

[SP21] Vivek Srivastava and Ravi Shankar Pandey. Load balancing for software-defined network: a review. International Journal of Computers and Applications, 44 (8):746–759, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2021.1919835.

Shukla:2019:EET

[SPC19] Alok Kumar Shukla, Sanjeev Kumar Pippal, and Sansar Singh Chauhan. An empirical evaluation of teaching– learning-based optimization, genetic algorithm and particle swarm optimization. International Journal of Computers and Applications, 45 (1):36–50, 2019. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1686562.

Swathi:2024:NHH

[SPKS24] Bankapalli Swathi, M. Sunil Prakash, B. T. Krishna, and M. Satyanarayana. A novel hybrid heuristic-based network parameter optimization for spectral and energy efficiency in dynamic spectrum access on wireless mesh network system. International Journal of Computers and Applications, 46(4):266–279, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2307089.

Succi:2002:DCC

[SPW02] G. Succi, W. Pedrycz, and R. Wong. Dynamic composition of components using WebCODS. International Journal of Computers and Applications, 24(1): 20–27, 2002. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2002.11441656.

Sharafeddine:2006:DTB

[SR06] S. Sharafeddine and A. Riedl. Dimensioning token bucket policers for various voiceover IP applications using realscenario measurements. International Journal of Computers and Applications, 28(4): 379–387, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441824.

Shanthi:2019:NSC

[SR19] S. Shanthi and N. Rajkumar. Non-small-cell lung cancer prediction using radiomic features and machine learning methods. International Journal of Computers and Applications, 44 (12):1161–1169, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1693723.

Sapna:2020:CAS

[SR20] S. Sapna and A. Renuka.

Computer-aided system for leukocyte nucleus segmentation and leukocyte classification based on nucleus characteristics. International Journal of Computers and Applications, 42 (6):622–633, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1726013.

Swaroop:2023:DSS

[SR23a] Chigurupati Ravi Swaroop and K. Raja. AT-densenet with salp swarm optimization for outlier prediction. International Journal of Computers and Applications, 45 (12):735–747, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2273015.

Swaroop:2023:COL

[SR23b] Chigurupati Ravi Swaroop and K. Raja. CGSA optimized LSTM auto encoder for outlier detection. International Journal of Computers and Applications, 45 (7–8):497–507, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2239551.

Song:2010:GEP

[SRCXLC10] D. Song, W. Ru-Chuan, F. Xiong, and Y. Le-Chan. Gene expression programming for attribution reduction in rough set. International Journal of Computers and Applications, 32(2): 226–231, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2010.11441979.

Sun:2008:JBH

[SRJS08] J.-Z. Sun, J. Riekki, M. Jurmu, and J. Sauvola. Javabased HTTP input channel for heterogeneous wireless networks. International Journal of Computers and Applications, 30(2):

97–104, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441887.

Salhieh:2004:PAM

[SS04] A. Salhieh and L. Schwiebert.

Power-aware metrics for wireless sensor networks. International Journal of Computers and Applications, 26 (2):1–7, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2004.11441734.

Sari:2005:CAS

[SS05] T. Sari and M. Sellami. Cursive Arabic script segmentation and recognition system. International Journal of Computers and Applications, 27 (3):161–168, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441771.

Saini:2010:OCA

[SS10] D. S. Saini and N. Sharma.

OVSF code assignment schemes at the forward link of WCDMA systems. International Journal of Computers and Applications, 32(3):253–260, 2010. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

/www.tandfonline.com/doi/
full/10.2316/Journal.202.
2010.3.202-2324.

Sundarrajan:2012:ENP

[SS12] Kalavathy Sundarrajan and Ramalingam Suresh. An efficient neighbourhood pixel filtering algorithm for waveletbased image denoising. International Journal of Computers and Applications, 34 (2):90–97, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2012.
2.202-3110.

Selvaraj:2015:SPA

[SS15] Alagumani Selvaraj and Subashini Sundararajan. Survey on public auditability to ensure data integrity in cloud storage. International Journal of Computers and Applications, 37(3–4): 102–110, 2015. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2016.1188561.

Shamshirband:2018:LEF

[SS18] Shahaboddin Shamshirband and Hossein Soleimani. LAAPS: an efficient file-based search in unstructured peer-to-peer networks using reinforcement algorithm. International Journal of Computers and Applications, 43(1):62–69, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1511319.

Sardar:2019:TFS

[SS19a] Vijay M. Sardar and S. D.

Shirbahadurkar. Timbre features for speaker identification of whispering speech: selection of optimal audio descriptors. International Journal of Computers and Applications, 43(10):1047–1053, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1652788.

Shojafar:2019:IEN

[SS19b] Mohammad Shojafar and Mehdi Sookhak. Internet of Everything, Networks, Applications, and Computing Systems (IoENACS). International Journal of Computers and Applications, 42 (3):213–215, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1575621.

Saxena:2020:AAL

[SS20a] Deepika Saxena and Ashutosh Kumar Singh. Auto-adaptive learning-based workload forecasting in dynamic cloud environment. International Journal of Computers and Applications, 44(6):541–551, 2020. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (electronic). URL http:/

/www.tandfonline.com/doi/

full/10.1080/1206212X.2020.
1830245.

Shirzad:2020:JFP

[SS20b] Ehsan Shirzad and Hamid Saadatfar. Job failure prediction in Hadoop based on log file analysis. International Journal of Computers and Applications, 44 (3):260–269, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1732081.

Singh:2020:ASM

[SS20c] Jagsir Singh and Jaswinder Singh. Assessment of supervised machine learning algorithms using dynamic API calls for malware detection. International Journal of Computers and Applications, 44 (3):270–277, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1732641.

Souadih:2020:EEC

[SS20d] Rebiha Souadih and Fouzi Semchedine. Energy-efficient coverage and connectivity of wireless sensor network in the framework of hybrid sensor and vehicular network. International Journal of Computers and Applications, 44 (5):444–454, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1808346.

Selvaganapathy:2021:AME

[SS21] Shymalagowri Selvaganapathy and Sudha Sadasivam. Anti-malware engines under adversarial attacks. International Journal of Computers and Applications, 44 (8):791–804, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2021.1940744.

Stanley:2024:ECC

[SS24] S. A. Norton Stanley and Shoney Sebastian. Elastic circuit de-constructor: a pattern to enhance resiliency in microservices. International Journal of Computers and Applications, 46 (10):921–932, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2397018.

Sharma:2006:SDS

[SSG06] D. K. Sharma, R. K. Sharma, and D. Ghosh. A spatial decision support system for land management. International Journal of Computers and Applications, 28(1): 50–58, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.

tandfonline.com/doi/full/
10.1080/1206212X.2006.11441788.

Sharma:2012:NFI

[SSM12] Ashish K. Sharma, Jitendra Sharma, and Ilesh C. Mehta. A novel fuzzy integrated technical requirements prioritization software system for quality function deployment. International Journal of Computers and Applications, 34 (4):241–248, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2012.
4.202-3440.

Santra:2016:HPA

[SSMM16] Dipankar Santra, Krishna Sarker, Anirban Mukherjee, and Arindam Mondal. Hybrid PSO-ACO technique to solve multi-constraint economic load dispatch problems for 6-generator system. International Journal of Computers and Applications, 38 (2–3):96–115, 2016. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2016.1218241.

Santos:2005:HRT

[SSO05] R. Santos, J. Santos, and J. Orozco. Hard real-time systems with stochastic execution times: Deterministic and probabilistic guarantees. International Journal of Computers and Applications, 27 (2):57–62, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441758.

Sarkar:2007:BRL

[SSP07] D. Sarkar, U. K. Sarkar, and G. Peng. Bandwidth requirement of links in a hierarchical caching network: a graph-based formulation, an algorithm and its performance evaluation. International Journal of Computers and Applications, 29(1): 70–78, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441834.

Sharma:2019:NRI

[SSP19] Urvashi Sharma, Meenakshi Sood, and Emjee Puthooran. A novel resolution independent gradient edge predictor for lossless compression of medical image sequences. International Journal of Computers and Applications, 43 (8):764–774, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1610994.

Sidhu:2020:MLA

[SSS20] Brahmaleen Kaur Sidhu, Kawaljeet Singh, and Neeraj Sharma. A machine learning approach to software model refactoring. International Journal of Computers and Applications, 44 (2):166–177, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2020.1711616.

Sangeetha:2021:DPM

[SSS21a] S. Sangeetha, G. Sudha Sadasivam, and Ayush Srikanth. Differentially private model release for healthcare applications. International Journal of Computers and Applications, 44(10):953–958, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2021.2024958.

Suganya:2021:CMF

[SSS21b] D. Suganya, K. Thirunadana Sikamani, and J. Sasikala. Copy–move forgery detection of medical images using golden ball optimization. International Journal of Computers and Applications, 44(8):729–737, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2021.1907905.

Siddiqui:2016:CRE

[ST16] Zeeshan Ali Siddiqui and Kirti Tyagi. A critical review on effort estimation techniques for service-oriented-architecture-based applications. International Journal of Computers and Applications, 38(4): 207–216, 2016. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2016.1237132.

Sheremetov:2003:DRL

[SV03] L. Sheremetov and R. Peredo Valderrama. Development of reusable learning materials for WBE using intelligent components and agents. International Journal of Computers and Applications, 25(3): 170–178, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2003.11441700.

Subashini:2023:MDL

[SV23] N. J. Subashini and K. Venkatesh.

Multimodal deep learning for chronic kidney disease prediction: leveraging feature selection algorithms and ensemble models. International Journal of Computers and Applications, 45 (10):647–659, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2262786.

Saltouros:2000:EHG

[SVM^{+}00] M. P. Saltouros, E. A. Verentziotis, M. E. Markaki, M. E. Theologou, and I. S.

Venieris. An efficient hybrid genetic algorithm for finding (near-) optimal Steiner trees: an approach to routing of multipoint connections. International Journal of Computers and Applications, 22(3): 159–165, 2000. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2000.11441619.

Sum:2007:PDB

[SWL07] J. Sum, J. Wu, and C.-S. Leung. On profit density based greedy algorithm for a resource allocation problem in Web services. International Journal of Computers and Applications, 29(2): 155–163, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2007.11441843.

Sum:2008:ASM

[SWLH08] J. Sum, J. Wu, C. S. Leung, and K. I. J. Ho. Analysis on a simulated model for Gnutella topology: Connectedness and extension. International Journal of Computers and Applications, 30 (4):279–288, 2008. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2008.11441907.

Shi:2012:FPM

[SX12] Hailiang Shi and Xiangjun Xin. Fusion of panchromatic and multispectral image based on PCA and NSCT. International Journal of Computers and Applications, 34(4): 223–228, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2012.
4.202-3130.

Singh:2011:DSA

[SYJ11] Ghanshyam Singh, Rajendra P. Yadav, and Vijay Janyani. Design of symmetric and asymmetric 2 × 2 all optical Ti:LiNbO_{3} Machzehnder interferometric switches. International Journal of Computers and Applications, 33 (1):36–40, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2011.
1.202-2977.

Shen:2010:DTD

[SYZS10] J. Shen, D. Yoon, D. Zhao, and Y. Song. Denoising of two-dimensional geometric discontinuities. International Journal of Computers and Applications, 32 (2):129–140, 2010. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2010.11441969.

Samadi:2018:DMM

[SZT18] Yassir Samadi, Mostapha Zbakh, and Claude Tadonki. DT-MG: many-to-one matching game for tasks scheduling towards resources optimization in cloud computing. International Journal of Computers and Applications, 43(3):233–245, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1519630.

Touzene:2023:USB

[TA23] Abderezak Touzene and Manar Al Moqbali. User satisfaction-based genetic algorithm for load shifting in smart grid. International Journal of Computers and Applications, 45 (6):444–451, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2232167.

Tajariol:2003:SEC

[TAD03] F. Tajariol, J.-M. Adam, and M. Dubois. A study on the effect of communication and monitoring tools in Web-based tutoring. International Journal of Computers and Applications, 25(3): 206–211, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2003.11441704.

Tan:2011:CTA

[Tan11] Zuowen Tan. Comments on a threshold authenticated encryption scheme. International Journal of Computers and Applications, 33(2): 132–136, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2011.
2.202-2858.

Tahir:2025:SRM

[TAUK25] Mahjabeen Tahir, Azizol Abdullah, Nur Izura Udzir, and Khairul Azhar Kasmiran. A systematic review of machine learning and deep learning techniques for anomaly detection in data mining. International Journal of Computers and Applications, 47 (2):169–187, 2025. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2025.2449999.

Tanterdtid:2000:OSH

[TBS00] S. Tanterdtid, W. Benjapolakul, and W. Steanputtanagul. Optimizing a singlehop virtual path system using a genetic algorithm. International Journal of Computers and Applications, 22 (2):62–72, 2000. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2000.11441602.

Tai:2014:ETR

[TCC^{+}14] Shen-Chuan Tai, Chia-Ying Chang, Bo-Jhih Chen, Yu-Yi Liao, and Yung-Gi Wu. Early termination for residual quadtree decision-making in HEVC. International Journal of Computers and Applications, 36(1):23–33, 2014. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

                                        /www.tandfonline.com/doi/
                                        full/10.2316/Journal.202.
                                        2014.1.202-3809.

Thiyam:2023:SMF

[TD23] Bidyapati Thiyam and Shouvik Dey. Statistical methods for feature selection: unlocking the key to improved accuracy. International Journal of Computers and Applications, 45(6):433–443, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2223795.

Trivedi:2024:AFR

[TG24] Himani Trivedi and Mahesh Goyani. Adapting face recognition to the masked world: leveraging deep attention networks with a custom dataset. International Journal of Computers and Applications, 46 (6):427–441, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2328499.

Tiancheng:2018:OSA

[Tia18]  Wang Tiancheng.  Online

student achievement prediction model based on data envelopment analysis (DEA) for multidimensional educational internet of data mining. International Journal of Computers and Applications, 43(4):315–319, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1537109.

Tyagi:2024:USF

[TK24] Gaurav Tyagi and Rahul Kumar. Understanding the security failures of four user authentication schemes for wireless sensor networks in IoT environment. International Journal of Computers and Applications, 46 (9):769–784, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2388266.

Tchendji:2023:TPB

[TKY23] Vianney Kengne Tchendji, Joëlle Kabdjou, and Yannick Florian Yankam. Traffic predictive-based flow splitting rerouting scheme for link failures in softwaredefined networks. International Journal of Computers and Applications, 45 (7–8):508–515, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2241185.

Tao:2018:MFM

[TLC18] Song Tao, Zhuang Lei, and Jing Chenkai. Multi-frame moving video detection algorithm for IOT based on Gauss Monte Carlo particle filter. International Journal of Computers and Applications, 42(3):260–265, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1534370.

Taibi:2004:TBS

[TN04] T. Taibi and D. C. L. Ngo. Towards a balanced specification language for distributed object computing patterns. International Journal of Computers and Applications, 26 (1):1–8, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2004.11441725.

Thulasidharan:2024:NBQ

[TN24] Pillai Praveen Thulasidharan and Keshab Nath. A novel binary quantizer for variational autoencoder-based image compressor. International Journal of Computers and Applications, 46 (8):604–620, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2380651.

Thomas:2005:BFJ

[TOG^{+}05] G. Thomas, F. Ogel, A. Galland, B. Folliot, and I. Piumarta. Building a flexible Java runtime upon a flexible compiler. International Journal of Computers and Applications, 27(1):27–34, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441753.

Trabelsi:2008:NMM

[TS08] Z. Trabelsi and K. Shuaib.

A novel man-in-the-middle intrusion detection scheme for switched lans. International Journal of Computers and Applications, 30(3): 234–243, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441903.

Trabelsi:2011:IES

[TS11] Zouheir Trabelsi and Khaled Shuaib. Implementation of an effective and secure biometrics-based student attendance system. International Journal of Computers and Applications, 33(2): 144–153, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/

                                        10.2316/Journal.202.2011.
                                        2.202-2928.

Thakurta:2015:JRM

[TS15] Parag Kumar Guha Thakurta and Sujoy Sett. Joint routing in mobile networks: a weighted optimization approach. International Journal of Computers and Applications, 37(2):73–81, 2015. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

                                        /www.tandfonline.com/doi/
                                        full/10.1080/1206212X.2015.
                                        1088212.

Tseng:2019:DEF

[TS19] Yi-Ju Tseng and Yu-Lien Shih. Developing epidemic forecasting models to assist disease surveillance for influenza with electronic health records. International Journal of Computers and Applications, 42 (6):616–621, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1633762.

Tsujita:2007:RMP

[Tsu07] Y. Tsujita. Remote MPI-I/O on a parallel virtual file system using a circular buffer for high throughput. International Journal of Computers and Applications, 29(3): 291–299, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.

tandfonline.com/doi/full/
10.1080/1206212X.2007.11441859.

Thiongane:2008:SPP

[TT08] B. Thiongane and D. L.

Truong. Shared path protection for bandwidth guaranteed connections in multidomain networks. International Journal of Computers and Applications, 30(4): 289–297, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441908.

Tai:2015:SID

[TTW15] Shen-Chuan Tai, Ting-Chou Tsai, and Jui-Chiang Wen. Single image dehazing based on vector quantization. International Journal of Computers and Applications, 37 (3–4):83–93, 2015. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2016.1139871.

Tokmakoff:2002:CEN

[TVV^{+}02] Andrew Tokmakoff, Daan Velthausz, Mark Van Setten, Erik Oltmans, Ernst-Jan Goedvolk, René Bal, and Paul Porskamp. Content engineering for the nextgeneration Internet. International Journal of Computers and Applications, 24(2): 58–69, 2002. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.

                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2002.11441661.

Tian:2019:RAM

[TW19] Guixian Tian and Jian Wang. Recommendation algorithm for mobile e-commerce based on cone depth learning. International Journal of Computers and Applications, 43 (9):897–902, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1649346.

Tao-Yun:2017:HPN

[TYBwY17] Zhou Tao-Yun, Lian Baowang, and Zhang Yi. High precision nodes localization algorithm based on optimal dualfrequency comparison ranging. International Journal of Computers and Applications, 42(4):415–421, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2017.1419603.

Ta:2007:CAP

[TZ07] D. N. B. Ta and S. Zhou. Client assignment problem in distributed virtual environments. International Journal of Computers and Applications, 29(1):89–95, 2007. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

                                        /www.tandfonline.com/doi/
                                        full/10.1080/1206212X.2007.
                                        11441836.

uddin:2022:RCS

[udQ22] Aafaq Mohi ud din and Shaima Qureshi. A review of challenges and solutions in the design and implementation of deep graph neural networks. International Journal of Computers and Applications, 45 (3):221–230, 2022. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2022.2133805.

Uskov:2003:RLI

[UU03] V. Uskov and M. Uskova.

Reusable learning and information atoms approach to Web-based education. International Journal of Computers and Applications, 25(3): 188–197, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2003.11441702.

Villanova:2002:PES

[VBM02] M. Villanova, N. Belkhatir, and H. Martin. A process environment supporting Web multimedia information systems. International Journal of Computers and Applications, 24 (2):70–76, 2002. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2002.11441662.

Valenti:2000:CPV

[VCP00] S. Valenti, A. Cucchiarelli, and M. Panti. A communication paradigm for voluntary collaboration. International Journal of Computers and Applications, 22(2): 109–117, 2000. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2000.11441607.

Veeravalli:2005:DPA

[Vee05] B. Veeravalli. Design and performance analysis of heuristic load-balancing strategies for processing divisible loads on Ethernet clusters. International Journal of Computers and Applications, 27(2): 97–107, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441762.

Venkataramanan:2014:MFB

[VG14] Chakrapani Venkataramanan and Selvaraj M. Girirajkumar. Markov fuzzy based MAC protocol for life time maximization of wireless sensor network. International Journal of Computers and Applications, 36(4):133–139, 2014. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

/www.tandfonline.com/doi/
full/10.2316/Journal.202.
2014.4.202-3934.

Vecchiola:2009:AFD

[VGPB09] C. Vecchiola, A. Grosso, A. Passadore, and A. Boccalatte. Agentservice: a framework for distributed multiagent system development. International Journal of Computers and Applications, 31 (3):204–210, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2009.11441942.

VanHilst:2011:PAW

[VHL11] Michael VanHilst, Shihong Huang, and Hugh Lindsay. Process analysis of a waterfall project using repository data. International Journal of Computers and Applications, 33(1):49–56, 2011. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2011.
1.202-2986.

Vashisht:2024:ESJ

[VJTG24] Rohit Vashisht, Abhinav Juneja, Gagan Thakral, and Sonam Gupta. An empirical study of just-in-time-defect prediction using various machine learning techniques. International Journal of Computers and Applications, 46 (6):397–406, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2328489.

Vaithyasubramanian:2019:EWS

[VLK19] S. Vaithyasubramanian, D. Lalitha, and C. K. Kirubhashankar. Enhancing website security against bots, spam and web attacks using l CAPTCHA. International Journal of Computers and Applications, 45 (1):63–69, 2019. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1702285.

Vargo:2003:LOE

[VNBA03] J. Vargo, J. C. Nesbit, K. Belfer, and A. Archambault. Learning object evaluation: Computermediated collaboration and inter-rater reliability. International Journal of Computers and Applications, 25(3): 198–205, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2003.11441703.

Venkateswarlu:2019:DSD

[VNR19] V. T. Venkateswarlu, P. V. Naganjaneyulu, and D. N. Rao. Delay sensitive data routing optimization using rendezvous agents in wireless sensor networks with mobile sink. International Journal of Computers and Applications, 43(5):445–452, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1576326.

Valiveti:2017:PAS

[VP17] Hima Bindu Valiveti and Trinatha Rao Polipalli. Performance analysis of SLTC-D2D handover mechanism in software-defined networks. International Journal of Computers and Applications, 41 (4):245–254, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1396414.

Veerashetty:2019:DRI

[VP19a] Sachinkumar Veerashetty and Nagaraj B. Patil. Design of rotation, illumination, and scale invariant Gabor texture descriptor for image texture analysis and retrieval. International Journal of Computers and Applications, 43 (9):940–948, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1658378.

Veerashetty:2019:MDB

[VP19b] Sachinkumar Veerashetty and Nagaraj B. Patil. Manhattan distance-based histogram of oriented gradients for contentbased medical image retrieval. International Journal of Computers and Applications, 43 (9):924–930, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1653011.

Vyas:2008:RCG

[VR08] V. S. Vyas and P. P. Rege. Real coded genetic algorithm for optimal parameters selection in circular Mellin feature extractors. International Journal of Computers and Applications, 30(2): 134–141, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2008.11441892.

Venkateswaramma:2019:NPS

[VR19] P. V. Venkateswaramma and I. Ravi Prakash Reddy. Node pattern state and trust-rate base route selection for reliable data transmission in mobile ad hoc networks. International Journal of Computers and Applications, 43 (9):874–880, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1645416.

Verma:2024:TAB

[VS24] Gurudatta Verma and Tirath Prasad Sahu. TOPSIS-ACO based feature selection for multilabel classification. International Journal of Computers and Applications, 46 (6):363–380, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2321683.

Vakali:2001:JBM

[VT01] A. I. Vakali and E. D. Terzi.

A Java-based model for I/O scheduling in tertiary storage subsystems. International Journal of Computers and Applications, 23(1): 45–50, 2001. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2001.11441633.

Vidyarthi:2006:CBL

[VTS06] D. Prakash Vidyarthi, A. Kumar Tripathi, and B. Kumer Sarker. Cluster-based load partitioning and allocation in distributed computing systems. International Journal of Computers and Applications, 28(4):301–307, 2006. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441814.

Virupaksha:2019:SBA

[VV19] Shashidhar Virupaksha and D. Venkatesulu. Subspacebased aggregation for enhancing utility, information measures, and cluster identification in privacy preserved data mining on high-dimensional continuous data. International Journal of Computers and Applications, 44 (12):1130–1139, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1686211.

Wang:2006:ECB

[Wan06] C. S. Wang. An efficient content-based retrieval system for 3D objects. International Journal of Computers and Applications, 28(4): 308–313, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2006.11441815.

Wang:2011:VAD

[Wan11] Kun-Ching Wang. Voice activity detector for noise spectrum estimation using a dynamic band-splitting entropy estimate. International Journal of Computers and Applications, 33(3):220–228, 2011. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

                                        /www.tandfonline.com/doi/
                                        full/10.2316/Journal.202.
                                        2011.3.202-2979.

Wang:2019:ARE

[Wan19] Wei Wang. An algorithm for rationality estimation of movement trend based on big data analysis. International Journal of Computers and Applications, 43 (9):868–873, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1645409.

Wee:2004:SSG

[WBV04] C. M. Wee, M. S. Beg, and B. Vaillant. Secure Setrequest and Getresponse for SNMP: APSSNMP. International Journal of Computers and Applications, 26 (1):1–7, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2004.11441723.

Wahlang:2024:DRU

[WC24] Roisan Wahlang and K. Chandrasekaran. Dimensionality reduction using neural networks for lattice-based cryptographic keys. International Journal of Computers and Applications, 46(10):889–910, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2396328.

Wu:2006:EIS

[WCC06] B.-F. Wu, Y.-L. Chen, and C.-C. Chiu. Efficient implementation of several multilevel thresholding algorithms using a combinatorial scheme. International Journal of Computers and Applications, 28(3): 259–269, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441811.

Wei:2017:SDS

[WDW17] Bingyang Wei, Harry S. Delugach, and Yi Wang. From state diagrams to sequence diagrams: a requirements acquisition approach. International Journal of Computers and Applications, 41 (2):89–111, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1408982.

Werghi:2009:ROR

[Wer09] N. Werghi. Robust 3D object registration based on pairwise matching of geometric distributions. International Journal of Computers and Applications, 31(1):1–8, 2009. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2009.11441917.

Winley:2006:AHF

[WF06] G. K. Winley and J. P.-O. Fan. Admissible heuristic functions for flow-shop problems. International Journal of Computers and Applications, 28(3): 222–226, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441806.

Wenguo:2017:ISW

[WG17] Liao Wenguo and Liao Guangping. Incremental semantic web retrieval model based on web service. International Journal of Computers and Applications, 42 (1):76–83, 2017. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1397386.

Wu:2019:AIR

[WG19] Desheng Wu and Yuhui Guan.

Artificial intelligence retrieval algorithm for text data from multiple data sources. International Journal of Computers and Applications, 43 (7):715–719, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1639353.

Wu:2017:AFG

[WGE17] Zhiqiang (John) Wu, Sampson Gholston, and Letha Etzkorn. An approach to finding good anchor nodes in ontologies. International Journal of Computers and Applications, 41(2):140–148, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1412166.

Welch:2001:SVD

[WHBS01] P. H. Welch, G. H. Hilderink, A. W. P. Bakkers, and G. S. Stiles. Safe and verifiable design of concurrent Java programs. International Journal of Computers and Applications, 23(3):

159–165, 2001. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2001.11441647.

Wu:2020:SMS

[WhP20] Shinq-Jen Wu and Van hung Pham. Speeding up multi-SVMs through modified working set selection. International Journal of Computers and Applications, 44 (5):426–432, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2020.1779978.

Williams:2007:FTA

[WL07] D. Williams and H. Lutfiyya. Fault-tolerant authentication services. International Journal of Computers and Applications, 29 (2):107–114, 2007. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2007.11441838.

Wang:2018:SPE

[WL18] Wenjun Wang and Wei-Ming Lin. System performance enhancement with thread suspension for simultaneous multi-threading processors. International Journal of Computers and Applications, 42(8):774–786, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1489572.

Wu:2005:ABA

[WLC05] S.-Y. Wu, S.-J. Leu, and R.- S. Chang. An agent-based architecture for resource reservation in mobile IP networks. International Journal of Computers and Applications, 27 (4):266–273, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441780.

Wu:2010:RTL

[WLC10] B. F. Wu, C.-T. Lin, and C.- J. Chen. Real-time lane and vehicle detection based on a single camera model. International Journal of Computers and Applications, 32(2): 149–159, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2010.11441971.

Wang:2018:IDB

[WLHZ18] Limin Wang, Mingyang Li, Xuming Han, and Kaiyue Zheng. An improved densitybased spatial clustering of application with noise. International Journal of Computers and Applications, 40 (3):1–7, 2018. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1424103.

Wang:2010:NPG

[WLL10] F. Wang, H. Li, and R. Li. Non-parametric generalized cross entropy estimator for BSS algorithm. International Journal of Computers and Applications, 32(3): 372–380, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2010.
                                        3.202-2822.

Wang:2012:DPM

[WLZZ12] Yanlong Wang, Jinhua Liu, Ting Zhang, and Jin Zhang. 3-D porous media reconstruction using a 2-D micro-CT image and MPS. International Journal of Computers and Applications, 34(1): 48–54, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2012.
                                        1.202-3141.

Wang:2002:PSP

[WMB02] H. Wang, D. Mamora, and L. N. Bhuyan. Ppctsi: a software package for rapid analysis of CT scan images. International Journal of Computers and Applications, 24 (3):103–111, 2002. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2002.11441668.

Wang:2008:WGB

[WPJZ08] L. Wang, S. Piao, Y. Jiang, and L. Zhang. On a Webgraph-based micronetwork architecture for SoCs. International Journal of Computers and Applications, 30 (1):1–8, 2008. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441879.

Wahaballa:2015:TSE

[WQX^{+}15] Abubaker Wahaballa, Zhen Qin, Hu Xiong, Zhiguang Qin, and Mohammed Ramadan. A taxonomy of secure electronic English auction protocols. International Journal of Computers and Applications, 37 (1):28–36, 2015. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2015.1063248.

Wu:2000:ICD

[WR00] Jie Wu and V. Rancov. Implementing context-driven parallel computations. International Journal of Computers and Applications, 22(1): 29–37, 2000. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2000.11441596.

Wu:2015:FED

[WT15] Yung-Gi Wu and Sheng-Lun Tsai. Fall event detection by gyroscopic and accelerometer sensors in smart phone. International Journal of Computers and Applications, 37 (2):60–66, 2015. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2015.1080446.

Wang:2007:IAQ

[WTLW07] Y.-H. Wang, C.-H. Tsai, H.- Z. Lin, and C.-A. Wang. Interference-aware QoS multipath routing for Ad Hoc wireless networks. International Journal of Computers and Applications, 29(4): 372–378, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441868.

Wu:2008:MFP

[Wu08]  Z. D. Wu.  Modelling

for a federated peer-to-peer MMOG architecture. International Journal of Computers and Applications, 30(4): 309–318, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441910.

Wu:2010:IEL

[Wu10] Yung-Gi Wu. Image enlargement by Lagrange interpolation. International Journal of Computers and Applications, 32(4):420–427, 2010. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

/www.tandfonline.com/doi/
full/10.2316/Journal.202.
2010.4.202-2470.

Wu:2018:LGI

[Wu18]  Ze Wu.  Local-global it-

erative linear embedding optimization-based positioning algorithm in sensor network. International Journal of Computers and Applications, 43(2):181–185, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1537094.

Wang:2007:PBL

[WVK07] Q. Wang, J. Vincent, and G. King. Prediction based link state update. International Journal of Computers and Applications, 29(4): 379–393, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441869.

Wang:2018:SSH

[WWHW18] Limin Wang, Nianbo Wang, Xuming Han, and Yizhang Wang. Semi-supervised hierarchical optimization-based affinity propagation algorithm and its applications. International Journal of Computers and Applications, 40 (3):1–10, 2018. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1424283.

Wahaballa:2014:MLS

[WWL^{+}14] Abubaker Wahaballa, Osman Wahballa, Fagen Li, Mohammed Ramadan, and Zhiguang Qin. Multiplelayered securities using steganography and cryptography. International Journal of Computers and Applications, 36 (3):93–100, 2014. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2014.
                                        3.202-3917.

Wei:2013:PMS

[WXYN13] Yuanwei Wei, Yali Xue, Jiongyu Yin, and Weidou Ni. Prediction of municipal solid waste generation in China by multiple linear regression method. International Journal of Computers and Applications, 35(3): 136–140, 2013. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2013.
                                        3.202-3898.

Wang:2019:IGC

[WYZ^{+}19] Lei Wang, Shanlin Yang, Aizhu Zhou, Rongjing Huang, Shuai Ding, Hao Wang, and Juncheng Hu. An intelligent gastric cancer screening method based on convolutional neural network and support vector machine. International Journal of Computers and Applications, 43 (7):720–725, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1640345.

Wang:2005:RVT

[WZ05] Z. Wang and J. Zhang. Routing with virtual transmission range in mobile AD HOC networks. International Journal of Computers and Applications, 27(4): 237–243, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441776.

Wang:2018:APC

[WZTH18] Limin Wang, Kaiyue Zheng, Xing Tao, and Xuming Han. Affinity propagation clustering algorithm based on large-scale data-set. International Journal of Computers and Applications, 40 (3):1–6, 2018. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1425184.

Wang:2020:RAB

[WZW20] Dan Wang, Kai Zhao, and Yi Wang. Retracted article: Based on deep learning in traffic remote sensing image processing to recognize target vehicle. International Journal of Computers and Applications, 45(2):180–186, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1080/1206212X.2020.1735764.

Xiong:2013:STP

[XCCL13] Hu Xiong, Yanan Chen, Zhong Chen, and Fagen Li. Simple three-party password-based key exchange protocol with provable security. International Journal of Computers and Applications, 35(1): 44–50, 2013. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2013.
                                        1.202-3540.

Xinxin:2018:KSM

[XF18] Hu Xinxin and Liang Fengshou. Keyword search method of distributed file in large data environment. International Journal of Computers and Applications, 42 (7):644–648, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1518663.

Xiong:2012:SPB

[XGC12] Hu Xiong, Zhi Guan, and Zhong Chen. On the security of a pairing-based signcryption scheme using selfcertified public keys. International Journal of Computers and Applications, 34(1): 55–57, 2012. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2012.
1.202-3156.

Xiaoming:2018:AIP

[Xia18]  Kou Xiaoming.  Analysis

on the Internet public opinions in the era of we media based on fast dual-cycle level set. International Journal of Computers and Applications, 43(4):305–309, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1537097.

Xiong:2009:PSI

[XLQ09] H. Xiong, F. Li, and Z. Qin.

Provably secure identity based threshold signature without random oracles. International Journal of Computers and Applications, 31(4): 290–295, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2009.11441953.

Xu:2023:ENA

[XM23] Yang Xu and Yongjie Ma.

Evolutionary neural architecture search based on group whitening residual ConvNet and large kernel subspace attention module. International Journal of Computers and Applications, 46 (4):195–207, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2023.2300221.

Xu:2009:DDR

[XXR09] H. Xu, L. Xing, and R. Robidoux. Drbd: Dynamic reliability block diagrams for system reliability modelling. International Journal of Computers and Applications, 31 (2):132–141, 2009. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2009.11441934.

Xia:2002:CMN

[XY02] Shundong Xia and Jinyuan You. A coordination model for network computing environments. International Journal of Computers and Applications, 24(3):153–159, 2002. ISSN 1206-212X (print), 1925-7074 (electronic). URL

                                        https://www.tandfonline.
                                        com/doi/full/10.1080/1206212X.
                                        2002.11441675.

Xiangdong:2010:MOC

[XYSX10] L. Xiangdong, Z. Yuelong, C. Songqiao, and Y. Xiaoli. A multiversion optimistic concurrency control protocol in mobile broadcast environments. International Journal of Computers and Applications, 32(3):

261–266, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2010.
3.202-2462.

Xiaoqiong:2018:ISA

[XZ18] Wei Xiaoqiong and Yin E.

Zhang. Image segmentation algorithm based on dynamic particle swarm optimization and K-means clustering. International Journal of Computers and Applications, 42 (7):649–654, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1521090.

Xu:2017:HBP

[XZCL17] Xiaofei Xu, Zeyu Zhang, Yanbin Chen, and Li Li. HMMbased predictive model for enhancing data quality in WSN. International Journal of Computers and Applications, 42 (4):351–359, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1395133.

Youssef:2006:LCT

[YA06] M. Youssef and A. Agrawala.

Location-clustering techniques for WLAN location determination systems. International Journal of Computers and Applications, 28(3):

278–284, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441813.

Yang:2017:ARC

[Yan17] Shiqi Yang. Analysis for the reliability of computer network by using intelligent cloud computing method. International Journal of Computers and Applications, 41 (4):306–311, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1417770.

Yang:2018:MNV

[Yan18] Fan Yang. Model of network video retrieval of dynamic Bayesian network optimization for the mobile terminal. International Journal of Computers and Applications, 43(4):310–314, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1537098.

Yang:2019:NNG

[Yan19] Ge Yang. Network node grouping algorithm and evaluation model based on clustering and Bayesian classifier. International Journal of Computers and Applications, 45(1):70–76, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1703337.

Yao:2017:ACD

[Yao17] Weihua Yao. Application of computer digital simulation technology in action structure image research on martial art. International Journal of Computers and Applications, 42(3):216–221, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1396431.

Yang:2018:PLO

[YB18] Hui Yang and Peter A.

Bath. Predicting loneliness in older age using two measures of loneliness. International Journal of Computers and Applications, 42 (6):602–615, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1562408.

Yang:2004:IAD

[YCYL04] C.-H. Yang, L.-Y. Chuang, C.- H. Yang, and C.-H. Luo. Internet access for disabled persons using Morse code. International Journal of Computers and Applications, 26 (1):10–16, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2004.11441719.

Ye:2017:ITF

[Ye17] Tao Ye. Internet of Things financial data capture technology based on improved particle swarm optimization FLFNN. International Journal of Computers and Applications, 42 (1):102–107, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2017.1397389.

Yi:2018:CMO

[YF18] Tong Yi and Chun Fang. A complexity metric for objectoriented software. International Journal of Computers and Applications, 42 (6):544–549, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1478194.

Yu:2010:CTF

[YHK^{+}10a] Jia Yu, Rong Hao, Fanyu Kong, Xiangguo Cheng, Huawei Zhao, and Yangkui Chen. Cryptanalysis of a type of forward secure signatures and multi-signatures. International Journal of Computers and Applications, 32(4): 476–481, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2010.
                                        4.202-2992.

Yu:2010:IBF

[YHK^{+}10b] Jia Yu, Rong Hao, Fanyu Kong, Xiangguo Cheng, Huawei Zhao, and Chen Yangkui. Identity-based forward secure threshold signature scheme based on mediated RSA. International Journal of Computers and Applications, 32 (4):469–475, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2010.
4.202-2927.

Yeo:2004:JBW

[YHL04] C. K. Yeo, S. C. Hui, and B. S. Lee. Java-based, WAPenabled unified messaging system. International Journal of Computers and Applications, 26(4):1–8, 2004. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2004.11441741.

Yindi:2018:MCA

[Yin18a] Dong Yindi. Multi-classification analysis of large data based on knowledge element in micro-blog short text. International Journal of Computers and Applications, 43 (4):377–382, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1554835.

Ying:2018:PAB

[Yin18b] Wang Ying. Positioning algorithm based on linear embedding optimization in internet of sensor network. International Journal of Computers and Applications, 43 (2):153–158, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1534373.

Yang:2007:IUC

[YLY07] F. W. Yang, H. J. Lin, and S. H. Yen. An improved unsupervised clustering algorithm based on population Markov chain. International Journal of Computers and Applications, 29(3): 253–258, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2007.11441855.

Yang:2015:DFP

[YLZ^{+}15] Yin Yang, Wenyi Li, Manning Zhang, Xiaomei Ding, and Jing Dai. Disk failure prediction model for storage systems based on disk SMART technology. International Journal of Computers and Applications, 37(3–4): 111–119, 2015. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2016.1188562.

Yang:2001:CP

[YM01] Simon X. Yang and Max Meng. Call for papers. International Journal of Computers and Applications, 23 (1):77, 2001. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2001.11441626.

Ya:2017:TAW

[YMRLJ17] Wang Ya, Zhou Meng-Ran, Niu Lei, and Zhao Jia. Trust analysis of WSN nodes based on fuzzy theory. International Journal of Computers and Applications, 42 (1):52–56, 2017. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1396430.

Yahi:2015:RTH

[YMTbB15] Amira Yahi, Kamel Messaoudi, Salah Toumi, and El bay Bourennane. Realtime hardware implementation of a speed FSBMA used in H.264/AVC. International Journal of Computers and Applications, 37(3–4): 134–142, 2015. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2016.1188554.

Yassine:2023:RAN

[YNA23] Sembati Yassine, Naja Najib, and Jamali Abdellah.

Routing approaches in named data network: a survey and emerging research challenges. International Journal of Computers and Applications, 46(1):32–45, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2279811.

Younis:2005:EPA

[YP05] M. Younis and S. Ponnusamy. Efficient and predictable approach for supporting spatial partitioning of real-time applications. International Journal of Computers and Applications, 27(4): 218–227, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441774.

Yogita:2020:DPP

[YPY20] Yogita Yogita, Vipin Pal, and Anju Yadav. DDC protocol to protract network lifetime of wireless sensor networks. International Journal of Computers and Applications, 44 (4):316–323, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2020.1737372.

Yuan:2006:CGR

[YTW^{+}06] C. Yuan, C. Tang, Y. Wen, J. Zuo, J. Peng, and J. Hu.

Convergency of genetic regression in data mining based on gene expression programming and optimized solution. International Journal of Computers and Applications, 28(4): 359–366, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441822.

Yu:2019:MDF

[Yu19] Jia Yu. Missing data filling algorithm in OA (office automation) cooperative office system. International Journal of Computers and Applications, 44(8):713–720, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1662172.

Yuting:2018:DGL

[Yut18] Lu Yuting. A 3-D geographic location routing protocol based on forward region adaptive. International Journal of Computers and Applications, 43(4):360–366, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1551757.

Yang:2008:PSI

[YW08] M. Yang and L. Wang.

Preface: Special issue on high-performance computing architectures. International Journal of Computers and Applications, 30(1): 62–63, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2008.11441882.

Yu:2024:NLB

[YWDC24] Xiaohan Yu, Haixia Wang, Keming Dong, and Chao Chen. A novel LDVP-based anomaly detection method for data streams. International Journal of Computers and Applications, 46 (6):381–389, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2323266.

Yuan:2019:ASC

[YWH19] Zhi Yuan, Weiqing Wang, and Shan He. Application of sustainable computing based advanced intelligent power electronic technology for smart grid systems. International Journal of Computers and Applications, 43(5):453–461, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1576993.

Yu:2006:DSG

[YWZ^{+}06] H. Yu, L. Wang, J. Zhang, J. Barksdale, and X. Yuan. Developing a secure geospatial visualization and collaboration system using software engineering technology. International Journal of Computers and Applications, 28(4): 350–358, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441821.

Yang:2006:COP

[YWZJ06] M. Yang, J. Wang, S. Q.

Zheng, and Y. Jiang. Code optimization of polynomial approximation functions on clustered instruction-level parallelism processors. International Journal of Computers and Applications, 28(4): 367–378, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441823.

Yong-xiong:2017:NAD

[YxLmLx17] Zhang Yong-xiong, Wang Liang-ming, and Yi Lu-xia. A network attack discovery algorithm based on unbalanced sampling vehicle evolution strategy for intrusion detection. International Journal of Computers and Applications, 42(1):84–92, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1397387.

Yuan:2010:ISA

[YYD10] H. Yuan, Y. Ye, and J. Deng.

Iterative sib algorithm based on simulated annealing. International Journal of Computers and Applications, 32(3): 309–317, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2010.
3.202-2717.

Yong:2017:DCA

[YYM17] Cao Yong, Mu Yongmin, and Shen Meie. Data change analysis based on function call path. International Journal of Computers and Applications, 40(3):1–10, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1413625.

Yang:2008:LLM

[YYZ08] Y. Yang, X. Yang, and C. Zhou. Lmcgrid: a low management cost grid computation model. International Journal of Computers and Applications, 30(1): 56–61, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441875.

Yu:2019:RAR

[YZ19a] Dahai Yu and Juan Zhu. Retracted article: Research on digital image wavelet transform filtering optimization processing method based on DSP Internet of Things. International Journal of Computers and Applications, 45 (1):77–87, 2019. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1080/1206212X.2019.1706031.

Zhang:2019:OBI

[yZ19b] Chun yu Zhang. The optimization of bibliographic information resource integration of digital library. International Journal of Computers and Applications, 43 (9):910–915, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1649838.

Yan:2021:PPC

[YZ21] Shiliang Yan and Xiaofeng Zhang. PCNet: partial convolution attention mechanism for image inpainting. International Journal of Computers and Applications, 44 (8):738–745, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2021.1909280.

Zanaty:2018:GHK

[ZA18] E. A. Zanaty and Ashraf Afifi.

Generalized Hermite kernel function for support vector machine classifications. International Journal of Computers and Applications, 42 (8):765–773, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1489571.

Zou:2008:NCK

[ZB08]  X. Zou and L. Bai.  A

new class of key management scheme for access control in dynamic hierarchies. International Journal of Computers and Applications, 30(4): 331–337, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441913.

Zhuo:2020:RAK

[ZB20] Rongqing Zhuo and Zhongxian Bai. Retracted article: Key technologies of cloud computing-based IoT data mining. International Journal of Computers and Applications, 45 (2):196–203, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1080/1206212X.2020.1738665.

Zheng:2007:CSB

[ZC07] Z. Zheng and T. K. Y. Chan. A client–server based view-dependent multiresolution mesh hierarchy. International Journal of Computers and Applications, 29(4): 362–371, 2007. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441867.

Zhang:2023:MPD

[ZDG^{+}23] Xinyu Zhang, Hang Dong, Liang Gong, Xin Cheng, Zhenghui Ge, and Liangchao Guo. Multiple paddy disease recognition methods based on deformable transformer attention mechanism in complex scenarios. International Journal of Computers and Applications, 45 (10):660–672, 2023. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2023.2263254.

Zhu:2021:VBF

[ZDW^{+}21] Huazheng Zhu, Jinglong Du, LuLu Wang, Baoru Han, and Yuanyuan Jia. A vision-based fall detection framework for the elderly in a room environment using motion features and DAG-SVM. International Journal of Computers and Applications, 44 (7):678–686, 2021. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2021.1886417.

Zanaty:2017:NAC

[ZEZ17] E. A. Zanaty and S. F. El-Zoghdy. A novel approach for color image segmentation based on region growing. International Journal of Computers and Applications, 39(3):123–139, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2017.1309218.

Zhang:2024:MVT

[ZFH24] Xuan Zhang, Mengqi Fan, and Mo Hou. Mobilenet V3-transformer, a lightweight model for image caption. International Journal of Computers and Applications, 46 (6):418–426, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2024.2328498.

Zara:2004:PSL

[ZFV04] F. Zara, F. Faure, and J.- M. Vincent. Parallel simulation of large dynamic system on a PC cluster: Application to cloth simulation. International Journal of Computers and Applications, 26 (3):1–8, 2004. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2004.11441739.

Zhang:2017:RRA

[Zha17a] Fengqiong Zhang. Research on reliability analysis of computer network based on intelligent cloud computing method.

International Journal of Computers and Applications, 41 (4):283–288, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1402622.

Zhang:2017:CSA

[Zha17b] Yan Zhang. Classified scheduling algorithm of big data under cloud computing. International Journal of Computers and Applications, 41 (4):262–267, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1397343.

Zheng:2019:FSA

[Zhe19] Xiaochang Zheng. Feature selection algorithm of network attack big data under the interference of fading noise. International Journal of Computers and Applications, 44(9):807–813, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2019.1703327.

Zhu:2017:RDM

[Zhu17] Jia Zhu. Research on data mining of electric power system based on Hadoop cloud computing platform. International Journal of Computers and Applications, 41 (4):289–295, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1402623.

Zhu:2017:FPE

[ZJG17] Honglei Zhu, Dahai Jin, and Yunzhan Gong. False positive elimination in suspected code fault automatic confirmation. International Journal of Computers and Applications, 40(3):1–9, 2017. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2017.1397342.

Zhang:2008:MNA

[ZJH08] J. A. Zhang, S. F. Jones, and A. S. Helal. Mobile network adaptation in the Ubidata mobile file system. International Journal of Computers and Applications, 30(2):65–72, 2008. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441883.

Zhang:2008:FTR

[ZJR08] L. Zhang, Y. Jiang, and E. E. Regentova. Fault tolerance routing for wavelength routed optical networks in ONoC. International Journal of Computers and Applications, 30 (1):23–35, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441880.

Zhu:2015:EPR

[ZJZ15] Zhu Zhu, Julang Jiang, and Xiaoguo Zhang. Edgepreserving regularized filter with visual perception outlier measurement. International Journal of Computers and Applications, 37(3–4): 120–126, 2015. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2016.1188567.

Zhang:2010:ADP

[ZKNL10] Weiyi Zhang, Jun Kong, Kendall Nygard, and Ming Li. Adaptive design of pervasive computing system under QoS constraints1. International Journal of Computers and Applications, 32(4): 482–492, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2010.
4.202-2837.

Zhen:2005:RTO

[ZKS05] B. Zhen, M. Kobayashi, and M. Shimizu. The reading of transmission-only active RFID tags. International Journal of Computers and Applications, 27(1):10–19, 2005. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2005.11441751.

Zhou:2018:QGA

[ZL18] Aiping Zhou and Hongbing Li. A QoS guarantee approach for multimedia software system. International Journal of Computers and Applications, 42(6):578–588, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2018.1486559.

Zhu:2006:IPC

[ZLC06] X. Zhu, J. Liao, and J. Chen. Improved protocol conversion methodology and its application. International Journal of Computers and Applications, 28(3):210–221, 2006. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:/

                                        /www.tandfonline.com/doi/
                                        full/10.1080/1206212X.2006.
                                        11441805.

Zeng:2019:PMK

[ZLF19] Min Zeng, Fenjuan Li, and Huiping Feng. Prediction model of key customer loss in e-commerce combined with cemo real time streaming data framework. International Journal of Computers and Applications, 43 (9):889–896, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1649345.

Zhou:2018:DPC

[ZLHW18] Ruihong Zhou, Qiaoming Liu, Xuming Han, and Limin Wang. Density Peak Clustering algorithm using knowledge learning-based fruit fly optimization. International Journal of Computers and Applications, 40(3):1–10, 2018. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2018.1440340.

Zhenyu:2000:AGA

[ZMY^{+}00] Chen Zhenyu, J. B. Mbede, Zhou Yan, Li Dehua, and Hu Hanping. Application of a genetic algorithm in triangulation of a 3-D object surface. International Journal of Computers and Applications, 22(2):73–77, 2000. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2000.11441603.

Zhang:2008:HPA

[ZNLA08] M. Z. Zhang, H. T. Ngo, A. R. Livingston, and V. K. Asari. A high performance architecture for implementation of 2-D convolution with quadrant symmetric kernels. International Journal of Computers and Applications, 30(4): 298–308, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441909.

Zhang:2019:FDA

[ZP19] Jie Zhang and Haibo Pu. Fast data acquisition algorithm for remote monitoring system of smart home. International Journal of Computers and Applications, 44 (8):705–712, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2019.1667092.

Zhang:2020:RAD

[ZSWF20] Hongxiang Zhang, Shaojie Shi, Yongkai Wu, and Tong Feng. Retracted article: Development of computer-based agricultural remote intelligent information monitoring system. International Journal of Computers and Applications, 45(2):151–160, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1080/1206212X.2020.1730567.

Zhou:2008:TCV

[ZTSL08] S. P. Zhou, S.-P. Ting, Z. Q. Shen, and L. B. Luo. Twilight City — a virtual environment for MOUT. International Journal of Computers and Applications, 30(2): 117–123, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2008.11441889.

Zheng:2008:SRP

[ZW08] J.-Y. Zheng and X. Wang.

Streaming route panorama for seamless city guidance. International Journal of Computers and Applications, 30(3): 192–200, 2008. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2008.11441898.

Zheng:2024:MCG

[ZWC24] Liwei Zheng, Yan Wang, and Zhanqi Cui. Modeling contextual goals and detecting context conflicts from BDDbased user stories. International Journal of Computers and Applications, 46 (12):1057–1068, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2387444.

Zhang:2014:RTD

[ZWLH14] Junqiang Zhang, Chonglong Wu, Junqi Liu, and Haihua Hu. The research of three-dimensional integrated framework of landslide disaster monitoring data. International Journal of Computers and Applications, 36(4): 148–154, 2014. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.2316/Journal.202.2014.
4.202-3994.

Zhang:2024:MLL

[ZX24] Hanyu Zhang and Hao Xu. Manifold learning localization based on local tangent space alignment for wireless sensor networks. International Journal of Computers and Applications, 46 (9):785–794, 2024. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/1206212X.2024.2388865.

Zhang:2007:AAN

[ZXF07] M. Zhang, S. Xu, and J. Fulcher. Anser: Adaptive neuron artificial neural network system for estimating rainfall. International Journal of Computers and Applications, 29 (3):215–222, 2007. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2007.11441850.

Zhang:2006:ASV

[ZXT^{+}06] X. Zhang, X. L. Xiao, J. W. Tian, J. Liu, and G. Y. Xu. Application of support vector machines in classification of magnetic resonance images. International Journal of Computers and Applications, 28 (2):122–128, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.
tandfonline.com/doi/full/
10.1080/1206212X.2006.11441795.

Zhao:2019:RAI

[ZY19a] Jiewen Zhao and Jiaojie Yuan.

Retracted article: An intelligent model to reduce the energy consumption of sensor network nodes. International Journal of Computers and Applications, 45 (1):105–113, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1080/1206212X.2019.1707436.

Zhao:2019:RAW

[ZY19b] Lei Zhao and Huang Yu.

Retracted article: A wireless network remote monitoring method driven by artificial intelligence. International Journal of Computers and Applications, 45 (2):115–123, 2019. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1080/1206212X.2019.1710664.

Zheng:2003:CSH

[ZYM03] S. Q. Zheng, M. Yang, and F. Masetti. Constructing schedulers for highspeed, high-capacity switches/ routers. International Journal of Computers and Applications, 25(4):264–271, 2003. ISSN 1206-212X (print), 1925-7074 (electronic). URL

https://www.tandfonline.
com/doi/full/10.1080/1206212X.
2003.11441714.

Zhiwei:2013:RSA

[ZYS13] Liang Zhiwei, Chen Yanyan, and Zhu Songhao. Resource service access mechanism for service robots. International Journal of Computers and Applications, 35 (1):51–57, 2013. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL https://www.
                                        tandfonline.com/doi/full/
                                        10.2316/Journal.202.2013.
                                        1.202-3641.

Zhu:2020:RVI

[ZZ20] Zhu Zhu and Xiaoguo Zhang. A random-valued impulse noise removal algorithm via just noticeable difference threshold detector and weighted variation method. International Journal of Computers and Applications, 44 (2):187–200, 2020. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/1206212X.2020.1719309.